A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

20 May 2025 Version 2.00

Title word cross-reference

$104.00 [Ryp13]. $105.00 [Cha13]. $149 [Pet15]. $167.99 [Fle16]. $189 [Pet15]. $209.95 [Fle16]. $25 [LDA^{+}13]. $35.00 [LDA^{+}13, Smi13]. $40.00 [Pra14]. $48 [Zyd11]. $55.30 [Chi13, Chr14, Nel14a]. $64.56 [HCD^{+}14]. $69 [Zyd11]. $72.80 [Ryp13]. $75 [Gid14]. $79.00 [Chi13, Chr14, Nel14a]. $85.00 [Var13]. $99.95 [Qui16, Shi13]. ^{+} [SZZ10]. ^{13} [GDPW10]. ^{18} [GDPW10]. $A130.00 [Ber13]. δ [GDPW10]. S^{4} [Dud19].

-ATPase [SZZ10].

. [Zyd11].

1* [HRR^{+}11]. 11-Ketotestosterone [MGK^{+}18]. 15-Year [DCJ^{+}13].

2010 [Yas16]. 20E [BTWB17]. 222 [BLG^{+}13, PET^{+}17].

3-Trifluoromethyl-4-Nitrophenol [CHB^{+}12]. 30-Year [CAB11].

8-mm [CHR^{+}17].

Abalone [BRS^{+}12]. Abandonment [LS10b]. Abilities [FCG^{+}17]. Ability [FMJ11, FWSW^{+}15, ZZD10]. Abiotic [DHA^{+}17]. Above [HP14, BLB^{+}13]. Above-Optimal [HP14]. Absorbable [JLA17]. Abundance [ABS^{+}11, Arc16, BPS11, BAKT12, BS15, BRS^{+}12, BB10b, CKC^{+}12, CYW^{+}19, FSP^{+}17, HPF^{+}16, HAG13, HW14, HJS^{+}16, HEW12, HMP18, KYK17, MAS10, MBG10, MHE12, MFH13, MOS13, MCO^{+}10, OI16, PT10, PB13, PPJ15, PMN19, PGS13, PDM^{+}14, PA10, RH14, RSB14, RSL11, RGM^{+}11, SP10, SWB^{+}16, TA13, WBMG10, WS15, ZBHW16, ZKB^{+}15]. Abundances [Han11]. Accelerating [EGAW12]. Access [AJFO10, AJF13a, AJF13b]. Acclimation [BRGR10]. Accounting [HRG12]. Accumulation [HSG10]. Accuracy [AHS11a, BFM10, HSF12, MM10, TL14]. Accurate [BSSP18]. Accurately [DYWY17, MGK^{+}18]. Acid [HWH10]. Acidification [BGLP19]. Acids [YSW^{+}14]. Acipenser [FCG^{+}17]. Acknowledgments [Ano18g, Ano19h]. Acoustic [Cha13, CHW^{+}13, DBC^{+}10, DBB^{+}12, EPT^{+}16, GZB12, HJM^{+}12, JDS15, KHF^{+}18, LBD^{+}16, MCSF14, MGW13, PSH^{+}19, REN17, RPVS15, SNP^{+}10, SFF14, SJFF15, SHH^{+}19]. Acoustic-Tagged [HJM^{+}12, REN17]. Acoustically [NBC15]. Acoustically-Tagged [NBC15]. Acoustics [DMR^{+}10]. Across [FFYI16, RPC14, SMM13, BBM^{+}14, CKM^{+}11, CJG13, DA16, FBF10, FWD10, FDWG13, FA10, PPH15, PDH^{+}18, WDW11]. Acrylic [BN16]. Act [WHSK19, ACDL19]. Action [BC12]. Actions [WCYH18]. Actively [ACMM^{+}15]. Activities [SJFF15, WRQH17]. Activity [BBZGB15, BMM10, BNW^{+}12, GZB12, MDP12, SB12a, SZZ10, WB15a]. acuta [SPCJ11]. Acute [MFAS^{+}12, SDOK11]. Adaptation [DZW^{+}12, TNOU11]. Adaptations [Fre19]. Adaptive [MHN11]. Additions [MWS10]. Additive [YJW11]. Addressing [TLM^{+}19]. Adfluvial [ACDL19, BP13, HSM12, IRJ11, JJNB17]. Adirondack [BGLP19, BSGB17, JK17, RJWK10]. Adjacent [TT17]. Admixture [CCKF11]. Adult [ASJ17, BHC^{+}18, BHS^{+}17, BWD^{+}17, CR10, CSDW10, CJL^{+}14, CTS^{+}15, CVB14, DCS^{+}17, DEP^{+}18, DH11, FSJ^{+}11, GCM12, GKG^{+}19, HKKS14, HN11, HBA^{+}15, JCK^{+}12, KST^{+}12, KCJ^{+}15, KHD^{+}14, LUC^{+}15, PM14, RSB14, RHR^{+}15, SGP16, Str10, Str12, TA13, TCBI11, VFB^{+}16, YSC10, YHWH10]. Adulthood [AHS10, EHC^{+}14, FBWZ19]. Adults [JCS14, KPS^{+}19, PPP^{+}13]. Advances [Chi13]. Advancing [Chr14]. Advantages [ACB^{+}14, DWE^{+}19]. Aerial [DFB15]. Affect [BVA11, FBF10, MDT10, NPW12, SW11, TL15, WHP14, ZSVK13]. Affected [CEFQ11]. Affecting [CT12, KLR^{+}15, OTR^{+}14, RSVAB19, RBH^{+}12]. Affects [JDS15, RFM^{+}15, SAF^{+}19]. AFS [Zyd11]. after [BAKT12, DSHG10, DEW^{+}14, HGB^{+}16, HCZ13, HS13, MBG10, SRP^{+}18, WCMP13]. Age [ASP^{+}18, AS17, BMGL12, BBF^{+}18, BMH19, BA13, BBS12, BSR12, BSSP18, CFHE11, CGA13, DLHF16, DL13, DMHO12, GN13, HNK^{+}12, HS17, HBS12, Hoe17, HKO^{+}18, HWHF14, JRBC13, JPHK10, KGW14, KSSK^{+}17, KQ11, KK16, KDS18, KKM11, LS18, Mat18, MM10, MGK^{+}18, Mic10, MFH13, MD11, MBPP18, MLC13, MS12, NAWM14, NTHR18, OWN11, OWP13, PW10b, PDH^{+}18, PA10, RL14, RMMGPB16, RDH11, SR10, SMO^{+}14, SDW^{+}17a, SW11, SS11, SB12b, SPR^{+}16, SSM16, SMH^{+}11, SKBS18, SHP^{+}19, SAW^{+}17, TRMC15, TP11, TA12, UWL12a, WDW11, ZZMC13]. Age-0 [BBF^{+}18, HBS12, HWHF14, JRBC13, KGW14, KKM11, LS18, Mic10, MFH13, MD11, MS12, NTHR18, PW10b, PA10, RDH11, SDW^{+}17a, SW11, SS11, SPR^{+}16, SMH^{+}11, WDW11]. Age-1 [MGK^{+}18]. Age-2 [MGK^{+}18]. Age-Selective [KK16]. Age-Specific [BA13, RL14, TP11]. Age-Structured [BBS12, SR10]. Ages [GBB^{+}14]. Aging [LSJN13, LS10a]. Air [CCB16]. Alabama [CHPG12, SYH^{+}15, YITI12]. Alarm [LRM^{+}10]. Alaska [GSA^{+}19, Yas16, AHS11a, BBC^{+}14a, BBC^{+}14b, BDM^{+}10, BCS19, CTS^{+}15, CHS^{+}11, DSHG10, DF11, DHRQ10, EBG18, FSN^{+}13, FRM18, FFYI16, KQ11, MHF12, MSS^{+}11, NRZ^{+}13, OWP^{+}17, WTB10]. Alaskan [BES15, MWS10]. Albert [Var13]. Aleutian [GRS12, Spi12]. Alevins [RB13]. Alewife [CBMR14, EBA^{+}19, MJS^{+}19, MCO^{+}10, SPR^{+}16, SRP^{+}18]. Alewives [FHCSN12, GGS10, GDG^{+}15, RBK^{+}11]. Alga [VFFP13]. Algorithm [MDP12]. Alkaline [BRGR10]. Alligator [BSR12, BSD13, DSM18, FDSB18, LS16, MSD18, RAB^{+}18, SDSB18]. Allometric [CTN11]. Alluvial [AM13, BHP^{+}15]. Along [ZHWR11, BT16, FWSW^{+}15, GRS12, HT13, RSL11, SGP16, TT17, WB15a, WBMG10, YYA^{+}16]. Also [Hoe17]. Altamaha [PB13, SP10]. Alter [CS16, LUC^{+}15]. Alteration [FPD12, SWS10]. Altered [DSRA12, LSSYL10]. Alternate [JDS15]. Alternative [BVA11, MNM^{+}17, NYCC19, RKEB13b]. Ambient [LRM^{+}10]. America [BDTD17, BWDM11, GMB^{+}19, PSGB13, Wei10, WBW^{+}13]. American [Chi13, Chr14, Gid14, HKO^{+}18, LDA^{+}13, Nel14a, Ryp13, SRMB^{+}15, Zyd11, ACK17, Ano13b, Ano14b, Ano16, Ano17b, Ano19i, BJA^{+}13, CPS^{+}15, CTP^{+}12, CEM^{+}19, DMCO11, DMHO12, EWSR16, FSW10, FBB^{+}13, GPBA18, GBZ14, HCR14, HTR13, HBB17, HHES13, HEW12, HMO14, MLC13, RH14, RLL^{+}19, SJGL16, SH12, SS15b, SAH^{+}14, WL13, ZZMC13]. amidst [KGC^{+}12]. Ammocoetes [HB17]. Among [HN13, PF11a, BLCN^{+}10, BSS18, BGB^{+}16, BEMF11, CN11, CR10, CGB19, CJG13, EFC^{+}11, FBWZ19, FBS^{+}14, HR15, HLB14, HRR^{+}11, HBTT11, HSE^{+}12, KDS18, LEM^{+}11, MHN11, MSWS10, MFAS^{+}12, MWSS13, MLBS^{+}13, PPF^{+}13, PHM17, PTPS10, PL13, SCC17, SF15, WSI14, WSQ^{+}15, WKS^{+}19, WCQD18, WLW16]. Among-Pass [PF11a]. Among-Stream [HN13]. Amplifier [KM10]. Anadromous [BBC14c, BWRA16, BCS19, GGS10, GDG^{+}15, GDST16, HCZ13, HS13, HMO14, MC14, PBW16, RHW15, SHM^{+}12, SR15, SHC^{+}16, SHCB16].

Anadromy [HGB^{+}16, NZF^{+}11, WAR12]. Anaerobic [SB12a]. Analog [MWS10]. Analogs [KPZ^{+}12]. Analyses [AHS11a, BL11, DWE^{+}18, DSRA12, WMW^{+}12, WMG^{+}14]. Analysis [BMP^{+}17, DCJ^{+}13, DYWY17, EPT^{+}16, FSN^{+}13, GvdHBC12, GKBG12, GRS12, GNV^{+}14, GWG^{+}19, HH11, HHJH17, HWH10, HDH16, HCNL10, KHHG13, KPZ^{+}12, MSW^{+}10, MVK12, PZS17, RB12, RHKD12, RMMGPB16, RBK^{+}11, SPM^{+}17, SBC^{+}16, VSC15, WB13b, ZM10]. Analytical [PSM12]. Analyzed [GSVS^{+}10]. Analyzing [RPSB14]. Ancestry [MMY12]. Ancient [DKS18]. Andrew [Chi13]. Anesthesia [CJL^{+}14]. Anesthetic [VSBB^{+}14]. Anesthetics [BLG^{+}13]. Angled [HBC15]. Angler [CML19, LSV^{+}13, RSBF14, SC13]. Angler-Caught [RSBF14]. Angling [ARC^{+}13, BDS10, BNW^{+}12, DKS18, HVD^{+}16, LUC^{+}15, PWP13, THL^{+}17, VWS13]. Angling-Induced [THL^{+}17]. Angling-Sized [BDS10]. anguillicaudatus [KU12, UK14]. Annala [Chr14]. Annual [CR10, GDST16, KHW^{+}14, MKH10, MBL^{+}18, PGWS15, REN17, SR10, ZKH12]. Annuli [DMCO11]. Anthropogenic [BDTD17, TW14]. Antipredator [SS15b]. Apache [PB17, RB13, RBF14]. Apalachicola [YITI12, SYH^{+}15]. Aperture [BDM^{+}10, TFM^{+}18]. Apex [HB18, KBG^{+}18, SGK16]. Appalachian [BMP^{+}17, FK18, FKF^{+}19a, FKF^{+}19b, KP14, PHHM12, TNS^{+}14, WK13, WCD^{+}11]. Apparent [HMSG^{+}11, PGWS15, WCQD18]. Applicable [DMCO11]. Application [BBG^{+}18, CVL^{+}19, CYW^{+}19, FSN^{+}13, PPF^{+}13, VSC15]. Applications [HS17, MH11]. Approach [ACRA10, Chr14, Gre12, GBF12, HJS^{+}16, JDS15, KP14, KRM13, OGM^{+}15, PSM12, THGA15, TNS^{+}14, WRP^{+}12]. Approaches [TBGM12]. Aquaculture [CEM^{+}19, DHB15]. Aquaculture-Based [CEM^{+}19]. Aquatic [HWH10, KBG^{+}18, MWS10, NTHR18]. AQUI [BTWB17]. AQUI-S [BTWB17]. Arancibia [Fle16]. Arber [Chr14]. Archipelago [BRS^{+}12, KGC^{+}12]. Archival [NB12, RMDA11]. Archived [OPS^{+}14]. Arctic [BCS19, BS13, BDM^{+}10, CHH^{+}16, CPP^{+}14, DHA^{+}17, DMW^{+}19, GDST16, RPC14, SS18b, TFM^{+}18, WRP^{+}12]. Area [EBG18, GWG^{+}19, HGS^{+}16, TBL^{+}17]. Areas [ASJ17, BDM^{+}10, FM17]. Arizona [DSAM12, ESD^{+}17, CYP11, SRMB^{+}14, Sto10]. Arkansas [FM17, GBF12, SJ11]. Aromatic [MHPB14]. Arrival [BA13, MM17, RCC^{+}13]. Arrowtooth [MHF12]. Arroyo [BMOF16]. Artificial [AIC13, DSB^{+}18, EWW13, PS14, SKP^{+}10, SHR^{+}12, SAW^{+}17, WZCT12]. Asian [CCL^{+}19]. Assay [CHO^{+}15, LSD^{+}14]. Assays [BSSM12, DPS14, GMB^{+}19]. Assemblage [BWHF17, BR11, BKC^{+}19, CT16, FQW^{+}10, HCD10, JZBP16, KGC^{+}12, KCZH14, KP14, Lan16, MRPW19, MSK15, MKK15, MBG^{+}11, PTWK11, RBZ^{+}17, WSQ^{+}15, WB15b]. Assemblages [AM13, BEWM10, BP10, CHPG12, DSBS14, DM11a, GBW16, GWG^{+}19, IB14, KWW14, MP10, MKH13, RW11, SFB16, SGB13, THNG14, TF13, WDR16, WCZ^{+}18, WK13, WBWSK12, vVO10]. Assess [BES15, GvdHBC12, JPFB13, KMKS13, MLS^{+}13, RAB^{+}18]. Assessed [SJFF15]. Assessing [ARC^{+}13, CCH15, Gre12, HMTC12, HARL18, JRBC13, Lis14, Lis15, MCH^{+}13, NDC^{+}14, PWP13, RNG^{+}13, WK13]. Assessment [BBAK19, BSCA17, BBS12, CGDL18, DHA^{+}17, DBC^{+}10, FSC^{+}13, HSS17, HCW14, HLH14, HKZ11, KCZH14, KVA^{+}19, LSJN13, MCW^{+}14, MT19, OSA^{+}17, RL14, SJMG14]. Assessments [LDP19, RJ11]. Assigning [GBB^{+}14]. Assist [LFB^{+}15]. Assistance [RHR^{+}15]. Associated [ARB^{+}18, ACRAM11, BNC^{+}15, BGV11, CM11b, DSBS14, GPBA18, GWG^{+}19, HHJH17, JC16, NZF^{+}11, VJGV11]. Association [CBW^{+}16, DHB15]. Associations [BQHR17, CAB11, DGMC17, DWGF11, FBWZ19, HGSG17, Lan16, SL16]. Asymmetric [WDDS13]. At-Risk [FBWZ19]. Atchafalaya [MKK15]. Atlantic [AKM12, DG19, GN13, Han18, JZBP16, NG13, NAWM14, RBE^{+}10, ASJ15, ASJ16, ASJ17, AHS10, AS17, BMF^{+}15, BP16, BHLK10, BMGL12, BGV^{+}12, BLG^{+}13, BK11, BMC^{+}13, DCS^{+}17, FZZ^{+}10, FBP11, FP19, FGW^{+}16, GvdHBC12, GCA^{+}18, Goo15, HKW^{+}14, HPF^{+}16, Han11, HCR14, HSS17, HKZ11, HMSG^{+}11, HWN15, HWHF14, KHW^{+}14, KSSK^{+}17, LKS^{+}12, LMPvdH18, LRM^{+}10, LUC^{+}15, LKP^{+}13, LSJN13, MSK15, MSP12, NCW^{+}17, PDSP^{+}11, RSG12, RJ11, SMJ14, SF15, SP10, SWB^{+}16, SFH15, SZZ10, TL17, TLKT17, VJGV11, VHPP19, VSC15, WNCF10, WCZT11, WZCT12, WS15, WMW^{+}12, ZHWR11, ZZD10, dHST13, vZNW12]. ATPase [SZZ10]. Attached [JBC^{+}12]. Attachment [JDS15]. Attacks [SS15a]. Attention [LMPvdH18]. Augmentation [CCKF11, HBTT11]. Australia [Ber13]. Australian [Ber13, HBC15]. Authors [Ano11b, Ano12b, Ano13b, Ano14b, Ano16, Ano17b, Ano19i, SRMB^{+}15]. Autosomal [MSWD15]. Availability [HSSS16, KGW14, KM14, MBH12, PW10a, RSL11, TES14, YCB^{+}11]. Available [Yas16]. Avian [EHR^{+}12, EPT^{+}16, EPC^{+}19, HERC12, HEC^{+}15, KKB^{+}10]. Avoid [AJFO10, AJF13a, HD16]. Avoidance [JBC^{+}12, SPH19].

B [LDA^{+}13]. Back [BSM13, CC16, PM13]. Back-Calculated [CC16]. Back-Calculation [PM13]. Back-Calculations [BSM13]. Backpack [PVJH15]. Bad [RKEB13b]. Baits [HBC15]. Baja [FRRCGA^{+}15, TRHHOGSJ14]. Balance [JPG^{+}18, TJB^{+}14]. Balancing [MDP12]. Baltimore [Qui16]. Bank [RBZ^{+}17, JPLS13, ZC13]. Banking [BMTP13]. Barcoding [PCFH17]. Barge [BWD^{+}17]. Barker [CBSB15]. Barotrauma [BS16]. Barracuda [BNS^{+}14]. Barrier [BWRA16, MCW^{+}14, WCMP13, WAR12]. Barriers [WZC^{+}14, WCN^{+}19]. Base [MP10, OGM^{+}15]. Based [CJL^{+}14, CEM^{+}19, Dud19, EHR^{+}12, EPT^{+}16, FBH^{+}13, FA10, GKBG12, HSA^{+}16, Hoe17, KR18, NCW^{+}17, PPH15, PMN19, RSB14, RKEB13a, RKEB13b, SDBB12, SDSB18, SL16, TL17, THGA15, UNTA10, WR14, WBWSK12, ZM10, WCN^{+}19, vPWT12, WRP^{+}12]. Baseline [RKEB13a]. Baselines [RKEB13b]. Basin [BDTD17, BAHP11, BHD^{+}16, BCF^{+}11, BBS12, CM11a, EPS^{+}15, EMS^{+}14, HGBE16, HHES13, KPZ^{+}12, KCMS11, MKK15, MFR15, OTR^{+}14, PTPS10, PDP12, PPC13, PPMJ14, PPW^{+}14, PPJ15, RL19, SK11, WDHW14, WMW^{+}12, ZBW10, ACRAM11, DWGF11, DWG14, EBF^{+}14, FWBT17, HLB14, HCW14, HN11, JRP12, JAA^{+}13, KBC^{+}18, KK16, KHHG13, MHN11, MWG18, NHM10, PGJ11, RDR^{+}10, SYH^{+}15, SJ11, SMY11, SS13, Str10, Str12, SWP^{+}18, WDR16, WAW11, WM12]. Basin-Scale [HGBE16]. Basins [KMKS13, VFFP13, ACRA10]. Bass [ARC^{+}11, ARC^{+}13, BLCN^{+}10, BLCB13, BN16, BNW^{+}12, BB10b, BAR15, CSDW10, CBW^{+}16, CDB10, CCB16, DF12, DA12, DMHM^{+}18, EFC^{+}11, EPW15, ETC18, ER18, FDWG13, FRBA13, GWB^{+}15, GDWD10, GA10, HHJH17, HBS12, HVD^{+}16, HDF12, HPHB10, KGC^{+}12, KTE^{+}14, KAH16, KHD^{+}14, LSJN13, LAR16, LV14, Lov11, LS10b, LCHB^{+}18, MS13, MFH13, NDC^{+}14, Nel14a, NLF13, NTHR18, NDW10, PW10a, PWP13, PMF^{+}10, PET^{+}17, RHKD12, RO16, SPC^{+}18, SA14, SA16, SM13a, SAS17, SH12, SMH^{+}11, SS15b, TRW10, TPA14, TPLA16, TAM15, THL^{+}17, TBGM12, VWS13, WCZT11, WZCT12, WSK11, Nel14a]. Basses [TA12]. Batch [GDG^{+}15]. Bathythermal [BAKT12]. Bay [ARL^{+}12, DHRQ10, FSW10, GCM12, Gre12, IHTF11, JM14, RSG12, TT17, WLLCF16, WMW^{+}12, BAHP11, BL11, CMM^{+}17, HWHF14, Lov11, LN18, OSA^{+}17, OGM^{+}15, RL14, SMJ14, TLKT17, TF13, WF11, WLW16, WSK11]. Bayesian [DL13, SL16]. Bays [DSM18]. Beach [TT17, WF11]. Bear [CHT13]. Beat [MBBM10]. Beaver [LRW13, KGC^{+}12]. Bed [FSFS17, JPSM12, MKH10, MLB18]. Beds [DSBS14, FWSW^{+}15]. Behavior [BGW^{+}17, BNS^{+}14, CEFQ11, DMR^{+}10, EWW13, EGAW12, FK18, FKF^{+}19b, KCJ^{+}15, LUC^{+}15, MCSF14, MBT10, MCH^{+}13, PDP12, RH14, RSG12, RGP17, SKP^{+}10, SS15b, SHP^{+}19, SPH19, TBL^{+}17, TPLA16, TKC^{+}18, WEW12, YSC10]. Behavioral [BCF^{+}13, BSCA17, GWB^{+}15, KBG^{+}18, LRM^{+}10, NPW12, PET^{+}17, SAF^{+}19, SHH^{+}19]. Behaviorally [MCP^{+}15]. Behaviors [CW11]. Behind [PJ11]. below [BLB^{+}13, WAR12]. Benefit [Lan16]. Benefits [DCBKR14, HWD14]. Benthic [HLB^{+}16, KS11, PSH^{+}19, SS11, SAH^{+}14, WDWC10]. Benthivorous [MLBS^{+}13]. Bering [BBL^{+}12, SSH^{+}11a, SS15a, SS18b, BCP^{+}11, BD15, GRS12, HSM^{+}10, PBW16, Spi12]. Berkeley [Var13]. Best [RPVS15]. Bethesda [Chi13, Chr14, LDA^{+}13, Nel14a, Ryp13, Zyd11]. Better [HUMG^{+}18, TP11]. Between [AJF13a, Ano11a, BP13, BWRA16, BGOQ13, CBVF15, CHO^{+}15, CHLN15, CM11a, DWN^{+}15, EHC^{+}14, FRBA13, Han18, HPHB10, JPLS13, KGC^{+}12, KRRW18, LBC10, LHHL17, MJS^{+}19, MOS13, NMH12, OPS^{+}14, PB16, QSJ12, RHW15, RZW^{+}12, RNG^{+}13, RMW16, RB18, SB12a, SRMB^{+}14, SPR^{+}16, SE11, STK12, TC11, WDDS13]. Between-Group [AJF13a]. Bias [AJFO10, ELC13, KDS18, LSJN13, RL14]. Biased [MJGV14, SDBB12]. Big [PDM^{+}14, DHA^{+}17]. Bighead [GGC17, HGSG17, LYX^{+}10, LLWW19, SPC^{+}18, ZRM^{+}16]. Bigheaded [DAA^{+}16]. Bile [YSW^{+}14]. Binomial [IWB^{+}13]. Bioassessments [FA10]. Bioclimatic [CM11b]. Biodiversity [BNS^{+}16, SJ11]. Bioelectric [GvdHBC12]. Bioelectrical [DYWY17, HH11, HHJH17, VSC15]. Bioenergetic [AMBM12]. Bioenergetics [CORM13, Har17, HD16, HBHS15, JMM^{+}11, KWS14, MDP12, MWC^{+}13, MBBM10, PM15, SSS^{+}12, UNTA10, WvPC12, WBH10, ZLHS10, vPWT12]. Biogeographic [MSBT10]. Biogeography [Var13]. Biological [RJH12, WAFL10]. Biologically [BNTT17]. Biology [ABM17, BCS19, BCMPDSZ10, CABNQ^{+}13, CTG14, Fle16, GSR^{+}11, Gol18, HL10, HAV^{+}19, KSSK^{+}17, MWSS13, Mil19, NG13, Pag19, Pra14, Shi13, Nel14a]. Biomass [MCO^{+}10, PT10, STZ13, ZM10]. Biomass-Based [ZM10]. Biometric [MSD18]. Biotelemetry [DMR^{+}10]. Biotic [MAK12, PKP^{+}17]. Black [GMB^{+}19, KAH16, MS13, SM13a]. Blackfoot [PPW^{+}14, EPS^{+}15, PDP12, PPC13, PPMJ14, PPJ15]. Blackgill [LF15]. Blackspotted [GMP^{+}11]. Blackwell [Fle16]. Bloater [CBMR14, EBA^{+}19, PB16, PRK^{+}16]. Bloaters [BMR^{+}12, PBM^{+}12]. Blood [BS16]. Blooms [JC16]. Blue [BT16, CTN11, DDMR10, DK18, Goo15, HAR18, OWS16, RL14, RBE^{+}10, SO15]. Blueback [MHHH10]. Bluefin [AGH18, Gol18]. Bluefish [MJGV14, MBL12, TT17]. Bluegill [CHLN15, MFH13]. Bluegills [OWN11, OWP13, PM10, SW11, SBW14]. Bluehead [FWBT17, WvPC12]. Bluntnose [WAFL10]. Boats [Lev18]. Boca [Shi13]. Bodied [AUDL18, Cla16, HW14, SRMB^{+}14, SFF14]. Bodies [LBBW17]. Body [Ano11a, BS13, CTN11, GvdHBC12, KM14, KHF^{+}18, Mic10, PW10b, PKL^{+}11, RR13, SW11, TRLG13, TB16, TC11, VSBC13]. Bomb [GDWD10, MLC13]. Bonneville [BDTD17, CR10, HK12, KST^{+}12]. Bonytails [OSDT18]. Book [Ano10a, Ano10b, Fle16, Nel14a, Qui16, Yas16, Pet15]. Boom [KYK17]. Boom-and-Bust [KYK17]. Bootstrapping [SMN^{+}10]. Borderlands [BB15]. Boreal [BDS10, BSHK^{+}11, LSPB11, TPBB14]. Borrow [MKH13]. Borrow-Pit [MKH13]. Both [CPS^{+}15]. Bottleneck [KWCI10, KCM14]. Bottom [KS11]. Bottomless [RC16]. Bottorff [Yas16]. Bradford [Chi13]. Branchiostegal [BSSP18]. Brassy [FBF10]. Brazilian [CBW^{+}16]. Brazos [RMW16, VFFP13]. Breder [RC16]. Breeders [BSS18]. Breeding [JCS14, MGWP10, SKP^{+}10]. Bristol [DHRQ10]. British [BNS^{+}16, MPHG10, NBC15, TA13, WAKH12, Bea10, BBC^{+}14a, BBC^{+}14b, MKH10, PGR13]. Broad [BDM^{+}10, HLR^{+}12]. Brood [LS10b]. Brook [AGKW12, BSGB17, BGLP19, BBAK19, BB10a, BMP^{+}17, DW16, DW15, HH11, HCNL10, JK17, KLR^{+}15, KHK^{+}15, KMS^{+}18, KRRW18, MCW^{+}14, MFAS^{+}12, Mit16, MWKS13, NB13, PHHM12, RMM11, RJWK10, SDW17b, TPBB14, TNS^{+}14, WDDS13, WKS^{+}19, WCD^{+}11]. Brown [AMM^{+}16, DW16, DH11, FWB^{+}15, Har17, JC16, JLA17, KRRW18, KRM13, Mil19, Mit16, OS11, SHR^{+}12, SDW17b, WDDS13, WMS15, WBH10, YHWH10]. Brunswick [UWL12a, UPT^{+}13, DCS^{+}17, TL17, UWL12b]. Budget [MDP12]. Buffers [FQW^{+}10]. Bulak [Nel14a]. Bulkley [WAKH12]. Bull [ACMM^{+}15, ACDL19, ADS^{+}11, BP13, BHD^{+}16, CAB11, FH17, HCS^{+}16, LB15, MWC^{+}13, NAWM14, PPP^{+}13, SK11, WBR11]. Bullfrogs [ACK17]. Bullhead [Har17]. Bumpers [GCA^{+}18]. Bundle [MWS10]. Buoyancy [PBC12]. Buoyant [JBC^{+}12]. Burbot [AIC13, BJA^{+}13, BQHR17, CVL^{+}19, CJGH13, CJG13, GRDO13, GS13, HP13, HUMG^{+}18, JWHJ13, KSB16, SM13b, SWC13, SNI^{+}13, TA13, TASC17]. Burrhead [PSGB13]. Bust [KYK17]. Butte [CNM^{+}18]. Butterfly [BN16]. Bypass [SO12, HSG10].

C [Chi13, LDA^{+}13, Nel14a, Shi13, GDPW10, VSBC13]. Cage [JKP10]. Cairns [Zyd11]. Calcified [KDS18]. CalCOFI [Pet15]. Calculate [SMN^{+}10]. Calculated [CC16, STZ13]. Calculation [PM13]. Calculations [BSM13]. California [Gid14, Pet15, TRHHOGSJ14, Var13, ARL^{+}12, BSC^{+}10, BGB^{+}16, BHP^{+}15, CT16, FSP^{+}17, FRRCGA^{+}15, GLM^{+}15, GGHS^{+}14, GHPS^{+}16, GNMM12, JM14, LHHL17, LSSYL10, MSD^{+}19, MDT10, NLF13, NR16, PHM17, RG16, RL19, SMO^{+}14, SCC17, SMS^{+}12, TVL^{+}12]. californiensis [MFR15]. Callinectes [DDMR10, DK18]. Calorimetry [GDWD10]. Cambridge [Cha13]. Camera [GWG^{+}19]. Campeche [JPLS13]. Can [BMF^{+}15, CGL^{+}11, FH17, HKR18, HGS^{+}16, MCP^{+}15]. Canada [BNS^{+}16, CHH^{+}16, Han18, JKP10, LMPvdH18, NBC15, UWL12a, UPT^{+}13, WMW^{+}12]. Canadian [BDS10, BCP^{+}11, DF11, WCYH18]. Canadian-Origin [DF11]. Canal [CKFQ11, GBW16, Esh14, MBT10]. Canals [SO12]. Candidate [NZF^{+}11]. Candy [DA16]. Canning [BCS19]. Canyon [CYP11, SRMB^{+}14, WvPC12, YCB^{+}11]. Capability [CJM^{+}15]. Cape [DCBKR14, SH12]. Captive [FS16, HL10, RNG^{+}13, SVKK18]. Captive-Reared [SVKK18]. Captivity [JPG^{+}18]. Capture [BMC^{+}13, HLK11, LSV^{+}13, MSH^{+}14, RCC^{+}13, RHR^{+}15, TL17]. Captured [DCS^{+}17]. Carbon [DAA^{+}16, QSJ12, SPH19]. Carcass [KPZ^{+}12, MWS10, RSB14]. Carcasses [WHC^{+}10]. Caribbean [HAV^{+}19]. Carmel [Smi13]. Carolina [RH14, THNG14, AAD14, BRA14, CTP^{+}12, CEM^{+}19, DDMR10, DK18, FBH^{+}13, HARL18, MKA10, MWT14, MVP15, RLL^{+}19, SH12, SFH15]. Carp [BBZGB15, BW10, GGC17, GMB^{+}19, HGSG17, LBBW17, LYX^{+}10, LLWW19, SS13, SWP^{+}18, WB13a, WB13b, ZRM^{+}16]. Carps [DAA^{+}16, SPC^{+}18]. Carrier [Shi13]. Carryover [AW19]. Cascade [THKG10]. Case [KCM14, MRPW19, PGW^{+}11, RJ11]. caspia [ARL^{+}12]. Caspian [ARL^{+}12, MGWP10, SCL^{+}13]. Casselman [Zyd11]. Catalina [LSSYL10]. Catch [AJFO10, AJF10, AJF13a, AJF13b, CML19, FM17, Goo15, HAG13, HJS^{+}16, HGP^{+}17, LUC^{+}15, Sto10, TP11, WAFL10, YJW11]. Catch-and-Release [FM17, LUC^{+}15]. Catch-Rate [YJW11]. Catchability [HJS^{+}16, KS11, PTPS10, RSVAB19]. Catchable [MHE12]. Catchable-Sized [MHE12]. Catches [CM11a, VIW^{+}17]. Catchment [YYA^{+}16]. catesbeianus [ACK17]. Catfish [BK11, FBP11, LS10a, MSB^{+}18, MBEvdB^{+}10, SO15, SSR^{+}17]. Cating [DMCO11]. Catostomid [Lan16]. Catostomids [BBF^{+}18, BFH14, SFF14]. Catskill [BEWM10, EBMV10]. Cattaraugus [HDBS14]. Caught [DDMR10, RSBF14]. Cause [Mit16]. Causes [FSRK11, TLM^{+}19]. Cautionary [ACWIK13]. Cedar [BLB^{+}13, TW14]. Census [MLS^{+}13]. Central [CM11b, DDMR10, NRZ^{+}13, PTWK11, RHKD12, SCC17, SHR^{+}12, SMS^{+}12, TVL^{+}12, DG14, JPHK10, MBPP18, PDH^{+}18, RL19, SDW17b]. Centrarchid [AM13, MKK15]. Cephalopod [PM19]. Ceratomyxosis [FMG^{+}11]. Ceratonova [CRC^{+}16]. cerebralis [FWB^{+}15, KKB^{+}10, PDP12]. cerebralis-Resistant [FWB^{+}15]. Cetaceans [Cha13]. CFT [ACK17]. Chain [BSC^{+}16, HRHS16]. Challenge [SKBS18]. Challenged [BRGR10]. Challenges [ACB^{+}14]. Champlain [Esh14, HML13, SPR^{+}16, SRP^{+}18]. Change [BAKT12, FSFS17, FBWT19, LSSYL10, MDT10, NB12, RA10, SNC^{+}10, SRP^{+}18, SWS10, WHSK19, ZTC12]. Changes [BCRP13, BS15, CTPM13, HSSS16, HK12, KGC^{+}12, KCS^{+}19, KR18, PT10, PRK^{+}16, RSL11, RSBF14, SWC13, SWS10, TJB^{+}14]. Changing [EBA^{+}19, Har17, JMB^{+}13, Mit16, SDOK19]. Channel [BEWM10, EBMV10, FSFS17, LS10a, PPJ15, SSR^{+}17, WSQ^{+}15]. Channels [MC14, RBZ^{+}17, WHC^{+}10]. Char [CPP^{+}14, GDST16, SDK^{+}10]. Character [FGW^{+}16]. Characteristics [CKC^{+}12, CJG13, DCS^{+}17, EPW15, HDH16, HERC12, MKH10, MP10, MK15, PSM12, PGW^{+}11, PGVG15, SDW15, SMS^{+}12, VB14, WSI14, WQD14, WAFL10]. Characterization [CTP^{+}12, KCJ^{+}15, MJHP18, WBM^{+}17]. Characterize [BDM^{+}10, DFB15, FWSW^{+}15, LMD^{+}17, WRP^{+}12]. Characterizing [ACWIK13, BES15, CHS^{+}11, GZB12, HCD10, RJ11]. Charles [Nel14a]. Chattahoochee [SYH^{+}15]. Chemical [LRM^{+}10, NYCC19, TBGM12, TL14, ZSVK13]. Chemistry [BBG^{+}18, CBVF15, CC16, NW14, SMJ14, SYH^{+}15, SBC17]. Chesapeake [BL11, WF11, CMM^{+}17, FSW10, HWHF14, Lov11, LN18, OSA^{+}17, OGM^{+}15, RL14, SMJ14, TF13, WSK11]. Chicago [Pra14, Smi13, LLWW19]. Chichester [Fle16]. Chignik [CHS^{+}11]. Chinese [FCG^{+}17]. Chinook [Ano11a, Bea10, BHS^{+}17, BS18, BNC^{+}15, BVA11, BBM^{+}14, BWD^{+}17, BRK^{+}18, BDBE11, BBS12, BEMF11, BCG^{+}12, BLB^{+}13, CJL^{+}14, CHO^{+}15, CEFQ11, CKFQ11, CFHE11, CC16, CTS^{+}15, CTPM13, CVB14, CNM^{+}18, CTK^{+}17, DF11, DEW^{+}14, DWE^{+}18, DWE^{+}19, DBC^{+}10, DML^{+}10, DBS^{+}10, EMS^{+}14, ETC18, FJP18, FWT^{+}14, FPM15, FMS10, FMG^{+}11, GCG^{+}18, GDM^{+}10, GDM^{+}11, GJC^{+}18, GA17, GSA^{+}19, HMTC12, HLB14, HBB17, HRR^{+}11, HSG10, HN11, HM17, JMB^{+}13, JBC^{+}12, JJC12, JRP12, JAA^{+}13, KQ11, KK16, KHHG13, KRM11, KRM13, LBC10, LHS^{+}13, LHHL17, MM17, MSBT10, MHN11, MBH16, MM10, MGK^{+}18, MCO^{+}10, MCH^{+}13, MCP^{+}15, MSWD15, NHM10, NBC15, PBG^{+}11, PJ11, PPF^{+}13, PPH15, PBC12, PM15, RSB14, RGM^{+}11, RT14, SJT^{+}16, SMO^{+}14, SCC17, SKP^{+}10, SCL^{+}13, SHH^{+}19, SBDW11, SE11, SFLH14, SLG^{+}14, SVKK18, SMN^{+}10]. Chinook [Str10, Str12, TRMC15, TM19, TC11, TKC^{+}12, TES14, TKC^{+}18, TTW^{+}11, VBJS10, WHF^{+}12, WM12, Wei10, WBD14, ZKB^{+}15, ZKH12]. Chiricahua [ACK17]. chiricahuensis [ACK17]. Chloride [SSH11b]. Choice [TL14]. Choptank [AS17]. Chromosomal [MGP11, MSWD15, MSS^{+}17]. Chronic [DMRK14]. Chronicling [OGM^{+}15]. Chronology [MCT17]. Chub [BB15, BMOF16, CCL^{+}19, DSRA12, GLM^{+}15, HMW19, LRH12, OMA13, PSGB13, RMW16, SPH^{+}15, WZC^{+}14, WMS15]. Chum [HT13, MSS^{+}17, PGS13, RCC^{+}13, STS^{+}14, WAD^{+}17, WRP^{+}12]. Cichlid [BMH19]. Cisco [BHG^{+}10, FJJ^{+}12, HDH16]. Ciscoes [BD15, PBW16]. Clair [FPR^{+}18]. clarkii [SPM^{+}17]. Class [CPP^{+}14, MBPA13, PW10a, SWP^{+}18]. Classification [WBWSK12]. Cleaner [WHSK19]. Cleaning [GNV^{+}14]. Clear [GZB12]. Clearwater [TKC^{+}18]. Climate [AMBM12, BBL^{+}12, FJJ^{+}12, FBWT19, Har17, MDT10, RR13, SNC^{+}10, SWS10, UACS^{+}19, WAD^{+}17, ZTC12]. Climate-Growth [RR13]. Climate-Related [UACS^{+}19]. Close [YSC10]. Closed [MSH17]. Closing [JLA17]. Closure [BHC^{+}13]. Closures [BDBE11, GA10]. Club [VWS13]. Clupeid [HMO14]. Cluster [Nel14b]. Co [PDP12]. Co-occurrence [PDP12]. Coancestry [BSS18]. Coanda [PPW^{+}14]. Coarse [DW13]. Coast [FRRCGA^{+}15, SMS^{+}12, Wei10, YYA^{+}16, ASJ16, GSVS^{+}10, KSSK^{+}17, TVL^{+}12, ZHWR11]. Coast-Wide [ASJ16]. Coastal [AGJ11, AGKW12, BGB^{+}16, BKC^{+}19, BAR15, BGOQ13, CHPG12, CYW^{+}19, DGMC17, DSM18, DF12, FHCSN12, GGHS^{+}14, HRB^{+}16, Han18, HBTT11, HCZ13, HCZG15, KHD^{+}14, KWW14, MSK15, MSD^{+}19, MP10, MPM^{+}14, NDW10, OSA^{+}17, OWS16, PGJ11, RHEF19, RSL11, RG16, RBH^{+}12, RRG^{+}17, RL19, SFB16, SGP16, SR15, SSM16, THNG14, TH15, TTM^{+}15, vVO10, SMO^{+}14]. Coastwide [ASJ15, WMG^{+}14]. Cobia [Ada18]. Cochlodinium [RGP17]. Cod [DCBKR14, Han11, MHF12, PDSP^{+}11, SS15a, Spi12]. Coded [HB17, HSA^{+}16, Wei10]. Coding [GPBA18, NB13]. Coexistence [CHT13]. Coho [AKK^{+}19, BWDM11, BNTT17, DSHG10, FSJ^{+}11, FWT^{+}14, FBS^{+}14, FRM18, GWOD18, GHPS^{+}16, HGBE16, HWH10, JPSM12, JADJ^{+}14, KBC^{+}18, LPM^{+}17, MSD^{+}19, MGW13, MSS^{+}17, NMH12, NRZ^{+}13, OPHD18, PKL^{+}11, QHS^{+}13, RRA^{+}15, RKF^{+}13, RBH^{+}12, WTB10]. Cohort [TA13, ZM10]. Cold [FBJR11, IB14]. Coldwater [MSH17]. Collected [DPS14]. Collection [PCT^{+}12]. Collections [BGV11]. Collectors [KVA^{+}19]. Collingwood [Ber13]. Colonial [CML19]. Colonization [Esh14, HHES13, KBC^{+}18, WCMP13]. Colony [ARL^{+}12, MGWP10]. Color [PSM12]. Colorado [DWGF11, DF14, FWBT17, MWG18, Sto10, VFFP13, ZBW10, CYP11, CRH10, DF14, GBF12, HNV^{+}18, SRMB^{+}14, SPH^{+}15, YCB^{+}11, ZBW10, ZBHW16]. Columbia [BNS^{+}16, EPC^{+}19, MC14, NBC15, RDR^{+}10, RT14, ACRA10, ACRAM11, Bea10, BBC^{+}14a, BBC^{+}14b, BSHK^{+}11, BCH^{+}11, EHR^{+}12, FWT^{+}14, HJM^{+}12, HLB14, HTR13, HBB17, HN11, HHES13, HS13, JAA^{+}13, KK16, KPZ^{+}12, MPHG10, MKH10, MGWP10, MHN11, MBH16, MOS13, NHM10, PGR13, SZGS12, SJT^{+}16, SCL^{+}13, TA13, WAKH12]. Combination [SS13]. Combined [BWD^{+}17]. Combining [GA17, HUMG^{+}18]. Comment [ARC^{+}13, CCH15, Lis15, PWP13, SKBS18]. Commercial [DCS^{+}17, KQ11, PGS13, WCYH18]. Common [BSS18, BBZGB15, BW10, GN13, LDLH15, LBBW17, MOS13, NG13, NLF13, RNG^{+}13, SS13, TBTS12, WB13a, WB13b, WBMG10, YYA^{+}16]. Communication [LDA^{+}13]. Communities [BGLP19, ERL15, GS16, NJ12, SJT^{+}16, WHSK19]. Community [BBL^{+}12, BES15, CGM15, DGMC17, FPD12, HCZG15, MWS10, MWT14, RSBF14, STK12, TJB^{+}14, HCD^{+}14]. Comparative [BLB^{+}13, CBMR14, FA10, KWS14, ZC13]. Compare [TPLA16]. Comparing [ARB^{+}18, MBL12, PPF^{+}13]. Comparison [BLG^{+}13, BB10a, BN16, BSC^{+}16, CM10, CBSB15, CNM^{+}18, DM11a, DEP^{+}18, DSB^{+}18, ELC13, GGC17, HLB^{+}16, HOGO10, Han11, HDBS14, HS13, KGC^{+}12, KCZH14, MSD18, NMH12, OWN11, PNMH^{+}12, PET^{+}17, QHS^{+}13, RC16, RNG^{+}13, SMO^{+}14, SAW^{+}17, THGA15, YJW11, HGS^{+}16]. Comparisons [BT16, HNV^{+}18, OPS^{+}14, PF11a, RBE^{+}10, TT17, WCQD18]. Compensation [SA16]. Compensatory [RBK^{+}11]. Competition [DMRK14, HWN15, JBV13, PKL^{+}11]. Complementary [CR10]. Complex [AFHF12, FFYI16, KWW14, PHHM12, PPW^{+}14]. Complexity [Con17, DWG14, DW13, JADJ^{+}14, LF15, WDWC10]. Components [IWB^{+}13, KAH16, PDSP^{+}11, SF15, YSW^{+}14]. Composition [BAKT12, BCM^{+}19, CHLN15, CBM15, EMS^{+}14, FDWG13, GJC^{+}18, JC16, JPHK10, KHD^{+}14, PPF^{+}16, PM14, PDH^{+}18, SJT^{+}16, SFLH14, ZSVK13]. Compressed [KHF^{+}18]. Computational [GKBG12]. Concentration [FWD10, SJGL16]. Concentrations [MHPB14, SARC13, WNCF10]. Concept [BYM19, PRW^{+}10, TH15]. Concern [PVJH15]. Concerns [Zyd11]. Condition [BMC^{+}13, BS13, CTN11, CT16, CM11a, DYWY17, EHC^{+}14, FKF^{+}19a, GvdHBC12, GA17, HOGO10, HHJH17, HML13, HER^{+}11, KHF^{+}18, MAK12, PRK^{+}16, RNG^{+}13, RT14, SF17, TRMC15, WDR16]. Conditions [AS17, BJA^{+}13, CM11b, CPS^{+}15, FS16, FRRCGA^{+}15, GA17, HR15, HERC12, JPSM12, KCJ^{+}15, Mit16, MWG18, SLG^{+}14, Str12, TFM^{+}18, WB13a]. Confer [DWE^{+}19]. Confidence [SMN^{+}10]. Confirm [ESD^{+}17, EBF^{+}14]. Conflict [LBC10]. Conflicts [GKG^{+}19]. Congeneric [PL13]. Connection [MC14]. Connections [JPLS13]. Connectivity [DM11a, ERL15, KWW14, Lan16, STK12, WCN^{+}19]. Consecutive [JPG^{+}18]. Consequences [ACMM^{+}15, BMC^{+}13, Goo19, JSW^{+}15, PET^{+}17, TRW10]. Consequent [TTM^{+}15]. Conservation [AGKW12, ADS^{+}11, BMF^{+}15, BYM19, BLCB13, BMOF16, BGV11, CVL^{+}19, DKS18, DF12, FPM13, FBS^{+}14, Fre19, LRH12, LFB^{+}15, MDD^{+}16, SDOK19, SDSB18, SPH^{+}15, WCYH18, Zyd11]. Conserved [GGP^{+}15]. Consider [ARC^{+}13, Hoe17, PWP13].

Consideration [EPS^{+}15]. Considering [MAH^{+}12]. Consistency [SBMH12]. Consistent [PHG^{+}16]. Conspecific [SJGL16]. Constant [GDM^{+}10, HR15, ZBC^{+}13]. Constraints [HBB13, WDHW14]. Construct [PMN19]. Constructed [FSFS17]. Consumer [BR14]. Consumers [FH17]. Consumption [FM17, HBB17, JBV13, KLS17, KSB16, MBH12, MCO^{+}10, PM15, PB16, RBF^{+}11, RJWK10, SSS^{+}12, SHCB16, TES14, vPWT12]. Contemporary [HSE^{+}12, KHHG13, WKC18]. Content [BL11, JPH17, MBL12, SF17, VSC15]. Contents [BSSM12]. Conterminous [HTH^{+}17]. Context [MAH^{+}12, RCD^{+}11]. Continental [GRS12, PDH^{+}18, MAS10]. Contingents [NW14, PMF^{+}10]. Continued [ZHWR11]. Continuous [Dud19]. Contraction [RG16]. Contrasting [CBM^{+}13, DWE^{+}18, NRZ^{+}13, PNM^{+}10]. Contrasts [SMS^{+}12]. Contribute [TES14]. Contributing [HM17, SMMK12]. Contribution [ASJ16, ASJ17, BNS^{+}16, GLM^{+}15, NDC^{+}14]. Contributions [Stu18, THL^{+}17, WBR11]. Control [BBC14c, CHB^{+}12, CYP11]. Controlled [CPS^{+}15, KKM11]. Controlling [LV14, NAG13]. Controls [BPS11, SS13]. Cool [BNC^{+}15]. Copepoda [MFR15]. Copepods [Mit16]. Copper [AHS11a]. Copy [MSS^{+}17]. Core [DKD^{+}15]. Coregonine [AHS^{+}11b, TFM^{+}18]. Cormack [CBSB15]. Cormorants [BDS10, DCJ^{+}13, SCL^{+}13]. Corrections [HHJH17]. Correlate [PSGB13]. Correlated [HN11, vS10]. Correlates [BS13, WL13]. Correlation [GDPW10]. Correspondence [CBVF15]. Corresponding [SCC17]. Corridor [HGG^{+}16, WB15a]. Corridors [Esh14]. Cortisol [BLG^{+}13]. Cost [FDSB18]. Coterminous [ADS^{+}11]. Count [IWB^{+}13]. Countergradient [DKD^{+}15]. Counts [DMCO11]. County [BHK^{+}10]. Coupling [STK12]. Coutant [Nel14a]. Covariation [CHLN15]. Cover [CSDW10, SWS10]. Covered [BDM^{+}10]. Cownose [CMM^{+}17]. Crab [BT16, CGDL18, HCR14, RL14, RMC^{+}17, SBDW11]. Crabs [DDMR10, DK18, MBP^{+}11, OWS16, RCM^{+}19, SFF^{+}10]. Crappie [MBH12, MS13, SM13a]. CRC [Shi13]. Create [SSH11b]. Creek [BCF^{+}11, MDT10, PPMJ14, PPW^{+}14, PPJ15, WAD^{+}17, CNM^{+}18, HDBS14, QHS^{+}13, SBDW11]. Creeks [FBH^{+}13, JPRS10, RLL^{+}19]. Crested [BDS10, DCJ^{+}13, SCL^{+}13]. Critical [NCW^{+}17, RL14, UK14]. Critically [AUDL18]. Croaker [GvdHBC12]. Croakers [TLKT17]. Cross [CHPG12, DA16, GA17]. Cross-Life-Stage [GA17]. Cross-Scale [DA16]. Cross-Shore [CHPG12]. Crossing [BP10]. Crowded [WB15b]. Cryopreservation [BMTP13]. CSIRO [Ber13]. Cues [LRM^{+}10, VHPP19]. Culture [JKP10, SM13b]. Cultured [SDOK11]. Culvert [ERL15, JPSM12]. Culverts [MCW^{+}14, TPBB14]. Cumulative [EPC^{+}19, WHF^{+}12]. Cure [RPSB14]. Current [MMY12, Pet15, RCD^{+}11]. Curves [TP11]. Cutthroat [BB10a, BRG19, BRGR10, BGOQ13, CKM^{+}11, CR10, CRH10, DZW^{+}12, DWN^{+}15, HBB13, HB18, HNV^{+}18, HK12, HMP18, JJNB17, MSW^{+}10, MPM^{+}14, MKHC17, MDD^{+}16, ODG^{+}11, Pag19, PRW^{+}10, PPW^{+}14, SPM^{+}17, SK11, UACS^{+}19, ZTC12, ZBC^{+}13]. Cycle [BBZGB15, BMM10, KYK17]. Cycles [CM10, LKP^{+}13, SSS^{+}12]. Cycling [ZKH12]. Cyprinid [BB15, HMW19, HKO^{+}18, LRH12, PCFH17]. Cyprinidae [HOWS10]. Cyprinids [DMRK14, NAG13].

D [Ryp13]. Dace [BMTP13]. Daily [LS18, TL15]. Dakota [LBBW17, PGW^{+}11, WBD14]. Dam [BHK^{+}10, BLB^{+}13, FBWZ19, GBZ14, HGB^{+}16, HEW12, HCZ13, HCZG15, HSM12, JZS^{+}19, KCJ^{+}15, QHS^{+}13, RH14, SDW15, SH12, WCZ^{+}18, WBM^{+}17, WZK^{+}15, YITI12, KST^{+}12, SMN^{+}10]. Dams [BW10, EWSR16, HUMG^{+}18, KVA^{+}19, LRW13, PCT^{+}12, THNG14]. Danio [HKP11]. Darter [AFHF12, DA16, GBF12, KLS17, PAG11, PL13]. Darters [PSB16]. Data [ASP^{+}18, AJFO10, CYW^{+}19, GWG^{+}19, GBB^{+}14, HUMG^{+}18, IWB^{+}13, KM10, KHHG13, MRPW19, PM13, PGWS15, SR10, SMO^{+}14, WMV^{+}16, WBR11]. Date [BQF^{+}19, CEFQ11, KGW14]. Dates [AS17, SPR^{+}16]. David [Qui16, Zyd11]. Day [AJF13b, MT19]. Debris [TVL^{+}12]. Decades [PPC13]. Decelerating [EGAW12]. Deciphering [BBGB15]. Decisions [AS12]. Declination [HAR18]. Decline [ABM17, BHG^{+}10, NR16, PHM17, SBO12]. Declines [HMSG^{+}11, MLC13]. Declining [DWGF11]. Decompression [PKPS11]. Decreases [MJH17]. Deep [HBC15]. Deepwater [Con17, JDS15, LF15, MAN^{+}12, MHPB14, SFB16]. Defining [JJC12]. Deformities [BBF^{+}18]. Deformity [SO15]. Degree [MT19]. Degree-Day [MT19]. Delaware [TLKT17, HPF^{+}16, TT17]. Delayed [PJ11]. Delays [VJGV11]. Deletion [MSS^{+}17]. Delineate [GMP^{+}11]. Delta [BSSM12, DEW^{+}14, DWE^{+}18, DWE^{+}19, FSP^{+}17, KR18, LFB^{+}15, NLF13, PMN19, RKEB13a, RKEB13b, SBC^{+}16, YSC10, PBB^{+}15]. Delta-Reared [DWE^{+}19]. Demand [SHCB16]. Demersal [AW19, RA10]. Demographic [MHN11, NDP16, SH10]. Demographics [FSW10, VB14]. Demography [DH11, GBZ14, HCS^{+}16]. Demonstration [BDM^{+}10]. Dennys [SZZ10]. densa [CBW^{+}16]. Densities [CM11b, EBA^{+}19, FRM18, KBG^{+}18, SAH^{+}14]. Density [BPS11, BHLK10, BT16, BJA^{+}13, BB10a, CTPM13, DA12, FSJ^{+}11, GvdHBC12, GGC17, GDWD10, HML13, JPH17, MK15, PM14, PBM^{+}12, RMC^{+}17, RT14, RAR^{+}16, SR10, SJT^{+}16, SF15, SF17, VFB^{+}16, WB13a, ZKH12]. Density-Dependence [WB13a]. Density-Dependent [BPS11, CTPM13, SR10]. Dependence [BHLK10, HGB^{+}16, WB13a]. Dependent [BPS11, CTPM13, DW16, DCBKR14, HD16, PPH15, PM15, RBZ^{+}17, SR10, SHM^{+}12]. Depleted [HPSA16]. Depletion [CKS^{+}19, HKR18, HJS^{+}16]. Deployment [KM10]. Deposition [FPR^{+}18, HEC^{+}15, PAG11]. Depression [DSHG10]. Depth [BBZGB15, CHPG12, PBC12, RSG12]. Depth-Related [CHPG12]. Depths [PCC^{+}13]. Derived [HWH10, HSE^{+}12, MCW^{+}14, WM12, WTB10]. Deriving [CYW^{+}19]. Descaling [ZZD10]. Describe [HUMG^{+}18, OSS^{+}17]. Description [RKEB13a]. Desert [BMOF16, CM10, DFB15, DYWY17, HB18, MCD17, SBDW11, SBC17, WB15a]. Desiccation [KU12]. Design [BEWM10, EBMV10, FMJ11, JDS15, Nel14b, PMN19, SC13, TF13]. Design-Based [PMN19]. Designed [KM10, MAN^{+}12]. Designs [BP10, CBSB15, FVK^{+}18, KCZH14, SFF^{+}10]. Detailed [HPG10]. Detect [BSGB17, LSD^{+}14, SFF14]. Detectability [BSC^{+}16, RG15, WF11]. Detecting [WMV^{+}16]. Detection [AGJ11, BSSM12, BR14, CCL^{+}19, FVK^{+}18, GLM^{+}15, HKR18, HBS13, JSW12, KO16, MH11, MLB18, PF11a, PVJH15, PDM^{+}14, SL16]. Deter [DAA^{+}16, MCP^{+}15]. Determinants [MBG^{+}11, TKC^{+}18]. Determinate [GDG^{+}15, HMO14]. Determination [GDWD10, HS17, KWCI10, MSD18, PB17, SB12b, TL14]. Determine [CW11, SYH^{+}15]. Determined [BWDM11, GZB12, Wei10, WMW^{+}12]. Determines [BJC12, DGMC17]. Determining [CHO^{+}15, PBW16, ZSVK13]. Detroit [FPR^{+}18]. Develop [MRPW19]. Developing [ABS^{+}11, Poe14]. Development [BS18, BMTP13, CVL^{+}19, CEM^{+}19, DKL^{+}17, DA16, DG19, HHJH17, HBHS15, IB14, LS16, LS18, MWC^{+}13, MR11, NMH12, PKP^{+}17, PM19, RAB^{+}18, SLG^{+}14, TKC^{+}18, WHP14, ZLHS10]. Developmental [ODG^{+}11]. Devices [BES15]. Diagnostic [GMP^{+}11]. DIDSON [BFM10]. Didymosphenia [JC16]. Diel [AHS^{+}11b, BBZGB15, BJR14, GZB12, LKP^{+}13, MCSF14, RW11, RKF^{+}13, SBW14, SJFF15, TASC17, WZCT12]. Diet [AMBM12, BDS10, BN16, BR14, BSC^{+}16, CBM15, DWE^{+}18, DCJ^{+}13, DBS^{+}10, FRBA13, FBB^{+}13, GCA^{+}18, GJC^{+}18, HSSS16, Han11, HW14, HML13, JC16, JK17, JPHK10, LSPB11, MKK15, RL19, SBC^{+}16, SRMB^{+}14, SRP^{+}18, WCD^{+}11]. Diets [AMM^{+}16, EBA^{+}19, MFH13, MLBS^{+}13, NTHR18, RSBF14, WBD17]. Differ [BSM13, SB12a]. Differences [BMR^{+}12, MT19, MWSS13, PM19, PVJH15, QSJ12, RGP17, SS18b, SWC13]. Different [BHS^{+}17, CPS^{+}15, HVD^{+}16, JMM^{+}11, KMKS13, MC14, PB15, SFF^{+}10]. Differential [BQF^{+}19, BNW^{+}12, BEMF11, HRR^{+}11, LRH12, MGP11, SGP16]. Differentiation [BYM19, CPP^{+}14]. Differing [OWN11, PTPS10]. Digestion [LJHL10]. Digital [MVK12, WTB10]. Diked [KWW14]. diluviana [AHS^{+}11b]. Dimensional [GKBG12]. Diminished [BCF^{+}11]. Dinoflagellate [RGP17]. Dinosaurs [DKS18]. Dioxide [DAA^{+}16, SPH19]. Diporeia [PB16]. Direct [KST^{+}12, MAK12, SS13]. Direct-Gradient [MAK12]. Disaggregating [NR16]. Discard [BC12]. Discharge [HHES13, PA10]. Discordance [CHO^{+}15]. Discovery [MSW^{+}10]. Disease [EPS^{+}15, PPMJ14]. Disease-Positive [PPMJ14]. Disjoined [DM11a]. Dispersal [BLB^{+}13, FWT^{+}14, GRS12, HBS12, HCNL10, HPHB10, RHKD12, SNI^{+}13, WAKH12, YYA^{+}16]. Displacement [PPP^{+}13]. Dissolved [AS12, GWB^{+}15, KM10]. Distance [SSR^{+}17]. Distant [HEW12]. Distinct [BRG19, FH17, KBG^{+}18, MBL^{+}18, BYM19]. Distinctiveness [SPM^{+}17]. Distinguishing [Ano11a, TC11]. Distributed [RPC14].

Distribution [AGKW12, BPS11, BGW^{+}17, BA13, BD15, BW10, CKM^{+}11, CZK^{+}17, CKS^{+}19, CTK^{+}17, CAB11, EMS^{+}14, FZZ^{+}10, FSP^{+}17, FRM18, FSRK11, FRRCGA^{+}15, GGP^{+}15, GCM12, HSM^{+}10, HW14, HCS^{+}16, LKS^{+}12, LKP^{+}13, MLS^{+}13, NCW^{+}17, OSS^{+}17, PBW16, PSGB13, PSB16, PDH^{+}18, PGVG15, QHS^{+}13, RSL11, RG15, SMO^{+}14, SMH^{+}11, SWC13, TPBB14, TTM^{+}15, VJGV11, VCP^{+}14, WDHW14, WLLCF16, WBMG10]. Distributions [BRG19, CGM15, DWGF11, HLR^{+}12, SGP16, SS18b, SWS10, TW14, Wei10, WTB10, ZC13]. Disturbances [TW14]. Ditches [JPRS10]. Diurnal [MCT17]. Divergence [CHS^{+}11, DBHV14, MHN11, PGJ11]. Diverse [MKK15, Str12]. Diversifying [AHS11a]. Diversion [MCP^{+}15]. Diversions [HUMG^{+}18, MCH^{+}13, PPW^{+}14]. Diversity [BKC^{+}19, CN11, CHMG^{+}18, EPR^{+}18, HRB^{+}16, HBTT11, HWD^{+}16, LMD^{+}17, LYX^{+}10, MSP12, MBZ^{+}14, OWP^{+}17, RMM11, SDK^{+}10, STS^{+}14, SCC17, SSR^{+}17, TKC^{+}18, TPBB14, WDR16, WSI14]. DNA [BSGB17, CKM^{+}11, FVK^{+}18, GMB^{+}19, HOWS10, LYX^{+}10, PCFH17, RBF^{+}11, SBC^{+}16, Whi12, WMW^{+}12, WMG^{+}14]. Do [BB10a, LRW13, NPW12]. Documentation [MPM^{+}14]. Does [DYWY17, HSM12, LUC^{+}15, SB12a, SBDW11, TL15, VSBB^{+}14, VSBC13, WHSK19]. Dogfish [DCBKR14, DG19]. Dolly [BCS19, WTB10]. Dolphinfish [BAYS16, CYW^{+}19, RBE^{+}10, TRHHOGSJ14]. Domain [NAW12]. Domain-Sensitive [NAW12]. Domesticated [NMH12]. Dominance [IMO^{+}14, TB16, vZNW12]. Dose [VSBB^{+}14]. Double [BDS10, DCJ^{+}13, HRG12, SCL^{+}13]. Double-Crested [BDS10, DCJ^{+}13, SCL^{+}13]. Downs [GS13]. Downstream [CTPM13, EWSR16, MJH17, TKC^{+}12]. Downstream-Migrating [MJH17]. DPS [BYM19]. Drainage [BAHP11, ERL15, HUMG^{+}18, WL13, WKC18]. Drainages [RG15, RG16]. Dramatic [HSSS16]. Drawdowns [GBW16]. Dredging [FPD12]. Dreissenid [HML13, KS11, MBG10]. Drift [GGC17, HGS^{+}16, UPT^{+}13]. Drift-Feeding [HGS^{+}16]. Drift-Flow [HGS^{+}16]. Driven [OSDT18]. Drivers [EPR^{+}18, KO16, MSB^{+}18]. Driving [NR16]. Drones [HNRB19]. Drought [PSGB13]. Drowned [BR11, JRBC13]. Drum [RR13]. Dry [CNM^{+}18, JPH17]. Drying [FBF10]. Dual [GBB^{+}14, MJHP18, MBBM10]. Dual-Frequency [GBB^{+}14, MJHP18, MBBM10]. due [JHK^{+}19]. Duration [MGW13]. During [BNTT17, CNM^{+}18, LUC^{+}15, Arc16, AJF13b, BP13, CKFQ11, CHW^{+}13, CSD^{+}14, GCG^{+}18, GA10, HPG10, HBA^{+}15, JCK^{+}12, KYK17, MR11, PM14, PKL^{+}11, PHG^{+}16, PBC12, PTG^{+}10, PGS13, PGR13, SWB^{+}16, SGB13, SGK16, TB14, TKC^{+}12, ZKH12]. Duskystripe [SB12b]. Dwelling [BS13, STZ13, WNCF10]. Dynamic [AHS^{+}11b, BWHF17, BCH^{+}11]. Dynamics [BGB^{+}16, BBS12, CBMR14, FM17, FMG^{+}11, GKBG12, GP12, HSS17, IHTF11, JRBC13, KGC^{+}12, KR18, KSB16, KRM13, LBBW17, NR16, PF11b, PM10, RKEB13a, RKEB13b, SFB16, SH10, SM13a, WvPC12, WRQH17, ZBHW16]. e-book [Fle16, Pet15]. Early [BB15, BNS^{+}16, CSED10, CDB10, CSD^{+}14, DWE^{+}19, DMHM^{+}18, FWT^{+}14, FPM15, GCG^{+}18, GDST16, HPF^{+}16, HSS17, JWHJ13, KKM11, LME16, LDLH15, LHS^{+}13, LJHL10, MNS^{+}19, MSWS10, MWKS13, MBT10, NMH12, PFG19, PKL^{+}11, PTG^{+}10, PGS13, PGR13, RRA^{+}15, SBC^{+}16, SSH^{+}11a, SWB^{+}16, SHP^{+}19, TD13, TKC^{+}12, WDW11]. Early-Life [BB15]. East [BSCA17, DG14, GSVS^{+}10, YYA^{+}16]. East-Central [DG14]. Eastern [BBL^{+}12, CTG14, CGDL18, DSBS14, DW15, SSH^{+}11a, SS15a, SS18b, SBDW11, SMMK12]. Echograms [MBBM10]. Ecological [GS16, OTR^{+}14]. Ecology [ACDL19, BGW^{+}17, BAYS16, BL11, DKS18, FK18, Fre19, GRDO13, Gol18, HRHS16, HB18, IRJ11, Mil19, PNM^{+}10, PSH^{+}19, PFG19, QSJ12, Qui16, Qui19, RCM^{+}19, RMM11, RBE^{+}10, SK11, Sec19, SM13b, SGK16, TA12, UWL12b, Ber13, Gid14, HCD^{+}14]. Ecomorphological [MBZ^{+}14]. Economical [DJ17]. Ecoregion [PKP^{+}17]. Ecoregions [KMKS13]. Ecosystem [Chr14, EBA^{+}19, HCR14, Han18, KGC^{+}12, KKB^{+}10, RA10]. Ecotype [HW14]. Ectoparasitic [Mit16]. Edge [CZK^{+}17, CKS^{+}19, RO16, VJGV11, Zyd11]. Edisto [CTP^{+}12]. Edited [Ber13, Chi13, Chr14, Fle16, LDA^{+}13, Nel14a, Ryp13, Shi13, Var13, Zyd11]. edition [Ryp13, Shi13]. eDNA [FVK^{+}18, GMB^{+}19]. edwardsii [Mit16]. Eel [CPS^{+}15, GPBA18, HEW12, SJGL16, ZZMC13]. Eel-Stage [GPBA18]. Eels [EWSR16, FSW10, FBB^{+}13, GPBA18, RLL^{+}19, SJGL16, SAH^{+}14, WL13, Zyd11]. Effect [CJM^{+}15, CT16, EPW15, GDM^{+}11, GCA^{+}18, HJM^{+}12, JRP12, LS10b, MGW13, OWN11, PBB^{+}15, SSS^{+}12, SHR^{+}12, SH12, UWL12b, WEW12, vZNW12]. Effective [BSS18, CYP11, HMSG^{+}11, JCS14, LV14, MSP12, ZC13]. Effectiveness [BGV11, BR14, JADJ^{+}14, LDP19]. Effects [ARL^{+}12, AW19, ACK17, AGKW12, BPS11, BB15, BJA^{+}13, BSC^{+}10, BBZGB15, BCRP13, BBL^{+}12, BWD^{+}17, BKC^{+}19, BHK^{+}10, CSDW10, Cla16, CDB10, CCB16, DW13, DWN^{+}15, EGAW12, EBMV10, EPC^{+}19, FWB^{+}15, FPM13, FKH15, FH12, FMS10, FMG^{+}11, GWOD18, GBW16, GJC^{+}18, GA17, Gre12, GIW^{+}13, GA10, HBS13, HB17, HTR13, HMW19, HWN15, HWHF14, HMP18, IBH12, JBC^{+}12, KH13, KHF^{+}18, KKM11, LDLH15, LJHL10, LBBW17, LSJN13, LSV^{+}13, MDP12, MHE12, MFH13, MK10, MS12, NDW10, NAG13, OPHD18, OPS^{+}14, PB15, PPF^{+}13, PGWS15, Poe14, PB16, QSVD11, RHW15, Rey16, RBF^{+}11, SDW^{+}17a, SO12, SDOK11, SLG^{+}14, SPCJ11, Sto10, SCB15, TASC17, TPA14, THNG14, TM19, UACS^{+}19, VSBB^{+}14, VSBC13, WHF^{+}12, WMS15, WZCT12, WR14, WK13, WDWC10, WF19, WCN^{+}19, WOM14, YCB^{+}11, YHWH10, vVO10]. Effects [KR18]. Efficacy [BSGB17, BDBE11, FDSB18]. Efficiency [CGL^{+}11, MJH17, SIKR17, SFF^{+}10]. Efficient [WNCF10]. Effort [HAG13, HCD10, PDM^{+}14, SA14]. Efforts [CCKF11, CBSB15, DF12]. Egeria [CBW^{+}16]. Egg [AGKH17, Bea10, BS18, BGV11, FPR^{+}18, FSC^{+}13, JRP12, MLBS^{+}13, PAG11, QSVD11, SS13, SVKK18, TM19, WBD14].

Egg-to-Fry [JRP12, SVKK18]. Eggs [BCR13, CHLN15, FH12, GGC17, RB13, RAR^{+}16]. Eight [SFLH14]. Elasmobranch [OPS^{+}14]. Elastomer [HB17]. Electric [ARB^{+}18]. Electrical [MJH17, Rey16, TBGM12]. Electricity [NAG13]. Electrode [HH11]. Electrodes [Rey16]. Electrofishing [HAG13, HKR18, HJS^{+}16, LDP19, MKHC17, MK10, PVJH15, Rey16]. Electroimmobilization [NYCC19]. Electromagnetic [BCF^{+}13]. Electronarcosis [BLG^{+}13]. Electronic [LEM^{+}11]. Electrosedation [ARB^{+}18, PET^{+}17, RFM^{+}15]. Electroshocking [MAN^{+}12]. Elevation [HNV^{+}18, WDHW14, WTB10]. Eliminates [JJNB17]. Elk [PSB16]. Elucidates [PCFH17]. Elusive [JLV^{+}10]. Elwha [HGB^{+}16, LPM^{+}17, MPM^{+}14, QHS^{+}13, WBM^{+}17]. Embryo [MYS^{+}11]. Embryonic [DG19, HDQ^{+}17, WHP14]. Embryos [BQF^{+}19, NAG13]. Emerge [BRG19]. Emergence [BS18, JSW^{+}15]. Emerging [AMM^{+}16]. Emigrant [BMF^{+}15]. Emigration [AHS10, GGS10, HLK11, PM14, PHG^{+}16, RRA^{+}15, TRLG13, WF19]. Empirical [BGV^{+}12, JPH17]. Encounter [HN13]. Encounters [YSC10]. Endangered [AUDL18, BB15, BH11, CCL^{+}19, DPB13, DYWY17, DSAM12, EBF^{+}14, GLM^{+}15, HKO^{+}18, JZS^{+}19, Lan16, LFB^{+}15, LSD^{+}14, PAG11, PPD16, SBC17, TTM^{+}15, WOM14, ACDL19]. Endemic [BYM19, MKA10]. Endocrinology [LSSYL10]. Endogenous [GGS10]. Energetics [OGM^{+}15]. Energy [CKS^{+}19, GvdHBC12, GDWD10, HR15, HGS^{+}16, HML13, JPG^{+}18, JPH17, MDP12, MBL12, PM14, PBM^{+}12, SF15, SF17]. England [RJ11, DKL^{+}17, FHCSN12, PMF^{+}10]. Enhance [BMF^{+}15]. Enhanced [Bea10, PK11]. Enhancement [CEM^{+}19, DF12, HGG^{+}16]. Enhances [ERL15]. Enrichment [KPZ^{+}12]. Entering [ZZD10]. Entrainment [KMM^{+}11, KR18, MBPA13, MCH^{+}13, MCP^{+}15, PBB^{+}15, SO12, WHF^{+}12]. Entrance [SFF^{+}10]. Entry [MJHP18]. Environment [BWD^{+}17, CSD^{+}14, DSRA12, FS16, GDST16, MSWS10, MOS13, WDW11]. Environmental [BSGB17, CM11a, CJG13, DSM18, GGS10, HMTC12, HBB13, JPSM12, KLR^{+}15, KCJ^{+}15, KO16, MPHG10, MSK15, Mit16, OSDT18, PTWK11, SM13a, Str12, UWL12b, VHPP19, WB13a, WL13, FVK^{+}18, GMB^{+}19]. Environments [BHS^{+}17, BCH^{+}11, HT13, MAH^{+}12, NRZ^{+}13, WSQ^{+}15]. Episodic [BAR15]. Erie [BBG^{+}18, CM11a, Esh14, HDBS14, PDM^{+}14, SWC13, VB14, YJW11, ZRM^{+}16]. Errata [Ano10c, Ano10d]. Erratum [Ano10e, Ano11a, Ano12a, Ano14a, Ano15, Ano17a, Ano19a, Ely12, Ano13a]. Error [AJF13a]. Escapements [MK16]. Esocids [EWW13]. Establish [SBC17]. Establishment [PW10a, SRP^{+}18]. Estimate [CGA13, DYWY17, HPF^{+}16, KHW^{+}14, PPH15, PHFN17, SL16, TS13, VSC15]. Estimated [dHST13]. Estimates [AJF13a, BSR12, BSSP18, BFM10, CN11, CC16, CBSB15, FBH^{+}13, HLB^{+}16, HJS^{+}16, HBS12, HSE^{+}12, JCS14, KDS18, MAS10, MCO^{+}10, RSB14, STZ13, SJMG14, UWL12a]. Estimating [CTN11, CKC^{+}12, HOGO10, HCD10, HGP^{+}17, IWB^{+}13, MH11, MBL^{+}18, PCT^{+}12, PM15, WBR11, ZM10, ZC13, vPWT12]. Estimation [AFHF12, ASP^{+}18, BC12, CCKF11, CEM^{+}19, DLHF16, DL13, DMHO12, HBTT11, HSA^{+}16, MBP^{+}11, MK16, PVJH15, SKBS18, SC13]. Estimations [ZZMC13]. Estimators [HSF12, Hoe17, STZ13, THGA15]. Estuaries [DK18, DSBS14, FWSW^{+}15, Fre19, LEM^{+}11, NW14, NDW10, WBMG10]. Estuarine [Ada18, CT16, DWE^{+}19, DHB15, FWD10, FDWG13, FBH^{+}13, LDP19, MJS^{+}19, NRZ^{+}13, NR16, PHG^{+}16, PE14, PA10, SF15, SMMK12, WF11]. Estuary [BR14, CBW^{+}16, DCS^{+}17, DHB15, FZZ^{+}10, HPF^{+}16, HSSS16, HJM^{+}12, MSH17, NCW^{+}17, OSS^{+}17, PMF^{+}10, RT14, SCL^{+}13, VBJS10, WBD17, BSSM12, BSC^{+}16, KR18, MBH16, NLF13, NR16, QHS^{+}13, RKEB13a, RKEB13b, SBC^{+}16, SMH^{+}11]. Ethanol [FH12, OPS^{+}14, VSBC13]. Ethanol-Stored [OPS^{+}14]. Etheostoma [DPB13]. Etowah [AFHF12]. Eugenol [BTWB17, CJL^{+}14]. Eugenol-Based [CJL^{+}14]. Eulachon [FSN^{+}13]. European [HMSG^{+}11, SHR^{+}12, VJGV11]. Evaluate [DF12, KST^{+}12, WCQD18]. Evaluated [HPHB10]. Evaluating [ACRA10, ARC^{+}11, GSD^{+}12, GBF12, KRRW18, KP14, MCW^{+}14, NAG13, PGWS15, PSB16, WCN^{+}19]. Evaluation [BSS18, BHC^{+}13, BSSP18, CCKF11, CC16, CGA13, CTP^{+}12, EHR^{+}12, FBJR11, FHCSN12, GDWD10, GSA^{+}19, HAG13, HB17, HJS^{+}16, JPH17, MP10, MWC^{+}13, RHEF19, SC13, UNTA10, WDDS13, WBH10, ZBHW16, ZLHS10]. Evaluations [AIC13]. Events [BS16, PGVG15]. Evidence [BGV^{+}12, BNTT17, CTS^{+}15, CT12, CTPM13, DKD^{+}15, DZW^{+}12, Gre12, HKW^{+}14, HDQ^{+}17, LSSYL10, PF11b, SS13, SSR^{+}17, Str10, TLKT17, ZTC12]. Evidenced [WvPC12]. Evolution [Bea10]. Evolutionarily [PRW^{+}10]. Evolutionary [AKM12, ADS^{+}11, Pag19]. Examination [HPG10, SBC^{+}16]. Examine [GA17]. Examined [FSJ^{+}11]. Examining [BR14, BCM^{+}19, LBC10, NHM10]. Example [AGH18, ACWIK13, Goo15, GA10]. Exclusion [CWP^{+}19]. Exercise [CJM^{+}15]. Exhaustive [CJM^{+}15]. Exhibit [BB10a, BNW^{+}12]. Exotic [TD13]. Expansion [LN18, SGK16]. Expectations [GKG^{+}19]. Experience [CGB19, WEW12]. Experiencing [HBTT11]. Experiment [JCK^{+}12, SLG^{+}14]. Experimental [BNW^{+}12, CTP^{+}12, DMRK14, FSC^{+}13, GS16, HJS^{+}16, HDQ^{+}17, JDS15, SS18a, SAF^{+}19, WHC^{+}10]. Explains [BRA14, PJ11, SFF^{+}10]. Exploit [FH17]. Exploitation [KQ11, PGWS15, SS18a, Fle16]. Exploring [BMR^{+}12, GA10, NLF13, NDW10, SZGS12, SM13a, WB15b]. Exposed [BDBE11, BCG^{+}12, DMRK14, GDM^{+}10, PET^{+}17, RGP17, YSC10]. Exposure [AW19, CRC^{+}16, CHB^{+}12, CCB16, MGW13, SMO^{+}14, TNS^{+}14]. Exposure-Sensitivity [TNS^{+}14]. Expression [BHP^{+}15, GPBA18]. Expulsion [IBH12]. Extending [LDP19]. Extensive [HUSST19, JHK^{+}19]. Extent [ACRAM11, RHW15, UPT^{+}13]. Extents [FSJ^{+}11]. External [DEP^{+}18, HER^{+}11, JDS15, MHPB14]. Externally [JBC^{+}12]. Extinction [BHG^{+}10]. Extinctions [FSRK11]. Extirpations [FPM13]. Extreme [GJC^{+}18, STK12, SGB13]. Extrinsic [MNB18].

F/SPO [Yas16]. Facilitate [BDTD17, DPS14, KBG^{+}18]. Facilitates [HBS12]. Factor [ABM17, CT12, JRP12]. Factors [ACRAM11, GGS10, KLR^{+}15, KVA^{+}19, MPHG10, MNB18, MD11, MK16, OTR^{+}14, OSDT18, PPF^{+}16, PTPS10, RSVAB19, RBH^{+}12, SAS17, SWP^{+}18, SMMK12, TTM^{+}15, UACS^{+}19, WBD17]. Failure [BHC^{+}18, Smi13]. Fall [Ano11a, BGV^{+}12, BWD^{+}17, FJP18, FMS10, FMG^{+}11, GDM^{+}10, GDM^{+}11, GGC17, HSM^{+}10, HGBE16, MPM^{+}14, MWKS13, PJ11, SCL^{+}13, SE11, SFH15, SMN^{+}10, SDW17b, TC11, TKC^{+}12, TKC^{+}18, WBD14, WRP^{+}12]. Fall-Run [BWD^{+}17, FMG^{+}11, WRP^{+}12]. Falling [PJ11]. Family [TB16]. Farms [JKP10]. Fast [JLA17, LSPB11]. Fast-Growing [LSPB11]. Fasting [SSS^{+}12]. Fat [DG14]. Fate [LUC^{+}15]. Fathead [AS12, SRMB^{+}14]. Fatty [HWH10]. Favor [MWG18]. Fear [SH12]. Feasibility [SNP^{+}10, SHCB16]. Feature [CYW^{+}19]. Features [BCH^{+}11, MWG18]. Fecundity [BEMF11, DG19, GDG^{+}15, HMO14, QSVD11]. Fed [PB15, WRP^{+}12]. Federally [CCL^{+}19]. Feeding [Bel12, BAYS16, DA12, DG14, GRDO13, Han18, HGS^{+}16, LME16, LSPB11, MCT17, NCW^{+}17, PFG19, RBE^{+}10, SSS^{+}12, SGK16, TT17, TRHHOGSJ14, TA12, UWL12b]. Female [BRK^{+}18, CTG14, CPS^{+}15, DK18, JPG^{+}18, OWS16, PSM12, SSS^{+}12]. Females [CHLN15]. Fertilizer [WHC^{+}10]. Fidelity [DPB13, GDST16, TLKT17, WLW16, YYA^{+}16]. Field [Ano11a, CJL^{+}14, HUMG^{+}18, LDP19, MT19, SS13, TC11, VSC15, WTB10, vPWT12]. Field-based [vPWT12]. Field-Surveyed [WTB10]. Fields [BCF^{+}13, Rey16]. Fin [BSR12, BSSP18, OPS^{+}14, SWH^{+}17, VSBC13]. Findings [RMW16]. Fine [BR11, HBA^{+}15, PS14, TPBB14, WAKH12]. Fine-Scale [BR11, HBA^{+}15, PS14, TPBB14, WAKH12]. Finger [BAKT12]. Finley [Smi13]. Fire [SMY11]. First [BNTT17, CKFQ11, KM14, SO15, SKP^{+}10, VFB^{+}16, WB13a]. First-Generation [SKP^{+}10]. First-Winter [KM14]. First-Year [VFB^{+}16, WB13a]. Fish [ABM17, ARB^{+}18, AUDL18, AW19, ACB^{+}14, BEWM10, BGLP19, BT16, BDTD17, BMOF16, BAKT12, BWHF17, BCF^{+}13, BSCA17, BR11, BBL^{+}12, BP10, BKC^{+}19, BTWB17, BES15, BDM^{+}10, BHK^{+}10, CCL^{+}19, CS16, CGM15, CJL^{+}14, Chi13, Cla16, CYP11, CGA13, CBM^{+}13, CWP^{+}19, DGMC17, DSBS14, DW13, DM11a, DYWY17, DHB15, ESD^{+}17, EB16, EPR^{+}18, EHC^{+}14, ERL15, FWD10, FBWZ19, FK18, FQW^{+}10, FPD12, FPR^{+}18, FBWT19, FA10, FH12, GKBG12, GBW16, GDWD10, GA17, GNV^{+}14, GWG^{+}19, GSD^{+}12, HLB^{+}16, HSSS16, HAG13, HR15, HS17, HN13, HARL18, HMW19, HKP11, HSA^{+}16, HCZG15, HD16, HCD10, HLK11, HER^{+}11, HERC12, HEC^{+}15, IRJ11, IB14, JZBP16, JPH17, JKP10, KGC^{+}12, KMKS13, KCS^{+}19, KMM^{+}11, KCZH14, KP14, KHF^{+}18, KVA^{+}19, KWW14, Lan16, LF15, LJHL10, LSD^{+}14, MDP12, MRPW19]. Fish [MSK15, MP10, MLS^{+}13, MH11, MKA10, MWT14, MKH13, MVP15, MWG18, MBG^{+}11, MCO^{+}10, NPW12, NDC^{+}14, NJ12, NYCC19, NR16, NDW10, NAG13, OSS^{+}17, PPF^{+}16, PW10b, PAB^{+}11, PTWK11, PK11, PE14, PCT^{+}12, PHFN17, PDM^{+}14, RHEF19, RZW^{+}12, RBZ^{+}17, RW11, RSBF14, RJ11, RGP17, RFM^{+}15, SARC13, SJT^{+}16, SDBB12, SAF^{+}19, SFB16, SF17, SJ11, STZ13, SCF16, Smi13, SR15, SHC^{+}16, STK12, SWS10, SGB13, SAH^{+}14, TF13, VFFP13, VIW^{+}17, VSBC13, WDR16, WB15a, WMS15, WR14, WSQ^{+}15, WRQH17, WCZ^{+}18, WK13, WBWSK12, WHSK19, WB15b, WBW^{+}13, YCB^{+}11, YSC10, vVO10, HLH14]. Fish-Based [FA10]. Fished [HVD^{+}16]. Fisheries [Ada18, Ano13b, Ano14b, Ano16, Ano17b, Ano19i, BGB^{+}16, CML19, Chi13, Chr14, CEM^{+}19, IWB^{+}13, Jac18, LS10a, LDA^{+}13, MPHG10, Nel14a, Nel14b, PM19, Pet15, RCC^{+}13, Ryp13, SRMB^{+}15, SA16, Smi13, TH15, TBC^{+}13, WMV^{+}16, WBR11, WCYH18, Yas16, Zyd11, Ryp13]. Fishers [AJF10, AJF13a, AJF13b, WCYH18]. Fishery [ABS^{+}11, BCM^{+}19, CYW^{+}19, CGDL18, DCS^{+}17, DDMR10, DPS14, DCBKR14, GKG^{+}19, GA10, HNRB19, KQ11, KAH16, KSB16, KYK17, PGW^{+}11, SMO^{+}14, VB14, WOM14]. Fishery-Dependent [DCBKR14]. Fishery-Independent [ABS^{+}11, GKG^{+}19]. Fishes [AHS^{+}11b, BYM19, BGB^{+}16, BS15, BSC^{+}16, CHR^{+}17, CT16, DWGF11, DWG14, DKS18, FMJ11, GP12, Gre12, HBC15, Han18, HLH14, HTH^{+}17, HUSST19, JDS15, JBV13, KS11, LBBW17, LRR^{+}15, Mat18, MK10, MLB18, MS12, MHPB14, MLBS^{+}13, PF11a, Poe14, PA10, RA10, SF15, SGP16, SRMB^{+}14, SHCB16, SPH^{+}15, SL16, Sto10, WHSK19, WF11, Ber13, Gid14, Var13, Qui16]. Fishing [AJF13a, AJF13b, BBL^{+}12, BNS^{+}14, KK16, LLWW19, PAA14, PGS13, RSVAB19, TS13, ZM10, dHST13]. Fishing-Related [BNS^{+}14]. Fishpass [CHH^{+}16]. Fishway [BMB^{+}10, FMJ11, JCK^{+}12, KCM14, RHEF19]. Fishways [FHCSN12, KCJ^{+}15, LMPvdH18]. FishXing [MCW^{+}14]. FISK [HLH14]. Fitness [DWN^{+}15, HWD14, PK11]. Fitness-Related [DWN^{+}15]. Five [EWSR16, PA10]. Fixation [FH12]. Fixed [BFH14, PNMH^{+}12]. Fixed-Frame [PNMH^{+}12]. Flannelmouth [CGB19, FWBT17, WvPC12]. Flatfishes [RB18]. Flathead [BK11, FBP11, MSB^{+}18, WZC^{+}14]. Flint [SYH^{+}15]. Flood [KKM11, MDT10]. Flooding [YHWH10]. Floodplain [AM13, DSRA12, GJC^{+}18, HSG10, Lan16, MWS10, RW11, RAB^{+}18]. Florida [Shi13, BLCN^{+}10, CGDL18, CAB11, DG14, FWSW^{+}15, Gre12, HAG13, MHHH10, PA10, RHKD12, SA14, SA16, SAS17, SGB13, TPLA16, TAM15, THL^{+}17, WBMG10, YITI12, YYA^{+}16]. Flounder [AKM12, BFK15, BC12, Bel12, BL11, FDWG13, HPSA16, MHF12, NW14, RJ11, RJH12, WMG^{+}14]. Flow [Arc16, BBZGB15, BWHF17, FBWT19, GP12, HGS^{+}16, HSE^{+}12, MP10, OPHD18, PBB^{+}15, PHHM12, RAB^{+}18, RMW16, STK12, WKC18, YHWH10]. Flow-Related [OPHD18]. Flow-Specific [RAB^{+}18]. Flowing [EMS^{+}14]. Flows [EGAW12, FKH15, KKM11, YSC10]. Fluctuating [GDM^{+}10, GDM^{+}11, HR15, KKM11, RB13, RBF14, ZBC^{+}13]. Fluctuation [TASC17]. Fluid [GKBG12]. Flume [MCH^{+}13]. Fluorescent [HKP11, HLH14]. Fluvial [HGG^{+}16, HCS^{+}16, LB15]. Following [BS16, Lan16, HML13, PPJ15, RH14, RHR^{+}15, SFB16, SFLH14, VWS13]. fonticola [DPB13]. Food [BCRP13, HTR13, JMM^{+}11, KM14, KPZ^{+}12, LB15, PK11, SW11, ZRM^{+}16, vPWT12]. Forage [EB16, KHF^{+}18, NR16, OGM^{+}15, PW10b]. Foraging [CW11, DEW^{+}14, HBB13, HRHS16, HDF12, JMB^{+}13, NPW12, NCW^{+}17, PMF^{+}10, SK11, WDWC10, WB13b]. Forced [PKPS11]. Forces [NR16]. Forebays [KVA^{+}19]. Forecasting [ZRM^{+}16]. Forest [TPBB14]. Forested [FSFS17]. Forests [DWE^{+}19]. Fork [ACDL19, HGBE16, SHCB16]. Formalin [FH12]. Found [LLWW19]. Foundations [Jac18]. Fountain [PAG11]. Four [BHC^{+}13, CN11, DW13, FSJ^{+}11, HML13, HWN15, MHF12, ODG^{+}11, SCC17, THKG10]. Four-Level [THKG10]. Fragmentation [BBAK19, OMA13, PSGB13]. Fragmented [ACMM^{+}15, SHR^{+}12]. Frame [PNMH^{+}12]. Frameworks [FA10]. Francis [PGW^{+}11]. Francisco [ARL^{+}12, BSSM12, BSC^{+}16, KR18, NLF13, NR16, RKEB13a, RKEB13b, SBC^{+}16, SMH^{+}11]. Fraser [BNS^{+}16, MPHG10, PGR13, REN17]. Free [BNC^{+}15, EMS^{+}14, Yas16]. Free-Flowing [EMS^{+}14]. Frequency [BBF^{+}18, GBB^{+}14, MJHP18, MDT10, MBBM10, ZZMC13]. Freshwater [Ber13, BCF^{+}13, BTWB17, CHW^{+}13, DWE^{+}19, DW13, EPR^{+}18, Gid14, HMTC12, HBC15, HTH^{+}17, MCC^{+}12, NDW10, Poe14, RR13, RDR^{+}10, SJT^{+}16, SGB13, SHP^{+}19, SPH19, TB14, Var13, WDW11, vVO10]. Fright [ELC13]. Frog [BSHK^{+}11]. Frogs [ACK17]. Frozen [OPS^{+}14]. Fry [BCR13, HT13, JRP12, JSW^{+}15, LME16, SVKK18, TB16]. Fuca [MBT10]. Function [CSD^{+}14, SJGL16, TASC17, TL14]. Functional [EPR^{+}18, MSWD15, PPF^{+}16]. Functioning [HCR14]. Functions [HD16]. Fundy [WMW^{+}12]. Future [DDMR10, DKS18, FJJ^{+}12, HP13, SWS10, ZTC12].

G [HCD^{+}14]. Gag [ABS^{+}11]. Gags [RMMGPB16, SMMK12]. Gain [AGH18]. Galaxias [AUDL18]. Game [SWS10]. Gar [BSR12, BSD13, BSSP18, FS16, GCM12, KDS18, LS16, MRB^{+}18, RAB^{+}18, SDSB18]. Gard [Yas16]. Gars [DSM18, DKD^{+}15, FVK^{+}18, FDSB18, LS18, MSD18, SSM16, SKBS18]. Gary [HCD^{+}14]. Gas [RMC^{+}17, RCM^{+}19, SDOK11]. Gastric [BN16, DEP^{+}18]. Gate [PBB^{+}15]. Gauging [BGLP19]. Gear [CCB16, PVJH15]. Gears [Poe14]. geminata [JC16]. Gene [CHO^{+}15, GPBA18, HSE^{+}12, MSWD15, NB13, WKC18]. Generalist [HSSS16, PHFN17, WDWC10]. Generalized [YJW11]. Generation [BSC^{+}16, LFB^{+}15, SKP^{+}10]. Generations [DSHG10, ODG^{+}11, SFLH14]. Genes [CPP^{+}14, GPBA18]. Genesis [LS18]. Genetic [ACB^{+}14, ADS^{+}11, BLCN^{+}10, BBAK19, BWRA16, BCM^{+}19, BSC^{+}16, BEMF11, BJC12, CN11, CKC^{+}12, CVL^{+}19, CMM^{+}17, CHS^{+}11, CTP^{+}12, CEM^{+}19, DSRA12, DSAM12, FPM13, FGW^{+}16, GRS12, GSVS^{+}10, HDBS14, HBTT11, HBS12, HUMG^{+}18, KSF13, KHK^{+}15, KHHG13, KCMS11, LFB^{+}15, LYX^{+}10, LCHB^{+}18, MBEvdB^{+}10, MSP12, MDD^{+}16, NHM10, NZF^{+}11, NDP16, OWP^{+}17, PZS17, PDSP^{+}11, PRW^{+}10, PGJ11, RSB14, RHKD12, RT14, SPM^{+}17, SJT^{+}16, SDK^{+}10, STS^{+}14, SMO^{+}14, SFLH14, SSR^{+}17, SBMH12, SS15c, TNOU11, TPBB14, VHJ^{+}15, WAW11, WBR11, WSI14, WKS^{+}19, WCN^{+}19, WAKH12, WKC18, WCQD18, WLW16, WBM^{+}17]. Genetic-Based [RSB14]. Genetically [HSE^{+}12]. Genetics [AGKW12, BMOF16, CBM^{+}13, MCD17, NB13, RPC14, SH10, Spi12]. Genome [BMTP13, MGP11]. Genomic [BYM19, BMP^{+}17, SPM^{+}17]. Genotoxic [VSBB^{+}14]. Genotypes [HRR^{+}11, SSH^{+}11a]. Genotypic [CHO^{+}15]. Genotyping [DPS14]. Genus [SDK^{+}10]. Geochemical [AW19]. Geographic [CTN11, CHS^{+}11, VCP^{+}14, WDHW14]. Geography [MSS^{+}11]. Geology [NJ12]. Geomorphic [MP10]. Georges [ZC13]. Georgia [NBC15, BP16, BP17, FP19, FGW^{+}16, JPFB13, KTE^{+}14, PB13, SP10, SJFF15]. Georgian [WLLCF16, WLW16]. Ghost [NB13]. Giant [BMH19]. Gila [MCD17]. Gill [BHC^{+}18, PTPS10, PNMH^{+}12, RB18, SZZ10]. Gill-Net [BHC^{+}18, PTPS10]. GIS [WRP^{+}12]. GIS-Based [WRP^{+}12]. Gizzard [FBJR11, Mic10]. Glacially [WBD17, WRP^{+}12]. Glaciated [BBM^{+}14]. Glaciation [CKM^{+}11]. Glacier [MDD^{+}16]. Glass [GPBA18, SJGL16]. Glass-Eel-Stage [SJGL16]. Global [ILH^{+}18, Lev18]. Gloves [ARB^{+}18]. Glue [JLA17]. Go [SCF16]. Goals [BGLP19, HP13]. Golden [VFFP13]. Goldfish [LCO^{+}15]. Gonad [BS16]. Good [MYS^{+}11, RKEB13b]. Government [Lev18]. Gradient [FWD10, FDWG13, Mic10, MAK12, RSL11]. Gradients [CJG13]. Grand [SRMB^{+}14, WvPC12, YCB^{+}11, CYP11]. Grande [Arc16, BMB^{+}10, HKO^{+}18, HMP18, ZTC12, ZBC^{+}13]. Granite [SMN^{+}10]. Grass [GGC17]. Grasse [TCBI11]. Gravel [FSFS17, TM19]. Gravel-Bed [FSFS17]. Gray [FWSW^{+}15, KSSK^{+}17, SS11]. Grayling [BS13, CHH^{+}16, DHA^{+}17, DMW^{+}19, RPC14]. Grays [RDR^{+}10]. Great [BNS^{+}14, MBG^{+}11, PAB^{+}11, PBM^{+}12, SS15c, WDHW14, BAKT12, BWHF17, BP10, CHT13, CHB^{+}12, FBF10, FMJ11, HML13, IRJ11, JWHJ13, KSF13, KWW14, LRR^{+}15, MM17, MBM10, MLB18, RAR^{+}16, SC13, TH15]. Greater [BB10a, KKB^{+}10, RGM^{+}11]. Green [LEM^{+}11, MBL^{+}18, PGVG15]. Greenland [PDSP^{+}11]. Groove [DMCO11]. Groundfish [DKL^{+}17, GDPW10, MHF12, ZC13]. Grounds [HSA^{+}16, MJS^{+}19, RCC^{+}13]. Group [AJF13a]. Grouper [CABNQ^{+}13, GWG^{+}19, RSVAB19]. Groupers [TA12]. Groups [SBMH12, SS15c, WCQD18]. Growing [LSPB11, MT19]. Growth [ASP^{+}18, AS17, BMGL12, BDS10, BMH19, BSC^{+}10, BS13, BVA11, CTN11, CM10, CT16, CGA13, CT12, CTPM13, CDB10, CPS^{+}15, CSD^{+}14, DEW^{+}14, DKD^{+}15, DWE^{+}19, DBHV14, DMHO12, DF14, EWW13, EBA^{+}19, Eva17, FBF10, FKF^{+}19a, FMS10, GWOD18, GCG^{+}18, GDM^{+}10, GDM^{+}11, GN13, Goo19, GSA^{+}19, HNK^{+}12, HSG10, HD16, HKO^{+}18, HWHF14, HMP18, IBH12, JC16, JPHK10, JSW^{+}15, JMM^{+}11, KGW14, KM14, KCS^{+}19, KSSK^{+}17, KHF^{+}18, KKM11, KYK17, LBD^{+}16, MSD^{+}19, MT19, MSB^{+}18, Mat18, MAH^{+}12, MNB18, MCC^{+}12, MHE12, MBH12, MBPP18, MLC13, MSWD15, NAWM14, NTHR18, OWN11, ODG^{+}11, PBG^{+}11, PK11, PJ11, PPH15, PRK^{+}16, PA10, QSVD11, RBF14, RR13, RBF^{+}11, RJWK10, RBH^{+}12, SDW^{+}17a, SB12b, SPR^{+}16, SSM16, SLG^{+}14, SAW^{+}17, SJ14, TRLG13, TASC17, TRW10, TB14, TES14, TL15, UACS^{+}19]. Growth [VFB^{+}16, WNCF10, WDW11, WB13a, WCD^{+}11, vPWT12, vZNW12]. Growth-Enhanced [PK11]. Growth-Mediated [DBHV14]. Guadalupe [BLCB13]. Guatemala [BMH19]. Guidance [MJH17]. Guide [Ano11b, Ano12b, Ano13b, Ano14b, Ano16, Ano17b, Ano19i, SRMB^{+}15]. Guilds [MRPW19]. Gulf [BAYS16, BCMPDSZ10, BJR14, CABNQ^{+}13, Chr14, CGDL18, DSBS14, DSB^{+}18, GGP^{+}15, GCA^{+}18, GSD^{+}12, Han18, HPG10, JPHK10, MBPP18, MLC13, MS12, PS14, PDH^{+}18, RCM^{+}19, RMMGPB16, RSVAB19, SFB16, SGP16, SS11, SAW^{+}17, SMMK12, SCB15, AKM12, ASP^{+}18, BBC^{+}14a, BBC^{+}14b, DF12, FWSW^{+}15, JPLS13, LKS^{+}12, MHF12, MHPB14, PHG^{+}16, TA12, WZK^{+}15]. Gulf-Strain [DF12]. Gun [GIW^{+}13, TLM^{+}19]. Gut [BSSM12].

H [Chr14, Fle16]. Habitat [AS12, ACRA10, ACRAM11, ACWIK13, AKK^{+}19, AGJ11, AMBM12, BEWM10, BQHR17, BAKT12, BBM^{+}14, BWRA16, BB10b, BGOQ13, BW10, CTK^{+}17, CAB11, DGMC17, DG13, DHA^{+}17, DWG14, DFB15, DW16, DMW^{+}19, DW13, DF12, DSB^{+}18, DA16, EWW13, ELC13, EBMV10, FBF10, FPD12, FKH15, FFYI16, FS16, GPBA18, GCM12, GKG^{+}19, GZB12, HPF^{+}16, HSSS16, HKKS14, HK12, HGG^{+}16, HGSG17, HPSA16, HM17, JZS^{+}19, JADJ^{+}14, KBO^{+}12, KWW14, KCMS11, LKP^{+}13, Lov11, MM17, MAH^{+}12, MNM^{+}17, MSS^{+}11, MLS^{+}13, MFH13, MKA10, MD11, MWKS13, MK15, NAW12, NDP16, NDW10, NCW^{+}17, PSH^{+}19, PKL^{+}11, PHG^{+}16, PB17, PTG^{+}10, PCC^{+}13, PPMJ14, PPW^{+}14, PGVG15, PL13, RB12, RHW15, RSL11, RA10, RMM11, RLL^{+}19, SHM^{+}12, SMY11, SS11, SR15, SMS^{+}12, STK12, SL16, SCB15, TD13, TRW10, TFM^{+}18]. Habitat [UNTA10, WDR16, WCZT11, WZCT12, WSQ^{+}15, WDWC10, WQD14, WS15, WZK^{+}15, WRP^{+}12]. Habitat-Specific [DSB^{+}18]. Habitats [BP13, BDM^{+}10, CGL^{+}11, CR10, DBS^{+}10, DHB15, Fre19, LDP19, MWS10, MJS^{+}19, NRZ^{+}13, RL14, RDH11, SJT^{+}16, SAW^{+}17, TES14, WDW11, WBMG10]. Habits [DG14, NCW^{+}17, TRHHOGSJ14]. Haddock [BSM13, BJR14]. Hake [Han11]. Hakes [Fle16]. Half [BSC^{+}16, HWD14, PHM17]. Half-Lives [BSC^{+}16]. Half-Pounder [HWD14, PHM17]. Halibut [MHF12, dHST13]. Haliotis [BRS^{+}12]. Hall [Chr14]. Hall-Arber [Chr14]. Hallock [Qui16]. Handling [ARB^{+}18, EB16]. Haplotypes [RHKD12]. Hard [TVL^{+}12]. hardcover [Ber13, Cha13, Fle16, Gid14, Nel14a, Pet15, Pra14]. Hardened [BT16, TT17]. Harmful [RGP17]. Harrison [BNS^{+}16]. Harvest [CS16, MSH^{+}14, SDSB18]. Harvest-Induced [CS16]. Harvesting [SAF^{+}19]. Hatch [AS17, KGW14, SPR^{+}16]. Hatchery [Ano11a, BBGB15, BSS18, Bea10, BHS^{+}17, BGV11, BCF^{+}11, BRK^{+}18, BCM^{+}19, CCH15, CEFQ11, CML19, CFHE11, CWP^{+}19, DWE^{+}18, DML^{+}10, EWW13, FPM15, HLB14, HSS17, HBTT11, HSF12, HSA^{+}16, HM17, JSW^{+}15, KMS^{+}18, KPS^{+}19, LHS^{+}13, LMD^{+}17, LFB^{+}15, Lis14, Lis15, MM17, MSWS10, MFAS^{+}12, MM10, MHE12, PB15, PT10, QSJ12, SKP^{+}10, SCF16, SFLH14, SLG^{+}14, SZZ10, SNI^{+}13, TB16, TPLA16, TC11, TLM^{+}19, WM12, WOM14, ZBW10]. Hatchery- [DML^{+}10, LHS^{+}13, MSWS10]. Hatchery-Enhanced [Bea10]. Hatchery-Origin [HSF12, HSA^{+}16, WM12]. Hatchery-Produced [QSJ12]. Hatchery-Reared [SZZ10, SNI^{+}13, TB16, TPLA16, TLM^{+}19, ZBW10]. Hatching [TM19]. Hawaii [HLB^{+}16]. Head [BW10, KVA^{+}19, SH12, SHC^{+}16]. Headwater [ACRA10, ACRAM11, BSGB17, CRH10, FRM18, HEW12, KLR^{+}15, MBEvdB^{+}10, WKS^{+}19]. Headwaters [RSL11, SHR^{+}12]. Healing [CHW^{+}13, IBH12, LBD^{+}16, PBG^{+}11]. Hearing [CJGH13]. Heat [DMRK14]. Heavily [WHF^{+}12]. Heavyweight [CCB16]. Heithaus [Shi13]. Held [CM10]. Hemorrhagic [EFC^{+}11]. Herring [BDB^{+}15, EB16, MHHH10, OSA^{+}17, RRG^{+}17, SNP^{+}10, TL14, TL15]. Hickory [MR11]. Hierarchical [GHPS^{+}16, HJS^{+}16, MWT14, SL16, WBR11]. High [HPG10, HNV^{+}18, KAH16, KVA^{+}19, MSW^{+}10, SRMB^{+}14, SHC^{+}16]. High-Elevation [HNV^{+}18]. High-Head [KVA^{+}19, SHC^{+}16]. High-Release [KAH16]. High-Resolution [HPG10, MSW^{+}10]. Higher [RMC^{+}17]. Highly [CBW^{+}16, DSRA12]. Hill [BBAK19]. Hills [PPF^{+}16, PKP^{+}17]. Hind [ZHWR11]. Histo [JLA17]. Histo-Glue [JLA17]. Histocompatibility [CPP^{+}14]. Histological [RNG^{+}13]. Historic [ZTC12]. Historical [CKM^{+}11, FSRK11, HB18, HSE^{+}12, KGC^{+}12, MNS^{+}19, WKC18, Var13]. Historically [Han11, HS13]. Histories [CBVF15, HCS^{+}16, JJC12]. History [AKM12, ADS^{+}11, BNS^{+}16, BCRP13, BBC14c, BHP^{+}15, BCM^{+}19, CKC^{+}12, CJG13, DBHV14, DMHM^{+}18, FBWZ19, FH17, GA17, HRB^{+}16, HDBS14, HP13, HWD14, HWD^{+}16, HS13, JWHJ13, KBG^{+}18, LDLH15, LHS^{+}13, LMD^{+}17, LCO^{+}15, MSBT10, MSWS10, PJ11, PHM17, PTG^{+}10, QSVD11, SCC17, SWH^{+}17, SMS^{+}12, Stu18, TKC^{+}18, TTW^{+}11, VBJS10, WAW11, WDW11, Yas16, ZSVK13, ZKB^{+}15, dHST13]. Hitch [BYM19]. Holes [GWG^{+}19]. Holiday [AFHF12]. Home [BAHP11, PS14, SK11]. Homing [BQF^{+}19, DML^{+}10]. Hood [CKFQ11, MBT10]. Hooking [HBC15]. Hoop [SL16, Sto10]. Hoop-Net [Sto10]. Hoop-net-Based [SL16]. Hopkins [Qui16]. Horizon [MHPB14, SFB16]. Horizontal [SFF^{+}10]. Hormone [GWOD18, GPBA18, MSWD15]. Hormone-Coding [GPBA18]. Hornyhead [HMW19]. Horseshoe [MBP^{+}11]. Host [PM10, SPCJ11]. Hosts [MOS13]. Householders [AJF10]. Hudson [BAHP11, PSH^{+}19]. Humpback [CCL^{+}19, SPH^{+}15, WMS15]. Humphries [Ber13]. Huron [GRDO13, JKP10, BAKT12, BBS12, CBMR14, IHTF11, OTR^{+}14, PRK^{+}16, RA10, RSBF14, SJ14, WLLCF16, WLW16]. Hurricane [vVO10]. Hybrid [BDB^{+}15, Nel14a, ODG^{+}11, TBGM12]. Hybridization [BDTD17, BGOQ13, DWN^{+}15, ER18, GLM^{+}15, JHK^{+}19, KCMS11, LCHB^{+}18, SDK^{+}10, SLM10]. Hybridizing [DSHG10]. Hybrids [MKHC17]. Hydraulic [BGV11]. Hydro [BCG^{+}12]. Hydro-Turbine [BCG^{+}12]. Hydroacoustic [BSCA17]. Hydroacoustics [OSS^{+}17]. Hydrocarbon [MHPB14]. Hydroelectric [EWSR16, MBPA13]. Hydrogeographic [BJC12]. Hydrographic [FRRCGA^{+}15]. Hydrokinetic [BES15]. Hydrologic [GJC^{+}18]. Hydrological [MC14, WBD17]. Hydrologically [STK12]. Hydrology [BKC^{+}19]. Hydropower [SZGS12]. Hydroprogne [ARL^{+}12]. Hydroturbine [PBC12]. Hypothesis [MNM^{+}17]. Hypoxia [AW19, BAR15, DMW^{+}19, GCA^{+}18, RBF^{+}11, SCB15]. Hypoxic [RCM^{+}19].

Ice [BDM^{+}10]. Ice-Covered [BDM^{+}10]. Icelandic [AHS10]. Ichthyophoniasis [ZKH12]. Ichthyoplankton [CHPG12, HPG10, SWB^{+}16]. Idaho [CML19, CM11b, KMS^{+}18, KU12, KCMS11, KPS^{+}19, SK11, UK14, WSQ^{+}15]. Identification [ACB^{+}14, BFK15, BCP^{+}11, CKC^{+}12, DPS14, FVK^{+}18, GDPW10, GGP^{+}15, GBB^{+}14, HWH10, JHK^{+}19, MJHP18, MSW^{+}10, MBBM10, PHFN17, SMO^{+}14]. Identifies [MGP11]. Identify [BDM^{+}10, GMP^{+}11, TFM^{+}18, WRP^{+}12]. Identifying [BBM^{+}14, CM11a, FJJ^{+}12, FSRK11, TLM^{+}19, VHPP19]. Identity [LS10a, TM19]. II [PM10, CPP^{+}14, RKEB13b]. III [KR18, YSC10]. Illinois [KMM^{+}11, Pra14, HCD10]. Image [GKBG12]. Image-Analysis-Based [GKBG12]. Imagery [DFB15]. Images [BFM10, MVK12]. Imaging [OSA^{+}17]. Immature [HSM^{+}10]. Immediate [TBC^{+}13]. Immediate-Release [TBC^{+}13]. Immigration [PHG^{+}16]. Immobilization [ARB^{+}18, NYCC19]. Immune [TASC17]. Imnaha [BCF^{+}11]. Impact [AS12, BHS^{+}17, BP10, EWSR16, MGWP10, MMY12, SSR^{+}17]. Impacts [AUDL18, BGLP19, BDTD17, BS18, CORM13, FPD12, FBB^{+}13, IRJ11, JBV13, MDT10, SBO12, TD13, THL^{+}17, VFFP13, WAD^{+}17, ZRM^{+}16]. Impaired [WDR16]. Impairment [CCB16]. Impairs [ZZD10]. Impassable [HSM12]. Impedance [DYWY17, GvdHBC12, HH11, HHJH17, VSC15]. Impede [LRW13, WZC^{+}14]. Impediments [HPSA16]. Imperiled [FK18, FWBT17, GS16, HMW19, MKA10]. Implant [HB17]. Implantation [DBB^{+}12, JDS15, MRB^{+}18, MGW13, SHH^{+}19]. Implanted [CHR^{+}17, DBC^{+}10, DM11b, FMS10, KHF^{+}18, NB12]. Implanting [RPVS15, SNP^{+}10, TPA14]. Implications [Ada18, BLCB13, BS18, BWRA16, BCM^{+}19, BMTP13, CHB^{+}12, CBSB15, CPS^{+}15, DPB13, DCM10, FMJ11, FBS^{+}14, FBWT19, HGG^{+}16, JAA^{+}13, JMM^{+}11, JBV13, LRH12, LPM^{+}17, MCO^{+}10, PBC12, PDM^{+}14, RRA^{+}15, RG15, SA16, SJMG14, SDSB18, SHC^{+}16, SPH^{+}15, SPCJ11, TBTS12, TJB^{+}14, TF13, WBM^{+}17, ZTC12]. Importance [BMM10, CVB14, MNS^{+}19, RL19, TAM15]. Important [HCR14, HGG^{+}16, WS15]. Impounded [BW10, GBW16, HSM12]. Impoundment [HS13]. Impoundments [HHES13, KTE^{+}14, SSR^{+}17]. Imprinting [HDQ^{+}17]. Improve [CGL^{+}11]. Improved [FWSW^{+}15, STZ13]. Improvement [BWRA16, WCZ^{+}18]. Inbreeding [BBAK19]. Incidental [BMC^{+}13]. Incision [BHC^{+}13, BDBE11, PBG^{+}11]. Incisions [JLA17]. Inconnu [Stu18]. Incorporating [HGS^{+}16, TP11]. Increase [WHC^{+}10]. Increased [AHS11a]. Increases [BRGR10, BRK^{+}18, HEW12, MBG10, PPJ15]. Increasing [IMO^{+}14, MJH17, TNS^{+}14, VJGV11]. Incubation [KKM11]. Incurred [RCC^{+}13]. Independent [ABS^{+}11, AJF13a, GKG^{+}19]. Indeterminate [GDG^{+}15, HMO14]. Index [ACRA10, CYW^{+}19, IB14, Lov11, MAK12, PKP^{+}17]. Indexing [HAG13]. Indian [CAB11]. Indiana [HDH16, MFH13]. Indicate [BH11, FDWG13, QSJ12]. Indications [ZHWR11]. Indicator [SMO^{+}14, SNC^{+}10]. Indicators [ARB^{+}18, BS16, KH13]. Indices [ABS^{+}11, DA16, LMD^{+}17, PAB^{+}11, PMN19, PDM^{+}14, SF17, vS10]. Indigenous [Gre12]. Individual [Dud19, HERC12, JK17, KR18, MGK^{+}18, RKEB13a, RKEB13b, SDBB12, UACS^{+}19, WR14]. Individual-Based [Dud19, KR18, RKEB13a, RKEB13b, SDBB12, WR14]. Individuals [ARC^{+}13, PWP13]. Indo [CORM13]. Indo-Pacific [CORM13]. Induce [VSBB^{+}14]. Induced [CS16, GWB^{+}15, MBP^{+}11, PM10, THL^{+}17]. Induction [TBGM12]. Inertia [SBO12]. Inexpensive [KM10]. Infection [MFR15]. Inference [CHLN15]. Inferior [vS10]. Inferred [BL11, HOWS10, STS^{+}14, VB14, WB13b]. Inferring [BA13]. Infers [PTG^{+}10]. Infestations [JZBP16]. Inflow [SGB13]. Influence [ARC^{+}13, AS17, BQF^{+}19, CCH15, CGB19, CEFQ11, FRRCGA^{+}15, HH11, HARL18, HUMG^{+}18, HERC12, HDF12, JC16, KGW14, KM14, KLS17, KS11, KRM11, KRM13, Lis14, Lis15, LCO^{+}15, MPHG10, MP10, MBPA13, MSS^{+}11, NTHR18, OWP13, PBG^{+}11, PW10b, PWP13, PK11, RPSB14, RJWK10, RHR^{+}15, SMY11, SCW^{+}13, SM13a, SHP^{+}19, SPH19, TRLG13, TRMC15, TW14, TPBB14, TF13, VHJ^{+}15, WDR16, WCZ^{+}18, WQD14]. Influenced [HRHS16, VSC15, WBD17]. Influences [BBGB15, BDS10, BHD^{+}16, CKM^{+}11, DSM18, HMTC12, JPFB13, KCMS11, MSK15, MBH12, PTWK11, PKL^{+}11, PTPS10, RR13, TB16]. Influencing [GGS10, KVA^{+}19, MNB18, MD11, MSP12, PPF^{+}16, SAS17, TTM^{+}15]. Inform [PSH^{+}19]. Information [Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g]. Informs [MCD17]. Inhibit [BNS^{+}14]. Initial [CTP^{+}12, TPLA16]. Injected [LBD^{+}16]. Injections [SSH11b]. Injuries [BHC^{+}18]. Injury [BCG^{+}12, ZZD10]. Inland [MWSS13, Nel14a, NDP16]. Inner [WMW^{+}12, PDH^{+}18]. Inorganic [WHC^{+}10]. InSAR [TFM^{+}18]. Insects [HWH10]. Insertion [DEP^{+}18]. Inshore [PDSP^{+}11]. Insight [TRW10]. Insights [ACDL19, BCS19, CMM^{+}17, HBB13, HLR^{+}12, HPSA16].

Instability [ODG^{+}11]. Instantaneous [BS13]. Instream [ACWIK13, JADJ^{+}14, MK15, PPMJ14, PPJ15, RMW16, STK12]. Instrument [VSC15]. Intact [BWHF17]. Intake [HGS^{+}16]. Integrated [BBS12, CHR^{+}17, Cla16, DWE^{+}18, DM11b, DMR^{+}10, ELC13, EHR^{+}12, EPT^{+}16, HB17, HLK11, MCC^{+}12, SDW^{+}17a]. Integrating [HEC^{+}15, MRPW19, WB15a]. Integration [HD16]. Integrity [PKP^{+}17]. Intensity [MDT10]. Intensive [BJA^{+}13]. Interaction [HB18]. Interactions [BES15, CW11, CBW^{+}16, DW13, FRA13, Han18, HBB17, LRH12, Mit16, PCFH17, SPR^{+}16, WDDS13, WDWC10]. Interactive [CCB16]. Interannual [IHTF11, RJWK10]. Interconnected [HRHS16]. Interferometric [TFM^{+}18]. Interior [ACRA10, ACRAM11, MKH10, PBB^{+}15]. Intermediate [SPCJ11]. Intermittency [Arc16]. Intermittent [HUSST19, OPHD18]. Intermittently [MSH17]. Internal [GWOD18, JDS15]. Interoceanic [MJGV14]. Interpreting [CBVF15, SS15a]. Interregional [PF11a]. Intersex [KTE^{+}14]. Interspecies [CM11b]. Interspecific [HNV^{+}18]. Interstitial [MYS^{+}11]. Intersystem [PW10a]. Intertidal [DHB15]. Intervals [SMN^{+}10]. Intra [PW10a]. Intra- [PW10a]. Intrabasin [CR10]. Intracoelomic [CHW^{+}13, MRB^{+}18]. Intraperitoneal [NB12]. Intraspecific [HNV^{+}18, MWSS13]. Intrinsic [MT19]. Introduced [BK11, LRH12, LYX^{+}10, MSB^{+}18, MMY12, SBO12, THKG10, VHJ^{+}15, WCD^{+}11]. Introduction [DKS18, JJNB17, SDOK19, SM13b, TNOU11]. Introductions [NB13]. Introgressive [BLCB13, GLM^{+}15, JHK^{+}19, SLM10]. Inundation [RAB^{+}18]. Invaded [SDOK19]. Invasion [ABM17, KRRW18, LDLH15, LCO^{+}15, MBG10, RO16]. Invasions [HML13]. Invasive [BSSM12, CORM13, CHLN15, DAA^{+}16, ER18, FRBA13, FRA13, GS16, GMB^{+}19, KU12, LV14, NAG13, SS13, TD13, UK14, ZBHW16]. Invasiveness [HLH14, JMM^{+}11]. Invertebrate [PHFN17]. Invertebrates [AMM^{+}16]. Investigate [BBG^{+}18]. Investigation [BBL^{+}12]. Investment [OWN11]. Irrigated [WHF^{+}12]. Irrigation [SO12]. Island [LSSYL10, MHF12]. Islands [GRS12, Spi12]. Isle [GSJS17, MBZ^{+}14]. Isolated [CRH10, OWP^{+}17, WKS^{+}19]. Isolation [KRRW18, SDK^{+}10, SHR^{+}12, SE11, SSR^{+}17]. Isotope [BL11, CHLN15, OPS^{+}14, WB13b]. Isotopes [PBW16, QSJ12, RMM11, VSBC13]. Isotopic [GDPW10, HUMG^{+}18]. Issue [Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g]. Issues [TBC^{+}13]. Iteroparity [CBAC19]. IVb [EFC^{+}11].

J [Chr14, Var13]. Jack [JJC12]. James [Nel14a, BMGL12, BGV^{+}12]. Jams [FSFS17]. Japanese [STS^{+}14]. Jeffrey [Shi13]. Jennings [LDA^{+}13]. Jeremy [Chi13]. Jim [YITI12]. Joaquin [FSP^{+}17, LHHL17, PBB^{+}15]. John [DCS^{+}17, Shi13, TL17, UWL12a, UWL12b, UPT^{+}13, Zyd11]. Johns [MHHH10, Qui16, TAM15]. Jolly [CBSB15]. Juan [BBF^{+}18, BRS^{+}12, CGM15, DF14, GP12, MBT10, PFG19]. Jumping [FMJ11]. June [KWCI10]. Juvenile [ABS^{+}11, ARL^{+}12, ASJ15, AAD14, AIC13, AS17, BMF^{+}15, BP16, BFK15, BLG^{+}13, BAHP11, BBC^{+}14a, BBC^{+}14b, BQHR17, BCRP13, BRS^{+}12, BDBE11, BGW^{+}17, BES15, BCG^{+}12, CJM^{+}15, CBVF15, CT16, CRC^{+}16, CC16, CHW^{+}13, CVB14, CBM15, DEW^{+}14, DWE^{+}18, DBC^{+}10, DBB^{+}12, DH11, DBS^{+}10, DJ17, DHB15, DF14, EBG18, EPW15, EGAW12, ETC18, EHR^{+}12, EHC^{+}14, EPT^{+}16, FCG^{+}17, FBS^{+}14, FRM18, FP19, FBH^{+}13, GGS10, GvdHBC12, GDM^{+}10, GDM^{+}11, GJC^{+}18, GNMM12, HPF^{+}16, HSSS16, HGBE16, HOGO10, HJM^{+}12, HTR13, HBB17, HWH10, HSG10, HER^{+}11, HERC12, HPHB10, JBC^{+}12, JPSM12, JAA^{+}13, JADJ^{+}14, JM14, LMD^{+}17, LRM^{+}10, LJHL10, LBD^{+}16, MGWP10, MC14, MSH17, MSS^{+}11, MLS^{+}13, MBH16, MCC^{+}12, MJH17, MFR15, MBL12, MGW13, MAN^{+}12, MCT17, MCH^{+}13, MCP^{+}15, MK15, NPW12, NRZ^{+}13, NBC15, NLF13, OPHD18, PBG^{+}11, PNM^{+}10]. Juvenile [PPF^{+}13, PBB^{+}15, PPH15, PKL^{+}11, PBC12, PPP^{+}13, PM15, PM10, QHS^{+}13, RSB14, RRA^{+}15, RGM^{+}11, RDR^{+}10, RBH^{+}12, SZGS12, SPC^{+}18, SJT^{+}16, SMJ14, SF15, SGP16, SP10, SSH^{+}11a, SSH11b, SBW14, SO12, SHH^{+}19, SCF16, SDOK11, SHC^{+}16, SS15b, SHP^{+}19, SMMK12, SCB15, TA13, TD13, TPA14, TPLA16, TCBI11, TTW^{+}11, TL15, VBJS10, WMS15, WM12, WB13b, WAKH12, WTB10, ZKB^{+}15, vZNW12]. Juveniles [LSPB11, SF17, TKC^{+}12, WCZT11].

kamtschatkana [BRS^{+}12, BRS^{+}12]. Kansas [FPD12]. Karluk [Yas16]. Katrina [vVO10]. Keith [Ber13]. Keller [Chi13]. Kelt [PM14]. Kennebec [WS15, WZK^{+}15]. Ketotestosterone [MGK^{+}18]. Key [CGA13, GPBA18, MWG18]. Killifish [Gre12]. King [SFF^{+}10]. Kirikuchi [SDK^{+}10]. Kit [HLH14]. Klamath [BHD^{+}16, PGJ11, CRC^{+}16, FMG^{+}11, HWD14, HWD^{+}16, KHHG13, SMO^{+}14, Str10, Str12]. Klickitat [NZF^{+}11]. Klimley [Pra14]. Kodiak [MHF12]. Kokanee [HGB^{+}16, OWP^{+}17, SBO12, WQD14, WCQD18]. Kootenai [WSQ^{+}15, BQHR17, HP13]. Kootenay [SNI^{+}13]. Kopf [Chi13]. Kuskokwim [Stu18].

L [Chr14, Ryp13]. L. [ACK17]. Laboratory [CM10, LDP19, MDP12, MT19, MAN^{+}12, PPH15, SS13, WBH10]. Laboratory-Based [PPH15]. Lack [BDS10, WBD14]. Lagoon [CAB11]. Lagoons [TFM^{+}18]. Lahontan [BRGR10, HB18, PRW^{+}10]. Lake [AHS^{+}11b, BBGB15, BDS10, BP13, BAHP11, BAKT12, BRGR10, CSED10, CHT13, CM11a, CSD^{+}14, DG13, DMW^{+}19, DCJ^{+}13, DM11a, DMR^{+}10, FPR^{+}18, FSC^{+}13, GRDO13, GSR^{+}11, GSJS17, GS13, GZB12, HNK^{+}12, HML13, HB18, HSE^{+}12, HDH16, IHTF11, JRBC13, JKP10, JBV13, KGC^{+}12, KCS^{+}19, KBG^{+}18, KWS14, KPS^{+}19, LME16, MNS^{+}19, MSWS10, MWSS13, MBPA13, MNM^{+}17, MNB18, MFH13, MD11, MBZ^{+}14, NB12, PNM^{+}10, PTPS10, RB12, RA10, RJWK10, SS18a, SBO12, SH10, SCW^{+}13, SDOK19, SBW14, SJMG14, SCF16, SIKR17, SBC17, SGK16, TA13, TCBI11, THKG10, VCP^{+}14, VFB^{+}16, WDW11, WLLCF16, BAKT12, BR11, BMM10, BBS12, BHG^{+}10, BBG^{+}18, BMR^{+}12, CHT13, DSAM12, ESD^{+}17, EBA^{+}19, Esh14, FBB^{+}13, GS13, HNK^{+}12, HDBS14, HML13, HSE^{+}12, JMB^{+}13, JRBC13, KWCI10, KRM11, KRM13, MBG10, MCO^{+}10, MLBS^{+}13]. Lake [OTR^{+}14, PGW^{+}11, PB16, PRK^{+}16, PDM^{+}14, RCD^{+}11, RMM11, RSBF14, SPR^{+}16, SRP^{+}18, SWC13, SJ14, THKG10, TJB^{+}14, VB14, WLW16, WBD14, YJW11, ZRM^{+}16, ZZMC13, vVO10]. Lake-Scale [MFH13]. Lake-Type [PNM^{+}10]. Lakes [AM13, BB10b, FJJ^{+}12, HAG13, HRHS16, HNV^{+}18, JK17, LSPB11, MSS^{+}11, MFH13, MKH13, PTPS10, PCC^{+}13, PM10, RBK^{+}11, SS13, TBL^{+}17, WSI14, WB13a, WBWSK12, BAKT12, CHB^{+}12, CBMR14, HML13, IRJ11, JWHJ13, KSF13, KWW14, LRR^{+}15, MM17, MBM10, PBM^{+}12, RAR^{+}16, SC13, SS15c, TH15]. Lakes- [BAKT12]. Lakes-Origin [BAKT12]. Lakewide [HBS12, MCO^{+}10]. Lampetra [JSW12]. Lamprey [CHB^{+}12, HB17, KCJ^{+}15, MSW^{+}14, MJHP18, MAN^{+}12, MOS13, RG15, RG16, SS15a, YSW^{+}14]. Lampreys [BMM10, CHB^{+}12, Esh14, Eva17, HJS^{+}16, HCZ13, HBA^{+}15, JCK^{+}12, JSW12, KCM14, MCC^{+}12, MJH17, NAW12, SS18b, SJ14]. Land [DWGF11, GNMM12, HDH16, RBK^{+}11, SWS10]. Land-Locked [RBK^{+}11]. Land-Use [HDH16]. Landings [CYW^{+}19]. Landlocked [HGB^{+}16, WBD14]. Landsburg [BLB^{+}13]. Landscape [ACMM^{+}15, BCH^{+}11, CTPM13, DFB15, FSJ^{+}11, Spi12, WDDS13, WBWSK12]. Landscape-Based [WBWSK12]. Landscape-Scale [WDDS13]. LaPlatte [PPD16]. Large [BAHP11, BBM^{+}14, CHMG^{+}18, CTPM13, CM11b, DK18, DEW^{+}14, DWE^{+}18, EMS^{+}14, GJC^{+}18, HUSST19, JCK^{+}12, JSW12, KCZH14, KCJ^{+}15, KKM11, MCH^{+}13, PCC^{+}13, PE14, PRK^{+}16, RHEF19, RPC14, RBZ^{+}17, RA10, SPH^{+}15, SFF14, WRQH17, WDW11]. Large-Bodied [SFF14]. Large-Dam [KCJ^{+}15]. Large-River [CTPM13, HUSST19, RBZ^{+}17, SPH^{+}15]. Large-Scale [CM11b, DK18, JCK^{+}12, PE14, PRK^{+}16, RA10]. Largemouth [ARC^{+}11, ARC^{+}13, BNW^{+}12, BAR15, CBW^{+}16, CCB16, EPW15, FDWG13, GWB^{+}15, GDWD10, GA10, HBS12, HVD^{+}16, HDF12, KTE^{+}14, Lov11, MFH13, NDC^{+}14, NTHR18, NDW10, PW10a, PWP13, PET^{+}17, SPC^{+}18, SA16, SS15b, TPA14, TAM15, VWS13]. Larger [MS12]. Larvae [ESD^{+}17, FRRCGA^{+}15, GNV^{+}14, KWCI10, LSPB11, RAR^{+}16, UPT^{+}13]. Larval [ABS^{+}11, BH11, CHB^{+}12, DA12, EBA^{+}19, Eva17, GRDO13, HJS^{+}16, HGSG17, JPLS13, JSW12, LJHL10, NYCC19, PM10, PDM^{+}14, SS13, TL17]. Latent [SZGS12, SBO12]. Lateral [PPP^{+}13, SMH^{+}11]. Laterally [KHF^{+}18]. Latitudinal [MT19]. Lauer [LDA^{+}13]. Laurel [BBAK19]. Lavage [BN16]. Lawrence [EFC^{+}11, Han18, ZZMC13, MLC13]. Lead [FS16]. Lean [GSR^{+}11, SJMG14]. Led [Lev18]. Lee [Yas16]. Legacy [AW19, PRW^{+}10, PSGB13]. Legumine [ACK17]. Length [ASP^{+}18, BSM13, BFM10, CGA13, DLHF16, DL13, FBWZ19, GBB^{+}14, HS17, KQ11, KK16, RGM^{+}11, SDSB18, TRMC15, THGA15, TB16, ZM10, ZC13, ZZMC13]. Length- [KK16]. Length-at-Age [ASP^{+}18, DLHF16]. Length-Based [SDSB18, THGA15]. Length-Cohort [ZM10]. Lengths [HGP^{+}17, MFH13]. Lentic [LRH12]. Leopard [ACK17]. Lepisosteus [SKBS18]. Lernaeopodidae [MFR15]. Lesions [MHPB14]. Less [AJF10, CHLN15]. Level [BMM10, FQW^{+}10, GA10, NDW10, PM19, SRP^{+}18, THKG10]. Levels [BLG^{+}13, BYM19, BH11, CKS^{+}19, GDM^{+}11, MC14, WHC^{+}10]. Lewis [ACDL19, SHCB16]. Lice [JZBP16]. Life [BB15, BNS^{+}16, BCRP13, BBC14c, BHP^{+}15, BCM^{+}19, CKC^{+}12, CBVF15, CSED10, CDB10, CJG13, CSD^{+}14, DWE^{+}19, DBHV14, DMHM^{+}18, FH17, GA17, HMTC12, HRB^{+}16, HWD14, HWD^{+}16, HS13, HCS^{+}16, JJC12, JWHJ13, KBG^{+}18, LDLH15, LHS^{+}13, LMD^{+}17, LCO^{+}15, MSWS10, PFG19, PJ11, PHM17, PTG^{+}10, SCC17, SBC^{+}16, SWB^{+}16, SCF16, SMS^{+}12, Stu18, TB14, TKC^{+}18, TTW^{+}11, VBJS10, WAW11, WDW11, ZSVK13, ZKB^{+}15, dHST13]. Life-History [PJ11]. Life-Skills [SCF16]. Life-Stage-Specific [HMTC12]. Lift [RC16]. Light [HBS13]. Like [CHH^{+}16, FHCSN12, LMPvdH18, RHEF19]. Likelihood [HSA^{+}16]. Limit [ARC^{+}11, AGKW12]. Limitation [PK11]. Limiting [GNMM12, JRP12]. Limits [AGH18, BGOQ13, Str10]. Limulus [MBP^{+}11]. Lineage [BRG19]. Lineages [BCH^{+}11, NHM10, SE11]. Linear [YJW11]. Linking [HBB13]. Lion [KST^{+}12]. Lionfish [CORM13]. Lipid [SF17, VSC15]. Lipids [HWH10]. Lithobates [ACK17]. Little [Nel14b, PDM^{+}14, BCF^{+}11, FRM18, Sto10]. Littoral [SBW14]. Live [LSV^{+}13]. Live-Release [LSV^{+}13]. Lived [HKO^{+}18, LF15]. Lives [BSC^{+}16]. Livestock [PPJ15]. Load [MKH10]. Loads [WM12]. Lobster [FRRCGA^{+}15, HCR14]. Local [DZW^{+}12, FSRK11, MFH13]. Localized [SA14]. Locate [MSH17]. Location [BQF^{+}19, BWD^{+}17, HH11, PBG^{+}11, TL17]. Locations [CGB19, WB15a]. Loci [HN11]. Lock [SDW15, SH12, YITI12]. Lock-and-Dam [SDW15, SH12]. Locked [RBK^{+}11]. Locomotory [BNW^{+}12]. Logger [KM10]. Logistic [TP11]. Long [AGH18, BGB^{+}16, DSM18, GP12, HSSS16, IHTF11, KM10, KRM13, LF15, MLC13, NDC^{+}14, NR16, OGM^{+}15, PPJ15, RA10, SWB^{+}16, SMH^{+}11, VB14, WHSK19, ZBHW16]. Long-Lived [LF15]. Long-Term [AGH18, BGB^{+}16, DSM18, GP12, HSSS16, IHTF11, KM10, KRM13, MLC13, NDC^{+}14, NR16, OGM^{+}15, PPJ15, RA10, SWB^{+}16, SMH^{+}11, VB14, WHSK19, ZBHW16]. Longfin [NR16]. Longitudinal [BSM13, FRM18, PPP^{+}13]. Longline [RSVAB19]. Longlines [GSD^{+}12]. Longnose [BSSP18, KDS18, MRB^{+}18, SSM16]. Loss [DM11b, HVD^{+}16, HRG12, Mit16, PBG^{+}11, RH14]. Lota [KSB16]. Louisiana [KMKS13, MKK15, RMC^{+}17, vVO10]. Low [AW19, AAD14, ACB^{+}14, BYM19, BH11, BRS^{+}12, BW10, FDSB18, GDST16, Han11, KM10, MSWD15, SDK^{+}10, SH12, UK14, VSBB^{+}14, WOM14]. Low-Cost [FDSB18]. Low-Head [BW10, SH12]. Low-Maintenance [KM10]. Low-Temperature [AAD14, UK14]. Lower [HML13, HM17, REN17, RDR^{+}10, BL11, BSD13, HKKS14, JSW12, JHK^{+}19, MKH13, MWG18, RDH11, RMW16, RT14, SBDW11, SIKR17, SMN^{+}10, SS15c, TKC^{+}12, TLKT17, WSQ^{+}15, WF11]. Lowers [BNC^{+}15]. Lumps [GWG^{+}19]. Lunar [BMM10, LKP^{+}13].

M [Chr14, Ryp13, Zyd11]. Mackenzie [HLR^{+}12]. Macro [MSB^{+}18]. Macro-Scale [MSB^{+}18]. Macrohabitat [BSD13]. Macroinvertebrate [DGMC17, SAH^{+}14]. Macroinvertebrates [KMKS13]. Macrophytes [NTHR18, PAG11]. Macrotidal [DMHM^{+}18]. Madison [BGW^{+}17]. Madtom [MKA10]. Magnitude [FSC^{+}13]. Main [BBS12, HPHB10, OTR^{+}14, WSQ^{+}15]. Main- [WSQ^{+}15]. Main-Stem [HPHB10]. Maine [BJR14, Chr14, FZZ^{+}10, GZB12, GBZ14, HCZ13, HCZG15, HKZ11, LKS^{+}12, NCW^{+}17, OSS^{+}17, RSG12, SZZ10, WCZ^{+}18, WS15, WZK^{+}15]. Maintained [SFLH14]. Maintenance [KM10]. Major [BCH^{+}11, CPP^{+}14]. Male [BVA11, JJC12, KMS^{+}18, LBC10, SKP^{+}10]. Males [ARC^{+}13, MGK^{+}18, PWP13]. Managed [GP12]. Management [ARL^{+}12, BLCB13, DPB13, DKS18, DCBKR14, HP13, LFB^{+}15, LPM^{+}17, MPHG10, Mil19, Nel14a, SDSB18, SM13b, ZTC12, Smi13]. Managing [ACMM^{+}15]. Manganese [TL15]. Manistee [BHK^{+}10, DHA^{+}17, DG13]. Many [SR10, SS13, WB15b]. Mapping [PSH^{+}19]. Margin [GRS12]. Marine [AJFO10, AJF10, BNTT17, CKFQ11, DGMC17, GCG^{+}18, GA17, GWG^{+}19, HMTC12, HWH10, MBT10, MOS13, PGR13, QSJ12, Qui16, SHP^{+}19, TB14, Wei10, Yas16]. Marine-Derived [HWH10]. Mark [Eva17, HLB^{+}16, HRG12, MK16, PGWS15]. Marked [DMHO12, RGM^{+}11]. Marker [AIC13]. Markers [CVL^{+}19, EBF^{+}14, NZF^{+}11, PDSP^{+}11, STS^{+}14, TL14]. Marking [BQF^{+}19, Chi13, HB17]. Marks [SSH11b]. Marlin [Goo15]. Marlins [RBE^{+}10]. Marsh [BT16, MBH16, JPRS10, RC16]. Maryland [Chi13, Chr14, LDA^{+}13, Nel14a, Qui16, Ryp13, Zyd11, AS17]. Mass [PPH15]. Mass- [PPH15]. Massachusetts [HCD^{+}14, DCBKR14, KHD^{+}14, RRG^{+}17]. Match [FH17]. Matching [SBC17]. Material [IBH12]. Materials [BHC^{+}13]. Maternal [BBC14c, SSH11b]. Mating [BSS18, OSDT18]. Matrix [KRM13, SR10]. Matter [JPH17, VSBC13]. Matthew [Chi13]. Maturation [BS16, LBC10, MGK^{+}18, OWN11]. Maturing [FPM15, PM14]. Maturity [CFHE11, Con17, CGDL18, DL13, HNK^{+}12, OWS16, OWP13, PM14, RMC^{+}17, TRMC15]. Maximum [Goo15, HSA^{+}16, Hoe17, PBC12, Smi13]. May [DWE^{+}19, MDT10, MVP15]. McClatchie [Pet15]. McCloud [SLM10]. McKenzie [Chi13]. Mead [DSAM12]. Mean [HGP^{+}17, TB16]. Means [PM19]. Measurements [HH11]. Measures [CEM^{+}19, PMN19, WCYH18]. Measuring [Eva17]. Mechanism [NJ12]. Mechanisms [BMR^{+}12, HM17, LRM^{+}10]. Mediated [DBHV14]. Mediates [SS15b]. Melting [MSW^{+}10]. members [Chi13, Chr14, LDA^{+}13, Nel14a, Ryp13, Zyd11]. Memorandum [Yas16].

Menhaden [ASJ15, ASJ16, ASJ17, AS17, HWHF14, SMJ14, SWB^{+}16]. Menippe [CGDL18, RMC^{+}17]. mercenaria [CGDL18]. Mercury [Ada18, FWD10, HNV^{+}18, SARC13, TA12, WNCF10]. Mesa [Chi13]. Mesohabitat [TCBI11]. Metabolism [CJM^{+}15]. Metamorphosis [ACK17, SJ14]. Metapopulation [PF11b]. Metapopulations [BCH^{+}11]. Method [DMCO11, DJ17, GKBG12, JPRS10, LV14, MH11, NYCC19, PSM12, Poe14]. Methods [DEP^{+}18, MSD18, MVK12, Poe14, PMN19, RMDA11, SC13]. Methylmercury [HSG10]. Metrics [MRPW19, WCN^{+}19, vS10]. Mexico [DF14, GP12, AKM12, BB15, BAYS16, BCMPDSZ10, CABNQ^{+}13, CGDL18, DSBS14, DSB^{+}18, FWSW^{+}15, GGP^{+}15, GCA^{+}18, GSD^{+}12, HPG10, JPHK10, JPLS13, MBEvdB^{+}10, MBPP18, MS12, MHPB14, PS14, PDH^{+}18, RCM^{+}19, RMMGPB16, RSVAB19, SFB16, SGP16, SS11, SAW^{+}17, SMMK12, SCB15, TRHHOGSJ14, TA12]. Michael [Shi13]. Michigan [CBMR14, BR11, BMR^{+}12, BHK^{+}10, DG13, DHA^{+}17, EBA^{+}19, ERL15, GRDO13, HSE^{+}12, JMB^{+}13, JRBC13, KGC^{+}12, KRM11, KRM13, MBG10, MD11, MLBS^{+}13, NAW12, PB16, PM10, RCD^{+}11, TJB^{+}14, WDW11]. Micro [LBD^{+}16, SHH^{+}19]. Micro-acoustic [SHH^{+}19]. Microacoustic [FMS10]. Microchemical [BFK15, FDWG13, GNV^{+}14]. Microchemistry [CBM^{+}13, HPHB10, SWH^{+}17, VBJS10]. Microhabitat [FJP18]. Microsatellite [BCP^{+}11, FSN^{+}13, KHHG13, MGP11, WMW^{+}12]. Microsatellites [BWDM11, GSVS^{+}10]. Microscale [HT13]. Microstructure [VBJS10]. Microtrolling [DJ17]. Mid [MGWP10, RJ11]. Mid-Atlantic [RJ11]. Mid-Columbia [MGWP10]. Midcoastal [FBS^{+}14]. Midcontinent [PAB^{+}11]. Middle [BH11, PTG^{+}10]. Midwest [GSVS^{+}10]. Midwestern [KCS^{+}19]. Migrant [RSB14]. Migrating [CKS^{+}19, CJL^{+}14, DEP^{+}18, HCW14, HKZ11, MGWP10, MJH17, SHP^{+}19]. Migration [BHC^{+}18, BBC^{+}14a, BWD^{+}17, CZK^{+}17, CHMG^{+}18, CTS^{+}15, CVB14, CNM^{+}18, DHRQ10, FP19, GDST16, GA17, HJM^{+}12, HT13, HSS17, HER^{+}11, LUC^{+}15, MPHG10, MJS^{+}19, MSS^{+}11, MJGV14, MR11, OS11, PM14, PPP^{+}13, PGR13, RHR^{+}15, RKF^{+}13, RBH^{+}12, RRG^{+}17, SZGS12, SHM^{+}12, SDBB12, SSH^{+}11a, SHR^{+}12, SH12, Str10, Str12, TTW^{+}11, VCP^{+}14, WZC^{+}14, WAD^{+}17, WL13, ZKH12, Qui16]. Migrations [AHS^{+}11b, HBA^{+}15, TBTS12, VHPP19]. Migratory [BMM10, CEFQ11, CBM^{+}13, HLR^{+}12, PMF^{+}10, RH14, SWH^{+}17, SDW15, SZZ10, TKC^{+}18, WBR11, WMW^{+}12]. Minas [WMW^{+}12]. Miniature [KM10]. Minijack [HLB14, JJC12, SLG^{+}14]. Minimum [AGH18, JLV^{+}10, UK14, SPCJ11]. Minnesota [SCW^{+}13, AMM^{+}16, BS15, DH11, FJJ^{+}12, MMY12, PCC^{+}13]. Minnow [AS12, Arc16, BYM19, BMB^{+}10, DSRA12, FBF10, HKO^{+}18, OMA13, SRMB^{+}14]. Misgurnus [KU12, UK14]. Missing [PVJH15]. Mississippi [BH11, HKKS14, JHK^{+}19, KMM^{+}11, MKH13, PTG^{+}10, SS13, AM13, BSSM12, DF12, DLHF16, KBO^{+}12, SWP^{+}18]. Missouri [ACRA10, EBF^{+}14, RDH11, Mic10, MS13, SM13a]. Mistassini [HNK^{+}12].

Mitigation [WRQH17]. Mitochondrial [CKM^{+}11, CMM^{+}17, HOWS10, LYX^{+}10, MSBT10, Whi12, WMW^{+}12]. Mittelbach [HCD^{+}14]. Mixed [AHS11a, FSN^{+}13, IWB^{+}13, MM10, WBR11]. Mixed-Stock [AHS11a, FSN^{+}13, MM10, WBR11]. Mixing [BAZ12]. Mixture [BEMF11, HRG12, SL16]. mm [CHR^{+}17]. Moapa [DM11b]. Mobility [BFH14]. Model [BMB^{+}10, BRG19, BEMF11, BZ12, Dud19, HBB13, HGS^{+}16, HBHS15, JPFB13, MDP12, MWC^{+}13, MWT14, NAW12, PPH15, RAB^{+}18, SR10, SARC13, SDBB12, TFM^{+}18, UNTA10, WRP^{+}12, YJW11, ZLHS10, vPWT12, RKEB13a]. Modeled [JPLS13, MRPW19, RHW15]. Modeling [AFHF12, DF12, GGC17, Goo19, HS17, HLK11, JBV13, KWS14, KR18, KRM13, MPHG10, MAH^{+}12, PCT^{+}12, PSB16, RKEB13a, RKEB13b, SJ11, SDSB18, UWL12b]. Models [CBSB15, FSJ^{+}11, HHJH17, HD16, IWB^{+}13, JPH17, KCM14, MBM10, MLS^{+}13, PM13, PM15, PVJH15, PHFN17, RPSB14, SL16, WDDS13, WB15a, WR14, WBH10, WTB10]. Modes [CW11]. Modified [CBW^{+}16, MAN^{+}12, MCP^{+}15]. Mohave [ESD^{+}17]. Molecular [BRG19, CCL^{+}19, CEM^{+}19, ESD^{+}17, LSD^{+}14, MCD17]. Momma [SSH11b]. Monitoring [CVL^{+}19, Cha13, LPM^{+}17, MNS^{+}19, NDP16, RRA^{+}15, SFF14, SJFF15, WAD^{+}17, ZC13]. Monkfish [RMDA11]. Montana [SMY11, BGW^{+}17, DMW^{+}19, EPS^{+}15, PDP12, PPC13, PPMJ14, PPW^{+}14, PPJ15]. Montane [BZ12, BAZ12, CM10]. Moon [SHR^{+}12]. Morbidity [TLM^{+}19]. Morice [WAKH12]. Morphological [GWOD18, JHK^{+}19, KHK^{+}15, LMD^{+}17, MWG18]. Morphologically [GGP^{+}15]. Morphology [Ano11a, LSSYL10, MBEvdB^{+}10, ODG^{+}11, PTPS10, RB18, TC11]. Morphometric [GLM^{+}15, HDH16]. Morphometrics [CBVF15]. Morphotypes [CHT13, CPP^{+}14, GSJS17, HNK^{+}12, KWS14]. Mortal [BCG^{+}12]. Mortality [BC12, BMH19, BNTT17, BNC^{+}15, BRK^{+}18, BMR^{+}12, CR10, CFHE11, CGA13, ETC18, FBJR11, FSC^{+}13, FBH^{+}13, GCG^{+}18, HBC15, Hoe17, HGP^{+}17, JPHK10, KAH16, KR18, KRM11, LAR16, LBD^{+}16, MPHG10, MBP^{+}11, PGWS15, PM10, PA10, RJWK10, SZGS12, SPR^{+}16, THGA15, TB14, TPLA16, TP11, TS13, TLM^{+}19, WM12, WF19, YHWH10, ZM10, dHST13]. Motupiko [YHWH10]. Mount [RNG^{+}13]. Mountain [BGW^{+}17, BCR13, FK18, FKF^{+}19a, FKF^{+}19b, PDP12, WK13, AUDL18, BEWM10, EBMV10]. Mountains [ACWIK13, BSGB17, BGLP19]. Mouth [BR11, JRBC13, RSL11]. Movement [BAHP11, BFH14, BSD13, CR10, CTPM13, CNM^{+}18, DPB13, DAA^{+}16, DF14, EBG18, FWB^{+}15, FP19, HGBE16, HUMG^{+}18, JCK^{+}12, JPSM12, KHD^{+}14, KBO^{+}12, LAR16, LEM^{+}11, LRW13, MCW^{+}14, MBPA13, MBH16, MWKS13, NCW^{+}17, PKL^{+}11, PHHM12, RKF^{+}13, SCW^{+}13, TKC^{+}12, TCBI11, WR14, WCN^{+}19, WAR12, WSK11, YHWH10, ZBW10]. Movements [BP13, BW10, CHH^{+}16, CGM15, CKFQ11, DK18, DCM10, EB16, FZZ^{+}10, FFYI16, FH17, GBZ14, PBW16, PPW^{+}14, PS14, REN17, SBW14, SIKR17, WZCT12]. Moving [GKBG12]. MS [BLG^{+}13, PET^{+}17]. MS-222 [BLG^{+}13, PET^{+}17]. mtDNA [MBEvdB^{+}10]. Much [GNV^{+}14]. Mudminnow [KO16]. Multi [KP14]. Multi-Scaled [KP14]. Multibeam [DMR^{+}10]. Multimethod [PVJH15]. Multimetric [IB14, PAB^{+}11, vS10]. Multiple [BW10, CW11, CHT13, FPM13, FQW^{+}10, HMW19, MBH16, MBL12, PPF^{+}16, PF11a, Poe14, SHM^{+}12, SR15, WHC^{+}10]. Multiple-Season [PF11a]. Multiscale [EPS^{+}15, MSK15, PTWK11, RB12]. Multispecies [CYW^{+}19, PSB16, SL16]. Multistage [PMN19]. Multistate [HLK11, PCT^{+}12]. Multivariate [MAK12]. Multiyear [dHST13]. Musick [Shi13]. Muskegon [SWS10]. Muskellunge [KSF13, MMY12, SHC^{+}16, WEW12, WLLCF16, WKC18, WLW16]. Mussel [WBW^{+}13]. Mussels [KS11, MBG10]. Mutagenic [VSBB^{+}14]. mykiss [BCRP13, BBC14c, BCH^{+}11, BWRA16, HBTT11, NZF^{+}11, SMS^{+}12, TRLG13, WBM^{+}17]. Mysis [AHS^{+}11b, PB16]. Myxobolus [FWB^{+}15, KKB^{+}10, PDP12].

N [VSBC13]. Naivety [HVD^{+}16]. Namakan [SCW^{+}13]. Natal [BFK15, BQF^{+}19, SBC17, TL14, WCMP13, ZSVK13]. National [BCS19, MDD^{+}16, SGK16, Yas16]. Native [ABM17, BDTD17, BBF^{+}18, BB10a, BHD^{+}16, BRG19, BGOQ13, CKM^{+}11, DWGF11, DW15, DWN^{+}15, ER18, FBWT19, GP12, HBC15, HB18, HUSST19, JKP10, JBV13, KRRW18, LRH12, LYX^{+}10, MSB^{+}18, MMY12, SRMB^{+}14, SMY11, SPR^{+}16, SBDW11, Sto10, SGK16, TD13, WB15a, WDHW14, WCD^{+}11, WLW16]. Natives [BB10a]. Natural [Ano11a, BEWM10, BGV11, CCH15, CWP^{+}19, DML^{+}10, DSB^{+}18, EBMV10, FPM15, FS16, HBC15, HAR18, Hoe17, HSE^{+}12, LHS^{+}13, Lis14, Lis15, LDA^{+}13, MSD^{+}19, MSP12, RB12, SAW^{+}17, TNOU11, TP11, TC11, TW14, TS13, dHST13]. Natural-Channel-Design [BEWM10, EBMV10]. Natural-Origin [CWP^{+}19, DML^{+}10, LHS^{+}13]. Naturalized [JKP10, MM17]. Naturally [JSW^{+}15, KBO^{+}12, SVKK18]. Nature [CHH^{+}16, FHCSN12, LMPvdH18, RHEF19]. Nature-Like [CHH^{+}16, FHCSN12, LMPvdH18, RHEF19]. Near [BSCA17, SFB16, MCH^{+}13, YSC10]. Near-Coastal [SFB16]. Nearby [QHS^{+}13]. Nearshore [DBS^{+}10, JRBC13, RMC^{+}17, TBL^{+}17]. Need [TBC^{+}13]. Needed [GNV^{+}14, KMKS13, SR10]. Needs [PTG^{+}10]. Negative [IWB^{+}13]. Nekton [FSP^{+}17, JPRS10, RC16]. Neotropical [CBM^{+}13, Var13]. nerka [WAW11]. Nest [LS10b, SAS17]. Nesting [ARC^{+}13, PWP13]. Net [BHC^{+}18, HGS^{+}16, PTPS10, Sto10, SL16]. Nets [PNMH^{+}12, RC16]. Network [ERL15, MLS^{+}13]. Neutral [KHK^{+}15, PDSP^{+}11, PBC12, WKS^{+}19]. Neutrally [JBC^{+}12]. Nevada [BRGR10, DSAM12, ESD^{+}17, PPW^{+}14]. Next [LFB^{+}15]. Next-Generation [LFB^{+}15]. Niche [FRBA13, FRA13, JK17, RB18, WB15a, WB15b]. Niches [DWE^{+}18]. Nile [HAR18, VSBB^{+}14]. Ninespine [MBG10]. Nitrogen [QSJ12]. Nitrophenol [CHB^{+}12]. NMFS [Yas16]. NMFS-F [Yas16]. NMFS-F/SPO-125 [Yas16].

No [TLKT17]. NOAA [Yas16]. Nonequilibrium [HGP^{+}17]. Nongame [MK10, WK13]. Nonindigenous [Gre12, HHES13]. Nonlethal [CCL^{+}19, HOGO10, MSD18]. Nonlethally [DJ17]. Nonmigrant [MH11]. Nonnatal [DF11]. Nonnative [BB10a, BHD^{+}16, BR14, CYP11, DWN^{+}15, ESD^{+}17, GP12, HBB17, HRHS16, HMW19, HGSG17, HWN15, JBV13, KRRW18, KSB16, MWG18, NB13, OMA13, SRMB^{+}14, SMY11, SDOK19, SPR^{+}16, SGK16, WB15a, WDHW14, vZNW12]. Nonneutral [PDSP^{+}11]. Nonsalmonid [LMPvdH18]. Nontidal [DK18]. Nonvisual [DA12]. North [BDTD17, BB10b, CTG14, GN13, Gid14, LKP^{+}13, NG13, NAWM14, SH12, SFH15, Wei10, ACDL19, AMBM12, BJA^{+}13, BWDM11, BRA14, CEM^{+}19, DDMR10, DK18, FBH^{+}13, GMB^{+}19, HARL18, HKO^{+}18, JPHK10, MWT14, MBPP18, PSGB13, PDH^{+}18, RB12, RH14, RBE^{+}10, RLL^{+}19, SHCB16, SDW17b, THNG14, WBW^{+}13]. North-Central [JPHK10, MBPP18, PDH^{+}18, SDW17b]. Northeast [Yas16, DHB15]. Northeastern [MAS10]. Northern [ACK17, CZK^{+}17, CKS^{+}19, FSP^{+}17, GRS12, GIW^{+}13, KGC^{+}12, KBG^{+}18, LDLH15, LN18, OI16, PTPS10, PCC^{+}13, SHC^{+}16, WCD^{+}11, ZBHW16, CHH^{+}16, GRDO13, GGP^{+}15, GCA^{+}18, GSJS17, GSD^{+}12, HPG10, JKP10, JMM^{+}11, MSD^{+}19, MS12, PS14, RCM^{+}19, SS18a, SFB16, SGP16, SS11, SCB15, WSI14]. Northern-Strain [WCD^{+}11]. Northumberland [HCR14]. Northwest [DG19, FJP18, JZBP16]. Northwestern [BJC12, ILH^{+}18, MBG^{+}11]. Notes [CTG14]. Novel [CBW^{+}16, DPS14, EBF^{+}14, GS16, MSW^{+}10, WCN^{+}19, WB15b]. Nuclear [WMG^{+}14]. Nucleotide [CHS^{+}11, EBF^{+}14, GMP^{+}11, HSM^{+}10, LFB^{+}15, MSW^{+}10, NB13, STS^{+}14, SSH^{+}11a, AHS11a, CVL^{+}19, HN11]. Number [BSS18, DBB^{+}12, GDG^{+}15, JCS14, JHK^{+}19, MSS^{+}17]. Numerical [HD16]. Nurseries [BFK15]. Nursery [ASJ16, ASJ17, DGMC17, EBG18, MNS^{+}19, WCZT11]. Nushagak [KQ11]. Nutrient [CS16, KPZ^{+}12, WM12]. Nutritional [SPH19].

O [GDPW10]. Oahe [WBD14]. Oakland [Gid14]. Objectively [GBB^{+}14]. Observation [FBP11]. Observations [FSFS17, HL10, JPH17, KST^{+}12, MNS^{+}19, MR11, WvPC12]. Observed [DMR^{+}10, SMMK12]. Obstacles [SHR^{+}12]. Occupancy [AFHF12, GBF12, JSW12, KO16, PSB16, PVJH15, QHS^{+}13, RG15, RG16, SMMK12, WDDS13]. Occurrence [DW15, KLR^{+}15, MWT14, PDP12]. Occurrences [PVJH15]. Occurring [Han11, KBO^{+}12]. Ocean [BNS^{+}16, BCM^{+}19, CNM^{+}18, FWT^{+}14, Lev18, Qui19, SMO^{+}14, SCC17, TTW^{+}11, AKM12, AMBM12, CTG14, DG19, GMP^{+}11, GN13, JZBP16, NG13, NAWM14, RBE^{+}10]. Oceanic [BA13]. Oceanography [Pet15]. Oceans [RBE^{+}10]. Odd [IMO^{+}14]. Odd-Year [IMO^{+}14]. Odor [SJGL16, SFF^{+}10]. off [FRRCGA^{+}15, KSSK^{+}17, QSVD11]. Offer [TRW10]. Offers [JZS^{+}19]. Office [Yas16]. Offs [BGV11, KRRW18]. Offshore [Ada18, DBS^{+}10, KBG^{+}18, PDSP^{+}11, RA10]. Offspring [BVA11, BBC14c, CGB19]. Ogeechee [JPFB13]. Ohio [PZS17, WKC18]. Oil [RMC^{+}17, RCM^{+}19, SFB16]. Old [RCD^{+}11]. Olfactory [HDQ^{+}17]. Oligohaline [SGB13]. Olympic [KO16]. Omission [Con17]. Omnivorous [CS16]. Oncorhynchus [BCRP13, BBC14c, BCH^{+}11, BWRA16, HBTT11, NZF^{+}11, SPM^{+}17, SMS^{+}12, TRLG13, WAW11, WBM^{+}17]. Oneida [DCJ^{+}13, VFB^{+}16]. Ontario [BMM10, Esh14, FBB^{+}13, GCM12, MCO^{+}10, SCW^{+}13, ZZMC13]. Ontogenetic [DBS^{+}10, HSSS16, Han11, LS16, MHF12, RGP17, RL19, SDBB12]. Ontogeny [LB15, SCB15]. Operating [BSCA17]. Operational [KVA^{+}19]. Operations [PBB^{+}15]. Opercular [BBF^{+}18]. Optimal [DBB^{+}12, HP14]. Options [TBC^{+}13]. Order [Poe14]. Oregon [AGJ11, BHD^{+}16, BCF^{+}11, DSRA12, FBS^{+}14, HGBE16, KBC^{+}18, MSW^{+}14, MFR15, RSL11, VBJS10]. Organic [JAA^{+}13]. Oriental [KU12, UK14]. Orientation [RFM^{+}15]. Origin [BAKT12, BRK^{+}18, BBG^{+}18, CCH15, CWP^{+}19, DF11, DML^{+}10, HSF12, HSA^{+}16, LHS^{+}13, Lis14, Lis15, MSWS10, MM10, OWP^{+}17, RGM^{+}11, RT14, SSH11b, SBC17, TL14, WM12, WHP14, WMW^{+}12, ZSVK13]. Origins [CPS^{+}15, SYH^{+}15]. Ornamental [HLH14, HTH^{+}17]. Oscillations [MJS^{+}19]. Osmoregulatory [ZZD10]. Other [PFG19, AJF10, PTPS10, RBE^{+}10]. Otolith [ASJ15, BBG^{+}18, CBVF15, CC16, CBM^{+}13, DMHO12, FDWG13, HS17, HLR^{+}12, HPHB10, LS10a, NW14, PBW16, RMMGPB16, SMJ14, SYH^{+}15, SBC17, TL14, VBJS10, ZSVK13]. Otoliths [BFK15, BSM13, BSR12, BSSP18, GDPW10, GNV^{+}14, LS16, LS18, TL15]. Otter [BMC^{+}13]. Out-Migrant [RSB14]. Out-Migrating [MGWP10]. Out-Migration [CNM^{+}18, FP19, HER^{+}11]. Outbreeding [DSHG10]. Outcome [CW11]. Outcomes [SDOK19]. Ovary [PSM12]. Overdispersion [MK16]. Overestimates [JHK^{+}19]. Overestimation [HD16]. Overexploitation [ZHWR11]. Overfishing [BHG^{+}10, Lev18]. Overlap [FRBA13, MLBS^{+}13, SRMB^{+}14, WCZT11]. Overlapping [SR15]. Overriding [Sto10]. Oversummer [GNMM12, OPHD18]. Overwinter [FBJR11, Mic10, PM10, SW11, TFM^{+}18]. Overwintering [Bel12, BDM^{+}10, UWL12a]. Ovigerous [DDMR10]. Owens [FPM13, GLM^{+}15]. Oxbow [DM11a]. Oxygen [AS12, AW19, GWB^{+}15, KM10, SSS^{+}12]. Oxytetracycline [DMHO12]. Oxytetracycline-Marked [DMHO12]. Oyster [DHB15, PE14].

Pacific [AMBM12, CTG14, FJP18, AGH18, CN11, CORM13, DJ17, DHB15, EB16, FBWZ19, GMP^{+}11, HB17, JCK^{+}12, JSW12, KST^{+}12, KCM14, KCJ^{+}15, LB15, MHF12, MSW^{+}14, MCC^{+}12, MCT17, MOS13, PKPS11, Qui19, RG15, RG16, SNP^{+}10, SS15a, SS18b, Spi12, WCYH18, WBD17, YSW^{+}14, ZSVK13]. Paddlefish [PGW^{+}11, PGWS15]. Paddlefishat [SDW15]. pages [Ber13, Cha13, Chi13, Chr14, Fle16, HCD^{+}14, LDA^{+}13, Nel14a, Pet15, Pra14, Qui16, Ryp13, Shi13, Smi13, Var13, Yas16, Zyd11]. Pain [AGH18]. Paiute [SPM^{+}17]. Pallid [BH11, EBF^{+}14, HKKS14, JHK^{+}19, KBO^{+}12, PTG^{+}10]. Pamunkey [KHW^{+}14]. Panel [LFB^{+}15]. paper [Chr14, HCD^{+}14, LDA^{+}13]. Paradigm [MJS^{+}19]. Parahucho [FSRK11]. Parameters [PPH15, PM15, SH10, UWL12b]. Parasite [CRC^{+}16, PM10, SPCJ11]. Parasite-Induced [PM10]. Parasitism [FSW10]. Parentage [HSA^{+}16]. Parentage-Based [HSA^{+}16]. Parental [BQF^{+}19, JRP12, NMH12, OWP13, TM19]. Parents [JSW^{+}15]. Park [MDD^{+}16, SGK16]. Parr [CT12, FJP18, JJC12, KRM11, KRM13, VSC15]. Parr-to-Smolt [CT12]. Parrish [Ryp13]. Parsons [Chi13]. Part [PF11b]. Partial [OS11, RKF^{+}13]. Partially [BBC14c, GP12]. Partitioning [MKK15]. parvum [VFFP13]. Pass [HKR18, PF11a]. Passage [BMB^{+}10, BDBE11, BCG^{+}12, EWSR16, FBWZ19, FHCSN12, LMPvdH18, PBC12, RHEF19, SDW15, WCZ^{+}18, YITI12]. Passing [BSCA17]. Passive [CHR^{+}17, Cla16, DM11b, EHR^{+}12, EPT^{+}16, HB17, HLK11, SDW^{+}17a, SFF14, SJFF15, Cha13, ELC13, MCC^{+}12]. Past [MSP12, NB13]. Patagonia [JMM^{+}11, JBV13]. Patagonian [OS11]. Path [ILH^{+}18]. Pathogen [BNC^{+}15, HER^{+}11]. Pathogen-Associated [BNC^{+}15]. Pathogen-Free [BNC^{+}15]. Pathogens [RPSB14]. Pathway [HJM^{+}12]. Pathways [BBC^{+}14a, HBA^{+}15]. Pattern [HGBE16]. Patterns [BFH14, BA13, BLB^{+}13, CBAC19, DHRQ10, EBG18, EPR^{+}18, FDWG13, FPR^{+}18, FWT^{+}14, FBS^{+}14, FRM18, GPBA18, JMB^{+}13, JJC12, KBG^{+}18, KHD^{+}14, LAR16, MSW^{+}14, MBPA13, MBH12, MVP15, MBBM10, NCW^{+}17, PET^{+}17, RG16, RKF^{+}13, RRG^{+}17, RO16, SMM13, SSH^{+}11a, SMS^{+}12, SJFF15, SMMK12, UACS^{+}19, VCP^{+}14, WMV^{+}16, WSQ^{+}15, WB13b, WSK11]. Paul [Ber13]. Peacock [BN16]. Pectoral [BSR12, BSSP18]. Pedigree [HBS12, HCNL10]. Pelagia [AHS^{+}11b]. Pelagic [BGB^{+}16, EB16, FSP^{+}17, GRDO13, OSS^{+}17, SGP16, SBW14]. Pellets [WHC^{+}10]. Peninsula [TRHHOGSJ14]. Pennsylvania [BBAK19, BMP^{+}17, KRRW18, SDW17b]. Penobscot [SZZ10, FZZ^{+}10, GBZ14, HKZ11, OSS^{+}17, RSG12, WCZ^{+}18]. Percent [BRK^{+}18]. Perceptions [WCYH18]. Perch [AGKH17, BPS11, BDS10, BRA14, CM11a, FRBA13, FRA13, GMP^{+}11, GSVS^{+}10, IHTF11, JRBC13, KGW14, LSPB11, MCSF14, MD11, RCD^{+}11, RBF^{+}11, SR10, SSS^{+}12, SB12a, SS15c, VFB^{+}16, WDW11, YJW11]. Performance [BMB^{+}10, Cla16, CHW^{+}13, DBC^{+}10, GSD^{+}12, HP14, JBC^{+}12, JPG^{+}18, JLA17, KVA^{+}19, LHHL17, MDP12, MGW13, PB15, PPF^{+}13, SDW^{+}17a, YSC10, vS10]. Periodicity [BJR14]. Periods [KGC^{+}12, SGB13]. Peripheral [DKD^{+}15]. Peritoneal [BDB^{+}15]. Permanent [HLK11]. perryi [FSRK11]. Persist [HSM12]. Persistence [BZ12, CRH10, GWG^{+}19, HDH16, ZTC12]. Persistent [JAA^{+}13, SE11]. Perspective [MBEvdB^{+}10, MVK12, RCD^{+}11]. Perspectives [CBM^{+}13, MVK12]. Perturbation [VIW^{+}17]. Perturbed [LSPB11].

Pervasive [Nel14b]. Peshtigo [CSED10]. Pesticide [CHB^{+}12]. Peter [Pra14]. pH [LRM^{+}10, WCD^{+}11]. Phase [BMM10, RLL^{+}19, SHR^{+}12, WL13]. Phased [AGH18]. Phased-In [AGH18]. Phelps [Chi13]. Phenological [CHMG^{+}18]. Phenology [BS18, LRR^{+}15]. Phenotypes [BVA11, BRG19]. Phenotypic [BEMF11, CHO^{+}15, DBHV14, LCO^{+}15, LCHB^{+}18, WKS^{+}19, WCQD18]. Pheromone [YSW^{+}14]. Phototaxic [TBL^{+}17]. Phylogeographic [GHPS^{+}16]. Physa [SPCJ11]. Physical [ACRA10, ACRAM11, MAH^{+}12]. Physiological [BMC^{+}13, DYWY17, DEP^{+}18, GWB^{+}15, KH13, LMD^{+}17, TKC^{+}18, TBGM12, VWS13, YSC10]. Physiology [Cla16, MGW13]. Piedmont [CHR^{+}17, THNG14]. Pigmentation [BDB^{+}15]. Pike [Gre12, GIW^{+}13, KBG^{+}18, LDLH15, PTPS10, PCC^{+}13, ZBHW16]. Pikeminnow [SHC^{+}16]. Pikeminnows [DF14, HOWS10]. Pine [BHK^{+}10]. Pink [IMO^{+}14, MSS^{+}17, RCC^{+}13]. Pinto [BRS^{+}12]. pipiens [ACK17]. Piscicide [BSHK^{+}11]. Piscicides [BAZ12]. Piscivore [KM14, RSBF14, SBO12, SRP^{+}18]. Piscivores [KKB^{+}10]. Piscivorous [HBS13, JBV13, LRH12]. Piscivory [EPW15, HBB13, YCB^{+}11]. PIT [ELC13, MCC^{+}12, MKH13, AUDL18, BFH14, SCL^{+}13, VSC15]. PIT-Tagged [SCL^{+}13]. Placement [PPJ15]. Plaice [MLC13]. Plain [MSK15, MP10, OSA^{+}17, THNG14]. Plains [BKC^{+}19, BWHF17, BP10, FBF10, FMJ11, MLB18, MBG^{+}11]. Planktivory [HTR13]. Plasma [BS16, MGK^{+}18]. Plastic [DBHV14]. Platforms [RMC^{+}17, RCM^{+}19]. Played [ER18]. Plume [SFF^{+}10]. Point [Yas16]. Points [RJH12, SR10]. Pollock [MHF12]. Pollutants [JAA^{+}13]. Polluted [WHSK19]. Polycyclic [MHPB14]. Polyhaline [FWSW^{+}15]. polykrikoides [RGP17]. Polymerase [BSC^{+}16]. Polymorphic [SSH^{+}11a]. Polymorphism [CVL^{+}19, CHT13, EBF^{+}14, HN11, LFB^{+}15, NB13, STS^{+}14]. Polymorphisms [AHS11a, CHS^{+}11, GMP^{+}11, HSM^{+}10, MSW^{+}10]. polyphemus [MBP^{+}11]. Polyphyly [HOWS10]. Ponds [BNW^{+}12, LLWW19]. Pontchartrain [vVO10]. Pool [DW16, LMPvdH18]. Pool-and-Weir [LMPvdH18]. Poor [NBC15]. Population [ACMM^{+}15, ACDL19, AKM12, ASP^{+}18, AGKW12, ASJ16, ASJ17, BP17, BYM19, BWDM11, BBAK19, BRA14, BCM^{+}19, BBS12, BMP^{+}17, BLB^{+}13, CS16, CBM^{+}13, CWP^{+}19, DCS^{+}17, DSAM12, FQW^{+}10, FSN^{+}13, FMG^{+}11, GGHS^{+}14, GA10, HLB^{+}16, HKR18, Han11, HSM12, HMSG^{+}11, HLK11, HPSA16, HCS^{+}16, HCNL10, JLV^{+}10, JPFB13, KGC^{+}12, KSF13, KR18, KHHG13, KU12, KRM13, LEM^{+}11, MSS^{+}11, MBL^{+}18, MSP12, NB13, NR16, NDW10, OMA13, OS11, PGJ11, PM13, PHG^{+}16, PGW^{+}11, PM10, RL14, RRA^{+}15, RPC14, RKEB13a, RKEB13b, SR10, SS18a, SMM13, SH10, SB12b, SBDW11, SFLH14, Spi12, SS15c, UWL12a, VHJ^{+}15, VB14, WvPC12, WAW11, WSI14, WRQH17, WF19, WCN^{+}19, WAKH12, WHP14, WOM14, ZBHW16]. Population-Level [FQW^{+}10, GA10, NDW10]. Populations [AMM^{+}16, AGKW12, BBGB15, BSGB17, BGLP19, BLCN^{+}10, Bea10, BSC^{+}10, BBAK19, BJC12, CVL^{+}19, CCH15, CHO^{+}15, CRH10, CVB14, CJG13, CHS^{+}11, DSHG10, DKD^{+}15, DW16, DSRA12, DHRQ10, DA16, Eva17, FWD10, GSD^{+}12, GSVS^{+}10, GMB^{+}19, HLB14, HBTT11, HVD^{+}16, HSE^{+}12, HMSG^{+}11, HEC^{+}15, HWN15, JPLS13, Lis14, Lis15, MSB^{+}18, MHN11, MWSS13, MNM^{+}17, MMY12, MSS^{+}17, NDC^{+}14, OWP^{+}17, OWN11, PRW^{+}10, PPF^{+}13, PPH15, RMC^{+}17, SDK^{+}10, STS^{+}14, SB12a, SMY11, SDSB18, SSM16, TNOU11, VFFP13, WBR11, WKS^{+}19, WLW16]. Portable [ARB^{+}18]. Portion [DHA^{+}17]. Position [RFM^{+}15, TRHHOGSJ14]. Positive [PPMJ14]. Possible [HM17]. Post [MHPB14]. Post-Deepwater [MHPB14]. Postcapture [DDMR10, RHR^{+}15]. Postcopulatory [DK18]. Posthatch [LS18]. Posthodiplostomum [SPCJ11]. Postrelease [HBC15, TLM^{+}19]. Postsmolts [LKS^{+}12, RSG12]. Postspawning [EB16]. Poststocking [HDBS14]. Postsurgical [CHW^{+}13]. Pot [DDMR10, SFF^{+}10]. Potamodromous [FFYI16]. Potential [ARL^{+}12, AUDL18, ARC^{+}11, ARC^{+}13, BDS10, BCRP13, BHP^{+}15, BES15, CHLN15, DDMR10, DEW^{+}14, DCBKR14, DSB^{+}18, FRA13, HWD14, MT19, MBBM10, NAG13, PWP13, PGW^{+}11, PM10, TKC^{+}18, WBW^{+}13]. Potholes [MGWP10]. Potomac [OI16, WL13]. Pounder [HWD14, PHM17]. Practices [CEFQ11, RPVS15]. Pre [HDBS14]. Pre- [HDBS14]. Precision [AHS11a, BFM10, HSF12, KM10, KDS18, VSC15]. Precocious [JJC12, LBC10]. Predation [AS12, ARL^{+}12, BDS10, CSED10, CML19, ESD^{+}17, ETC18, EHR^{+}12, EPT^{+}16, EPC^{+}19, FBP11, HERC12, HEC^{+}15, JJNB17, KST^{+}12, KRM11, KRM13, LSD^{+}14, LB15, MLBS^{+}13, NLF13, PW10b, PK11, RBK^{+}11, SPC^{+}18, SBC^{+}16, SCL^{+}13, SS13, SHC^{+}16, WMS15]. Predator [BPS11, CW11, DW13, EPW15, GS13, HN13, HB18, HEC^{+}15, JBC^{+}12, MAH^{+}12, NPW12, NLF13, OGM^{+}15, SDOK19, SGK16, TJB^{+}14, WDWC10, ZBHW16]. Predator-Specific [HEC^{+}15]. Predators [AGKH17, CW11, HKP11, KBG^{+}18, OMA13, OWN11, WEW12]. Predatory [HSSS16, SBO12]. Predict [BHC^{+}18, GvdHBC12, HGS^{+}16, LSSYL10, WTB10]. Predicting [DW15, PSM12, PM19, PCT^{+}12, PPD16, SR10, SWS10]. Prediction [EPS^{+}15]. Predictive [SARC13]. Predictor [MYS^{+}11]. Predictors [BC12, MBL12, SGP16]. Predicts [FBWT19, MGK^{+}18]. Preferences [CTK^{+}17]. Preliminary [RMW16]. Preparation [GDWD10]. Preparations [RNG^{+}13]. Prerestoration [HKZ11]. Presence [PDSP^{+}11]. Present [MSP12]. Preservation [FH12, OPS^{+}14, VSBC13]. Preserving [PRW^{+}10]. Presmolt [MGW13]. Prespawn [BNC^{+}15, BRK^{+}18]. Press [Cha13, Gid14, Pra14, Qui16, Shi13, Smi13, Var13]. Pressure [GIW^{+}13]. Prevalence [EFC^{+}11, HER^{+}11, MHPB14]. Prey [BAKT12, CW11, CSDW10, DW13, DCJ^{+}13, DA12, EPW15, GJC^{+}18, HBS13, HN13, HB18, JC16, JBV13, KGW14, KLS17, KBG^{+}18, KR18, LJHL10, MBH12, MCO^{+}10, NLF13, NCW^{+}17, PW10b, PW10a, PB16, PHFN17, PGR13, SDOK19, SHCB16, TES14, TAM15, TJB^{+}14, WDWC10, YCB^{+}11]. Primary [JRP12]. Primer [FVK^{+}18]. Principles [Mat18]. Prior [QHS^{+}13, WBM^{+}17, GBZ14]. Prioritization [EPS^{+}15]. Prioritize [WB15a]. Private [AJF10, AJF13a]. Probabilities [CGL^{+}11, HEC^{+}15, MH11, PF11a, PVJH15, PDM^{+}14, SL16]. Probability [MLB18, PCT^{+}12]. Process [NAW12]. Processes [HGS^{+}16, MSP12]. Processing [GNV^{+}14]. Produced [BSS18, QSJ12, THKG10]. Production [BMF^{+}15, BB10a, HGB^{+}16, HTR13, MC14, MVP15, PM19, SFLH14, TTM^{+}15]. Productivity [CCH15, KPS^{+}19, Lis14, Lis15, Mic10, WOM14, WHC^{+}10]. Professionals [LDA^{+}13]. Profiles [DEP^{+}18]. Progenitor [MSWD15]. Progeny [BQF^{+}19, BSS18]. Program [BCF^{+}11, CYP11, CTP^{+}12, LLWW19, MNS^{+}19, ZBHW16, Pet15]. Programs [BGV11, CPS^{+}15, SWB^{+}16]. Progress [TH15]. Projection [SR10]. Promote [HDH16]. Prone [DMW^{+}19]. Pronounced [WLW16]. Propellers [KMM^{+}11]. Properties [PPF^{+}16]. Propofol [VSBB^{+}14]. Proportion [HSF12, HSA^{+}16]. Proportional [WBR11]. Proportions [CCKF11]. Protect [BYM19]. Protected [GWG^{+}19, SPCJ11]. Protocol [BMTP13]. Protocols [BSC^{+}16]. Provide [AHS11a, BFK15]. Provides [CRC^{+}16]. Proximate [Mit16, PM14, SF15]. Proximity [MS12]. Prymnesium [VFFP13]. Pseudogene [MSWD15]. Public [AJFO10, AJF13a, AJF13b]. Publication [Yas16]. Publishing [Ber13]. Puget [RGM^{+}11, CEFQ11, DBS^{+}10, GCG^{+}18, RKF^{+}13, WAFL10]. Pughead [SO15]. Pulse [GIW^{+}13]. Pulsed [BN16, CBM15, SBO12]. Pulses [FH17]. Pupfish [FPM13, SPCJ11]. Purebred [BDB^{+}15]. Purgatory [ILH^{+}18]. Purifying [MKHC17]. Putative [BWRA16, Esh14, YSW^{+}14].

Quality [ARC^{+}13, BPS11, EBMV10, HK12, LHHL17, PWP13, PKL^{+}11, TES14, UNTA10]. Quantification [ELC13, TH15]. Quantify [BSGB17, PVJH15, VIW^{+}17]. Quantifying [ACRAM11, BCG^{+}12, HEC^{+}15, LSJN13, MH11, MLB18, SBO12, SA16, WHF^{+}12, WM12]. Quantile [CTN11]. Quantitative [BSC^{+}16, HCW14, MCT17]. Quantitatively [JPRS10]. Quantity [MNM^{+}17, RB12]. Quebec [HNK^{+}12]. Queen [HAV^{+}19]. Quinton [Chi13].

R [Chi13, Chr14, Shi13, Var13]. Radar [BDM^{+}10, TFM^{+}18, WTB10]. Radar-Derived [WTB10]. Radio [JCK^{+}12, KCM14, TPA14]. Radio-Tagged [JCK^{+}12, KCM14]. Radiocarbon [MLC13]. Radiotelemetry [GCM12, KST^{+}12, TPLA16]. Rain [GWB^{+}15]. Rain-Induced [GWB^{+}15]. Rainbow [BS13, CML19, DWN^{+}15, FWB^{+}15, FFYI16, HP14, IBH12, JJNB17, JMM^{+}11, KKM11, KYK17, LMPvdH18, MGP11, MSW^{+}10, MHE12, MKHC17, OTR^{+}14, ODG^{+}11, PB15, PGJ11, PT10, RL19, SPR^{+}16, TD13, WMS15, WAR12]. Rajidae [MAS10]. Raker [RB18]. Ramp [RHEF19]. Random [SDBB12]. Range [BAHP11, CKM^{+}11, DW15, PF11b, RG16, SK11, SMM13, VCP^{+}14, ZTC12]. Ranges [PS14]. Rangewide [BLCB13]. Ranking [TNS^{+}14]. Rapid [HAR18, HMSG^{+}11, WNCF10]. Rare [NLF13]. Rate [BDS10, Eva17, HLB14, HGS^{+}16, LJHL10, SLG^{+}14, TRLG13, TL15, YJW11]. Rates [BMH19, BVA11, EBA^{+}19, HMTC12, HN13, HBB17, HAR18, HVD^{+}16, HSE^{+}12, HGP^{+}17, KMM^{+}11, KKM11, MH11, NB12, PGWS15, SPR^{+}16, Sto10, WAFL10, ZBW10, dHST13]. Ratio [CPS^{+}15, VSBC13]. Ration [GDM^{+}11]. Rations [PB15]. Ratios [RBF^{+}11]. Raton [Shi13]. Ray [SWH^{+}17]. Rays [BSR12, BSSP18, CMM^{+}17, Pra14]. Razorback [DSAM12, ESD^{+}17, PFG19, ZBW10]. Re [PM15]. Re-Estimating [PM15]. Reach [JZS^{+}19]. Reaches [DW15, REN17]. Reaction [BSC^{+}16]. Reappraisal [GDG^{+}15]. Reared [BJA^{+}13, DWE^{+}19, EWW13, MSWS10, SZZ10, SVKK18, SNI^{+}13, TB16, TPLA16, TLM^{+}19, ZBW10]. Rearing [BCP^{+}11, BHS^{+}17, BWD^{+}17, CEFQ11, CSD^{+}14, DF11, JADJ^{+}14, NRZ^{+}13, PB15, PPP^{+}13, SLG^{+}14, TES14]. Rebuild [AGH18]. Recapture [ASP^{+}18, Eva17, HLB^{+}16, HLK11, HRG12, MK16, PGWS15, WvPC12]. Recaptured [DMHO12]. Recognized [Nel14b]. Recognizing [SSH11b]. Recolonization [LPM^{+}17, PKL^{+}11, WBM^{+}17]. Recolonize [HCZ13]. Recolonizing [BLB^{+}13]. Recompression [BS16]. Reconditioned [JPG^{+}18]. Reconnected [RDR^{+}10]. Reconnection [NDP16]. Reconstruct [SWH^{+}17]. Reconstructing [FSRK11, VBJS10]. Reconstruction [HBS12, HCNL10, PPJ15, SMN^{+}10]. Record [SO15]. Recoveries [EHR^{+}12, Wei10]. Recovery [ARB^{+}18, BGLP19, CGL^{+}11, CHB^{+}12, FCG^{+}17, HGB^{+}16, HEC^{+}15, HPSA16, JAA^{+}13, PB15, PTG^{+}10, PET^{+}17, RFM^{+}15, TBGM12, VB14, YSC10]. Recreational [AJFO10, AJF10, AJF13a, AJF13b, BGB^{+}16, DPS14, KQ11, KAH16, WBR11]. Recruiting [DSAM12, ZZMC13]. Recruitment [ARC^{+}11, ARC^{+}13, BP16, CBMR14, CVB14, DMRK14, FBF10, FWSW^{+}15, GSA^{+}19, JRBC13, JM14, KWCI10, KYK17, MHE12, MS13, MS12, OTR^{+}14, PW10a, PWP13, RCD^{+}11, RAB^{+}18, RMW16, SP10, SA16, SM13a, SS13, SS11, SMMK12, THL^{+}17, VFB^{+}16, WB13a]. Recruits [ASJ16]. Red [BCMPDSZ10, DSB^{+}18, GWG^{+}19, JPLS13, MS12, PS14, PDH^{+}18, RSVAB19, SJ11, SFF^{+}10, SAW^{+}17, SCB15, TS13]. Redband [CM10, DFB15, HSM12, HS13, KCMS11, PGJ11, SLM10]. Redd [Dud19]. Redds [BGV11]. Redfish [KPS^{+}19]. Redhorse [FK18, FKH15, SJFF15]. Redside [BMTP13]. Reduce [MVK12]. Reduced [CRC^{+}16, LRM^{+}10, MWG18]. Reduces [WNCF10]. Reduction [Arc16, LS10b, PPMJ14]. Reductions [FBF10, GWB^{+}15]. Reed [LDLH15]. Reef [GSD^{+}12, JDS15, PE14, SDBB12]. Reefs [DSB^{+}18, PS14, RB12]. Reference [RJH12]. Reflect [RSBF14]. Reflex [BC12, CCB16]. Refuge [BNC^{+}15, BCS19, BNS^{+}14, CRC^{+}16, FJJ^{+}12, JJNB17]. Refuge-Seeking [BNS^{+}14]. Refugia [MSH17, PHHM12]. Regarding [TBC^{+}13]. Regime [BWHF17, GP12]. Regimes [BHD^{+}16, GJC^{+}18, HR15, MSD^{+}19, RB13, RBF14]. Region [BMP^{+}17].

Regional [AGJ11, EPR^{+}18, OWS16, SJ11, VJGV11, Pet15]. Regional-Scale [EPR^{+}18]. Regionalizing [FA10]. Regions [MGP11, MSS^{+}17]. Regression [CTN11]. Regulate [WB13a]. Regulated [FKH15, HK12, KKM11, WRQH17]. Regulating [SWP^{+}18]. Regulation [MRPW19]. Regulations [SDSB18]. Rehabilitation [KWW14]. Reintroducing [SHCB16]. Reintroduction [MNS^{+}19]. Reintroductions [SHC^{+}16]. Reis [Var13]. Reject [AGKH17]. Relate [VFB^{+}16]. Related [BNS^{+}14, CHPG12, DWN^{+}15, HK12, KQ11, MK16, MR11, OPHD18, PT10, SZGS12, TB14, UACS^{+}19, WBD17]. Relation [DM11a, JPSM12, KCJ^{+}15, LKP^{+}13, MSD^{+}19, MK15, PA10, RT14, RB18]. Relations [BK11, FKF^{+}19a, RHW15]. Relationship [BWRA16, DLHF16, EHC^{+}14, HER^{+}11, RZW^{+}12, RMW16, WBW^{+}13]. Relationships [BLCN^{+}10, CM11a, LHHL17, MOS13, RR13, SF15, SM13a, TA12, WSI14]. Relative [BS15, BMM10, BA13, CN11, HHJH17, JSW^{+}15, KDS18, MSWD15, PFG19, PA10, RL19, SCL^{+}13, SWB^{+}16, SIKR17, WF19, WBMG10]. Relatives [Shi13]. Release [BQF^{+}19, CS16, CEFQ11, CFHE11, FM17, KAH16, LUC^{+}15, LSV^{+}13, RHR^{+}15, TBC^{+}13]. Released [BNS^{+}14, KPS^{+}19, SVKK18, SNI^{+}13]. Reliable [CYW^{+}19]. Relief [KCM14]. Relocation [LPM^{+}17]. Remain [CEFQ11]. Remnant [AGKW12]. Remote [DFB15, WRP^{+}12, WTB10]. Remote-Sensing [WRP^{+}12]. Removal [BHK^{+}10, FWB^{+}15, GBZ14, HGB^{+}16, HEW12, HCZ13, HCZG15, JZS^{+}19, PPJ15, QHS^{+}13, STZ13, SDOK19, WCZ^{+}18, WCMP13, WBM^{+}17, WZK^{+}15, ZBHW16]. Removals [RH14]. Removing [LV14, MKHC17]. Repeat [JPG^{+}18]. Replace [BB10a, HKR18]. Replacement [ER18, ERL15]. Reported [AJFO10]. Representative [BCF^{+}13, DA16, FRA13]. Reproduction [BH11, FBWT19, GDG^{+}15, GSJS17, HGSG17, MBPP18, PW10a, RJWK10, SSM16, THL^{+}17]. Reproductive [BVA11, BCF^{+}11, BS16, BGW^{+}17, BCMPDSZ10, BJR14, CABNQ^{+}13, CTG14, Con17, DDMR10, DSB^{+}18, FK18, GSR^{+}11, HAV^{+}19, HM17, JCS14, JPG^{+}18, KSSK^{+}17, KMS^{+}18, LF15, LRR^{+}15, MR11, NDC^{+}14, NG13, OWN11, OSDT18, PSM12, PCFH17, QSVD11, RNG^{+}13, SCW^{+}13, SA14, SE11]. Required [CGA13, HCD10, MNM^{+}17]. Requirements [HGS^{+}16, PB17]. Research [Nel14b, Yas16]. Reservoir [ETC18, FRBA13, FRA13, LAR16, MGWP10, MBPA13, SCW^{+}13, SHC^{+}16, TES14, WF19, WB15b]. Reservoirs [BRA14, HBB17, HARL18, Mic10, MS13, MFR15, PNMH^{+}12, RHKD12, SM13a, SHCB16, SL16, TRW10, TKC^{+}12, VFFP13]. Residence [BNTT17, GCG^{+}18]. Residency [FDWG13, MBH16, WSK11]. Resident [BWRA16, CMM^{+}17, CM10, FRBA13, FPM15, MS12, SHCB16, TAM15, WAR12]. Residualization [MH11]. Resilience [PAA14, RCC^{+}13]. Resistance [Rey16]. Resistant [FWB^{+}15]. Resolution [ACB^{+}14, HPG10, MSW^{+}10, WR14]. Resolving [MHN11]. Resorption [HCW14]. Resource [BMTP13, CHLN15, CBM15, FH17, LDA^{+}13]. Resources [BCRP13, GJC^{+}18]. Respond [SJGL16]. Response [ACDL19, ARC^{+}13, AKK^{+}19, AMBM12, CWP^{+}19, DEP^{+}18, DBHV14, EMS^{+}14, EB16, FPD12, FPR^{+}18, HCZG15, JKP10, JMM^{+}11, LRM^{+}10, Lis15, PHHM12, PPC13, PE14, PPP^{+}13, PKPS11, SHH^{+}19, Str12, TASC17, WRQH17, WCYH18, WZK^{+}15]. Responses [BEWM10, BCF^{+}13, BSCA17, BAR15, FQW^{+}10, GWB^{+}15, HBS13, LRM^{+}10, MWS10, MT19, NDP16, OGM^{+}15, RBZ^{+}17, RBK^{+}11, SR10, SS18a, SDSB18, TBGM12, VWS13, VIW^{+}17, YSC10]. Responsible [CEM^{+}19]. Restoration [AKK^{+}19, BEWM10, DEW^{+}14, EPS^{+}15, EBMV10, FPR^{+}18, Lan16, PPC13, PPMJ14, PE14, WB15a]. Restore [JADJ^{+}14]. Restored [JZS^{+}19, MWS10, PPW^{+}14]. Restraint [CJL^{+}14]. Restricted [GRS12, SR15]. Results [AJF10, RKEB13a]. Resumption [HGB^{+}16]. Retention [LBD^{+}16, PBG^{+}11]. Retrospective [WS15]. Return [BQF^{+}19, MNS^{+}19]. Returning [EMS^{+}14, IMO^{+}14]. Reveal [DWE^{+}18, DBHV14, MLBS^{+}13, SSH^{+}11a]. Revealed [HSM^{+}10, HCNL10, KHHG13, MCSF14, NW14, PDSP^{+}11, RMM11, ZKH12]. Reveals [BQF^{+}19, BBAK19, GRS12, LEM^{+}11, RHKD12, SPM^{+}17, Spi12]. Review [Ber13, Cha13, Chi13, Chr14, Fle16, Gid14, HCD^{+}14, LDA^{+}13, Nel14a, Pet15, Pra14, Qui16, Ryp13, Shi13, Smi13, Var13, Yas16, Zyd11]. Reviewer [Ano18g, Ano19h]. Reviews [Ano10a, Ano10b]. Revisited [ETC18]. Revisiting [Bea10]. Rice [Nel14a]. Richard [Yas16]. Richness [HCD10]. Ring [LS18]. Rings [Rey16]. Rio [Arc16, BMB^{+}10, HKO^{+}18, HMP18, ZTC12, ZBC^{+}13]. Riparian [FQW^{+}10]. Riprap [BT16, TT17]. Riprap-Hardened [TT17]. Riprap-Sill-Hardened [BT16]. Rise [PGS13]. Risk [AS12, ABM17, BHG^{+}10, EPS^{+}15, FBWZ19, HTH^{+}17, HMSG^{+}11, MCH^{+}13, NPW12, PW10b]. Risk-Taking [NPW12]. Risks [HLH14]. River [AHS11a, ACRA10, ACRAM11, ACDL19, AFHF12, Arc16, AS17, BP16, BP17, BMGL12, BGV^{+}12, BP13, BQF^{+}19, BAHP11, BK11, BNS^{+}16, BHD^{+}16, BDB^{+}15, BSCA17, BR11, BWD^{+}17, BHP^{+}15, BKC^{+}19, BGW^{+}17, BDM^{+}10, BCS19, BSD13, BHK^{+}10, BW10, CSED10, CGM15, CTS^{+}15, CTPM13, CAB11, CTP^{+}12, DCS^{+}17, DHA^{+}17, DF11, DWGF11, DWG14, DEW^{+}14, DWE^{+}18, DHRQ10, DF14, EPS^{+}15, EBF^{+}14, EMS^{+}14, FJP18, FZZ^{+}10, FKH15, FRM18, FPM15, FWBT17, FGW^{+}16, GPBA18, GBW16, GP12, GJC^{+}18, GA17, GBZ14, HGBE16, HGB^{+}16, HLB14, HSS17, HBTT11, HCW14, HN11, HK12, HHES13, HGG^{+}16, HCZ13, HCZG15, HKZ11, HUMG^{+}18, HUSST19, HERC12, JPFB13, JRBC13, JRP12, JAA^{+}13, JZS^{+}19, JSW12, KHW^{+}14, KBC^{+}18, KQ11, KK16, KHHG13, KCZH14, KPZ^{+}12, KKM11, KCMS11, LHS^{+}13, LHHL17, LCO^{+}15, MPHG10, MHN11]. River [MSW^{+}14, MHHH10, MSS^{+}11, MPM^{+}14, MBH16, MKK15, MVP15, MFR15, MWG18, NHM10, NCW^{+}17, OS11, OSA^{+}17, PSH^{+}19, PFG19, PBB^{+}15, PB13, PHM17, PDP12, PPC13, PPMJ14, PSB16, PPD16, PA10, QHS^{+}13, RHEF19, RPC14, RBZ^{+}17, RW11, RMW16, RDR^{+}10, RRG^{+}17, SYH^{+}15, SK11, SP10, SJ11, SRMB^{+}14, SS13, SE11, SH12, SFH15, SIKR17, SHCB16, SPH^{+}15, SGB13, Sto10, SJFF15, Str10, Str12, SWP^{+}18, TL17, TB14, TW14, TCBI11, TAM15, TBTS12, TL14, TL15, UWL12a, UWL12b, UPT^{+}13, VFFP13, VBJS10, WDR16, WB15a, WAW11, WM12, WSQ^{+}15, WRQH17, WCZ^{+}18, WAKH12, WS15, WRP^{+}12, Yas16, YHWH10, YITI12, ZBW10, ZBHW16, ZKB^{+}15, ACRA10, Ano11a, BBF^{+}18, BNS^{+}16, BQHR17, BNC^{+}15, BCH^{+}11, BH11, BWD^{+}17, BEMF11, BD15, BLB^{+}13, CKC^{+}12, CCKF11, CRC^{+}16, CFHE11, CYP11]. River [CRH10, CDB10, CBAC19, DML^{+}10, DM11b, DMHM^{+}18, EFC^{+}11, ETC18, EHR^{+}12, EPC^{+}19, EWSR16, FPD12, FPR^{+}18, FWT^{+}14, GDM^{+}10, GDM^{+}11, HMTC12, HKW^{+}14, HPF^{+}16, HP13, HJM^{+}12, HLR^{+}12, HTR13, HBB17, HSG10, HKKS14, HWD14, HWD^{+}16, HS13, JSW12, JHK^{+}19, KBO^{+}12, KPS^{+}19, LPM^{+}17, MGWP10, MC14, MKH13, MOS13, NZF^{+}11, OI16, PZS17, PJ11, PTG^{+}10, PGVG15, PGS13, PGR13, RH14, RDH11, REN17, RT14, SZGS12, SJT^{+}16, SMO^{+}14, SCL^{+}13, SLM10, SFLH14, SZZ10, SPH^{+}15, SWS10, SNI^{+}13, Stu18, TC11, TKC^{+}12, TKC^{+}18, TLM^{+}19, TTM^{+}15, VCP^{+}14, VHJ^{+}15, VFFP13, VHPP19, WL13, WKC18, WBM^{+}17, YCB^{+}11, ZZMC13, ZKH12]. River-Specific [GPBA18, HSS17]. Riverine [CR10, HBA^{+}15, KM10, LB15, PNM^{+}10, SJ11, SSR^{+}17, TES14]. Rivers [BES15, BDM^{+}10, DF12, HBC15, HR15, HPHB10, ILH^{+}18, Lov11, MLB18, NAW12, PAB^{+}11, PDM^{+}14, RHKD12, GSA^{+}19, KMM^{+}11]. Riverscape [FBF10, FBS^{+}14, FFYI16, PHHM12]. RNA [RBF^{+}11]. Road [BP10, ILH^{+}18, MCW^{+}14, TPBB14]. Roanoke [ER18, SFH15]. Robust [FKH15, OS11, SJFF15]. Rock [ER18, HCR14, RHEF19]. Rockfish [BS16, Con17, GMP^{+}11, GRS12, JM14, LF15, PKPS11]. Rocky [ACWIK13]. Rods [Rey16]. Role [BHG^{+}10, DWG14, ER18, Esh14, GNMM12, HHES13, MKH10, MHF12, OMA13, PRW^{+}10]. Roles [Mic10]. Rondeau [GCM12]. Ross [Gid14]. Rotenone [ACK17, BSHK^{+}11, BZ12]. Rougheye [GMP^{+}11]. Roundtail [LRH12]. Route [HT13]. Routine [BNW^{+}12]. Royale [MBZ^{+}14, GSJS17]. Ruin [ILH^{+}18]. Rules [TRW10]. Run [BWD^{+}17, FMG^{+}11, HN11, KHW^{+}14, LMD^{+}17, MHHH10, MBL^{+}18, SCC17, SMN^{+}10, VJGV11, WRP^{+}12]. Runge [Chr14]. Runs [OSA^{+}17, Str12].

S [Nel14a, Var13, BTWB17]. Sablefish [CBM15, EBG18]. Saco [NCW^{+}17]. Sacramento [FSP^{+}17, PBB^{+}15, HSG10, PGVG15]. Safety [BTWB17]. Sagavanirktok [BDM^{+}10]. Sagehen [MDT10]. Saginaw [IHTF11]. Sagittal [BSSP18]. Saint [DCS^{+}17, TL17, UWL12a, UWL12b, UPT^{+}13]. Sakhalin [FSRK11]. Saline [BRGR10]. Salinity [CDB10, FWD10, FDWG13, NDW10, SPCJ11]. Salmincola [Mit16, MFR15]. Salmon [AHS11a, AKK^{+}19, Ano11a, AHS10, BMF^{+}15, BHLK10, BQF^{+}19, BHC^{+}18, Bea10, BWDM11, BCP^{+}11, BBC^{+}14a, BBC^{+}14b, BNTT17, BNS^{+}16, BHS^{+}17, BS18, BNC^{+}15, BVA11, BGV11, BBM^{+}14, BWD^{+}17, BRK^{+}18, BDBE11, BA13, BBS12, BEMF11, BCG^{+}12, BLB^{+}13, BFM10, CBVF15, CZK^{+}17, CKS^{+}19, CHMG^{+}18, CJL^{+}14, CHO^{+}15, CEFQ11, CKFQ11, CRC^{+}16, CFHE11, CC16, CTS^{+}15, CHW^{+}13, CTPM13, CTG14, CVB14, CNM^{+}18, CTK^{+}17, CHS^{+}11, DSHG10, DF11, DEW^{+}14, DWE^{+}18, DWE^{+}19, DBC^{+}10, DEP^{+}18, DML^{+}10, DHRQ10, DBS^{+}10, DJ17, DHB15, EMS^{+}14, ETC18, FBWZ19, FJP18, FSJ^{+}11, FWT^{+}14, FBS^{+}14, FRM18, FPM15, FMS10, FMG^{+}11, FH17, GWOD18, GCG^{+}18, GDM^{+}10, GDM^{+}11, GHPS^{+}16, GJC^{+}18, GA17, GSA^{+}19, HSM^{+}10, HMTC12, HGBE16, HBS13, HGB^{+}16, HLB14, HT13, HBB17, HDQ^{+}17, HSS17, HRR^{+}11, HWH10, HSG10, HN11, HKZ11, HMSG^{+}11]. Salmon [HWN15, HM17, IMO^{+}14, ILH^{+}18, JMB^{+}13, JBC^{+}12, JJC12, JRP12, JPSM12, JAA^{+}13, JADJ^{+}14, KBC^{+}18, KQ11, KK16, KHHG13, KPZ^{+}12, KPS^{+}19, KRM11, KRM13, LKS^{+}12, LBC10, LHS^{+}13, LRM^{+}10, LHHL17, LUC^{+}15, LPM^{+}17, LBD^{+}16, LB15, MPHG10, MKH10, MM17, MWS10, MSBT10, MSD^{+}19, MHN11, MNS^{+}19, MSS^{+}11, MBH16, MM10, MGK^{+}18, MK16, MGW13, MCT17, MCO^{+}10, MCH^{+}13, MCP^{+}15, MSWD15, MSS^{+}17, NHM10, NMH12, NRZ^{+}13, NBC15, OPHD18, PBG^{+}11, PNM^{+}10, PT10, PJ11, PPF^{+}13, PBB^{+}15, PPH15, PKL^{+}11, PBC12, PM15, PGS13, PGR13, QHS^{+}13, Qui19, RCC^{+}13, RSB14, RPSB14, RRA^{+}15, RSG12, RGM^{+}11, RHR^{+}15, RT14, RKF^{+}13, RBH^{+}12, SJT^{+}16, STS^{+}14, SMO^{+}14, SCC17, SK11, SKP^{+}10, SCL^{+}13, SSH^{+}11a, SHH^{+}19, SBDW11, SE11, SFLH14, SLG^{+}14, SZZ10, SVKK18, SMN^{+}10, SHP^{+}19, Str10, Str12, TRMC15, TM19]. Salmon [TC11, TKC^{+}12, TES14, TKC^{+}18, TW14, TLM^{+}19, TTW^{+}11, TTM^{+}15, VJGV11, VBJS10, VSC15, WHF^{+}12, WNCF10, WM12, WCZT11, WZCT12, WCYH18, WAD^{+}17, Wei10, WHP14, WOM14, WBD14, WHC^{+}10, WRP^{+}12, WTB10, Yas16, ZSVK13, ZKB^{+}15, ZKH12, ZZD10, vZNW12, VBJS10, WAD^{+}17]. Salmon-Derived [WM12]. Salmonid [AGJ11, CM11b, Dud19, EGAW12, GBB^{+}14, HWN15, JKP10, LMPvdH18, MYS^{+}11, MC14, MLS^{+}13, PT10, RHW15, RDR^{+}10, SMY11, SDOK19, TNOU11, TD13]. Salmonids [ARL^{+}12, ACWIK13, CN11, DBB^{+}12, EHR^{+}12, EPT^{+}16, Fre19, HOGO10, HJM^{+}12, HGS^{+}16, KST^{+}12, LRH12, MGWP10, MFR15, RSL11, SZGS12, SHM^{+}12, SHC^{+}16, SHCB16, SGK16, TBL^{+}17, UNTA10, vZNW12]. Salmonine [TJB^{+}14]. Salmonines [BB10a]. Salt [JPRS10, RC16, QHS^{+}13]. Salt-Marsh [JPRS10, RC16]. Salvelinus [DPS14, SDK^{+}10]. Sam [Pet15]. Sample [CGA13, DJ17, GDWD10, Hoe17, JPRS10, SC13, ZC13]. Sampled [WBWSK12]. Samples [Goo19, OPS^{+}14]. Sampling [BN16, BR14, BSC^{+}16, CGL^{+}11, CBSB15, FWSW^{+}15, GBF12, HCD10, KCZH14, MAN^{+}12, Nel14b, Poe14, PNMH^{+}12, PDM^{+}14, RC16, SIKR17]. San [FSP^{+}17, KR18, PBB^{+}15, RKEB13a, RKEB13b, ARL^{+}12, BSSM12, BBF^{+}18, BRS^{+}12, BSC^{+}16, CGM15, DF14, GP12, LHHL17, NLF13, NR16, PFG19, SBC^{+}16, SMH^{+}11]. Sand [FPD12, MLB18, PPF^{+}16, PKP^{+}17, RMC^{+}17, Yas16]. Sand-Bed [MLB18]. Sandbanks [RW11]. Sands [SPCJ11]. Sandy [TT17]. Santa [LSSYL10]. sapidus [DDMR10, DK18]. SAR [BDM^{+}10]. Saskatchewan [VCP^{+}14]. Satilla [FGW^{+}16]. Saugers [Whi12]. Saugeye [ZLHS10]. Savannah [BP16, BP17, VHPP19]. Scale [BR11, CBVF15, CM11b, DK18, DW16, DA16, EPR^{+}18, HGBE16, HCW14, HBA^{+}15, HMP18, JJC12, JCK^{+}12, MSB^{+}18, MFH13, PE14, PS14, PRK^{+}16, RA10, SLG^{+}14, TPBB14, TL14, WDDS13, WAKH12]. Scale-Dependent [DW16]. Scaled [KP14]. Scales [BHLK10, BSM13, BSR12, FQW^{+}10, KMKS13, MBH16, MM10, PPF^{+}16]. Scan [FDSB18]. Scaphirhynchus [HGG^{+}16]. Scenarios [FJJ^{+}12, SWS10]. School [SCF16]. Science [HNRB19, Jac18, Zyd11]. Scientific [Yas16, LDA^{+}13]. Scientists [LS10a]. Screen [HTH^{+}17, MGP11, YSC10]. Screened [SO12]. Screening [CCL^{+}19, HLH14]. Sculpin [WCN^{+}19]. Sculpins [WBD17]. sdY [CHO^{+}15]. Sea [BMM10, CHB^{+}12, DJ17, Esh14, HCZ13, HBA^{+}15, JZBP16, KST^{+}12, MJHP18, MJH17, NAW12, Smi13, SJ14, TA12, VJGV11, BCP^{+}11, BBL^{+}12, GRS12, HSM^{+}10, SSH^{+}11a, SS15a, SS18b, Spi12]. Sea-Run [VJGV11]. Seabass [SDOK11]. Seabed [DGMC17]. Seagrass [DSBS14, DHB15, FWSW^{+}15]. Search [TLM^{+}19]. Seascape [Sec19]. Season [BP13, PF11a, WLLCF16]. Seasonal [BHLK10, BAHP11, CHPG12, CAB11, DW16, FWD10, FDWG13, FZZ^{+}10, FFYI16, GBW16, GZB12, HN13, LAR16, LB15, MP10, MBH12, MVP15, PMF^{+}10, QHS^{+}13, RSL11, RL19, SCW^{+}13, TCBI11, TTW^{+}11, WSK11]. Seasonality [HPG10, SCB15]. Seasonally [AHS^{+}11b, FH17, SB12a, TAM15]. Seatrout [AAD14, DLHF16, RPVS15]. Seattle [Yas16]. Seaward [HT13]. Seawater [CHW^{+}13, MCC^{+}12, MGW13, ZZD10]. Seber [CBSB15]. second [Shi13]. Secor [Qui16]. Section [DKS18, SM13b, TNOU11]. Sections [GBW16]. Sedation [TBGM12]. Sedative [BTWB17]. Sedatives [TBC^{+}13]. Sediment [BB15, MKH10]. Seeking [BNS^{+}14]. Segment [ASP^{+}18, BYM19, MBL^{+}18]. Segments [DM11a, HK12, KLR^{+}15]. Segregation [BGOQ13, DCBKR14, MCD17, WCQD18]. Seine [WF11]. Seining [MLB18]. Seismic [GIW^{+}13]. Seitz [Chi13]. Selected [BNW^{+}12, FA10]. Selecting [Goo19]. Selection [AS12, AHS11a, AIC13, BCM^{+}19, BB10b, DWG14, DCJ^{+}13, DML^{+}10, EWW13, GPBA18, HKKS14, KQ11, KBO^{+}12, MGP11, MAH^{+}12, RB12, TRW10, WCZT11]. Selective [BNTT17, CFHE11, GCG^{+}18, KK16, SAF^{+}19, TB14]. Selectivity [CSDW10, MSH^{+}14, PGR13, TP11, WDWC10]. seleniris [SPM^{+}17]. Selenium [HNV^{+}18]. Self [MNM^{+}17]. Self-Sustaining [MNM^{+}17]. Semi [BJA^{+}13]. Semi-Intensive [BJA^{+}13]. Semichronic [KH13]. Sensing [WRP^{+}12]. Sensitive [EB16, NAW12]. Sensitivity [CJGH13, SH10, TNS^{+}14]. Sensor [KM10]. Sensory [ABM17]. Sentinel [DCM10]. Separate [HLK11]. Separation [PL13, RB18]. Septicemia [EFC^{+}11]. Sequence [Whi12]. Sequences [HOWS10]. Sequencing [LFB^{+}15]. Serendipitous [CYP11]. Series [HPG10, WMV^{+}16]. Service [Yas16]. Several [BS15]. Severity [Mic10]. Sex [BS16, BMR^{+}12, CHO^{+}15, CPS^{+}15, LSSYL10, MSD18, MJGV14, MSH^{+}14, SCW^{+}13, YSW^{+}14].

Sex-Biased [MJGV14]. Sex-Determining [CHO^{+}15]. Sex-Specific [BMR^{+}12, MSH^{+}14]. Sexes [SB12a]. Sexual [DCBKR14, PM14, SWC13]. Shad [CTP^{+}12, CEM^{+}19, DMCO11, DMHO12, FBJR11, GBZ14, HTR13, HBB17, HHES13, HMO14, Mic10, MR11, PNMH^{+}12, RH14, SYH^{+}15, SH12, YITI12]. Shallow [CHT13, LBBW17, SJT^{+}16, WB13a]. Shallow-Water [CHT13]. Shape [MVK12, RMMGPB16]. Shapes [PNM^{+}10]. Shaping [CKM^{+}11]. Shared [JBV13]. Shark [GN13, NG13, NAWM14]. Sharks [CTG14, CAB11, GGP^{+}15, Pra14, Shi13, WAFL10]. shasta [CRC^{+}16]. Sheep [BCF^{+}11]. Sheephead [LSSYL10]. Sheepshead [RCM^{+}19]. Shelf [MAS10, PDH^{+}18]. Shenandoah [EWSR16]. Shift [PB16]. Shifting [OGM^{+}15]. Shifts [BCM^{+}19, CS16, DBS^{+}10, HSSS16, KGC^{+}12, LCO^{+}15, MD11, SRP^{+}18, SMH^{+}11]. Shiner [SB12b]. Shoal [DSBS14, RMC^{+}17, RMW16]. Shoal- [DSBS14]. Shore [CHPG12, TLKT17, WQD14]. Shore-Spawning [WQD14]. Shoreline [BT16, DSBS14, WBMG10]. Shoreline-Associated [DSBS14]. Shorelines [TT17]. Short [AGH18, CHB^{+}12, DEP^{+}18, FWB^{+}15, FH12, HKO^{+}18, MBP^{+}11, MGW13]. Short-Duration [MGW13]. Short-Lived [HKO^{+}18]. Short-Term [AGH18, CHB^{+}12, FWB^{+}15, FH12, MBP^{+}11, DEP^{+}18]. Shortjaw [BHG^{+}10]. Shortnose [BP17, FZZ^{+}10, JPFB13, JZS^{+}19, KDS18, PSH^{+}19, PF11b, PB13, UWL12a, UWL12b, UPT^{+}13, VHPP19, WS15, WZK^{+}15]. Should [ARC^{+}13, Hoe17, PWP13]. Shovelnose [PTG^{+}10, SDW^{+}17a]. Shows [CHO^{+}15]. Shrimp [RW11]. Shubenacadie [CDB10, DMHM^{+}18]. Shutdowns [EWSR16]. Siberian [BHC^{+}13, CJM^{+}15]. Sibling [RB18]. Sibship [WCN^{+}19]. Sibship-based [WCN^{+}19]. Sicklefin [FK18]. Side [FDSB18, MC14, RBZ^{+}17, SFF^{+}10, WSQ^{+}15]. Side-Channel [WSQ^{+}15]. Side-Entrance [SFF^{+}10]. Side-Scan [FDSB18]. Signatures [ASJ15, BFK15]. Significance [AMM^{+}16]. Significant [BNTT17, PRW^{+}10]. Sill [BT16]. Silver [EWSR16, GGC17, HGSG17, LLWW19, SPC^{+}18, SWP^{+}18, ZRM^{+}16]. Silverside [BSSM12]. Silvery [Arc16, BMB^{+}10, HKO^{+}18]. Similar [BNW^{+}12, KBG^{+}18]. Similarity [HCD10]. Simple [ZM10]. Simulated [BRGR10, BDBE11, BCG^{+}12, BAZ12, LSV^{+}13, RCC^{+}13]. Simulating [HR15, WR14]. Simulation [GKBG12, RJH12, THGA15]. Simulations [CM10, LV14, MCH^{+}13]. Sinauer [HCD^{+}14]. since [BBS12]. sinensis [FCG^{+}17]. Single [BDBE11, CHS^{+}11, EBF^{+}14, GMP^{+}11, HSM^{+}10, HKR18, HRG12, LFB^{+}15, MSW^{+}10, NB13, PF11a, STS^{+}14, SSH^{+}11a, AHS11a, CVL^{+}19, HN11]. Single- [PF11a]. Single-Nucleotide [CHS^{+}11, EBF^{+}14, GMP^{+}11, HSM^{+}10, LFB^{+}15, MSW^{+}10, STS^{+}14, SSH^{+}11a, AHS11a, CVL^{+}19, HN11]. Single-Pass [HKR18]. Single-Suture [BDBE11]. Siscowet [AHS^{+}11b, GSR^{+}11, SJMG14]. Site [DPB13, DML^{+}10, HCD10, TLKT17, TNS^{+}14, WLW16, YYA^{+}16].

Site-to-Site [HCD10]. Situ [DMR^{+}10, TM19]. Situations [HGP^{+}17]. Six [PKPS11]. Sixgill [WAFL10]. Size [AGH18, ARC^{+}11, BAHP11, Bea10, BBC^{+}14b, BNTT17, BBL^{+}12, CS16, CEFQ11, CFHE11, CC16, CTPM13, CPS^{+}15, EPW15, FH12, GCG^{+}18, GDG^{+}15, GGC17, Goo15, Goo19, HBS13, Hoe17, HMSG^{+}11, IBH12, JLV^{+}10, KM14, LJHL10, MBPA13, Mic10, MBL^{+}18, MSP12, MSH^{+}14, OWS16, OWN11, OWP13, PW10b, PKL^{+}11, PB13, QSVD11, QHS^{+}13, RZW^{+}12, RMC^{+}17, SJT^{+}16, SMO^{+}14, SCC17, SAF^{+}19, SK11, SW11, SAW^{+}17, TRLG13, TB14, TM19, TA12, UWL12a, VSBC13, WMS15, WBD14, ZC13]. Size- [MSH^{+}14]. Size-Class [MBPA13]. Size-Selective [BNTT17, CFHE11, GCG^{+}18, SAF^{+}19, TB14]. Sized [BDS10, MHE12]. Sizes [BGB^{+}16, CGA13]. Skagit [TB14, ZKB^{+}15]. Skates [MAS10]. Skeena [PGS13]. Skeins [AGKH17]. Skill [DBC^{+}10]. Skills [SCF16]. Skin [MHPB14]. Skip [JPG^{+}18]. Skipped [SJMG14, TBTS12]. Slimy [WCN^{+}19]. Small [ACMM^{+}15, AUDL18, AKK^{+}19, BRA14, BAZ12, Cla16, EB16, FMJ11, HW14, HCZG15, JPRS10, KHF^{+}18, MKH10, MBPA13, PCC^{+}13, SRMB^{+}14, SBW14, SAH^{+}14, THNG14, VWS13, vVO10]. Small-Bodied [AUDL18, Cla16, HW14, SRMB^{+}14]. Small-Stream [HCZG15]. Smallmouth [BB10b, CSDW10, EFC^{+}11, ETC18, HPHB10, KGC^{+}12, LV14, LS10b, RO16, WCZT11, WZCT12]. Smallscale [DG14]. Smelt [BSSM12, KR18, LMPvdH18, LFB^{+}15, NLF13, NR16, OTR^{+}14, PMN19, RKEB13a, RKEB13b, SBC^{+}16, SPR^{+}16, YSC10]. Smith [HGBE16]. Smoking [TLM^{+}19]. Smolt [AHS10, BSC^{+}10, BHS^{+}17, CHMG^{+}18, CT12, HGB^{+}16, HDQ^{+}17, SLG^{+}14, SHP^{+}19]. Smolt-to-Adult [BHS^{+}17]. Smoltification [HSM12]. Smolting [LBC10, TB16]. Smolts [CNM^{+}18, HSS17, HKZ11, MBT10, SZZ10, TRMC15, TLM^{+}19, ZZD10]. Smoothhound [GGP^{+}15]. Snake [KCMS11, TKC^{+}12, WDR16, Ano11a, BWD^{+}17, CKC^{+}12, CBAC19, DWG14, ETC18, GDM^{+}10, GDM^{+}11, HMTC12, HCW14, KPS^{+}19, PJ11, TC11, TLM^{+}19, TTM^{+}15, WAW11, WM12]. Snakehead [LN18]. Snakeheads [OI16]. Snapper [BCMPDSZ10, DSB^{+}18, FWSW^{+}15, JPHK10, JPLS13, MBPP18, MS12, PS14, PDH^{+}18, SAW^{+}17, SCB15, TS13]. Snook [DG14, RNG^{+}13, TBTS12, WBMG10, YYA^{+}16]. Snorkeling [ELC13]. SNP [CVL^{+}19, HN11, DPS14]. SNPs [AHS11a]. Social [SAF^{+}19, vZNW12]. Society [Ano13b, Ano14b, Ano16, Ano17b, Ano19i, Chi13, Chr14, LDA^{+}13, Nel14a, Ryp13, Zyd11, SRMB^{+}15]. Sockeye [AHS11a, BQF^{+}19, BHC^{+}18, BCP^{+}11, BBC^{+}14a, BBC^{+}14b, BNS^{+}16, CZK^{+}17, CKS^{+}19, CHW^{+}13, CHS^{+}11, DEP^{+}18, DHRQ10, FH17, HSM^{+}10, HGB^{+}16, HDQ^{+}17, KPS^{+}19, MPHG10, MKH10, MNS^{+}19, MSS^{+}11, PNM^{+}10, PGR13, RHR^{+}15, SSH^{+}11a, SHP^{+}19, TW14, TLM^{+}19, TTM^{+}15, WHP14, Yas16]. Software [MCW^{+}14]. Somatic [MNB18]. Some [JPH17]. Sonar [BFM10, FDSB18, GBB^{+}14, MJHP18, MBBM10, OSA^{+}17]. Sonic [SNI^{+}13]. Sonic-Tagged [SNI^{+}13]. Sound [MVP15, CEFQ11, DBS^{+}10, GCG^{+}18, RGM^{+}11, RKF^{+}13, WAFL10].

Source [BLB^{+}13]. Sources [AJF13a, FSC^{+}13, LLWW19]. South [CTP^{+}12, MVP15, NRZ^{+}13, TRHHOGSJ14, WBD14, AAD14, CEM^{+}19, LBBW17, MSK15, PGW^{+}11]. South-Central [NRZ^{+}13]. Southeast [DSHG10, EBG18, GSA^{+}19]. Southeastern [FRBA13, Ada18, ABS^{+}11, AMM^{+}16, DH11, EPR^{+}18, KSSK^{+}17, LAR16, PPF^{+}16, PKP^{+}17, RSVAB19, SDW15, ZHWR11]. Southern [AKM12, AGKW12, FDWG13, HL10, MBL^{+}18, NW14, PF11b, RG16, TVL^{+}12, VJGV11, WCD^{+}11, BHP^{+}15, BCMPDSZ10, CABNQ^{+}13, CML19, FK18, FKF^{+}19a, FKF^{+}19b, GSR^{+}11, Han18, HMSG^{+}11, KP14, MLC13, RMMGPB16, RJ11, SJMG14, TNS^{+}14, WK13]. Southern-Strain [WCD^{+}11]. Southernmost [SDK^{+}10]. Southwest [HKO^{+}18, OWP^{+}17, PCFH17, SGB13]. Southwestern [BJR14, MD11, WDW11]. Space [WB15b]. Spanning [Mic10]. Spartina [BT16]. Spatial [BHLK10, BR11, BGOQ13, CTK^{+}17, DHRQ10, DCM10, EPT^{+}16, FSJ^{+}11, FQW^{+}10, FM17, HGBE16, HCNL10, HMP18, IWB^{+}13, KMKS13, KHK^{+}15, MCD17, MSB^{+}18, MAH^{+}12, MBH16, MVP15, PZS17, PPF^{+}16, PSH^{+}19, PGVG15, RHW15, RL19, SJT^{+}16, SMO^{+}14, SMJ14, SA14, SS18b, SWC13, SBMH12, SMMK12, TL14, WR14, WBD17, YJW11]. Spatially [Dud19, HBTT11, RBZ^{+}17, VB14]. Spatiotemporal [HGSG17, RO16, WB13b, PHG^{+}16]. Spawn [SVKK18]. Spawner [CCH15, Lis14, Lis15, RSB14]. Spawners [HSF12]. Spawning [BGV^{+}12, BP13, BHC^{+}18, BMM10, BGW^{+}17, BD15, BGOQ13, CGB19, CM11a, Con17, CTK^{+}17, DML^{+}10, EBF^{+}14, FPR^{+}18, FPM15, FS16, GBZ14, GA10, HKW^{+}14, HDBS14, HCW14, HSA^{+}16, HBA^{+}15, HUSST19, HM17, JPG^{+}18, JZS^{+}19, JSW^{+}15, KHW^{+}14, MPHG10, MKH10, MM17, MSW^{+}14, MHHH10, MJS^{+}19, MSS^{+}11, MPM^{+}14, MBL^{+}18, MR11, OSA^{+}17, PDP12, PPP^{+}13, PGVG15, RB12, RCC^{+}13, RHW15, RRG^{+}17, RO16, SNC^{+}10, SKP^{+}10, SJMG14, SHR^{+}12, SH12, SFH15, SBMH12, SFF14, SJFF15, SS15c, SJ14, TL17, TW14, TBTS12, VHPP19, WZC^{+}14, WLLCF16, WQD14, WCQD18, WZK^{+}15, WRP^{+}12, YITI12, YYA^{+}16, ZKH12]. Spawning-Phase [BMM10]. Special [DKS18, LMPvdH18, SM13b, TNOU11]. Specialization [JK17]. Species [ACDL19, AFHF12, BBC14c, BCF^{+}13, CT16, CBW^{+}16, DPB13, DW13, DPS14, DYWY17, DCM10, FBWZ19, FRA13, GS16, GBB^{+}14, HARL18, HRHS16, HCD10, HWN15, LSD^{+}14, MHF12, MCD17, MSW^{+}10, MDT10, Mit16, MWG18, MBBM10, NDW10, PPF^{+}16, PFG19, PL13, PVJH15, PKPS11, SF17, SR15, Sto10, TD13, WDDS13, WB15a, ZSVK13]. Species-Specific [Sto10]. Specific [BBC^{+}14a, BBC^{+}14b, BA13, BMR^{+}12, DSB^{+}18, GPBA18, HMTC12, HSS17, HEC^{+}15, MSH^{+}14, RL14, RAB^{+}18, Sto10, TP11, TTW^{+}11]. Speckled [ZHWR11]. Spectra [BBL^{+}12]. Sperm [BMTP13]. Spheres [Rey16]. Spill [SFB16]. Spiny [DCBKR14, DG19, FRRCGA^{+}15]. SPO-125 [Yas16]. Sport [HARL18, KSB16, LBBW17, PGW^{+}11]. Spot [FBH^{+}13, TLKT17]. Spotted [AAD14, BSHK^{+}11, BSSP18, DKD^{+}15, DLHF16, FS16, GCM12, KDS18, LS18, LCHB^{+}18, RPVS15]. spp [JSW12, RMC^{+}17]. spp. [SKBS18]. Spring [Arc16, BHS^{+}17, DML^{+}10, FJP18, GCM12, HMTC12, HM17, JRP12, LBC10, LHS^{+}13, MGK^{+}18, PPMJ14, PPW^{+}14, PPJ15, SKP^{+}10, TKC^{+}12]. Springer [Pet15]. Springfish [DM11b]. Springs [SFLH14]. sSOD [HRR^{+}11]. sSOD-1* [HRR^{+}11]. St [EFC^{+}11, Han18, MLC13, ZZMC13]. St. [FPR^{+}18, MHHH10, TAM15]. Stability [BEWM10, CJG13, PPD16]. Stabilization [RBZ^{+}17]. Stable [BL11, CHLN15, OPS^{+}14, QSJ12, RMM11, VSBC13, WB13b]. Stage [BB15, CDB10, GPBA18, GA17, HMTC12, KRM13, LV14, PSM12, PFG19, SJGL16, SBC^{+}16, TD13, TB16]. Stage-Structured [LV14]. Stages [CSED10, CSD^{+}14, HDQ^{+}17, PKL^{+}11, TA13, TB14]. Staghorn [WBD17]. Standardization [YJW11]. Standardized [CEM^{+}19, PSM12, Poe14]. Starvation [FBJR11]. State [BBG^{+}18, EPR^{+}18, SHM^{+}12, BRS^{+}12, TBL^{+}17]. State- [EPR^{+}18]. State-Dependent [SHM^{+}12]. State-Straying [BBG^{+}18]. States [Ada18, ABS^{+}11, ADS^{+}11, DW15, FA10, HLH14, HTH^{+}17, MAS10, PPF^{+}16, PKP^{+}17, ZHWR11]. Static [RB13, RBF14]. Statistical [SMN^{+}10]. Statistically [CYW^{+}19]. Status [ACRA10, BP17, BLCB13, BS16, DKS18, FBB^{+}13, FGW^{+}16, HGG^{+}16, MBEvdB^{+}10, MGK^{+}18, MDD^{+}16, PSB16, SCW^{+}13, SPH19, Zyd11]. Steamboat [GWG^{+}19]. Steelhead [AMBM12, BSC^{+}10, BGV11, BCF^{+}11, BHP^{+}15, BCM^{+}19, BBG^{+}18, BGOQ13, CKC^{+}12, CCH15, CJL^{+}14, CBSB15, CBAC19, CWP^{+}19, DBHV14, EHC^{+}14, EPC^{+}19, GGHS^{+}14, GNMM12, HMTC12, HRB^{+}16, HCW14, HWD14, HWD^{+}16, HER^{+}11, HERC12, JPG^{+}18, JSW^{+}15, KH13, LMD^{+}17, Lis14, Lis15, MGP11, MSH17, MBT10, MK15, PM14, PHM17, QSVD11, QSJ12, RL19, SSH11b, SO12, TVL^{+}12, TB14, TB16, VHJ^{+}15, WCMP13, WAKH12]. Steelhead/Rainbow [RL19]. Stem [HPHB10]. Stephen [Gid14]. Stephenson [Chr14]. Stickleback [HBHS15, MBG10]. Stingray [HL10]. Stock [AHS11a, ARC^{+}11, ACB^{+}14, BBGB15, BBC^{+}14a, BBC^{+}14b, CKC^{+}12, CTS^{+}15, EMS^{+}14, FSN^{+}13, HP13, HS13, JAA^{+}13, KHD^{+}14, LSJN13, MM10, NMH12, PDSP^{+}11, RL14, RMMGPB16, RGM^{+}11, RT14, RJ11, SJT^{+}16, SMO^{+}14, SM13a, SJMG14, TTW^{+}11, VFB^{+}16, WBR11, WMW^{+}12, WMG^{+}14]. Stock-Recruitment [SM13a]. Stock-Specific [BBC^{+}14a, BBC^{+}14b, TTW^{+}11]. Stocked [CML19, FKF^{+}19a, FKF^{+}19b, HT13, KM14, KMS^{+}18, NDC^{+}14, WK13]. Stocking [AGKW12, BMF^{+}15, BJA^{+}13, CPS^{+}15, CTP^{+}12, MHE12, MMY12, PGW^{+}11]. Stocks [AGH18, AHS10, ACB^{+}14, CKC^{+}12, DMCO11, GDPW10, HSM^{+}10, RNG^{+}13, SZZ10, VHJ^{+}15]. Stomach [BN16, BL11]. Stone [CGDL18, RMC^{+}17, RCM^{+}19]. Stonecats [PPD16]. Stored [OPS^{+}14]. Strain [DF12, WCD^{+}11]. Strains [BAKT12, HP14, MFAS^{+}12]. Strait [HCR14, MBT10, NBC15]. Strategies [BSS18, HS13, JWHJ13, Str12, ZKB^{+}15]. Strategy [Bel12, HLK11, NPW12, NAG13]. Straying [BWD^{+}17, BBG^{+}18, HSE^{+}12]. Stream [ACRA10, ACRAM11, AKK^{+}19, BS13, BWHF17, CHR^{+}17, Cla16, CVB14, DWG14, DW15, ERL15, FQW^{+}10, FSFS17, HJS^{+}16, HN13, HCZG15, HUSST19, HLK11, HCNL10, JADJ^{+}14, KLR^{+}15, KPZ^{+}12, LBC10, MRPW19, MCD17, MWS10, MSD^{+}19, MJHP18, MP10, MBM10, MLS^{+}13, MKHC17, MKA10, MWT14, NJ12, OSA^{+}17, PPF^{+}16, PM14, PF11a, PCFH17, PKL^{+}11, PPC13, PPMJ14, PPP^{+}13, PHFN17, RSL11, SKP^{+}10, STZ13, UNTA10, UACS^{+}19, WDR16, WNCF10, WZCT12, WCMP13, WHC^{+}10, ZSVK13]. Stream-Dwelling [BS13, STZ13, WNCF10]. Stream-Maturing [PM14]. Stream-Rearing [PPP^{+}13]. Stream-Type [CVB14, LBC10]. Streamflow [FBF10, GNMM12]. Streams [ACRA10, ACRAM11, AGJ11, BEWM10, BSGB17, BGLP19, BP10, BZ12, BAZ12, CRH10, DF11, DMRK14, DH11, EBMV10, FKF^{+}19a, FKF^{+}19b, FRM18, FHCSN12, GNMM12, HEW12, HCD10, KMKS13, KMS^{+}18, KRRW18, KP14, MKH10, MSK15, MBM10, MFR15, MBG^{+}11, OPHD18, PTWK11, PSGB13, RMM11, SPH^{+}15, SAH^{+}14, THNG14, TNS^{+}14, WCZT11, WK13, WTB10]. Streamside [WHC^{+}10]. Strength [CCB16, SWP^{+}18]. Stress [ARB^{+}18, CKS^{+}19, CM10, DMRK14, FBJR11, KH13, RCC^{+}13, TASC17, YSC10]. Stressors [BNS^{+}14]. Stretch [SGB13]. Striped [CDB10, DF12, DA12, DMHM^{+}18, HHJH17, KHD^{+}14, LSJN13, Nel14a, NLF13, PMF^{+}10, SB12b, SH12, SMH^{+}11, TRW10, TBGM12, WSK11]. Stronach [BHK^{+}10]. Strontium [HLR^{+}12, PBW16, SSH11b]. Structural [KVA^{+}19]. Structure [AKM12, AGKW12, ADS^{+}11, BWDM11, BRA14, BR11, BKC^{+}19, BJC12, CHPG12, CS16, CC16, CTS^{+}15, CBM^{+}13, FSN^{+}13, GGHS^{+}14, GHPS^{+}16, HCNL10, KSF13, KHK^{+}15, KHHG13, KP14, KCMS11, LEM^{+}11, LCHB^{+}18, MSK15, MSS^{+}11, MBG^{+}11, PGJ11, PTWK11, PB13, RB12, STS^{+}14, SMM13, SB12b, STK12, SAW^{+}17, SS15c, VHJ^{+}15, VIW^{+}17, WAW11, WSQ^{+}15, WAKH12, WKC18, WLW16, WMG^{+}14]. Structured [BCH^{+}11, BBS12, LV14, SR10, SS11]. Structures [JMM^{+}11, KDS18, MCP^{+}15, SH12]. Structuring [NJ12]. Studies [BSC^{+}16, ELC13, HEC^{+}15, MBBM10, MSH^{+}14, VSC15]. Study [BEMF11, DKS18, DCJ^{+}13, HRG12, KCM14, MRPW19, RJH12, TF13, VB14, ZC13]. Sturgeon [ASP^{+}18, BP16, BP17, BMGL12, BGV^{+}12, BLG^{+}13, BAHP11, BMC^{+}13, BH11, BHC^{+}13, CJM^{+}15, CSED10, CSD^{+}14, DCS^{+}17, DG13, EBF^{+}14, FCG^{+}17, FZZ^{+}10, FPR^{+}18, FBP11, FSC^{+}13, FP19, FGW^{+}16, HKW^{+}14, HPF^{+}16, HKKS14, HSE^{+}12, JLV^{+}10, JPFB13, JCS14, JZS^{+}19, JHK^{+}19, KHW^{+}14, KBO^{+}12, LEM^{+}11, MBPA13, MNM^{+}17, MNB18, MBL^{+}18, MSP12, NCW^{+}17, PSH^{+}19, PF11b, PB13, PHG^{+}16, PTG^{+}10, PGVG15, RDH11, REN17, SMM13, SP10, SH10, SDW^{+}17a, SWH^{+}17, SCW^{+}13, SCF16, SFH15, SIKR17, TL17, TCBI11, UWL12a, UWL12b, UPT^{+}13, VCP^{+}14, VHPP19, WS15, WZK^{+}15, WMW^{+}12]. Sturgeons [HGG^{+}16]. Sub [WRP^{+}12]. Sub-Arctic [WRP^{+}12]. Subadult [BNS^{+}14, SIKR17]. Subarctic [DGMC17]. Subdivision [Spi12]. Subestuary [WSK11]. Subpopulations [WLLCF16]. Subsequent [AIC13].

Subsidies [CBM15, Lev18]. Substrate [FPR^{+}18, NAW12]. Subtropical [BR14, NW14, PSGB13]. Subyearling [FMS10, GCG^{+}18, HBB17, RT14, SCL^{+}13, TBL^{+}17, TES14]. Subyearlings [Ano11a, CTPM13, TC11]. Success [ABM17, BCF^{+}11, BMB^{+}10, FPM15, GBF12, HM17, JCS14, KMS^{+}18, KKM11, LSPB11, LCO^{+}15, MWG18, OSDT18, PGW^{+}11, RHW15, RAB^{+}18, RHR^{+}15, SKP^{+}10, TM19]. Successful [EBF^{+}14]. Sucker [CGB19, DCM10, ESD^{+}17, KWCI10, PFG19, SBC17, WvPC12]. Suckers [CTN11, DSAM12, FWBT17, ZBW10]. Suggest [LV14, RA10]. Suitability [ACWIK13, DF12, DA16, FJP18, FKH15, Lov11, MKA10, TFM^{+}18]. Suitable [BBM^{+}14, JZS^{+}19, TRW10]. Sulfated [YSW^{+}14]. Summer [BL11, CMM^{+}17, CM10, CKFQ11, DG13, GCM12, HSM^{+}10, HMTC12, MBM10, RJWK10, RJH12, TRW10, UACS^{+}19, WAD^{+}17, BFH14]. Summer-Resident [CMM^{+}17]. Sunderland [HCD^{+}14]. Sunfish [THNG14]. Superimposition [Dud19]. Superior [AHS^{+}11b, BHG^{+}10, GSR^{+}11, GSJS17, GS13, MBZ^{+}14, RMM11, SJMG14]. Supersaturation [SDOK11]. Supplementation [BCF^{+}11, PT10]. Supplemented [DML^{+}10]. Supply [SHCB16]. Support [MNM^{+}17, TH15]. Supports [BWHF17, HB18, HUSST19]. Surf [SGP16]. Surface [RGM^{+}11]. Surgeon [DBC^{+}10]. Surgical [BHC^{+}13, DBB^{+}12, SHH^{+}19]. Surgically [DBC^{+}10, FMS10, KHF^{+}18, RPVS15, SNP^{+}10, TPA14]. Surrogate [AUDL18]. Surrogates [SF17]. Survey [AJFO10, AJF10, AJF13a, AJF13b, BLCB13, DKL^{+}17, GSD^{+}12, HLB^{+}16, KTE^{+}14, MAS10, Nel14b, RL14, TF13, WF11]. Surveyed [WTB10]. Surveying [FDSB18]. Surveys [BH11, DCBKR14, GKG^{+}19, HCD10, MK16, RSB14, RG15, SC13]. Survival [ACK17, AHS10, BB15, BJA^{+}13, BMC^{+}13, BHS^{+}17, Bel12, BRGR10, CHR^{+}17, CHW^{+}13, CBSB15, CT12, CDB10, CNM^{+}18, CSD^{+}14, DDMR10, DMW^{+}19, DMRK14, DM11b, EHC^{+}14, EWSR16, FBF10, FBWZ19, FWB^{+}15, FKF^{+}19b, FMS10, GDM^{+}10, GA17, GSA^{+}19, GNMM12, HMTC12, HNK^{+}12, HJM^{+}12, HRR^{+}11, HAR18, HKZ11, HLK11, HER^{+}11, IBH12, JRP12, JSW^{+}15, KGW14, KM14, KMS^{+}18, KHF^{+}18, KKM11, KYK17, KRM11, LSV^{+}13, MYS^{+}11, MCC^{+}12, MHE12, Mic10, MBT10, NBC15, OPHD18, PBG^{+}11, PKL^{+}11, PBC12, PHFN17, RH14, RPSB14, RB13, RBF14, RSG12, RBH^{+}12, RLL^{+}19, RAR^{+}16, SDW^{+}17a, SW11, SWB^{+}16, SAS17, SS15b, SVKK18, SNI^{+}13, SHP^{+}19, SDW17b, SJ14, TM19, WEW12, WDW11, WAD^{+}17, WCD^{+}11, ZBW10, ZKB^{+}15]. Susceptibility [HERC12, MBPA13]. Susitna [FRM18]. Suspended [BB15]. Sustainability [MCO^{+}10]. Sustainable [Smi13]. Sustaining [MNM^{+}17]. Sutton [Ryp13]. Suture [BHC^{+}13, BDBE11, DBC^{+}10, IBH12, PBG^{+}11]. Sutures [DBB^{+}12, JLA17]. Suturing [DBB^{+}12]. Suwannee [PA10]. Swimming [BMB^{+}10, CJM^{+}15, Cla16, CHW^{+}13, DMR^{+}10, FCG^{+}17, FMJ11, JBC^{+}12, KCJ^{+}15, LHHL17, MGW13, PPF^{+}13, RSG12, RGP17, SDW^{+}17a]. Swordfish [LKP^{+}13]. Sympatric [BGOQ13, CHT13, CPP^{+}14, DW16, RB18, SE11]. Symposium [Chi13, Chr14, Nel14a]. Synchronous [ZKH12]. Synchrony [CM11b, JRBC13, MS13]. Syndromes [NPW12]. Synthesis [CAB11, HP13]. Synthetic [TFM^{+}18, BDM^{+}10]. System [ARB^{+}18, BR11, DKL^{+}17, DHRQ10, FPR^{+}18, FPM15, GWB^{+}15, GBW16, HKW^{+}14, HLR^{+}12, KM10, MPM^{+}14, MAN^{+}12, OSDT18, Pet15, Rey16, SZGS12, WS15, Yas16, AFHF12, WZK^{+}15]. Systematically [BSM13]. Systems [KM10, LRH12, RPC14, SDW15]. Systemwide [EHR^{+}12].

T [Gid14, LDA^{+}13, Ryp13]. Tactics [CVB14]. Tadpole [ACK17]. Tadpoles [BSHK^{+}11]. Tag [DM11b, DJ17, FMS10, HEC^{+}15, HRG12, IBH12, MBP^{+}11, MGW13, VB14, WvPC12, Wei10]. Tag-Induced [MBP^{+}11]. Tagged [BDBE11, DEP^{+}18, HJM^{+}12, HLK11, JCK^{+}12, KCM14, KHD^{+}14, MH11, MCC^{+}12, NBC15, REN17, SNI^{+}13, SCL^{+}13]. Tagging [AUDL18, AIC13, Chi13, CHW^{+}13, DEP^{+}18, EMS^{+}14, EB16, HSA^{+}16, LEM^{+}11, MSH^{+}14, RMDA11, dHST13]. Tags [CHR^{+}17, Cla16, DM11b, EHR^{+}12, EPT^{+}16, HB17, HRG12, JDS15, KHF^{+}18, NB12, SDW^{+}17a, SNP^{+}10, VSC15]. Tail [MBBM10]. Tail-Beat [MBBM10]. Tailwater [IB14, KYK17]. Tailwaters [FM17]. Taimen [FSRK11]. Taiwanese [CYW^{+}19]. Taken [AJF13b, BFM10]. Taking [NPW12]. Tale [RBK^{+}11]. Tales [RCD^{+}11]. Tampa [Gre12]. Tandem [SL16]. Tapeworm [CCL^{+}19]. TaqMan [BSSM12]. Targeted [FWSW^{+}15]. Tautogs [MT19]. Taxonomy [Pag19]. Technical [FHCSN12, Yas16]. Technique [DBB^{+}12, MCW^{+}14, RNG^{+}13]. Techniques [ESD^{+}17, ELC13, Eva17, FH12, GDWD10, HOGO10, Mat18, Ryp13]. Technology [GIW^{+}13, Chi13]. Telemetry [BFH14, FBH^{+}13, GZB12, HBB13, MCSF14, PSH^{+}19, TL17, TS13]. Telemetry-Based [FBH^{+}13]. Teleost [SPH19]. Teleostei [HOWS10]. Teleosts [PSM12]. Telephone [AJFO10]. Temperate [BB10b, MFH13, RB12, RW11]. Temperature [AAD14, BS18, BCRP13, BMM10, BRK^{+}18, CM10, CDB10, DBC^{+}10, DBHV14, FBJR11, FBWT19, HHJH17, HAR18, HK12, HHES13, HD16, HWHF14, HMP18, KGW14, KWCI10, LJHL10, MRPW19, MT19, MBM10, NB12, PPH15, PHHM12, PPMJ14, PM15, RB13, RBF14, RT14, SW11, TASC17, TVL^{+}12, UK14, VSC15, WMS15, WHP14, YCB^{+}11, YHWH10]. Temperature-Dependent [HD16, PPH15, PM15]. Temperature-Related [HK12]. Temperatures [GDM^{+}10, GDM^{+}11, HP14, KWCI10, MRPW19, MBH12, RJWK10, TNS^{+}14, VJGV11, ZBC^{+}13]. Temporal [ACRAM11, BSC^{+}10, BR11, CBM15, CHS^{+}11, DHRQ10, DCM10, EFC^{+}11, EPT^{+}16, FM17, GWG^{+}19, IWB^{+}13, MHF12, OWS16, PZS17, PGVG15, RRG^{+}17, SJT^{+}16, SMJ14, SA14, SWC13, SBMH12, SJFF15, SS15c, SMMK12, VWS13, WMV^{+}16, WR14, WBD17, WBD14]. Tendency [CEFQ11]. Tennessee [IB14, PSB16]. Term [AGH18, BGB^{+}16, CHB^{+}12, DSM18, FWB^{+}15, FH12, GP12, HSSS16, IHTF11, KM10, KRM13, MBP^{+}11, MLC13, NDC^{+}14, NR16, OGM^{+}15, PPJ15, RA10, SWB^{+}16, SMH^{+}11, VB14, WHSK19, ZBHW16, DEP^{+}18]. Terminal [GGC17]. Tern [ARL^{+}12, MGWP10]. Terns [SCL^{+}13]. Tessellated [KLS17]. Test [CJL^{+}14, GS16, JPSM12, WHC^{+}10]. Testing [MAN^{+}12]. Texan [BKC^{+}19]. Texas [BSD13, DSM18, LCHB^{+}18, MBEvdB^{+}10, OWS16, PTWK11, RHKD12, RMW16, VFFP13]. Their [AGKW12, BHD^{+}16, BP10, BDM^{+}10, CZK^{+}17, CHLN15, CNM^{+}18, GS13, MOS13, PF11b, Rey16, Shi13, TJB^{+}14, AS12, BNS^{+}16, CKS^{+}19, DW15, MBBM10]. There [BNTT17, CTS^{+}15]. Thermal [ACWIK13, AMBM12, BHD^{+}16, BHP^{+}15, BCR13, CRC^{+}16, DMRK14, HR15, HP14, HAR18, KH13, MRPW19, MFAS^{+}12, MWSS13, PHHM12, PCC^{+}13, RZW^{+}12, Str10, UK14, ZBC^{+}13]. Thiamine [FBB^{+}13]. third [Ryp13]. Threadfin [PNMH^{+}12]. Threatened [BSSM12, BMC^{+}13, CKC^{+}12, DSRA12, DYWY17, GCM12, NDP16]. Threatens [Arc16]. Three [CT16, Eva17, FMJ11, FWSW^{+}15, HKR18, HP14, HWN15, OWP^{+}17, PCC^{+}13, SF17, SHCB16, PM10]. Three-Pass [HKR18]. Threespine [HBHS15]. Thresher [GN13, NG13]. Thresholds [Sto10]. Throughout [SWP^{+}18, DW15]. Throw [CGL^{+}11]. Tibet [LCO^{+}15]. Tidal [BSCA17, DK18, DEW^{+}14, DWE^{+}19, JPRS10, Lov11, MBH16, RDR^{+}10, RLL^{+}19, SJT^{+}16, TBTS12, WF11]. Tides [PBB^{+}15]. Tiger [CABNQ^{+}13, SHC^{+}16, WEW12]. Tilapia [HARL18, HAR18, VSBB^{+}14]. Time [HJM^{+}12, HPG10, KGC^{+}12, KCS^{+}19, OWP^{+}17, RFM^{+}15, WMV^{+}16, vPWT12]. Time-Varying [vPWT12]. Times [PHFN17]. Timing [BQF^{+}19, BA13, CHMG^{+}18, CC16, CTS^{+}15, CYP11, EMS^{+}14, HN11, JSW^{+}15, LSSYL10, MM17, MSS^{+}11, MDT10, OMA13, PW10a, RBH^{+}12, SHM^{+}12, SNC^{+}10, TRLG13, TL17, UPT^{+}13, VJGV11, WAD^{+}17, WHP14]. Timings [SCC17]. Tissue [OPS^{+}14, SARC13]. Tissues [HNV^{+}18]. Toad [BSHK^{+}11]. Togiak [CTS^{+}15]. Tolerance [AAD14, BCR13, DMW^{+}19, DMRK14, HAR18, KWCI10, KU12, MRPW19, MFAS^{+}12, RZW^{+}12, SPH19, UK14, WCD^{+}11]. Tolerances [ZBC^{+}13]. Too [WB15b]. Tool [BR14, CCL^{+}19, CEM^{+}19, DAA^{+}16, PSB16, SWH^{+}17]. Top [MAH^{+}12, SFF^{+}10]. Top- [SFF^{+}10]. Topographic [WDHW14]. Total [CGA13, GvdHBC12, HGP^{+}17, THGA15]. Total-Body [GvdHBC12]. Tournaments [LSV^{+}13, VWS13]. Towboat [KMM^{+}11]. Towed [GWG^{+}19]. Toxicity [BSHK^{+}11]. Tracers [AW19]. Tracking [AW19, ELC13]. Trade [BGV11, KRRW18, QSVD11]. Trade-off [QSVD11]. Trade-Offs [BGV11, KRRW18]. Traditional [WCN^{+}19]. Training [SCF16]. Trait [BFH14]. Traits [BEMF11, DBHV14, DWN^{+}15, LCO^{+}15, PK11]. Trajectories [BCRP13]. Transactions [Ano13b, Ano14b, Ano16, Ano17b, Ano19i, SRMB^{+}15]. Transformation [BSC^{+}10]. Transgenerational [SSH11b]. Transgenesis [GWOD18]. Transgenic [HKP11, HLH14, PK11]. Transient [TAM15]. Transition [CHW^{+}13, KWCI10]. Translocated [PRW^{+}10]. Translocation [SPH^{+}15, SPCJ11]. Translocations [FPM13, GBF12]. Transmitter [IBH12, LBD^{+}16, PBG^{+}11]. Transmitters [DBC^{+}10, DBB^{+}12, JBC^{+}12, LBD^{+}16, MRB^{+}18, PPF^{+}13, RPVS15, SHH^{+}19, TPA14]. Transplanted [WAR12]. Transponder [CHR^{+}17, Cla16, DM11b, ELC13, EHR^{+}12, EPT^{+}16, HLK11, SDW^{+}17a]. Transponders [HB17, MCC^{+}12]. Transport [JPLS13, MKH10]. Transportation [BWD^{+}17, PB15]. Transverse [BAZ12, DMCO11]. Trapping [SFF^{+}10, WBR11]. Traps [CGL^{+}11, RC16, RSB14]. Travel [HJM^{+}12]. Trawl [BMC^{+}13, DKL^{+}17]. Trawls [KS11, PNMH^{+}12]. Treatments [JADJ^{+}14]. Trematodes [PM10]. Trend [HKR18]. Trends [ACRA10, AGJ11, Arc16, Bea10, CZK^{+}17, DSM18, KQ11, KYK17, LRR^{+}15, MS13, OI16, PM19, RA10, SWB^{+}16, SJFF15]. Trial [AUDL18]. Tributaries [BBG^{+}18, CGM15, CGB19, IRJ11, LPM^{+}17, MSW^{+}14, OI16, PDM^{+}14, REN17, TF13, WF11]. Tributary [BQHR17, CBSB15, DF11, DWG14, FWBT17, HCZ13, HCZG15, HPHB10, KRM11, MM17, RMM11, RAR^{+}16, SPH^{+}15, vVO10]. Trifluoromethyl [CHB^{+}12]. Triggerfish [HAV^{+}19, KSSK^{+}17, SS11]. Trinidad [JM14]. Trinity [BSD13, PHM17]. Trip [AJF10]. Trophic [BK11, BL11, CORM13, DWE^{+}18, FRBA13, FRA13, FKF^{+}19a, HBB17, HB18, KSB16, LBBW17, LB15, MHF12, RCM^{+}19, RMM11, RB18, SRP^{+}18, SHCB16, TRHHOGSJ14, THKG10, WB15b, WHC^{+}10]. Trophy [BGB^{+}16]. Tropical [HTH^{+}17, PNMH^{+}12]. Trout [AHS^{+}11b, ACMM^{+}15, ACDL19, AMM^{+}16, AGKW12, ADS^{+}11, BBGB15, BSGB17, BGLP19, BP13, BBAK19, BB10a, BHD^{+}16, BS13, BAKT12, BRG19, BRGR10, BGOQ13, BMP^{+}17, CKM^{+}11, CR10, CM10, CHT13, CML19, CRH10, DFB15, DW16, DW15, DH11, DZW^{+}12, DWN^{+}15, DMR^{+}10, FWB^{+}15, FKF^{+}19a, FM17, FKF^{+}19b, FFYI16, FH17, GSR^{+}11, GSJS17, GS13, HH11, HNK^{+}12, HBB13, HBS13, HP14, HB18, HNV^{+}18, HK12, HSM12, HS13, HCS^{+}16, HCNL10, HMP18, ILH^{+}18, IBH12, JC16, JLA17, JK17, JJNB17, JMM^{+}11, KLR^{+}15, KHK^{+}15, KBG^{+}18, KMS^{+}18, KWS14, KRRW18, KKM11, KYK17, KCMS11, KRM13, LME16, LRW13, LB15, MCW^{+}14, MGP11, MSWS10, MFAS^{+}12, MWSS13, MSW^{+}10, MPM^{+}14, MWC^{+}13, MHE12, MKHC17, MDT10, Mil19, Mit16, MWKS13, MDD^{+}16, MBZ^{+}14, NB12, NB13, NDP16, OS11, ODG^{+}11, Pag19, PB15]. Trout [PRW^{+}10, PGJ11, PT10, PB17, PHHM12, PPC13, PPW^{+}14, PPJ15, PPP^{+}13, Qui19, RB13, RBF14, RMM11, RJWK10, RL19, SPM^{+}17, SK11, SBO12, SLM10, SJMG14, SHR^{+}12, SDW17b, TD13, TPBB14, THKG10, TNS^{+}14, UACS^{+}19, WDDS13, WMS15, WBR11, WDHW14, WKS^{+}19, WK13, WCD^{+}11, WBH10, WAR12, YCB^{+}11, YHWH10, ZTC12, ZBC^{+}13]. True [HLK11]. Tsangpo [LCO^{+}15]. Tubes [BN16]. Tui [GLM^{+}15]. Tuna [AGH18, Gol18, RBE^{+}10]. Turbid [BES15]. Turbidity [CSDW10, HBS13, HDF12, KLS17, MSD^{+}19, MWG18, Sto10, YCB^{+}11]. Turbine [BSCA17, BDBE11, BCG^{+}12, EWSR16]. Turnover [RW11]. Twelve [HNV^{+}18]. Twenty [MMY12]. Twist [WBW^{+}13]. Two [AHS10, BSC^{+}10, CGM15, DSHG10, DM11a, FWD10, FBWZ19, FHCSN12, FBH^{+}13, GKBG12, HBC15, HRB^{+}16, HS13, IRJ11, KBG^{+}18, KWS14, KCZH14, LRH12, LPM^{+}17, MCD17, MSD18, PB15, PPC13, PDM^{+}14, RBH^{+}12, RBK^{+}11, SGP16, SDW15, SSM16, STK12, TBL^{+}17, THGA15, WDDS13, WBH10, YSC10, LRR^{+}15]. Two-Dimensional [GKBG12]. Two-Species [WDDS13]. Two-Vector [YSC10]. Type [CVB14, DBC^{+}10, EFC^{+}11, HH11, HKP11, LBC10, PW10b, PNM^{+}10, PK11, RR13]. Types [SHM^{+}12, SAF^{+}19].

U.S. [DF11, FJP18, HAV^{+}19, ILH^{+}18, KCS^{+}19, LKS^{+}12, LAR16, SDW15]. UK [Fle16]. Ultralight [CCB16]. Ultraoligotrophic [JBV13]. Ultrasonic [HBB13, IBH12, TL17, TS13]. Umpqua [OMA13]. Uncertainty [HRG12, PMN19, RJ11]. Undergoing [HSSS16]. Underlying [BMR^{+}12]. Understanding [Goo15]. Unfished [HVD^{+}16]. Unintended [MK10]. Unique [HB18]. Unit [HAG13, PRW^{+}10]. United [Ada18, ABS^{+}11, DW15, MAS10, PPF^{+}16, PKP^{+}17, ZHWR11, ADS^{+}11, FA10, HLH14, HTH^{+}17]. Units [ADS^{+}11]. University [Cha13, Gid14, Pra14, Qui16, Smi13, Var13]. unlike [HCR14]. Unmarked [RGM^{+}11]. Unobstructed [RPC14]. Unregulated [HK12]. Unsampled [WBWSK12]. Unscreened [MCH^{+}13]. Unstratified [RJWK10]. Untagged [SMO^{+}14]. Unusual [MPM^{+}14, WMV^{+}16]. upon [AS12, HH11, RCC^{+}13]. Upper [BCR13, KWCI10, RZW^{+}12, Str10, ZBC^{+}13, ACRA10, BHD^{+}16, CHR^{+}17, DWGF11, EFC^{+}11, EBF^{+}14, EPC^{+}19, FWBT17, JRP12, KBC^{+}18, KMM^{+}11, KR18, KCMS11, MC14, PGJ11, PBM^{+}12, RKEB13a, RKEB13b, SK11, SS13, SLM10, VFFP13, WDR16, ZBW10, ZZMC13]. Upriver [DHRQ10, LUC^{+}15]. Ups [GS13]. Upstate [GBW16]. Upstream [GDST16, HSM12, JPSM12, MCW^{+}14, PM14, PPP^{+}13, RO16, WZC^{+}14, WL13]. Uptake [TL15]. Urban [BMOF16, GWB^{+}15, HSSS16, LLWW19, TBL^{+}17]. Urge [SZZ10]. USA [BGLP19, BB15, CGM15, EPR^{+}18, FSW10, FP19, HNV^{+}18, HHES13, KTE^{+}14, KSSK^{+}17, KU12, KPZ^{+}12, PAB^{+}11, PSH^{+}19, TT17, UK14, Yas16]. Usage [TCBI11]. Use [BBZGB15, BAKT12, BDM^{+}10, BSD13, BW10, CGL^{+}11, CHLN15, CBM15, DG13, DWGF11, DWG14, DW16, DMW^{+}19, DCM10, ESD^{+}17, ELC13, FJP18, FFYI16, FWBT17, GvdHBC12, GGC17, GCM12, GKG^{+}19, GZB12, GNMM12, HPF^{+}16, HR15, HJS^{+}16, HNRB19, HKKS14, HK12, HGG^{+}16, HDH16, HPSA16, KST^{+}12, LMD^{+}17, LKP^{+}13, LSD^{+}14, LSSYL10, MM17, MWKS13, MBBM10, PBW16, PMF^{+}10, PHG^{+}16, PTG^{+}10, PAG11, PL13, RA10, RMM11, RDR^{+}10, SHM^{+}12, SMO^{+}14, SR15, SMN^{+}10, SCB15, TD13, TS13, TBC^{+}13, WZCT12, WSQ^{+}15, WZK^{+}15]. Useable [HGS^{+}16]. Used [HBA^{+}15, LS10a, PCC^{+}13]. Useful [EPS^{+}15]. Using [AGH18, ACRA10, AW19, ASP^{+}18, Ano11a, BSSM12, BC12, BSS18, BYM19, BEMF11, CKC^{+}12, CCL^{+}19, CC16, CPP^{+}14, CHS^{+}11, DFB15, DMR^{+}10, FDWG13, GCM12, GDWD10, Gre12, GA10, HLH14, HSA^{+}16, HOWS10, HPHB10, IWB^{+}13, JJC12, LFB^{+}15, MAS10, MJHP18, MLS^{+}13, MM10, MLBS^{+}13, OSS^{+}17, PSM12, PM13, PPH15, PGWS15, PCT^{+}12, PMN19, PVJH15, PHFN17, RPSB14, RMMGPB16, SR10, SDBB12, SBC^{+}16, STZ13, SSH11b, SL16, SJFF15, TPLA16, TC11, VIW^{+}17, VBJS10, WDDS13, WBR11, WCN^{+}19, WCQD18, WMG^{+}14, BFH14, CBSB15, HBC15, HS17, KCM14, MCSF14, OSA^{+}17, TL14]. Utah [DF14, GP12, KWCI10, PFG19, ZBW10]. Utility [BSR12, FA10, ZSVK13].

V [Ryp13]. Validates [MLC13]. Validating [PM13, SWH^{+}17]. Validation [DMHO12, DA16, HOGO10, SARC13]. Valley [AM13, SCC17, SMS^{+}12]. Valuable [KSB16]. Values [OPS^{+}14]. Varden [BCS19, WTB10]. Variability [ACRAM11, AJF13b, BSC^{+}10, BS18, CPS^{+}15, CTK^{+}17, DSAM12, MHF12, MSB^{+}18, RL19, SJT^{+}16, TA13, TRHHOGSJ14, WAW11, WDHW14]. Variable [BEWM10, GDST16, HBTT11, RJWK10]. Variables [CM11a, OPHD18, SM13a, WTB10]. Variance [OSDT18, STZ13, VIW^{+}17]. Variation [AS12, AMBM12, BRA14, BR11, CTN11, CKM^{+}11, CMM^{+}17, CBM15, DKD^{+}15, DSRA12, EFC^{+}11, FDWG13, GSVS^{+}10, HGBE16, HLR^{+}12, HLB14, HN13, IWB^{+}13, KHK^{+}15, LHS^{+}13, LYX^{+}10, MSBT10, MT19, MSWS10, MFAS^{+}12, MVK12, MSWD15, MSS^{+}17, MK15, OWS16, PW10a, PJ11, RL19, SMJ14, SJ11, SA14, SLG^{+}14, SGB13, TES14, VWS13, VSBC13, WKS^{+}19, WB13b, Whi12, WBD17, WBD14]. Various [SDSB18]. Vary [SMM13]. Varying [BHLK10, LDLH15, MBH12, VB14, vPWT12]. Vector [YSC10]. Vectors [KKB^{+}10]. Vegetated [CGL^{+}11]. Vegetation [EWW13]. Velocities [PB15]. Velocity [GGC17, JCK^{+}12, MYS^{+}11, MJH17]. Ventilation [RHR^{+}15]. Verification [LS10a]. Vermilion [JPHK10, MBPP18]. Vermont [BBGB15, PPD16]. Versus [HMO14, CJL^{+}14, CNM^{+}18, ELC13, GDG^{+}15, GDPW10, LMPvdH18, PF11a, PCC^{+}13, Poe14, RKEB13b, SMO^{+}14, SGP16, WF19]. Vertical [AHS^{+}11b, GKG^{+}19, GZB12, GSD^{+}12, LKP^{+}13, SFF^{+}10]. via [GDWD10, GSVS^{+}10, JPLS13, MKHC17]. Viability [JPFB13, KPS^{+}19, SVKK18]. Viable [JLV^{+}10]. Vicariance [BJC12]. Victoria [Ber13]. Video [DKL^{+}17, MAS10]. Viral [EFC^{+}11]. Virgin [CCKF11]. Virginia [BMGL12, BGV^{+}12, KHW^{+}14, OI16, PCFH17, TNS^{+}14]. Virus [EFC^{+}11]. Visible [BHC^{+}18]. Visual [HLB^{+}16, HBB13, HBS13, HB17]. Volcanism [CKM^{+}11]. volume [Ano11a]. Vondracek [LDA^{+}13]. Vulnerability [BNW^{+}12, SPC^{+}18, SCL^{+}13, TNS^{+}14, WMS15]. Vulnerable [HKP11, OMA13].

Wadeable [HJS^{+}16, HCD10, KMKS13]. Wahoos [RBE^{+}10]. Walk [SDBB12]. Walker [Ber13, BRGR10]. Walleye [BJC12, FPR^{+}18, HRHS16, IHTF11, JJNB17, LSV^{+}13, MHF12, RBK^{+}11, RAR^{+}16, SS18a, SNC^{+}10, SBMH12, WSI14, WF19]. Walleyes [GKG^{+}19, HDBS14, KRM13, MSH^{+}14, PZS17, RB12, VB14]. Walter [Cha13]. Warm [SFLH14]. Warming [ILH^{+}18]. Washington [BRS^{+}12, TBL^{+}17, Yas16, ACDL19, BLB^{+}13, CKFQ11, HRB^{+}16, HGB^{+}16, LHS^{+}13, MPM^{+}14, PT10, RBH^{+}12, SBDW11, SE11, SHCB16, TB14, TW14, WAD^{+}17, WBM^{+}17, ZKB^{+}15]. Water [BPS11, BBZGB15, BRGR10, BMM10, BRK^{+}18, CHT13, DBC^{+}10, GIW^{+}13, LJHL10, LHHL17, LBBW17, MRPW19, MBM10, MBH12, MJH17, MCH^{+}13, MCP^{+}15, PB15, RR13, RJWK10, RT14, SS15b, WMS15, WHP14, WHSK19, YHWH10]. Waterbird [CML19]. Waterfalls [TPBB14]. Waterfront [AJF10]. Waters [Ada18, BBC^{+}14a, BBC^{+}14b, CKFQ11, HAV^{+}19, KHD^{+}14, OWS16, PDSP^{+}11, RGM^{+}11, TRHHOGSJ14]. Watershed [BBM^{+}14, CHMG^{+}18, CTS^{+}15, CHS^{+}11, Lov11, LN18, MK15, PT10, SLM10, SR15, SWS10, SBC17, SDW17b, TVL^{+}12, TPBB14, WDR16, WHF^{+}12, WHSK19, WBM^{+}17]. Watersheds [HRB^{+}16, NRZ^{+}13, RBH^{+}12]. Waterweed [CBW^{+}16]. Way [Yas16]. Weakfish [TT17]. Weakly [TB16]. Weatherfish [KU12, UK14]. Web [JMM^{+}11, KPZ^{+}12, LB15, ZRM^{+}16]. Webs [HTR13]. Weight [FH12, HS17, RH14]. Weighted [Gre12, HGS^{+}16]. Weighted-Evidence [Gre12]. Weir [LMPvdH18]. West [FRRCGA^{+}15, HGBE16, Wei10]. Western [ASP^{+}18, BDTD17, PHG^{+}16, RG15, AHS^{+}11b, AKM12, BGLP19, BAYS16, CM11a, DSB^{+}18, GN13, LKP^{+}13, NG13, NAWM14, RA10, SMY11, SAW^{+}17, TBL^{+}17]. Westslope [DZW^{+}12, DWN^{+}15, MDD^{+}16, PPW^{+}14, SK11]. Wet [CNM^{+}18, RNG^{+}13]. Wetland [KWW14, TH15]. Wetlands [DWE^{+}19, RDR^{+}10]. Whirling [EPS^{+}15, PPMJ14]. White [BRA14, DM11b, DCM10, FRBA13, FRA13, Han11, JLV^{+}10, JCS14, LAR16, MCSF14, MBH12, REN17, SMM13, SJ11, SWH^{+}17, SDOK11, SPCJ11, VFB^{+}16, SE11]. Whitefish [BGW^{+}17, BCR13, BDM^{+}10, FPR^{+}18, GZB12, HLR^{+}12, HML13, PDP12]. Whitefishes [TFM^{+}18]. Who [SSH11b]. Wide [ASJ16]. Width [JK17]. Wild [BMF^{+}15, BBGB15, CKC^{+}12, CHO^{+}15, CBAC19, CNM^{+}18, CPS^{+}15, DW16, DWE^{+}18, EMS^{+}14, GSVS^{+}10, HKP11, JZBP16, JSW^{+}15, KU12, MSWS10, MHE12, MWKS13, PK11, PPC13, QSJ12, RNG^{+}13, SKP^{+}10, TRMC15, TPLA16]. Wild-Origin [MSWS10]. Wild-Type [HKP11, PK11]. Wildlife [BCS19]. Wiley [Fle16]. Wiley-Blackwell [Fle16]. Willamette [JSW12, KBC^{+}18, BNC^{+}15, CFHE11, MSW^{+}14, MFR15, VHJ^{+}15]. Willingness [PW10b]. Willow [SS15b]. Wind [HUMG^{+}18]. Windowpane [HW14]. Winter [AMM^{+}16, AS17, BFK15, BNTT17, Bel12, DMW^{+}19, FJP18, HR15, HPSA16, JADJ^{+}14, KM14, LMD^{+}17, Mic10, MWKS13, RJ11, SDW17b, TKC^{+}12, WMG^{+}14]. Winter-Emerging [AMM^{+}16]. Winter-Run [LMD^{+}17]. Winterkill [DMW^{+}19]. Winterkill-Prone [DMW^{+}19]. Wire [HB17, HSA^{+}16, Wei10]. Wisconsin [CSED10, SS18a, SIKR17, WSI14]. Within [AJF13b, BQF^{+}19, HCNL10, KP14, SYH^{+}15, ACRA10, AFHF12, CHMG^{+}18, CGB19, CEFQ11, EBG18, HPF^{+}16, HBTT11, KAH16, KCJ^{+}15, MFAS^{+}12, PPF^{+}13, PHHM12, WKS^{+}19, WSK11, ZTC12]. Within-Day [AJF13b]. Within-River [BQF^{+}19, SYH^{+}15]. Within-Stream [HCNL10]. without [SAH^{+}14, TRW10]. Wolf [SIKR17]. Wood [FSFS17, JADJ^{+}14, MWS10, PPJ15, DHRQ10, MSS^{+}11]. Woodruff [YITI12]. Woody [DW13, TVL^{+}12]. World [Fre19, PM19]. Wound [CHW^{+}13, IBH12, LBD^{+}16]. Woundfin [CCKF11]. Wyoming [CRH10, HUMG^{+}18, MRPW19].

Y-Chromosomal [MSWD15, MSS^{+}17]. Yakima [JRP12, DML^{+}10, LHS^{+}13]. Yampa [ZBHW16]. Yaqui [BB15]. Yarlung [LCO^{+}15]. Year [Bel12, BB10b, CNM^{+}18, CAB11, DCJ^{+}13, IMO^{+}14, LV14, PW10a, SWP^{+}18, VFB^{+}16, WB13a]. Year-Class [PW10a, SWP^{+}18]. Yearling [CKFQ11, CFHE11, SLG^{+}14]. Years [FBS^{+}14, RKEB13b, STK12]. Yellow [AGKH17, BPS11, BDS10, CM11a, GSVS^{+}10, IHTF11, JRBC13, KGW14, LSPB11, MD11, RCD^{+}11, RBF^{+}11, RLL^{+}19, SR10, SSS^{+}12, SB12a, SS15c, WDW11, WL13, YJW11]. Yellow-Phase [RLL^{+}19, WL13]. Yelloweye [BS16]. Yellowfin [RBE^{+}10]. Yellowstone [SGK16, CKM^{+}11, KKB^{+}10, MKHC17, ODG^{+}11, SGK16, THKG10]. Yellowtail [BC12]. Yield [BJA^{+}13, Smi13, vS10]. Yolo [HSG10]. York [BGLP19, Cha13, PSH^{+}19, Pet15, BSGB17, BSCA17, DCJ^{+}13, EBMV10, GBW16, HKW^{+}14, MBM10, TCBI11, VFB^{+}16]. Young [Bel12, BB10b, LV14, MT19, RBF^{+}11]. Young-of-the-Year [Bel12]. Young-of-Year [BB10b]. Yukon [BEMF11, BD15, DF11, ZKH12]. YY [KMS^{+}18].

Zale [Ryp13]. Zealand [YHWH10]. Zebra [HKP11]. Zimmer [Cha13]. Zone [RCM^{+}19, SGP16, TLKT17]. Zooplankton [HTR13].

References

Anweiler:2014:LTT

[AAD14] Katie V. Anweiler, Stephen A. Arnott, and Michael R. Denson. Low-temperature tolerance of juvenile spotted seatrout in South Carolina. Transactions of the American Fisheries Society, 143(4):999–1010, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abrahams:2017:SBR

[ABM17] M. V. Abrahams, D. K. Bassett, and J. C. Montgomery. Sensory biology as a risk factor for invasion success and native fish decline. Transactions of the American Fisheries Society, 146(6):1238–1244, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adamski:2011:DFI

[ABS^{+}11] Kyle M. Adamski, Jeffrey A. Buckel, Kyle W. Shertzer, Gretchen Bath Martin, and J. Christopher Taylor. Developing fishery-independent indices of larval and juvenile gag abundance in the Southeastern United States. Transactions of the American Fisheries Society, 140(4):973–983, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Araujo:2014:ACG

[ACB^{+}14] H. Andres Araujo, John R. Candy, Terry D. Beacham, Bruce White, and Colin Wallace. Advantages and challenges of genetic stock identification in fish stocks with low genetic resolution. Transactions of the American Fisheries Society, 143 (2):479–488, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Al-Chokhachy:2019:NIE

[ACDL19] Robert Al-Chokhachy, Jeremiah Doyle, and James S. Lamperth. New insights into the ecology of adfluvial bull trout and the population response to the Endangered Species Act in the North Fork Lewis River, Washington. Transactions of the American Fisheries Society, 148(6):1102–1116, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alvarez:2017:ECL

[ACK17] Guillermo Alvarez, Colleen A. Caldwell, and Carter G. Kruse. Effects of CFT legumine (5% rotenone) on tadpole survival and metamorphosis of Chiricahua leopard frogs Lithobates chiricahuensis, northern leopard frogs L. pipiens, and American bullfrogs L. catesbeianus. Transactions of the American Fisheries Society, 146(3):512–522, May 2017. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Al-Chokhachy:2015:CAM

[ACMM^{+}15] Robert Al-Chokhachy, Sean Moran, Peter A. McHugh, Shana Bernall, Wade Fredenberg, and Joseph M. DosSantos. Consequences of actively managing a small bull trout population in a fragmented landscape. Transactions of the American Fisheries Society, 144(3):515–531, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Al-Chokhachy:2010:EST

[ACRA10] Robert Al-Chokhachy, Brett B. Roper, and Eric K. Archer. Evaluating the status and trends of physical stream habitat in headwater streams within the interior Columbia River and Upper Missouri River Basins using an index approach. Transactions of the American Fisheries Society, 139(4):1041–1059, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Al-Chokhachy:2011:QEF

[ACRAM11] Robert Al-Chokhachy, Brett B. Roper, Eric K. Archer, and Scott Miller. Quantifying the extent of and factors associated with the temporal variability of physical stream habitat in headwater streams in the interior Columbia River Basin. Transactions of the American Fisheries Society, 140(2):399– 414, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Al-Chokhachy:2013:CTS

[ACWIK13] Robert Al-Chokhachy, Seth J. Wenger, Daniel J. Isaak, and Jeffrey L. Kershner. Characterizing the thermal suitability of instream habitat for salmonids: a cautionary example from the rocky mountains. Transactions of the American Fisheries Society, 142(3):793–801, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:2018:MCE

[Ada18] Douglas H. Adams. Mercury in cobia from estuarine and offshore waters of the Southeastern United States: Fisheries implications. Transactions of the American Fisheries Society, 147(2):363–369, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ardren:2011:GSE

[ADS^{+}11] William R. Ardren, Patrick W. DeHaan, Christian T. Smith, Eric B. Taylor, Robb Leary, Christine C. Kozfkay, Lindsay Godfrey, Matthew Diggs, Wade Fredenberg, Jeffrey Chan, C. William Kilpatrick, Maureen P. Small, and Denise K. Hawkins. Genetic structure, evolutionary history, and conservation units of bull trout in the coterminous United States. Transactions of the American Fisheries Society, 140(2):506– 525, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:2012:OME

[AFHF12] Gregory B. Anderson, Mary C. Freeman, Megan M. Hagler, and Byron J. Freeman. Occupancy modeling and estimation of the holiday darter species complex within the Etowah River System. Transactions of the American Fisheries Society, 141 (1):34–45, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ailloud:2018:STP

[AGH18] Lisa E. Ailloud, Todd Gedamke, and John M. Hoenig. Shortterm pain and long-term gain: Using phased-in minimum size limits to rebuild stocks — the Pacific bluefin tuna example. Transactions of the American Fisheries Society, 147(6):1015– 1029, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anlauf:2011:DRT

[AGJ11] Kara J. Anlauf, William Gaeuman, and Kim K. Jones. Detection of regional trends in salmonid habitat in coastal streams, Oregon. Transactions of the American Fisheries Society, 140 (1):52–66, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Almeida:2017:PRY

[AGKH17] L. Zoe Almeida, Samuel C. Guffey, Tyler A. Krieg, and Tomas O. Höök. Predators reject yellow perch egg skeins. Transactions of the American Fisheries Society, 146(1):173– 180, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Annett:2012:CGR

[AGKW12] Brendan Annett, Gabriele Gerlach, Timothy L. King, and Andrew R. Whiteley. Conservation genetics of remnant coastal brook trout populations at the southern limit of their distribution: Population structure and effects of stocking. Transactions of the American Fisheries Society, 141(5):1399–1410, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Antonsson:2010:SES

[AHS10] Thorolfur Antonsson, Thorkell Heidarsson, and Sigurdur S. Snorrason. Smolt emigration and survival to adulthood in two Icelandic stocks of Atlantic salmon. Transactions of the American Fisheries Society, 139(6):1688–1698, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ackerman:2011:SNP

[AHS11a] Michael W. Ackerman, Christopher Habicht, and Lisa W. Seeb. Single-Nucleotide Polymorphisms (SNPs) under diversifying selection provide increased accuracy and precision in mixed-stock analyses of sockeye salmon from the Copper River, Alaska. Transactions of the American Fisheries Society, 140(3):865–881, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ahrenstorff:2011:SDD

[AHS^{+}11b] Tyler D. Ahrenstorff, Thomas R. Hrabik, Jason D. Stockwell, Daniel L. Yule, and Greg G. Sass. Seasonally dynamic diel vertical migrations of Mysis diluviana, coregonine fishes, and Siscowet lake trout in the pelagia of Western Lake Superior. Transactions of the American Fisheries Society, 140(6):1504– 1520, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashton:2013:AMS

[AIC13] Neil K. Ashton, Susan C. Ireland, and Kenneth D. Cain. Artificial marker selection and subsequent tagging evaluations with juvenile burbot. Transactions of the American Fisheries Society, 142(6):1688–1698, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashford:2010:PWH

[AJF10] Julian Ashford, Cynthia Jones, and Lynn Fegley. Private waterfront householders catch less per trip than other fishers: Results of a marine recreational survey. Transactions of the American Fisheries Society, 139(4):1083–1090, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashford:2013:IEC

[AJF13a] Julian R. Ashford, Cynthia M. Jones, and Lynn Fegley. Independent estimates of catch by private and public access fishers avoid between-group sources of error in a recreational fishing survey. Transactions of the American Fisheries Society, 142 (2):422–429, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashford:2013:WDV

[AJF13b] Julian R. Ashford, Cynthia M. Jones, and Lynn Fegley. Within-day variability in catch taken by public access fishers during a recreational fishing survey. Transactions of the American Fisheries Society, 142(4):974–978, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashford:2010:CDR

[AJFO10] Julian Ashford, Cynthia Jones, Lynn Fegley, and Robert O’Reilly. Catch data reported by telephone avoid public access bias in a marine recreational survey. Transactions of the American Fisheries Society, 139(6):1751–1757, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:2019:CSH

[AKK^{+}19] Joseph H. Anderson, Kirk L. Krueger, Clayton Kinsel, Timothy Quinn, William J. Ehinger, and Robert Bilby. Coho salmon and habitat response to restoration in a small stream. Transactions of the American Fisheries Society, 148(5):1024– 1038, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:2012:PSE

[AKM12] Joel D. Anderson, William J. Karel, and Allison C. S. Mione. Population structure and evolutionary history of southern flounder in the Gulf of Mexico and Western Atlantic Ocean. Transactions of the American Fisheries Society, 141(1):46–55, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alfermann:2013:CAF

[AM13] T. J. Alfermann and L. E. Miranda. Centrarchid assemblages in floodplain lakes of the Mississippi alluvial valley. Transactions of the American Fisheries Society, 142(2):323–332, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atcheson:2012:BRS

[AMBM12] Margaret E. Atcheson, Katherine W. Myers, David A. Beauchamp, and Nathan J. Mantua. Bioenergetic response by steelhead to variation in diet, thermal habitat, and climate in the North Pacific Ocean. Transactions of the American Fisheries Society, 141(4):1081–1096, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:2016:WDB

[AMM^{+}16] A. M. Anderson, E. Mittag, B. Middleton, B. Vondracek, and L. C. Ferrington, Jr. Winter diets of brown trout populations in Southeastern Minnesota and the significance of winteremerging invertebrates. Transactions of the American Fisheries Society, 145(1):206–220, January 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2010:BRa

[Ano10a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 139(2):635–639, March 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2010:BRb

[Ano10b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 139(3):925–928, May 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2010:Ea

[Ano10c] Anonymous. Errata. Transactions of the American Fisheries Society, 139(1):316–317, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2010:Eb

[Ano10d] Anonymous. Errata. Transactions of the American Fisheries Society, 139(4):1261, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2010:Ec

[Ano10e] Anonymous. Erratum. Transactions of the American Fisheries Society, 139(6):1913, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2011:EDB

[Ano11a] Anonymous. Erratum: Distinguishing between natural and hatchery Snake River fall Chinook salmon subyearlings in the field using body morphology (volume 140:21–30). Transactions of the American Fisheries Society, 140(6):1683, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2011:GA

[Ano11b] Anonymous. Guide for authors. Transactions of the American Fisheries Society, 140(1):201–206, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2012:E

[Ano12a] Anonymous. Erratum. Transactions of the American Fisheries Society, 141(2):570, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2012:GA

[Ano12b] Anonymous. Guide for authors. Transactions of the American Fisheries Society, 141(1):257–262, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2013:E

[Ano13a] Anonymous. ERRATUM. Transactions of the American Fisheries Society, 142(4):1156, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2013:TAF

[Ano13b] Anonymous. Transactions of the American fisheries society: Guide for authors. Transactions of the American Fisheries Society, 142(1):317–322, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2014:E

[Ano14a] Anonymous. Erratum. Transactions of the American Fisheries Society, 143(2):564, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2014:TAF

[Ano14b] Anonymous. Transactions of the American fisheries society: Guide for authors. Transactions of the American Fisheries Society, 143(2):310–315, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2015:E

[Ano15] Anonymous. Erratum. Transactions of the American Fisheries Society, 144(1):227, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2016:TAF

[Ano16] Anonymous. Transactions of the American fisheries society guide for authors. Transactions of the American Fisheries Society, 145(1):221–226, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2017:E

[Ano17a] Anonymous. Erratum. Transactions of the American Fisheries Society, 146(6):1315, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2017:TAF

[Ano17b] Anonymous. Transactions of the American fisheries society guide for authors. Transactions of the American Fisheries Society, 146(1):194–200, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Transactions of the American Fisheries Society, 147(1):1–2, January 2018. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Transactions of the American Fisheries Society, 147(2):245–246, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Transactions of the American Fisheries Society, 147(3):417–418, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Transactions of the American Fisheries Society, 147(4):621–622, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Transactions of the American Fisheries Society, 147(5):773–774, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Transactions of the American Fisheries Society, 147(6):1013–1014, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2018:RA

[Ano18g] Anonymous. Reviewer acknowledgments. Transactions of the American Fisheries Society, 147(6):1216–1218, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:E

[Ano19a] Anonymous. Erratum. Transactions of the American Fisheries Society, 148(5):1039, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:IIa

[Ano19b] Anonymous. Issue information. Transactions of the American Fisheries Society, 148(1):1–2, January 2019. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:IIb

[Ano19c] Anonymous. Issue information. Transactions of the American Fisheries Society, 148(2):241–242, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:IIc

[Ano19d] Anonymous. Issue information. Transactions of the American Fisheries Society, 148(3):507–508, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:IId

[Ano19e] Anonymous. Issue information. Transactions of the American Fisheries Society, 148(4):685–686, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:IIe

[Ano19f] Anonymous. Issue information. Transactions of the American Fisheries Society, 148(5):873–874, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:IIf

[Ano19g] Anonymous. Issue information. Transactions of the American Fisheries Society, 148(6):1041–1042, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:RA

[Ano19h] Anonymous. Reviewer acknowledgments. Transactions of the American Fisheries Society, 148(6):1146–1147, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2019:TAF

[Ano19i] Anonymous. Transactions of the American fisheries society: Guide for authors. Transactions of the American Fisheries Society, 148(1):233–239, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abrams:2018:CIR

[ARB^{+}18] Alice E. I. Abrams, Andrew M. Rous, Jill L. Brooks, Michael J. Lawrence, Jonathan D. Midwood, Susan E. Doka, and Steven J. Cooke. Comparing immobilization, recovery, and stress indicators associated with electric fish handling gloves and a portable electrosedation system. Transactions of the American Fisheries Society, 147(2):390–399, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:2011:EPS

[ARC^{+}11] Micheal S. Allen, Mark W. Rogers, Matthew J. Catalano, Daniel G. Gwinn, and Stephen J. Walsh. Evaluating the potential for stock size to limit recruitment in largemouth bass.

Transactions of the American Fisheries Society, 140(4):1093– 1100, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:2013:APA

[ARC^{+}13] Michael S. Allen, Mark W. Rogers, Matthew J. Catalano, Daniel C. Gwinn, and Stephen J. Walsh. Assessing the potential for the angling of nesting males to influence largemouth bass recruitment should consider the quality of nesting individuals: Response to comment. Transactions of the American Fisheries Society, 142(2):419–421, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [PWP13].

Archdeacon:2016:RSF

[Arc16] Thomas P. Archdeacon. Reduction in spring flow threatens Rio Grande silvery minnow: Trends in abundance during river intermittency. Transactions of the American Fisheries Society, 145(4):754–765, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adrean:2012:PEM

[ARL^{+}12] Lindsay J. Adrean, Daniel D. Roby, Donald E. Lyons, Ken Collis, and Allen F. Evans. Potential effects of management on Caspian tern Hydroprogne caspia predation on juvenile salmonids at a colony in San Francisco Bay, California. Transactions of the American Fisheries Society, 141(6):1682–1696, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abrahams:2012:RPV

[AS12] Mark V. Abrahams and Jennifer Sloan. Risk of predation, variation in dissolved oxygen, and their impact upon habitat selection decisions by fathead minnow. Transactions of the American Fisheries Society, 141(3):580–584, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkinson:2017:IWC

[AS17] Alexandra N. Atkinson and David H. Secor. Influence of winter conditions on the age, hatch dates, and growth of juvenile Atlantic menhaden in the Choptank River, Maryland. Transactions of the American Fisheries Society, 146(6):1126–1136, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anstead:2015:COS

[ASJ15] Kristen A. Anstead, Jason J. Schaffler, and Cynthia M. Jones. Coastwide otolith signatures of juvenile Atlantic menhaden, 2009–2011. Transactions of the American Fisheries Society, 144(1):96–106, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anstead:2016:CWN

[ASJ16] Kristen A. Anstead, Jason J. Schaffler, and Cynthia M. Jones. Coast-wide nursery contribution of new recruits to the population of Atlantic menhaden. Transactions of the American Fisheries Society, 145(3):627–636, May 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anstead:2017:CNA

[ASJ17] Kristen A. Anstead, Jason J. Schaffler, and Cynthia M. Jones. Contribution of nursery areas to the adult population of Atlantic menhaden. Transactions of the American Fisheries Society, 146(1):36–46, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andres:2018:GEW

[ASP^{+}18] Michael J. Andres, William T. Slack, Mark S. Peterson, Kayla D. Kimmel, Bradley R. Lewis, and Paul O. Grammer. Growth estimation of western population segment Gulf sturgeon using length-at-age and mark–recapture data. Transactions of the American Fisheries Society, 147(1):139–150, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allan:2018:PIP

[AUDL18] Hugh Allan, Peter Unmack, Richard P. Duncan, and Mark Lintermans. Potential impacts of PIT tagging on a critically endangered small-bodied fish: a trial on the surrogate Mountain Galaxias. Transactions of the American Fisheries Society, 147(6):1078–1084, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Altenritter:2019:LHT

[AW19] Matthew E. Altenritter and Benjamin D. Walther. The legacy of hypoxia: Tracking carryover effects of low oxygen exposure in a demersal fish using geochemical tracers. Transactions of the American Fisheries Society, 148(3):569–583, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bracis:2013:IRO

[BA13] C. Bracis and J. J. Anderson. Inferring the relative oceanic distribution of salmon from patterns in age-specific arrival timing. Transactions of the American Fisheries Society, 142(2): 556–567, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barth:2011:HRS

[BAHP11] Cameron C. Barth, W. Gary Anderson, L. M. Henderson, and Stephan J. Peake. Home range size and seasonal movement of juvenile lake sturgeon in a large river in the Hudson Bay drainage basin. Transactions of the American Fisheries Society, 140(6):1629–1641, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bergstedt:2012:BHU

[BAKT12] Roger A. Bergstedt, Ray L. Argyle, Charles C. Krueger, and William W. Taylor. Bathythermal habitat use by strains of Great Lakes- and Finger Lakes-origin lake trout in Lake Huron after a change in prey fish abundance and composition. Transactions of the American Fisheries Society, 141(2):263– 274, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2015:RCL

[BAR15] Daniel T. Brown, D. Derek Aday, and James A. Rice. Responses of coastal largemouth bass to episodic hypoxia. Transactions of the American Fisheries Society, 144(4):655–666, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brewton:2016:FED

[BAYS16] Rachel A. Brewton, Matthew J. Ajemian, Peter C. Young, and Gregory W. Stunz. Feeding ecology of dolphinfish in the Western Gulf of Mexico. Transactions of the American Fisheries Society, 145(4):839–853, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2012:TMS

[BAZ12] Peter J. Brown, Jenifer L. Ard, and Alexander V. Zale. Transverse mixing of simulated piscicides in small montane streams. Transactions of the American Fisheries Society, 141(2):353– 356, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benjamin:2010:DNS

[BB10a] Joseph R. Benjamin and Colden V. Baxter. Do nonnative salmonines exhibit greater density and production than the natives they replace? A comparison of nonnative brook trout with native cutthroat trout. Transactions of the American Fisheries Society, 139(3):641–651, May 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2010:HSA

[BB10b] Peter James Brown and Michael Anthony Bozek. Habitat selection and abundance of young-of-year smallmouth bass in north temperate lakes. Transactions of the American Fisheries Society, 139(4):1247–1260, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barkalow:2015:ESS

[BB15] Stephani L. Clark Barkalow and Scott A. Bonar. Effects of suspended sediment on early-life stage survival of Yaqui chub, an endangered USA–Mexico borderlands cyprinid. Transactions of the American Fisheries Society, 144(2):345–351, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beer:2019:GAR

[BBAK19] Stephanie Dowell Beer, Meredith L. Bartron, David G. Argent, and William G. Kimmel. Genetic assessment reveals population fragmentation and inbreeding in populations of brook trout in the Laurel Hill of Pennsylvania. Transactions of the American Fisheries Society, 148(3):620–635, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2014:SSM

[BBC^{+}14a] Terry D. Beacham, Richard J. Beamish, John R. Candy, Colin Wallace, Strahan Tucker, Jamal H. Moss, and Marc Trudel. Stock-specific migration pathways of juvenile sockeye salmon in British Columbia waters and in the Gulf of Alaska. Transactions of the American Fisheries Society, 143(6):1386–1403, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2014:SSS

[BBC^{+}14b] Terry D. Beacham, Richard J. Beamish, John R. Candy, Colin Wallace, Strahan Tucker, Jamal H. Moss, and Marc Trudel. Stock-specific size of juvenile sockeye salmon in British Columbia waters and the Gulf of Alaska. Transactions of the American Fisheries Society, 143(4):876–889, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berejikian:2014:MCO

[BBC14c] Barry A. Berejikian, Richard A. Bush, and Lance A. Campbell. Maternal control over offspring life history in a partially anadromous species, Oncorhynchus mykiss. Transactions of the American Fisheries Society, 143(2):369–379, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barkstedt:2018:FOD

[BBF^{+}18] Judith M. Barkstedt, Stephani L. Clark Barkalow, Michael A. Farrington, Jennifer L. Kennedy, and Steven P. Platania. Frequency of opercular deformities in age-0 native catostomids in the San Juan River from 1998 to 2012. Transactions of the American Fisheries Society, 147(6):1115–1123, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budnik:2018:AOC

[BBG^{+}18] Richard R. Budnik, Christopher T. Boehler, Joel E. Gagnon, John R. Farver, and Jeffrey G. Miner. Application of otolith chemistry to investigate the origin and state-straying of steelhead in Lake Erie tributaries. Transactions of the American Fisheries Society, 147(1):16–30, January 2018. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baillie:2015:DHS

[BBGB15] Shauna M. Baillie, Craig Blackie, Leonard Gerardi, and Paul Bentzen. Deciphering hatchery stock influences on wild populations of Vermont lake trout. Transactions of the American Fisheries Society, 144(1):124–139, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boldt:2012:IFC

[BBL^{+}12] Jennifer L. Boldt, Shannon C. Bartkiw, Pat A. Livingston, Gerald R. Hoff, and Gary E. Walters. Investigation of fishing and climate effects on the community size spectra of Eastern Bering Sea fish. Transactions of the American Fisheries Society, 141(2):327–342, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bidlack:2014:ISH

[BBM^{+}14] Allison L. Bidlack, Lee E. Benda, Tom Miewald, Gordon H. Reeves, and Gabriel McMahan. Identifying suitable habitat for Chinook salmon across a large, glaciated watershed. Transactions of the American Fisheries Society, 143(3):689–699, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brenden:2012:ASI

[BBS12] Travis O. Brenden, James R. Bence, and Emily B. Szalai. An age-structured integrated assessment of Chinook salmon population dynamics in Lake Huron’s main basin since 1968. Transactions of the American Fisheries Society, 141(4):919– 933, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benito:2015:DCE

[BBZGB15] Josep Benito, Lluís Benejam, Lluís Zamora, and Emili García-Berthou. Diel cycle and effects of water flow on activity and use of depth by common carp. Transactions of the American Fisheries Society, 144(3):491–501, May 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barkley:2012:DME

[BC12] Adam S. Barkley and Steven X. Cadrin. Discard mortality estimation of yellowtail flounder using reflex action mortality predictors. Transactions of the American Fisheries Society, 141(3):638–644, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berntson:2011:DRS

[BCF^{+}11] Ewann A. Berntson, Richard W. Carmichael, Michael W. Flesher, Eric J. Ward, and Paul Moran. Diminished reproductive success of steelhead from a hatchery supplementation program (Little Sheep Creek, Imnaha Basin, Oregon). Transactions of the American Fisheries Society, 140(3):685–698, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bevelhimer:2013:BRR

[BCF^{+}13] Mark S. Bevelhimer, Glenn F. Cada, Allison M. Fortner, Peter E. Schweizer, and Kristina Riemer. Behavioral responses of representative freshwater fish species to electromagnetic fields. Transactions of the American Fisheries Society, 142(3):802– 813, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2012:QMI

[BCG^{+}12] Richard S. Brown, Thomas J. Carlson, Andrew J. Gingerich, John R. Stephenson, Brett D. Pflugrath, Abigail E. Welch, Mike J. Langeslay, Martin L. Ahmann, Robert L. Johnson, John R. Skalski, Adam G. Seaburg, and Richard L. Townsend. Quantifying mortal injury of juvenile Chinook salmon exposed to simulated hydro-turbine passage. Transactions of the American Fisheries Society, 141(1):147–157, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blankenship:2011:MLM

[BCH^{+}11] Scott M. Blankenship, Matt R. Campbell, Jon E. Hess, Maureen A. Hess, Todd W. Kassler, Christine C. Kozfkay, Andrew P. Matala, Shawn R. Narum, Melanie M. Paquin, Maureen P. Small, Jeff J. Stephenson, Kenneth I. Warheit, and Paul Moran. Major lineages and metapopulations in Columbia River Oncorhynchus mykiss are structured by dynamic landscape features and environments. Transactions of the American Fisheries Society, 140(3):665–684, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowersox:2019:ELH

[BCM^{+}19] Brett J. Bowersox, Matthew P. Corsi, Joshua L. McCormick, Timothy Copeland, and Matthew R. Campbell. Examining life history shifts and genetic composition in a hatchery steelhead population, with implications for fishery and ocean selection. Transactions of the American Fisheries Society, 148(6):1056– 1068, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brule:2010:RSR

[BCMPDSZ10] Thierry Brulé, Teresa Colás-Marrufo, Esperanza Pérez-Díaz, and Juan Carlos Sámano-Zapata. Red snapper reproductive biology in the Southern Gulf of Mexico. Transactions of the American Fisheries Society, 139(4):957–968, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2011:MIC

[BCP^{+}11] Terry D. Beacham, John R. Candy, Erin Porszt, Shunpei Sato, and Shigehiko Urawa. Microsatellite identification of Canadian sockeye salmon rearing in the Bering Sea. Transactions of the American Fisheries Society, 140(2):296–306, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brinkman:2013:UTT

[BCR13] Stephen F. Brinkman, Harry J. Crockett, and Kevin B. Rogers. Upper thermal tolerance of mountain whitefish eggs and fry. Transactions of the American Fisheries Society, 142 (3):824–831, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benjamin:2013:PEC

[BCRP13] Joseph R. Benjamin, Patrick J. Connolly, Jason G. Romine, and Russell W. Perry. Potential effects of changes in temperature and food resources on life history trajectories of juvenile Oncorhynchus mykiss. Transactions of the American Fisheries Society, 142(1):208–220, January 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2019:NIB

[BCS19] Randy J. Brown, Michael B. Courtney, and Andrew C. Seitz. New insights into the biology of anadromous Dolly Varden in the Canning River, Arctic National Wildlife Refuge, Alaska. Transactions of the American Fisheries Society, 148(1):73–87, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2015:SDB

[BD15] Randy J. Brown and David W. Daum. Spawning distribution of Bering ciscoes in the Yukon River. Transactions of the American Fisheries Society, 144(2):292–299, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berlinsky:2015:PPP

[BDB^{+}15] David L. Berlinsky, Matthew A. DiMaggio, Timothy S. Breton, Jennifer Walsh, and Adrienne I. Kovach. Peritoneal pigmentation in purebred and hybrid river herring. Transactions of the American Fisheries Society, 144(4):717–723, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:2011:ESS

[BDBE11] James W. Boyd, Katherine A. Deters, Richard S. Brown, and M. Brad Eppard. Efficacy of single-suture incision closures in tagged juvenile Chinook salmon exposed to simulated turbine passage. Transactions of the American Fisheries Society, 140(5):1186–1192, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2010:USA

[BDM^{+}10] Richard S. Brown, Claude R. Duguay, Robert P. Mueller, Larry L. Moulton, Peter J. Doucette, and Jerry D. Tagestad. Use of Synthetic Aperture Radar (SAR) to identify and characterize overwintering areas of fish in ice-covered Arctic rivers: a demonstration with broad whitefish and their habitats in the Sagavanirktok River, Alaska. Transactions of the American Fisheries Society, 139(6):1711–1722, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barks:2010:LAS

[BDS10] Patrick M. Barks, Jennifer L. Doucette, and Christopher M. Somers. Lack of angling-sized yellow perch in a Canadian boreal lake: Potential influences of growth rate, diet, and predation by double-crested cormorants. Transactions of the American Fisheries Society, 139(4):1029–1040, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangs:2017:AIF

[BDTD17] Max R. Bangs, Marlis R. Douglas, Paul Thompson, and Michael E. Douglas. Anthropogenic impacts facilitate native fish hybridization in the Bonneville Basin of Western North America. Transactions of the American Fisheries Society, 146 (1):16–21, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2010:RTE

[Bea10] Terry D. Beacham. Revisiting trends in the evolution of egg size in hatchery-enhanced populations of Chinook salmon from British Columbia. Transactions of the American Fisheries Society, 139(2):579–585, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:2012:WFO

[Bel12] Richard J. Bell. Winter feeding as an overwintering survival strategy in young-of-the-year winter flounder. Transactions of the American Fisheries Society, 141(4):855–871, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bromaghin:2011:UGM

[BEMF11] Jeffrey F. Bromaghin, Danielle F. Evenson, Thomas H. McLain, and Blair G. Flannery. Using a genetic mixture model to study phenotypic traits: Differential fecundity among Yukon River Chinook salmon. Transactions of the American Fisheries Society, 140(2):235–249, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berra:2013:REA

[Ber13] Tim M. Berra. A review of “Ecology of Australian Freshwater Fishes” Edited by Paul Humphries and Keith Walker. CSIRO Publishing, Collingwood, Victoria, Australia. 2013. 423 pages.

$A130.00 (hardcover). Transactions of the American Fisheries Society, 142(5):1477–1480, September 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bradley:2015:CJF

[BES15] Parker T. Bradley, Mark D. Evans, and Andrew C. Seitz. Characterizing the juvenile fish community in turbid Alaskan rivers to assess potential interactions with hydrokinetic devices. Transactions of the American Fisheries Society, 144(5): 1058–1069, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldigo:2010:VRF

[BEWM10] Barry P. Baldigo, Anne G. Ernst, Dana R. Warren, and Sarah J. Miller. Variable responses of fish assemblages, habitat, and stability to natural-channel-design restoration in Catskill Mountain streams. Transactions of the American Fisheries Society, 139(2):449–467, March 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Booth:2014:MFT

[BFH14] Michael T. Booth, Alexander S. Flecker, and Nelson G. Hairston, Jr. Is mobility a fixed trait? Summer movement patterns of catostomids using PIT telemetry. Transactions of the American Fisheries Society, 143(4):1098–1111, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bailey:2015:MSJ

[BFK15] David S. Bailey, Elizabeth A. Fairchild, and Linda H. Kalnejais. Microchemical signatures in juvenile winter flounder otoliths provide identification of natal nurseries. Transactions of the American Fisheries Society, 144(1):173–183, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burwen:2010:APS

[BFM10] Debby L. Burwen, Steven J. Fleischman, and James D. Miller. Accuracy and precision of salmon length estimates taken from DIDSON sonar images. Transactions of the American Fisheries Society, 139(5):1306–1314, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bellquist:2016:LTD

[BGB^{+}16] Lyall F. Bellquist, Jeffrey B. Graham, Amanda Barker, Jason Ho, and Brice X. Semmens. Long-term dynamics in “Trophy” sizes of pelagic and coastal pelagic fishes among California recreational fisheries (1966–2013). Transactions of the American Fisheries Society, 145(5):977–989, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldigo:2019:AIG

[BGLP19] Barry P. Baldigo, Scott D. George, Gregory B. Lawrence, and Eric A. Paul. Acidification impacts and goals for gauging recovery of brook trout populations and fish communities in streams of the Western Adirondack Mountains, New York, USA. Transactions of the American Fisheries Society, 148 (2):373–392, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buehrens:2013:SSS

[BGOQ13] T. W. Buehrens, J. Glasgow, C. O. Ostberg, and T. P. Quinn. Spatial segregation of spawning habitat limits hybridization between sympatric native steelhead and coastal cutthroat trout. Transactions of the American Fisheries Society, 142(1): 221–233, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berejikian:2011:ETO

[BGV11] Barry A. Berejikian, Joshua T. Gable, and Dmitri T. Vidergar. Effectiveness and trade-offs associated with hydraulic egg collections from natural salmon and steelhead redds for conservation hatchery programs. Transactions of the American Fisheries Society, 140(3):549–556, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Balazik:2012:EEF

[BGV^{+}12] Matthew T. Balazik, Greg C. Garman, Joel P. Van Eenennaam, Jerre Mohler, and L. Curry Woods III. Empirical evidence of fall spawning by Atlantic sturgeon in the James River, Virginia. Transactions of the American Fisheries Society, 141 (6):1465–1471, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyer:2017:RES

[BGW^{+}17] Jan K. Boyer, Christopher S. Guy, Molly A. H. Webb, Travis B. Horton, and Thomas E. McMahon. Reproductive ecology, spawning behavior, and juvenile distribution of mountain whitefish in the Madison River, Montana. Transactions of the American Fisheries Society, 146(5):939–954, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boley:2011:LSI

[BH11] Ryan M. Boley and Edward J. Heist. Larval surveys indicate low levels of endangered pallid sturgeon reproduction in the Middle Mississippi River. Transactions of the American Fisheries Society, 140(6):1604–1612, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boone:2013:EFS

[BHC^{+}13] S. Shaun Boone, Sonia M. Hernandez, Alvin C. Camus, Douglas L. Peterson, Cecil A. Jennings, James L. Shelton, and Stephen J. Divers. Evaluation of four suture materials for surgical incision closure in Siberian sturgeon. Transactions of the American Fisheries Society, 142(3):649–659, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bass:2018:VGN

[BHC^{+}18] Arthur Bass, Scott G. Hinch, Matthew T. Casselman, Nolan N. Bett, Nicholas J. Burnett, Collin T. Middleton, and David A. Patterson. Visible gill-net injuries predict migration and spawning failure in adult sockeye salmon. Transactions of the American Fisheries Society, 147(6):1085–1099, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benjamin:2016:TRN

[BHD^{+}16] Joseph R. Benjamin, Jeannie M. Heltzel, Jason B. Dunham, Michael Heck, and Nolan Banish. Thermal regimes, nonnative trout, and their influences on native bull trout in the Upper Klamath River Basin, Oregon. Transactions of the American Fisheries Society, 145(6):1318–1330, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bronte:2010:DSC

[BHG^{+}10] Charles R. Bronte, Michael H. Hoff, Owen T. Gorman, Wayne E. Thogmartin, Philip J. Schneeberger, and Thomas N. Todd. Decline of the shortjaw cisco in Lake Superior: The role of overfishing and risk of extinction. Transactions of the American Fisheries Society, 139(3):735–748, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burroughs:2010:ESD

[BHK^{+}10] Bryan A. Burroughs, Daniel B. Hayes, Kristi D. Klomp, Jonathan F. Hansen, and Jessica Mistak. The effects of the stronach dam removal on fish in the Pine River, Manistee County, Michigan. Transactions of the American Fisheries Society, 139(5):1595–1613, September 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bailey:2010:SDD

[BHLK10] Michael M. Bailey, Gregg E. Horton, Benjamin H. Letcher, and Michael T. Kinnison. Seasonal density dependence in Atlantic salmon over varying spatial scales. Transactions of the American Fisheries Society, 139(6):1642–1656, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boughton:2015:TPS

[BHP^{+}15] David A. Boughton, Lee R. Harrison, Andrew S. Pike, Juan L. Arriaza, and Marc Mangel. Thermal potential for steelhead life history expression in a Southern California alluvial river. Transactions of the American Fisheries Society, 144(2):258– 273, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:2017:IDH

[BHS^{+}17] Brian R. Beckman, Deborah L. Harstad, Dina K. Spangenberg, Ryan S. Gerstenberger, Chris V. Brun, and Donald A. Larsen. The impact of different hatchery rearing environments on smolt-to-adult survival of spring Chinook salmon. Transactions of the American Fisheries Society, 146(3):539–555, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barron:2013:ESD

[BJA^{+}13] James M. Barron, Nathan R. Jensen, Paul J. Anders, Joshua P. Egan, Susan C. Ireland, and Kenneth D. Cain. Effects of stocking density on survival and yield of North American burbot reared under semi-intensive conditions. Transactions of the American Fisheries Society, 142(6):1680–1687, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burke:2012:HVD

[BJC12] Lindsey A. Burke, Richard M. Jobin, and David W. Coltman. Hydrogeographic vicariance determines the genetic structure of northwestern walleye populations. Transactions of the American Fisheries Society, 141(3):697–706, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burchard:2014:DRP

[BJR14] Katie A. Burchard, Francis Juanes, and Rodney A. Rountree. Diel reproductive periodicity of haddock in the Southwestern Gulf of Maine. Transactions of the American Fisheries Society, 143(2):451–466, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baumann:2011:TRI

[BK11] Jessica R. Baumann and Thomas J. Kwak. Trophic relations of introduced flathead catfish in an Atlantic river. Transactions of the American Fisheries Society, 140(4):1120–1134, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:2019:EHF

[BKC^{+}19] Luke M. Bower, Friedrich W. Keppeler, Eduardo R. Cunha, Yasmin Quintana, David E. Saenz, Edwin O. Lopez-Delgado, Thethela Bokhutlo, Caroline C. Arantes, Marcelo C. Andrade, Clinton R. Robertson, Kevin B. Mayes, and Kirk O. Winemiller. Effects of hydrology on fish diversity and assemblage structure in a Texan coastal plains river. Transactions of the American Fisheries Society, 148(1):207–218, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buchheister:2011:TES

[BL11] Andre Buchheister and Robert J. Latour. Trophic ecology of summer flounder in Lower Chesapeake Bay inferred from stomach content and stable isotope analyses. Transactions of the American Fisheries Society, 140(5):1240–1254, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burton:2013:CDP

[BLB^{+}13] Karl D. Burton, Larry G. Lowe, Hans B. Berge, Heidy K. Barnett, and Paul L. Faulds. Comparative dispersal patterns for recolonizing Cedar River Chinook salmon above Landsburg Dam, Washington, and the source population below the dam. Transactions of the American Fisheries Society, 142(3):703– 716, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:2013:RSI

[BLCB13] Preston T. Bean, Dijar J. Lutz-Carrillo, and Timothy H. Bonner. Rangewide survey of the introgressive status of guadalupe bass: Implications for conservation and management. Transactions of the American Fisheries Society, 142(3):681–689, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barthel:2010:GRA

[BLCN^{+}10] Brandon L. Barthel, Dijar J. Lutz-Carrillo, Kristen E. Norberg, Wesley F. Porak, Michael D. Tringali, Todd W. Kassler, William E. Johnson, Anne M. Readel, Richard A. Krause, and David P. Philipp. Genetic relationships among populations of Florida bass. Transactions of the American Fisheries Society, 139(6):1615–1641, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Balazik:2013:CME

[BLG^{+}13] Matthew T. Balazik, Briana C. Langford, Greg C. Garman, Michael L. Fine, Jennifer K. Stewart, Robert J. Latour, and Stephen P. McIninch. Comparison of MS-222 and electronarcosis as anesthetics on cortisol levels in juvenile Atlantic sturgeon. Transactions of the American Fisheries Society, 142(6): 1640–1643, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bestgen:2010:SPF

[BMB^{+}10] Kevin R. Bestgen, Brent Mefford, Jay M. Bundy, Cameron D. Walford, and Robert I. Compton. Swimming performance and fishway model passage success of Rio Grande silvery minnow. Transactions of the American Fisheries Society, 139(2):433– 448, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beardsall:2013:CIO

[BMC^{+}13] Jeffrey W. Beardsall, Montana F. McLean, Steven J. Cooke, Brian C. Wilson, Michael J. Dadswell, Anna M. Redden, and Michael J. W. Stokesbury. Consequences of incidental otter trawl capture on survival and physiological condition of threatened Atlantic sturgeon. Transactions of the American Fisheries Society, 142(5):1202–1214, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bacon:2015:CCS

[BMF^{+}15] P. J. Bacon, I. A. Malcolm, R. J. Fryer, R. S. Glover, C. P. Millar, and A. F. Youngson. Can conservation stocking enhance juvenile emigrant production in wild Atlantic salmon? Transactions of the American Fisheries Society, 144(3):642– 654, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Balazik:2012:AGA

[BMGL12] Matthew T. Balazik, Stephen P. McIninch, Greg C. Garman, and Robert J. Latour. Age and growth of Atlantic sturgeon in the James River, Virginia, 1997–2011. Transactions of the American Fisheries Society, 141(4):1074–1080, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barrientos:2019:AGM

[BMH19] Christian A. Barrientos, Debra J. Murie, and Jeffrey E. Hill. Age, growth, and mortality rates of the giant cichlid in Guatemala. Transactions of the American Fisheries Society, 148(1):176–190, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Binder:2010:RIW

[BMM10] Thomas R. Binder, Robert L. McLaughlin, and D. Gordon McDonald. Relative importance of water temperature, water level, and lunar cycle to migratory activity in spawning-phase sea lampreys in Lake Ontario. Transactions of the American Fisheries Society, 139(3):700–712, May 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benjamin:2016:CGU

[BMOF16] Alyssa Benjamin, Bernie May, John O’Brien, and Amanda J. Finger. Conservation genetics of an urban desert fish, the arroyo chub. Transactions of the American Fisheries Society, 145(2):277–286, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buonaccorsi:2017:PGA

[BMP^{+}17] V. P. Buonaccorsi, J. Malloy, M. Peterson, K. Brubaker, and C. J. Grant. Population genomic analysis of brook trout in Pennsylvania’s Appalachian region. Transactions of the American Fisheries Society, 146(3):485–494, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bunnell:2012:EMU

[BMR^{+}12] David B. Bunnell, Charles P. Madenjian, Mark W. Rogers, Jeffrey D. Holuszko, and Linda J. Begnoche. Exploring mechanisms underlying sex-specific differences in mortality of Lake Michigan bloaters. Transactions of the American Fisheries Society, 141(1):204–214, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butts:2013:DSC

[BMTP13] Ian A. E. Butts, Ali Mokdad, Edward A. Trippel, and Trevor E. Pitcher. Development of a sperm cryopreservation protocol for redside dace: Implications for genome resource banking. Transactions of the American Fisheries Society, 142 (3):671–680, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bies:2016:CPG

[BN16] Jason M. Bies and J. Wesley Neal. Comparison of pulsed gastric lavage and acrylic stomach tubes for sampling the diet of butterfly peacock bass. Transactions of the American Fisheries Society, 145(4):854–859, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benda:2015:CPF

[BNC^{+}15] Susan E. Benda, George P. Naughton, Christopher C. Caudill, Michael L. Kent, and Carl B. Schreck. Cool, pathogenfree refuge lowers pathogen-associated prespawn mortality of Willamette River Chinook salmon. Transactions of the American Fisheries Society, 144(6):1159–1172, November 2015. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brownscombe:2014:FRS

[BNS^{+}14] Jacob W. Brownscombe, Liane Nowell, Emma Samson, Andy J. Danylchuk, and Steven J. Cooke. Fishing-related stressors inhibit refuge-seeking behavior in released subadult great barracuda. Transactions of the American Fisheries Society, 143(3):613–617, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:2016:EOL

[BNS^{+}16] Richard J. Beamish, Chrys M. Neville, Ruston M. Sweeting, Terry D. Beacham, Joy Wade, and Lingbo Li. Early ocean life history of Harrison River sockeye salmon and their contribution to the biodiversity of sockeye salmon in the Fraser River, British Columbia, Canada. Transactions of the American Fisheries Society, 145(2):348–362, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2017:TEB

[BNTT17] Terry D. Beacham, Chrys M. Neville, Strahan Tucker, and Marc Trudel. Is there evidence for biologically significant sizeselective mortality of Coho salmon during the first winter of marine residence? Transactions of the American Fisheries Society, 146(3):395–407, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Binder:2012:LBS

[BNW^{+}12] Thomas R. Binder, Michael A. Nannini, David H. Wahl, Robert Arlinghaus, Thomas Klefoth, David P. Philipp, and Steven J. Cooke. Largemouth bass selected for differential vulnerability to angling exhibit similar routine locomotory activity in experimental ponds. Transactions of the American Fisheries Society, 141(5):1252–1259, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouska:2010:RCD

[BP10] Wesley W. Bouska and Craig P. Paukert. Road crossing designs and their impact on fish assemblages of Great Plains streams. Transactions of the American Fisheries Society, 139 (1):214–222, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barnett:2013:MAB

[BP13] Heidy K. Barnett and Dwayne K. Paige. Movements by adfluvial bull trout during the spawning season between lake and river habitats. Transactions of the American Fisheries Society, 142(3):876–883, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bahr:2016:RJA

[BP16] Derek L. Bahr and Douglas L. Peterson. Recruitment of juvenile Atlantic sturgeon in the Savannah River, Georgia. Transactions of the American Fisheries Society, 145(6):1171–1178, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bahr:2017:SSS

[BP17] Derek L. Bahr and Douglas L. Peterson. Status of the shortnose sturgeon population in the Savannah River, Georgia. Transactions of the American Fisheries Society, 146(1):92–98, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bacheler:2011:CAD

[BPS11] Nathan M. Bacheler, Tammie J. Paoli, and Garret M. Schacht. Controls on abundance and distribution of yellow perch: Predator, water quality, and density-dependent effects. Transactions of the American Fisheries Society, 140(4):989–1000, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barnett:2019:DME

[BQF^{+}19] Heidy K. Barnett, Thomas P. Quinn, Kurt L. Fresh, Michele E. Koehler, Karl D. Burton, Mary Bhuthimethee, and N. Phil Peterson. Differential marking of embryos by location and date of release reveals within-river natal homing and parental influence on progeny return timing in sockeye salmon. Transactions of the American Fisheries Society, 148 (2):393–405, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beard:2017:HAJ

[BQHR17] Zachary S. Beard, Michael C. Quist, Ryan S. Hardy, and Tyler J. Ross. Habitat associations of juvenile burbot in a tributary of the Kootenai River. Transactions of the American Fisheries Society, 146(5):1008–1015, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bhagat:2011:TFS

[BR11] Yakuta Bhagat and Carl R. Ruetz III. Temporal and fine-scale spatial variation in fish assemblage structure in a drowned river mouth system of Lake Michigan. Transactions of the American Fisheries Society, 140(6):1429–1440, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boucek:2014:EEC

[BR14] Ross E. Boucek and Jennifer S. Rehage. Examining the effectiveness of consumer diet sampling as a nonnative detection tool in a subtropical estuary. Transactions of the American Fisheries Society, 143(2):489–494, March 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bethke:2014:WPS

[BRA14] Bethany J. Bethke, James A. Rice, and D. Derek Aday. White perch in small North Carolina reservoirs: What explains variation in population structure? Transactions of the American Fisheries Society, 143(1):77–84, January 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bestgen:2019:DPN

[BRG19] Kevin R. Bestgen, Kevin B. Rogers, and Robert Granger. Distinct phenotypes of native cutthroat trout emerge under a molecular model of lineage distributions. Transactions of the American Fisheries Society, 148(2):442–463, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bigelow:2010:ASL

[BRGR10] John P. Bigelow, Wendy M. Rauw, and Luis Gomez-Raya. Acclimation in simulated lake water increases survival of Lahontan cutthroat trout challenged with saline, alkaline water from Walker Lake, Nevada. Transactions of the American Fisheries Society, 139(3):876–887, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowerman:2018:PMF

[BRK^{+}18] Tracy Bowerman, Adrienne Roumasset, Matthew L. Keefer, Cameron S. Sharpe, and Christopher C. Caudill. Prespawn mortality of female Chinook salmon increases with water temperature and percent hatchery origin. Transactions of the American Fisheries Society, 147(1):31–42, January 2018. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouma:2012:LJP

[BRS^{+}12] Joshua V. Bouma, Don P. Rothaus, Kristina M. Straus, Brent Vadopalas, and Carolyn S. Friedman. Low juvenile pinto abalone Haliotis kamtschatkana kamtschatkana abundance in the San Juan Archipelago, Washington State. Transactions of the American Fisheries Society, 141(1):76–83, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bentley:2013:BCC

[BS13] Kale T. Bentley and Daniel E. Schindler. Body condition correlates with instantaneous growth in stream-dwelling rainbow trout and Arctic grayling. Transactions of the American Fisheries Society, 142(3):747–755, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bethke:2015:CRA

[BS15] Bethany J. Bethke and David F. Staples. Changes in relative abundance of several Minnesota fishes from 1970 to 2013. Transactions of the American Fisheries Society, 144(1):68–80, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blain:2016:RSB

[BS16] Brittany J. Blain and Trent M. Sutton. Reproductive status and blood plasma indicators of sex and gonad maturation status for yelloweye rockfish following barotrauma and recompression events. Transactions of the American Fisheries Society, 145(6):1234–1240, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beer:2018:IIT

[BS18] W. Nicholas Beer and E. Ashley Steel. Impacts and implications of temperature variability on Chinook salmon egg development and emergence phenology. Transactions of the American Fisheries Society, 147(1):3–15, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beakes:2010:STT

[BSC^{+}10] Michael P. Beakes, William H. Satterthwaite, Erin M. Collins, David R. Swank, Joseph E. Merz, Robert G. Titus, Susan M. Sogard, and Marc Mangel. Smolt transformation in two California steelhead populations: Effects of temporal variability in growth. Transactions of the American Fisheries Society, 139(5):1263–1275, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandl:2016:GQP

[BSC^{+}16] Scott C. Brandl, Brian M. Schreier, J. Louise Conrad, Bernie May, and Melinda R. Baerwald. Generation of quantitative polymerase chain reaction detectability half-lives and comparison of sampling protocols for genetic diet studies of San Francisco Estuary fishes. Transactions of the American Fisheries Society, 145(3):441–449, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bevelhimer:2017:HAB

[BSCA17] Mark Bevelhimer, Constantin Scherelis, Jonathan Colby, and Mary Ann Adonizio. Hydroacoustic assessment of behavioral responses by fish passing near an operating tidal turbine in the East River, New York. Transactions of the American Fisheries Society, 146(5):1028–1042, September 2017. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckmeier:2013:AGM

[BSD13] David L. Buckmeier, Nathan G. Smith, and Daniel J. Daugherty. Alligator gar movement and macrohabitat use in the Lower Trinity River, Texas. Transactions of the American Fisheries Society, 142(4):1025–1035, July 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldigo:2017:EED

[BSGB17] Barry P. Baldigo, Lee Ann Sporn, Scott D. George, and Jacob A. Ball. Efficacy of environmental DNA to detect and quantify brook trout populations in headwater streams of the Adirondack Mountains, New York. Transactions of the American Fisheries Society, 146(1):99–111, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Billman:2011:TPR

[BSHK^{+}11] Hilary G. Billman, Sophie St-Hilaire, Carter G. Kruse, Teri S. Peterson, and Charles R. Peterson. Toxicity of the piscicide rotenone to Columbia spotted frog and boreal toad tadpoles. Transactions of the American Fisheries Society, 140(4):919– 927, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baumann:2013:LLB

[BSM13] Hannes Baumann, Sandra J. Sutherland, and Richard S. McBride. Longitudinal length back-calculations from otoliths and scales differ systematically in haddock. Transactions of the American Fisheries Society, 142(1):184–192, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckmeier:2012:UAG

[BSR12] David L. Buckmeier, Nathan G. Smith, and Kerry S. Reeves. Utility of alligator gar age estimates from otoliths, pectoral fin rays, and scales. Transactions of the American Fisheries Society, 141(6):1510–1519, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartron:2018:EEN

[BSS18] Meredith L. Bartron, Nicholas M. Sard, and Kim T. Scribner. Evaluation of effective number of breeders and coancestry among progeny produced using common hatchery mating strategies. Transactions of the American Fisheries Society, 147(1):185–194, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baerwald:2012:DTD

[BSSM12] Melinda R. Baerwald, Brian M. Schreier, Gregg Schumer, and Bernie May. Detection of threatened delta smelt in the gut contents of the invasive Mississippi silverside in the San Francisco Estuary using TaqMan assays. Transactions of the American Fisheries Society, 141(6):1600–1607, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckmeier:2018:AEL

[BSSP18] David L. Buckmeier, Richard Snow, Nathan G. Smith, and Clayton Porter. Are age estimates for longnose gar and spotted gar accurate? An evaluation of sagittal otoliths, pectoral fin rays, and branchiostegal rays. Transactions of the American Fisheries Society, 147(4):639–648, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See comment [SKBS18].

Balouskus:2016:FBC

[BT16] Richard G. Balouskus and Timothy E. Targett. Fish and blue crab density along a riprap-sill-hardened shoreline: Comparisons with Spartina marsh and riprap. Transactions of the American Fisheries Society, 145(4):766–773, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowker:2017:SAE

[BTWB17] James D. Bowker, Jesse T. Trushenski, Niccole Wandelear, and Molly P. Bowman. Safety of AQUI-S 20E (10% eugenol) as a sedative for freshwater fish. Transactions of the American Fisheries Society, 146(3):384–394, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berejikian:2011:AMR

[BVA11] Barry A. Berejikian, Donald M. Van Doornik, and Jeffrey J. Atkins. Alternative male reproductive phenotypes affect offspring growth rates in Chinook salmon. Transactions of the American Fisheries Society, 140(5):1206–1212, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:2010:CCD

[BW10] Steven E. Butler and David H. Wahl. Common carp distribution, movements, and habitat use in a river impounded by multiple low-head dams. Transactions of the American Fisheries Society, 139(4):1121–1135, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bond:2017:CEB

[BWD^{+}17] Morgan H. Bond, Peter A. H. Westley, Andrew H. Dittman, Dean Holecek, Tiffani Marsh, and Thomas P. Quinn. Combined effects of barge transportation, river environment, and rearing location on straying and migration of adult Snake River fall-run Chinook salmon. Transactions of the American Fisheries Society, 146(1):60–73, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2011:CSP

[BWDM11] Terry D. Beacham, Michael Wetklo, Langtuo Deng, and Cathy MacConnachie. Coho salmon population structure in North America determined from microsatellites. Transactions of the American Fisheries Society, 140(2):253–270, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bestgen:2017:DFR

[BWHF17] Kevin R. Bestgen, Cameron T. Wilcox, Angela A. Hill, and Kurt D. Fausch. A dynamic flow regime supports an intact Great Plains stream fish assemblage. Transactions of the American Fisheries Society, 146(5):903–916, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowersox:2016:GRB

[BWRA16] Brett Bowersox, Thea Wickersham, Laura Redfield, and Michael W. Ackerman. The genetic relationship between anadromous and resident Oncorhynchus mykiss at a putative barrier with implications for habitat improvement. Transactions of the American Fisheries Society, 145(2):305–318, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baumsteiger:2019:UDP

[BYM19] Jason Baumsteiger, Matthew Young, and Peter B. Moyle. Using the Distinct Population Segment (DPS) concept to protect fishes with low levels of genomic differentiation: Conservation of an endemic minnow (hitch). Transactions of the American Fisheries Society, 148(2):406–416, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2012:RPM

[BZ12] Peter J. Brown and Alexander V. Zale. Rotenone persistence model for montane streams. Transactions of the American Fisheries Society, 141(2):560–569, March 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Curtis:2011:SDH

[CAB11] Tobey H. Curtis, Douglas H. Adams, and George H. Burgess. Seasonal distribution and habitat associations of bull sharks in the Indian River Lagoon, Florida: a 30-year synthesis. Transactions of the American Fisheries Society, 140(5):1213–1226, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caballero-Arango:2013:RBT

[CABNQ^{+}13] Doralice Caballero-Arango, Thierry Brulé, Virginia Nóh-Quiñones, Teresa Colás-Marrufo, and Esperanza Pérez-Díaz. Reproductive biology of the tiger grouper in the Southern Gulf of Mexico. Transactions of the American Fisheries Society, 142(1):282–299, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Copeland:2019:PIW

[CBAC19] Timothy Copeland, Brett J. Bowersox, Michael W. Ackerman, and Carlos Camacho. Patterns of iteroparity in wild Snake River steelhead. Transactions of the American Fisheries Society, 148(5):926–937, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:2013:PSN

[CBM^{+}13] Sarah M. Collins, Nate Bickford, Peter B. McIntyre, Aurélie Coulon, Amber J. Ulseth, Donald C. Taphorn, and Alexander S. Flecker. Population structure of a neotropical migratory fish: Contrasting perspectives from genetics and otolith microchemistry. Transactions of the American Fisheries Society, 142(5):1192–1201, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutre:2015:TVD

[CBM15] K. M. Coutré, A. H. Beaudreau, and P. W. Malecha. Temporal variation in diet composition and use of pulsed resource subsidies by juvenile sablefish. Transactions of the American Fisheries Society, 144(4):807–819, July 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collingsworth:2014:CRD

[CBMR14] Paris D. Collingsworth, David B. Bunnell, Charles P. Madenjian, and Stephen C. Riley. Comparative recruitment dynamics of alewife and bloater in Lakes Michigan and Huron. Transactions of the American Fisheries Society, 143(2):294– 309, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conner:2015:CTS

[CBSB15] Mary M. Conner, Stephen N. Bennett, W. Carl Saunders, and Nicolaas Bouwes. Comparison of tributary survival estimates of steelhead using Cormack–Jolly–Seber and Barker models: Implications for sampling efforts and designs. Transactions of the American Fisheries Society, 144(1):34–47, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2015:CBS

[CBVF15] Lance A. Campbell, Daniel L. Bottom, Eric C. Volk, and Ian A. Fleming. Correspondence between scale morphometrics and scale and otolith chemistry for interpreting juvenile salmon life histories. Transactions of the American Fisheries Society, 144(1):55–67, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conrad:2016:NSI

[CBW^{+}16] J. Louise Conrad, Andrew J. Bibian, Kelly L. Weinersmith, Denise De Carion, Matthew J. Young, Patrick Crain, Erin L. Hestir, Maria J. Santos, and Andrew Sih. Novel species interactions in a highly modified estuary: Association of largemouth bass with Brazilian waterweed Egeria densa. Transactions of the American Fisheries Society, 145(2):249–263, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Claiborne:2016:EBC

[CC16] Andrew M. Claiborne and Lance A. Campbell. Evaluation of back-calculated size and timing estimates for juvenile Chinook salmon using otolith structure and chemistry. Transactions of the American Fisheries Society, 145(3):493–501, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooke:2016:UHI

[CCB16] S. J. Cooke, C. J. A. Cooke, and J. W. Brownscombe. Ultralight or heavyweight? The interactive effects of gear strength and air exposure on reflex impairment in largemouth bass. Transactions of the American Fisheries Society, 145(6):1202– 1207, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:2015:CNP

[CCH15] Richard W. Carmichael, Mark W. Chilcote, and Charles W. Huntington. Comment: Natural productivity in steelhead populations of natural and hatchery origin — assessing hatchery spawner influence. Transactions of the American Fisheries Society, 144(6):1292–1297, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [Lis15].

Chen:2011:EWA

[CCKF11] Yongjiu Chen, Michael R. Childs, and Connie Keeler-Foster. Evaluation of woundfin augmentation efforts in the Virgin River by estimation of admixture proportions. Transactions of the American Fisheries Society, 140(3):598–604, May 2011.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2019:NDA

[CCL^{+}19] Meredith C. Campbell, Colleen A. Caldwell, Teresa D. Lewis, Wade D. Wilson, and Charlotte C. Gard. Nonlethal detection of Asian fish tapeworm in the federally endangered humpback chub using a molecular screening tool. Transactions of the American Fisheries Society, 148(4):832–842, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:2010:TSE

[CDB10] A. M. Cook, J. Duston, and R. G. Bradford. Temperature and salinity effects on survival and growth of early life stage Shubenacadie River striped bass. Transactions of the American Fisheries Society, 139(3):749–757, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chamberlin:2011:IHR

[CEFQ11] Joshua W. Chamberlin, Timothy E. Essington, John W. Ferguson, and Thomas P. Quinn. The influence of hatchery rearing practices on salmon migratory behavior: Is the tendency of Chinook salmon to remain within Puget Sound affected by size and date of release? Transactions of the American Fisheries Society, 140(5):1398–1408, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cushman:2019:DSM

[CEM^{+}19] Elizabeth L. Cushman, Heather K. Evans, Gregory R. Moyer, Morgan E. Raley, Ashantye S. Williams, and Tanya L. Darden. Development of a standardized molecular tool and estimation of genetic measures for responsible aquaculture-based fisheries enhancement of American shad in North and South Carolina. Transactions of the American Fisheries Society, 148(1):148– 162, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Claiborne:2011:SRS

[CFHE11] Andrew M. Claiborne, Joseph P. Fisher, Sean A. Hayes, and Robert L. Emmett. Size at release, size-selective mortality, and age of maturity of Willamette River hatchery yearling Chinook salmon. Transactions of the American Fisheries Society, 140 (4):1135–1144, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coggins:2013:EAL

[CGA13] Lewis G. Coggins, Jr., Daniel C. Gwinn, and Micheal S. Allen. Evaluation of age–length key sample sizes required to estimate fish total mortality and growth. Transactions of the American Fisheries Society, 142(3):832–840, May 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cathcart:2019:SLW

[CGB19] Charles N. Cathcart, Keith B. Gido, and W. Howard Brandenburg. Spawning locations within and among tributaries influence flannelmouth sucker offspring experience. Transactions of the American Fisheries Society, 148(5):963–977, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crowley:2018:AME

[CGDL18] Claire E. Crowley, Ryan L. Gandy, Kendra L. Daly, and Erin Hoerl Leone. Assessment of maturity in the Eastern Gulf of Mexico Florida stone crab Menippe mercenaria fishery. Transactions of the American Fisheries Society, 147(6): 1124–1145, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Camp:2011:URP

[CGL^{+}11] Edward V. Camp, Daniel C. Gwinn, Matthew V. Lauretta, William E. Pine, and Thomas K. Frazer. Use of recovery probabilities can improve sampling efficiency for throw traps in vegetated habitats. Transactions of the American Fisheries Society, 140(1):164–169, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cathcart:2015:FCD

[CGM15] C. Nathan Cathcart, Keith B. Gido, and Mark C. McKinstry. Fish community distributions and movements in two tributaries of the San Juan River, USA. Transactions of the American Fisheries Society, 144(5):1013–1028, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:2013:RPA

[Cha13] Ross Chapman. A review of “Passive Acoustic Monitoring of Cetaceans” By Walter Zimmer. Cambridge University Press, New York, 2011. 366 pages. $105.00 (hardcover). Transactions of the American Fisheries Society, 142(2):578–579, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clifford:2012:RLS

[CHB^{+}12] Alexander M. Clifford, Matthew Henry, Roger Bergstedt, D. Gordon McDonald, Adam S. Smits, and Michael P. Wilkie. Recovery of larval sea lampreys from short-term exposure to the pesticide 3-trifluoromethyl-4-nitrophenol: Implications for sea lamprey control in the Great Lakes. Transactions of the American Fisheries Society, 141(6):1697–1710, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cahill:2016:AGM

[CHH^{+}16] Christopher L. Cahill, Kimberly L. Howland, Mark F. Hulsman, Fred Noddin, William M. Tonn, Gregory Courtice, and David Z. Zhu. Arctic grayling movements through a naturelike fishpass in Northern Canada. Transactions of the American Fisheries Society, 145(5):951–963, September 2016. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chittenden:2013:RAF

[Chi13] Cedar Chittenden. A review of “Advances in Fish Tagging and Marking Technology” Edited by Jeremy R. McKenzie, Bradford Parsons, Andrew C. Seitz, R. Keller Kopf, Matthew Mesa, and Quinton Phelps. American Fisheries Society, Symposium 76, Bethesda, Maryland. 560 pages. $79.00 ($55.30 for American Fisheries Society members). Transactions of the American Fisheries Society, 142(3):884, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colborne:2015:PLI

[CHLN15] Scott F. Colborne, Timothy J. A. Hain, Fred J. Longstaffe, and Bryan D. Neff. The potential for less invasive inference of resource use: Covariation in stable isotope composition between females and their eggs in bluegill. Transactions of the American Fisheries Society, 144(2):283–291, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carr-Harris:2018:PDS

[CHMG^{+}18] Charmaine N. Carr-Harris, Jonathan W. Moore, Allen S. Gottesfeld, Jennifer A. Gordon, William M. Shepert, James D. J. Henry, Jr., Harvey J. Russell, Wade N. B. Helin, David J. Doolan, and Terry D. Beacham. Phenological diversity of salmon smolt migration timing within a large watershed. Transactions of the American Fisheries Society, 147(5): 775–790, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cavileer:2015:SDG

[CHO^{+}15] Timothy D. Cavileer, Samuel S. Hunter, Jeffery Olsen, John Wenburg, and James J. Nagler. A sex-determining gene (sdY) assay shows discordance between phenotypic and genotypic sex in wild populations of Chinook salmon. Transactions of the American Fisheries Society, 144(2):423–430, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carassou:2012:CSS

[CHPG12] Laure Carassou, Frank J. Hernandez, Sean P. Powers, and William M. Graham. Cross-shore, seasonal, and depth-related structure of ichthyoplankton assemblages in coastal Alabama. Transactions of the American Fisheries Society, 141(4):1137– 1150, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Christensen:2014:RAE

[Chr14] Villy Christensen. A review of “Advancing an Ecosystem Approach in the Gulf of Maine” Edited by R. L. Stephenson, J. H. Annala, J. A. Runge, and M. Hall-Arber. American Fisheries Society, Symposium 79, Bethesda, Maryland. 2012. 415 pages. $79.00 ($55.30 for American Fisheries Society members) (paper). Transactions of the American Fisheries Society, 143(3): 832, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cary:2017:SUP

[CHR^{+}17] Joshua B. Cary, Jessica L. Holbrook, Morgan E. Reed, Thomas B. Austin, Matthew S. Steffensen, Seoghyun Kim, Kasey C. Pregler, and Yoichiro Kanno. Survival of Upper Piedmont stream fishes implanted with 8-mm passive integrated transponder tags. Transactions of the American Fisheries Society, 146(6):1223–1232, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Creelman:2011:TGG

[CHS^{+}11] Elisabeth K. Creelman, Lorenz Hauser, Ryan K. Simmons, William D. Templin, and Lisa W. Seeb. Temporal and geographic genetic divergence: Characterizing sockeye salmon populations in the chignik watershed, Alaska, using singlenucleotide polymorphisms. Transactions of the American Fisheries Society, 140(3):749–762, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chavarie:2013:SPL

[CHT13] Louise Chavarie, Kimberly L. Howland, and William M. Tonn. Sympatric polymorphism in lake trout: The coexistence of multiple shallow-water morphotypes in Great Bear Lake. Transactions of the American Fisheries Society, 142 (3):814–823, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:2013:IAT

[CHW^{+}13] Alison L. Collins, Scott G. Hinch, David W. Welch, Steven J. Cooke, and Timothy D. Clark. Intracoelomic acoustic tagging of juvenile sockeye salmon: Swimming performance, survival, and postsurgical wound healing in freshwater and during a transition to seawater. Transactions of the American Fisheries Society, 142(2):515–523, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cott:2013:SLH

[CJG13] Peter A. Cott, Tom A. Johnston, and John M. Gunn. Stability in life history characteristics among burbot populations across environmental gradients. Transactions of the American Fisheries Society, 142(6):1746–1756, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cott:2013:HSB

[CJGH13] Peter A. Cott, Tom A. Johnston, John M. Gunn, and Dennis M. Higgs. Hearing sensitivity of the burbot. Transactions of the American Fisheries Society, 142(6):1699–1704, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caudill:2014:FTE

[CJL^{+}14] Christopher C. Caudill, Michael A. Jepson, Steven R. Lee, Travis L. Dick, George P. Naughton, and Matthew L. Keefer. A field test of eugenol-based anesthesia versus fish restraint in migrating adult Chinook salmon and steelhead. Transactions of the American Fisheries Society, 143(4):856–863, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cai:2015:EEE

[CJM^{+}15] L. Cai, D. Johnson, P. Mandal, M. Gan, X. Yuan, Z. Tu, and Y. Huang. Effect of exhaustive exercise on the swimming capability and metabolism of juvenile Siberian sturgeon. Transactions of the American Fisheries Society, 144(3):532–538, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2012:EAL

[CKC^{+}12] Matthew R. Campbell, Christine C. Kozfkay, Timothy Copeland, William C. Schrader, Michael W. Ackerman, and Shawn R. Narum. Estimating abundance and life history characteristics of threatened wild Snake River steelhead stocks by using genetic stock identification. Transactions of the American Fisheries Society, 141(5):1310–1327, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chamberlin:2011:MYC

[CKFQ11] Joshua W. Chamberlin, Anna N. Kagley, Kurt L. Fresh, and Thomas P. Quinn. Movements of yearling Chinook salmon during the first summer in marine waters of Hood Canal, Washington. Transactions of the American Fisheries Society, 140(2):429–439, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2011:HIV

[CKM^{+}11] Matthew R. Campbell, Christine C. Kozfkay, Kevin A. Meyer, Madison S. Powell, and Richard N. Williams. Historical influences of volcanism and glaciation in shaping mitochondrial DNA variation and distribution in Yellowstone cutthroat trout across its native range. Transactions of the American Fisheries Society, 140(1):91–107, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carey:2019:EDS

[CKS^{+}19] Michael P. Carey, Kevin D. Keith, Merlyn Schelske, Charlie Lean, Stephen D. McCormick, Amy Regish, and Christian E. Zimmerman. Energy depletion and stress levels in sockeye salmon migrating at the northern edge of their distribution. Transactions of the American Fisheries Society, 148(4):785– 797, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:2016:EPI

[Cla16] Scott R. Clark. Effects of passive integrated transponder tags on the physiology and swimming performance of a smallbodied stream fish. Transactions of the American Fisheries Society, 145(6):1179–1192, November 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cassinelli:2010:CGS

[CM10] John D. Cassinelli and Christine M. Moffitt. Comparison of growth and stress in resident redband trout held in laboratory simulations of montane and desert summer temperature cycles. Transactions of the American Fisheries Society, 139 (2):339–352, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collingsworth:2011:IRB

[CM11a] Paris D. Collingsworth and Elizabeth A. Marschall. Identifying relationships between catches of spawning condition yellow perch and environmental variables in the Western Basin of Lake Erie. Transactions of the American Fisheries Society, 140(1):31–36, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Copeland:2011:ISS

[CM11b] Timothy Copeland and Kevin A. Meyer. Interspecies synchrony in salmonid densities associated with large-scale bioclimatic conditions in central Idaho. Transactions of the American Fisheries Society, 140(4):928–942, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chiaramonte:2019:CWP

[CML19] Luciano V. Chiaramonte, Kevin A. Meyer, and James A. Lamansky, Jr. Colonial waterbird predation and angler catch of hatchery rainbow trout stocked in Southern Idaho fisheries. Transactions of the American Fisheries Society, 148(6):1088– 1101, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carney:2017:IMG

[CMM^{+}17] Susan L. Carney, Doreen M. McVeigh, Jeanette B. Moss, M. Drew Ferrier, and John F. Morrissey. Insights on mitochondrial genetic variation in Chesapeake Bay summer-resident cownose rays. Transactions of the American Fisheries Society, 146(3):478–484, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2011:RGD

[CN11] Nathan R. Campbell and Shawn R. Narum. Relative genetic diversity estimates among four Pacific salmonids. Transactions of the American Fisheries Society, 140(3):822–828, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cordoleani:2018:MSW

[CNM^{+}18] Flora Cordoleani, Jeremy Notch, Alex S. McHuron, Arnold J. Ammann, and Cyril J. Michel. Movement and survival of wild Chinook salmon smolts from Butte Creek during their outmigration to the ocean: Comparison of a dry year versus a wet year. Transactions of the American Fisheries Society, 147 (1):171–184, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conrath:2017:MSO

[Con17] Christina Lyn Conrath. Maturity, spawning omission, and reproductive complexity of deepwater rockfish. Transactions of the American Fisheries Society, 146(3):495–507, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cerino:2013:BTI

[CORM13] David Cerino, Anthony S. Overton, James A. Rice, and James A. Morris, Jr. Bioenergetics and trophic impacts of the invasive Indo-Pacific lionfish. Transactions of the American Fisheries Society, 142(6):1522–1534, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conejeros:2014:DSA

[CPP^{+}14] Pablo Conejeros, Anna Phan, Michael Power, Michael O’Connell, Sergey Alekseyev, Iván Salinas, and Brian Dixon. Differentiation of sympatric Arctic char morphotypes using major histocompatibility class II genes. Transactions of the American Fisheries Society, 143(3):586–594, May 2014. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cote:2015:GFS

[CPS^{+}15] Caroline L. Côté, Scott A. Pavey, Joshua A. Stacey, Thomas C. Pratt, Martin Castonguay, Céline Audet, and Louis Bernatchez. Growth, female size, and sex ratio variability in American eel of different origins in both controlled conditions and the wild: Implications for stocking programs. Transactions of the American Fisheries Society, 144(2):246– 257, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:2010:AIM

[CR10] Andrew J. Carlson and Frank J. Rahel. Annual intrabasin movement and mortality of adult Bonneville cutthroat trout among complementary riverine habitats. Transactions of the American Fisheries Society, 139(5):1360–1371, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chiaramonte:2016:KRT

[CRC^{+}16] Luciano V. Chiaramonte, R. Adam Ray, R. Alex Corum, Toz Soto, Sascha L. Hallett, and Jerri L. Bartholomew. Klamath River thermal refuge provides juvenile salmon reduced exposure to the parasite Ceratonova shasta. Transactions of the American Fisheries Society, 145(4):810–820, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:2010:PCR

[CRH10] Nathan Cook, Frank J. Rahel, and Wayne A. Hubert. Persistence of Colorado River cutthroat trout populations in isolated headwater streams of Wyoming. Transactions of the American Fisheries Society, 139(5):1500–1510, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Catalano:2016:HIS

[CS16] Matthew J. Catalano and Maynard H. Schaus. Harvestinduced size structure shifts alter nutrient release by a population of omnivorous fish. Transactions of the American Fisheries Society, 145(3):637–648, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crossman:2014:SGL

[CSD^{+}14] James A. Crossman, Kim T. Scribner, Christin A. Davis, Patrick S. Forsythe, and Edward A. Baker. Survival and growth of lake sturgeon during early life stages as a function of rearing environment. Transactions of the American Fisheries Society, 143(1):104–116, January 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carter:2010:ETC

[CSDW10] Mark W. Carter, Daniel E. Shoup, John M. Dettmers, and David H. Wahl. Effects of turbidity and cover on prey selectivity of adult smallmouth bass. Transactions of the American Fisheries Society, 139(2):353–361, March 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caroffino:2010:PEL

[CSED10] David C. Caroffino, Trent M. Sutton, Robert F. Elliott, and Michael C. Donofrio. Predation on early life stages of lake sturgeon in the Peshtigo River, Wisconsin. Transactions of the American Fisheries Society, 139(6):1846–1856, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connor:2012:EPG

[CT12] William P. Connor and Kenneth F. Tiffan. Evidence for parr growth as a factor affecting parr-to-smolt survival. Transactions of the American Fisheries Society, 141(5):1207–1218, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chase:2016:ESA

[CT16] D. A. Chase and A. E. Todgham. Effect of species assemblage on juvenile growth and condition in three California estuarine fishes. Transactions of the American Fisheries Society, 145 (1):94–107, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conrath:2014:NRB

[CTG14] Christina L. Conrath, Cindy A. Tribuzio, and Kenneth J. Goldman. Notes on the reproductive biology of female salmon sharks in the Eastern North Pacific Ocean. Transactions of the American Fisheries Society, 143(2):363–368, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cram:2017:SVC

[CTK^{+}17] Jeremy M. Cram, Christian E. Torgersen, Ryan S. Klett, George R. Pess, Darran May, Todd N. Pearsons, and Andrew H. Dittman. Spatial variability of Chinook salmon spawning distribution and habitat preferences. Transactions of the American Fisheries Society, 146(2):206–221, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cade:2011:EGV

[CTN11] Brian S. Cade, James W. Terrell, and Ben C. Neely. Estimating geographic variation in allometric growth and body condition of blue suckers with quantile regression. Transactions of the American Fisheries Society, 140(6):1657–1669, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cushman:2012:GCA

[CTP^{+}12] Elizabeth Cushman, Carolyn Tarpey, Bill Post, Kent Ware, and Tanya Darden. Genetic characterization of American shad in the Edisto River, South Carolina, and initial evaluation of an experimental stocking program. Transactions of the American Fisheries Society, 141(5):1338–1348, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connor:2013:EDD

[CTPM13] William P. Connor, Kenneth F. Tiffan, John M. Plumb, and Christine M. Moffitt. Evidence for density-dependent changes in growth, downstream movement, and size of Chinook salmon subyearlings in a large-river landscape. Transactions of the American Fisheries Society, 142(5):1453–1468, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:2015:MTA

[CTS^{+}15] Sydney C. Clark, Theresa L. Tanner, Suresh A. Sethi, Kale T. Bentley, and Daniel E. Schindler. Migration timing of adult Chinook salmon into the Togiak River, Alaska, watershed: Is there evidence for stock structure? Transactions of the American Fisheries Society, 144(4):829–836, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Copeland:2014:IJM

[CVB14] Timothy Copeland, David A. Venditti, and Bruce R. Barnett. The importance of juvenile migration tactics to adult recruitment in stream-type Chinook salmon populations. Transactions of the American Fisheries Society, 143(6):1460–1475, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2019:DAS

[CVL^{+}19] Matthew R. Campbell, Ninh V. Vu, Amanda P. LaGrange, Ryan S. Hardy, Tyler J. Ross, and Shawn R. Narum. Development and application of Single-Nucleotide Polymorphism (SNP) genetic markers for conservation monitoring of burbot populations. Transactions of the American Fisheries Society, 148(3):661–670, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carey:2011:FMP

[CW11] Michael P. Carey and David H. Wahl. Foraging modes of predators and behaviors of prey determine the outcome of multiple predator interactions. Transactions of the American Fisheries Society, 140(4):1015–1022, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Courter:2019:NOS

[CWP^{+}19] Ian I. Courter, Garth J. Wyatt, Russell W. Perry, John M. Plumb, Forrest M. Carpenter, Nicklaus K. Ackerman, Robert B. Lessard, and Peter F. Galbreath. A natural-origin steelhead population’s response to exclusion of hatchery fish. Transactions of the American Fisheries Society, 148(2):339– 351, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coggins:2011:NFC

[CYP11] Lewis G. Coggins, Jr., Michael D. Yard, and William E. Pine III. Nonnative fish control in the Colorado River in Grand Canyon, Arizona: an effective program or serendipitous timing? Transactions of the American Fisheries Society, 140 (2):456–470, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chang:2019:DSR

[CYW^{+}19] Shui-Kai Chang, Tzu-Lun Yuan, Sheng-Ping Wang, Yi-Jay Chang, and Gerard DiNardo. Deriving a statistically reliable abundance index from landings data: an application to the Taiwanese coastal dolphinfish fishery with a multispecies feature. Transactions of the American Fisheries Society, 148(1): 106–122, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carey:2017:MTS

[CZK^{+}17] Michael P. Carey, Christian E. Zimmerman, Kevin D. Keith, Merlyn Schelske, Charlie Lean, and David C. Douglas. Migration trends of sockeye salmon at the northern edge of their distribution. Transactions of the American Fisheries Society, 146(4):791–802, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duston:2012:PDN

[DA12] J. Duston and T. Astatkie. Prey density and nonvisual feeding by larval striped bass. Transactions of the American Fisheries Society, 141(3):610–614, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunn:2016:DHS

[DA16] Corey G. Dunn and Paul L. Angermeier. Development of habitat suitability indices for the candy darter, with cross-scale validation across representative populations. Transactions of the American Fisheries Society, 145(6):1266–1281, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:2016:CDT

[DAA^{+}16] Michael R. Donaldson, Jon Amberg, Shivani Adhikari, Aaron Cupp, Nathan Jensen, Jason Romine, Adam Wright, Mark Gaikowski, and Cory D. Suski. Carbon dioxide as a tool to deter the movement of invasive bigheaded carps. Transactions of the American Fisheries Society, 145(3):657–670, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deters:2012:OST

[DBB^{+}12] Katherine A. Deters, Richard S. Brown, James W. Boyd, M. Brad Eppard, and Adam G. Seaburg. Optimal suturing technique and number of sutures for surgical implantation of acoustic transmitters in juvenile salmonids. Transactions of the American Fisheries Society, 141(1):1–10, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deters:2010:PAS

[DBC^{+}10] Katherine A. Deters, Richard S. Brown, Kathleen M. Carter, James W. Boyd, M. Brad Eppard, and Adam G. Seaburg. Performance assessment of suture type, water temperature, and surgeon skill in juvenile Chinook salmon surgically implanted with acoustic transmitters. Transactions of the American Fisheries Society, 139(3):888–899, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doctor:2014:GML

[DBHV14] Katy Doctor, Barry Berejikian, Jeffrey J. Hard, and Don Van-Doornik. Growth-mediated life history traits of steelhead reveal phenotypic divergence and plastic response to temperature. Transactions of the American Fisheries Society, 143 (2):317–333, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duffy:2010:ODS

[DBS^{+}10] Elisabeth J. Duffy, David A. Beauchamp, Ruston M. Sweeting, Richard J. Beamish, and James S. Brennan. Ontogenetic diet shifts of juvenile Chinook salmon in nearshore and offshore habitats of Puget Sound. Transactions of the American Fisheries Society, 139(3):803–823, May 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DellApa:2014:SSS

[DCBKR14] Andrea Dell’Apa, Jennifer Cudney-Burch, David G. Kimmel, and Roger A. Rulifson. Sexual segregation of spiny dogfish in fishery-dependent surveys in Cape Cod, Massachusetts: Potential management benefits. Transactions of the American Fisheries Society, 143(4):833–844, July 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeBruyne:2013:APS

[DCJ^{+}13] Robin L. DeBruyne, Jeremy T. H. Coleman, James R. Jackson, Lars G. Rudstam, and Anthony J. VanDeValk. Analysis of prey selection by double-crested cormorants: a 15-year diet study in Oneida Lake, New York. Transactions of the American Fisheries Society, 142(2):430–446, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doherty:2010:STM

[DCM10] Chad A. Doherty, R. Allen Curry, and Kelly R. Munkittrick. Spatial and temporal movements of white sucker: Implications for use as a sentinel species. Transactions of the American Fisheries Society, 139(6):1818–1827, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dadswell:2017:PCA

[DCS^{+}17] Michael J. Dadswell, Cornel Ceapa, Aaron D. Spares, Nathan D. Stewart, R. Allen Curry, Rodney G. Bradford, and Michael J. W. Stokesbury. Population characteristics of adult Atlantic sturgeon captured by the commercial fishery in the Saint John River Estuary, New Brunswick. Transactions of the American Fisheries Society, 146(2):318–330, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Darnell:2010:PSF

[DDMR10] M. Zachary Darnell, Kelly M. Darnell, Ruth E. McDowell, and Dan Rittschof. Postcapture survival and future reproductive potential of ovigerous blue crabs Callinectes sapidus caught in the central North Carolina pot fishery. Transactions of the American Fisheries Society, 139(6):1677–1687, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dick:2018:STP

[DEP^{+}18] M. Dick, E. J. Eliason, D. A. Patterson, K. A. Robinson, S. G. Hinch, and S. J. Cooke. Short-term physiological response profiles of tagged migrating adult sockeye salmon: a comparison of gastric insertion and external tagging methods. Transactions of the American Fisheries Society, 147(2):300– 315, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

David:2014:FGP

[DEW^{+}14] Aaron T. David, Christopher S. Ellings, Isa Woo, Charles A. Simenstad, John Y. Takekawa, Kelley L. Turner, Ashley L. Smith, and Jean E. Takekawa. Foraging and growth potential of juvenile Chinook salmon after tidal restoration of a large river delta. Transactions of the American Fisheries Society, 143(6):1515–1529, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Daum:2011:COC

[DF11] David W. Daum and Blair G. Flannery. Canadian-origin Chinook salmon rearing in nonnatal U.S. tributary streams of the Yukon River, Alaska. Transactions of the American Fisheries Society, 140(2):207–220, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dieterich:2012:HSM

[DF12] Jay W. Dieterich and Richard S. Fulford. Habitat suitability modeling to evaluate conservation and enhancement efforts for Gulf-strain striped bass in Mississippi coastal rivers. Transactions of the American Fisheries Society, 141(3):731–746, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Durst:2014:MGJ

[DF14] Scott L. Durst and Nathan R. Franssen. Movement and growth of juvenile Colorado pikeminnows in the San Juan River, Colorado, New Mexico, and Utah. Transactions of the American Fisheries Society, 143(2):519–527, March 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauwalter:2015:UAI

[DFB15] Daniel C. Dauwalter, Kurt A. Fesenmyer, and Robin Bjork. Using aerial imagery to characterize redband trout habitat in a remote desert landscape. Transactions of the American Fisheries Society, 144(6):1322–1339, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Damstra:2013:SHU

[DG13] Richard A. Damstra and Tracy L. Galarowicz. Summer habitat use by lake sturgeon in Manistee Lake, Michigan. Transactions of the American Fisheries Society, 142(4):931–941, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dutka-Gianelli:2014:FHS

[DG14] Jynessa Dutka-Gianelli. Feeding habits of the smallscale fat snook from East-Central Florida. Transactions of the American Fisheries Society, 143(5):1199–1203, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dutton:2019:FED

[DG19] Jessica Dutton and Kaitlyn A. Gioia. Fecundity and embryonic development of spiny dogfish in the Northwest Atlantic Ocean. Transactions of the American Fisheries Society, 148 (1):48–57, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dalley:2017:SHD

[DGMC17] Kate L. Dalley, Robert S. Gregory, Corey J. Morris, and David Cote. Seabed habitat determines fish and macroinvertebrate community associations in a subarctic marine coastal nursery. Transactions of the American Fisheries Society, 146(6):1115– 1125, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dieterman:2011:DJA

[DH11] Douglas J. Dieterman and R. John H. Hoxmeier. Demography of juvenile and adult brown trout in streams of Southeastern Minnesota. Transactions of the American Fisheries Society, 140(6):1642–1656, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Danhoff:2017:AHA

[DHA^{+}17] Brian M. Danhoff, Casey J. Huckins, Nancy A. Auer, Cameron W. Goble, Stephanie A. Ogren, and J. Marty Holtgren. Abiotic habitat assessment for Arctic grayling in a portion of the Big Manistee River, Michigan. Transactions of the American Fisheries Society, 146(4):645–662, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dumbauld:2015:AJS

[DHB15] Brett R. Dumbauld, Geoffrey R. Hosack, and Katelyn M. Bosley. Association of juvenile salmon and estuarine fish with intertidal seagrass and oyster aquaculture habitats in a Northeast Pacific estuary. Transactions of the American Fisheries Society, 144(6):1091–1110, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doctor:2010:STP

[DHRQ10] K. K. Doctor, R. Hilborn, M. Rowse, and T. Quinn. Spatial and temporal patterns of upriver migration by sockeye salmon populations in the Wood River System, Bristol Bay, Alaska. Transactions of the American Fisheries Society, 139(1):80–91, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

denHeyer:2013:FNM

[dHST13] Cornelia E. den Heyer, Carl James Schwarz, and M. Kurtis Trzcinski. Fishing and natural mortality rates of Atlantic halibut estimated from multiyear tagging and life history. Transactions of the American Fisheries Society, 142(3):690–702, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duguid:2017:MEM

[DJ17] William D. P. Duguid and Francis Juanes. Microtrolling: an economical method to nonlethally sample and tag juvenile Pacific salmon at sea. Transactions of the American Fisheries Society, 146(2):359–369, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Darnell:2018:LSM

[DK18] M. Zachary Darnell and Adam A. Kemberling. Large-scale movements of postcopulatory female blue crabs Callinectes sapidus in tidal and nontidal estuaries of North Carolina. Transactions of the American Fisheries Society, 147(4):716– 728, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

David:2015:ECV

[DKD^{+}15] Solomon R. David, Richard S. Kik IV, James S. Diana, Edward S. Rutherford, and Michael J. Wiley. Evidence of countergradient variation in growth of spotted gars from core and peripheral populations. Transactions of the American Fisheries Society, 144(4):837–850, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeCelles:2017:DVT

[DKL^{+}17] Gregory R. DeCelles, Emily F. Keiley, Travis M. Lowery, Nicholas M. Calabrese, and Kevin D. E. Stokesbury. Development of a video trawl survey system for New England groundfish. Transactions of the American Fisheries Society, 146(3):462–477, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

David:2018:ISS

[DKS18] Solomon R. David, Sarah M. King, and Jeffrey A. Stein. Introduction to a special section: Angling for dinosaurs — status and future study of the ecology, conservation, and management of ancient fishes. Transactions of the American Fisheries Society, 147(4):623–625, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doll:2013:BEA

[DL13] Jason C. Doll and Thomas E. Lauer. Bayesian estimation of age and length at 50% maturity. Transactions of the American Fisheries Society, 142(4):1012–1024, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dippold:2016:ELA

[DLHF16] David A. Dippold, Robert T. Leaf, J. Read Hendon, and James S. Franks. Estimation of the length-at-age relationship of Mississippi’s spotted seatrout. Transactions of the American Fisheries Society, 145(2):295–304, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dembkowski:2011:CFA

[DM11a] Daniel J. Dembkowski and L. E. Miranda. Comparison of fish assemblages in two disjoined segments of an oxbow lake in relation to connectivity. Transactions of the American Fisheries Society, 140(4):1060–1069, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dixon:2011:STL

[DM11b] Christopher J. Dixon and Matthew G. Mesa. Survival and tag loss in Moapa White River springfish implanted with passive integrated transponder tags. Transactions of the American Fisheries Society, 140(5):1375–1379, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duffy:2011:CMT

[DMCO11] William J. Duffy, Richard S. McBride, Steven X. Cadrin, and Kenneth Oliveira. Is Cating’s method of transverse groove counts to annuli applicable for all stocks of American shad? Transactions of the American Fisheries Society, 140(4):1023– 1034, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duston:2018:SBE

[DMHM^{+}18] J. Duston, J. Manríquez-Hernández, G. M. MacInnis, C. M. Reesor, and T. Astatkie. Striped bass early life history in the macrotidal Shubenacadie River. Transactions of the American Fisheries Society, 147(5):919–938, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duffy:2012:OAV

[DMHO12] William J. Duffy, Richard S. McBride, Michael L. Hendricks, and Kenneth Oliveira. Otolith age validation and growth estimation from oxytetracycline-marked and recaptured American shad. Transactions of the American Fisheries Society, 141(6):1664–1671, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dittman:2010:HSS

[DML^{+}10] Andrew H. Dittman, Darran May, Donald A. Larsen, Mary L. Moser, Mark Johnston, and David Fast. Homing and spawning site selection by supplemented hatchery- and naturalorigin Yakima River spring Chinook salmon. Transactions of the American Fisheries Society, 139(4):1014–1028, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunlop:2010:SSB

[DMR^{+}10] Erin S. Dunlop, Scott W. Milne, Mark S. Ridgway, Jeff Condiotty, and Ian Higginbottom. In situ swimming behavior of lake trout observed using integrated multibeam acoustics and biotelemetry. Transactions of the American Fisheries Society, 139(2):420–432, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dekar:2014:TTS

[DMRK14] Matthew P. Dekar, Cagney McCauley, Jesse W. Ray, and Ryan S. King. Thermal tolerance, survival, and recruitment of cyprinids exposed to competition and chronic heat stress in experimental streams. Transactions of the American Fisheries Society, 143(4):1028–1036, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:2019:WSH

[DMW^{+}19] Michael N. Davis, Thomas E. McMahon, Molly A. H. Webb, Jason E. Ilgen, Alan T. Hitch, Matthew E. Jaeger, and Kyle A. Cutting. Winter survival, habitat use, and hypoxia tolerance of Montana Arctic grayling in a winterkill-prone lake. Transactions of the American Fisheries Society, 148(4):843–856, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dammeyer:2013:SFM

[DPB13] Nathan T. Dammeyer, Catherine T. Phillips, and Timothy H. Bonner. Site fidelity and movement of Etheostoma fonticola with implications to endangered species management. Transactions of the American Fisheries Society, 142(4):1049–1057, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeHaan:2014:NSG

[DPS14] Patrick W. DeHaan, Carita E. Pascal, and James E. Seeb. Novel SNP genotyping assays facilitate species identification of Salvelinus collected in a recreational fishery. Transactions of the American Fisheries Society, 143(1):164–172, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dowling:2012:GVR

[DSAM12] Thomas E. Dowling, Melody J. Saltzgiver, Deborah Adams, and Paul C. Marsh. Genetic variability in a recruiting population of endangered razorback suckers from Lake Mead, Arizona–Nevada. Transactions of the American Fisheries Society, 141(4):990–999, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Downey:2018:HSR

[DSB^{+}18] Charles H. Downey, Matthew K. Streich, Rachel A. Brewton, Matthew J. Ajemian, Jennifer J. Wetz, and Gregory W. Stunz. Habitat-specific reproductive potential of red snapper: a comparison of artificial and natural reefs in the Western Gulf of Mexico. Transactions of the American Fisheries Society, 147 (6):1030–1041, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeAngelo:2014:FAS

[DSBS14] Jacquelyn A. De Angelo, Philip W. Stevens, David A. Blewett, and Theodore S. Switzer. Fish assemblages of shoal- and shoreline-associated seagrass beds in Eastern Gulf of Mexico estuaries. Transactions of the American Fisheries Society, 143 (4):1037–1048, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dann:2010:ODA

[DSHG10] Tyler H. Dann, William W. Smoker, Jeffrey J. Hard, and Anthony J. Gharrett. Outbreeding depression after two generations of hybridizing Southeast Alaska Coho salmon populations? Transactions of the American Fisheries Society, 139(5):

1292–1305, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Daugherty:2018:AGT

[DSM18] Daniel J. Daugherty, J. Warren Schlechte, and Dusty L. Mc-Donald. Alligator gars in Texas coastal bays: Long-term trends and environmental influences. Transactions of the American Fisheries Society, 147(4):653–664, July 2018. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeHaan:2012:AGV

[DSRA12] Patrick W. DeHaan, Paul D. Scheerer, Ron Rhew, and William R. Ardren. Analyses of genetic variation in populations of Oregon chub, a threatened floodplain minnow in a highly altered environment. Transactions of the American Fisheries Society, 141(2):533–549, March 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dudley:2019:SCI

[Dud19] Peter N. Dudley. S^{4}: a spatially continuous, individual-based model of salmonid redd superimposition. Transactions of the American Fisheries Society, 148(2):352–372, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeBoom:2013:ECW

[DW13] Corey S. DeBoom and David H. Wahl. Effects of coarse woody habitat complexity on predator–prey interactions of four freshwater fish species. Transactions of the American Fisheries Society, 142(6):1602–1614, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeWeber:2015:PBT

[DW15] Jefferson T. DeWeber and Tyler Wagner. Predicting brook trout occurrence in stream reaches throughout their native range in the Eastern United States. Transactions of the American Fisheries Society, 144(1):11–24, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:2016:SDS

[DW16] Lori A. Davis and Tyler Wagner. Scale-dependent seasonal pool habitat use by sympatric wild brook trout and brown trout populations. Transactions of the American Fisheries Society, 145(4):888–902, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:2018:IDA

[DWE^{+}18] Melanie J. Davis, Isa Woo, Christopher S. Ellings, Sayre Hodgson, David A. Beauchamp, Glynnis Nakai, and Susan E. W. De La Cruz. Integrated diet analyses reveal contrasting trophic niches for wild and hatchery juvenile Chinook salmon in a large river delta. Transactions of the American Fisheries Society, 147(5):818–841, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:2019:FTF

[DWE^{+}19] Melanie J. Davis, Isa Woo, Christopher S. Ellings, Sayre Hodgson, David A. Beauchamp, Glynnis Nakai, and Susan E. W. De La Cruz. Freshwater tidal forests and estuarine wetlands may confer early life growth advantages for delta-reared Chinook salmon. Transactions of the American Fisheries Society, 148(2):289–307, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauwalter:2014:RCH

[DWG14] Daniel C. Dauwalter, Seth J. Wenger, and Peter Gardner. The role of complexity in habitat use and selection by stream fishes in a Snake River Basin tributary. Transactions of the American Fisheries Society, 143(5):1177–1187, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauwalter:2011:LUA

[DWGF11] Daniel C. Dauwalter, Seth J. Wenger, Kevin R. Gelwicks, and Kurt A. Fesenmyer. Land use associations with distributions of declining native fishes in the Upper Colorado River Basin. Transactions of the American Fisheries Society, 140(3):646– 658, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drinan:2015:EHB

[DWN^{+}15] Daniel P. Drinan, Molly A. H. Webb, Kerry A. Naish, Steven T. Kalinowski, Matthew C. Boyer, Amber C. Steed, Bradley B. Shepard, and Clint C. Muhlfeld. Effects of hybridization between nonnative rainbow trout and native westslope cutthroat trout on fitness-related traits. Transactions of the American Fisheries Society, 144(6):1275–1291, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dibble:2017:DBI

[DYWY17] Kimberly L. Dibble, Michael D. Yard, David L. Ward, and Charles B. Yackulic. Does bioelectrical impedance analysis accurately estimate the physiological condition of threatened and endangered desert fish species? Transactions of the American Fisheries Society, 146(5):888–902, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drinan:2012:ELA

[DZW^{+}12] Daniel P. Drinan, Alexander V. Zale, Molly A. H. Webb, Mark L. Taper, Bradley B. Shepard, and Steven T. Kalinowski. Evidence of local adaptation in westslope cutthroat trout. Transactions of the American Fisheries Society, 141 (4):872–880, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eiler:2016:TRP

[EB16] John H. Eiler and Mary A. Bishop. Tagging response and postspawning movements of Pacific herring, a small pelagic forage fish sensitive to handling. Transactions of the American Fisheries Society, 145(2):427–439, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eppehimer:2019:DDG

[EBA^{+}19] Drew E. Eppehimer, David B. Bunnell, Patricia M. Armenio, David M. Warner, Lauren A. Eaton, David J. Wells, and Edward S. Rutherford. Densities, diets, and growth rates of larval alewife and bloater in a changing Lake Michigan ecosystem. Transactions of the American Fisheries Society, 148(4):755– 770, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eichelberger:2014:NSN

[EBF^{+}14] Jennifer S. Eichelberger, Patrick J. Braaten, David B. Fuller, Matthew S. Krampe, and Edward J. Heist. Novel singlenucleotide polymorphism markers confirm successful spawning of endangered pallid sturgeon in the Upper Missouri River Basin. Transactions of the American Fisheries Society, 143 (6):1373–1385, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ehresmann:2018:MPJ

[EBG18] Rhea K. Ehresmann, Anne H. Beaudreau, and Kristen M. Green. Movement patterns of juvenile sablefish within a nursery area in Southeast Alaska. Transactions of the American Fisheries Society, 147(6):1052–1066, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ernst:2010:ENC

[EBMV10] Anne G. Ernst, Barry P. Baldigo, Christiane I. Mulvihill, and Mark Vian. Effects of natural-channel-design restoration on habitat quality in Catskill Mountain streams, New York. Transactions of the American Fisheries Society, 139(2):468– 482, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eckerlin:2011:TVP

[EFC^{+}11] Geofrey E. Eckerlin, John M. Farrell, Rufina N. Casey, Kristine M. Hope, Geoffrey H. Groocock, Paul R. Bowser, and James Casey. Temporal variation in prevalence of viral hemorrhagic septicemia virus type IVb among Upper St. Lawrence River smallmouth bass. Transactions of the American Fisheries Society, 140(3):529–536, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Enders:2012:EDA

[EGAW12] Eva C. Enders, Michael H. Gessel, James J. Anderson, and John G. Williams. Effects of decelerating and accelerating flows on juvenile salmonid behavior. Transactions of the American Fisheries Society, 141(2):357–364, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:2014:RBJ

[EHC^{+}14] A. F. Evans, N. J. Hostetter, K. Collis, D. D. Roby, and F. J. Loge. Relationship between juvenile fish condition and survival to adulthood in steelhead. Transactions of the American Fisheries Society, 143(4):899–909, July 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:2012:SEA

[EHR^{+}12] Allen F. Evans, Nathan J. Hostetter, Daniel D. Roby, Ken Collis, Donald E. Lyons, Benjamin P. Sandford, Richard D. Ledgerwood, and Scott Sebring. Systemwide evaluation of avian predation on juvenile salmonids from the Columbia River based on recoveries of passive integrated transponder tags. Transactions of the American Fisheries Society, 141 (4):975–989, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:2013:PIT

[ELC13] Theoren R. Ellis, Tommi Linnansaari, and Richard A. Cunjak. Passive Integrated Transponder (PIT) tracking versus snorkeling: Quantification of fright bias and comparison of techniques in habitat use studies. Transactions of the American Fisheries Society, 142(3):660–670, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ely:2012:E

[Ely12] Patrick C. Ely. Erratum. Transactions of the American Fisheries Society, 141(6):1711, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eiler:2014:DSC

[EMS^{+}14] John H. Eiler, Michele M. Masuda, Ted R. Spencer, Richard J. Driscoll, and Carl B. Schreck. Distribution, stock composition and timing, and tagging response of wild Chinook salmon returning to a large, free-flowing river basin. Transactions of the American Fisheries Society, 143(6):1476–1507, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:2019:CEA

[EPC^{+}19] Allen F. Evans, Quinn Payton, Bradley M. Cramer, Ken Collis, Nathan J. Hostetter, Daniel D. Roby, and Curtis Dotson. Cumulative effects of avian predation on Upper Columbia River steelhead. Transactions of the American Fisheries Society, 148(5):896–913, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Epstein:2018:SRS

[EPR^{+}18] Joshua M. Epstein, William E. Pine III, Christina M. Romagosa, Mark C. Scott, Catherine T. Phillips, Cathy A. Marion, and Benjamin Baiser. State- and regional-scale patterns and drivers of freshwater fish functional diversity in the Southeastern USA. Transactions of the American Fisheries Society, 147(6):1179–1198, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eby:2015:MPW

[EPS^{+}15] L. A. Eby, R. Pierce, M. Sparks, K. Carim, and C. Podner. Multiscale prediction of whirling disease risk in the Blackfoot River Basin, Montana: a useful consideration for restoration prioritization? Transactions of the American Fisheries Society, 144(4):753–766, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:2016:APJ

[EPT^{+}16] Allen F. Evans, Quinn Payton, Aaron Turecek, Bradley Cramer, Ken Collis, Daniel D. Roby, Peter J. Loschl, Leah Sullivan, John Skalski, Mark Weiland, and Curtis Dotson. Avian predation on juvenile salmonids: Spatial and temporal analysis based on acoustic and passive integrated transponder tags. Transactions of the American Fisheries Society, 145(4):860– 877, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Einfalt:2015:EPS

[EPW15] Lisa M. Einfalt, Joseph J. Parkos III, and David H. Wahl. Effect of predator size and prey characteristics on piscivory of juvenile largemouth bass. Transactions of the American Fisheries Society, 144(4):682–692, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschenroeder:2018:WRH

[ER18] Jackman C. Eschenroeder and James H. Roberts. What role has hybridization played in the replacement of native Roanoke bass with invasive rock bass? Transactions of the American Fisheries Society, 147(3):497–513, May 2018. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:2015:CRE

[ERL15] Nathan T. Evans, Christopher W. Riley, and Gary A. Lamberti. Culvert replacement enhances connectivity of stream fish communities in a Michigan drainage network. Transactions of the American Fisheries Society, 144(5):967–976, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ehlo:2017:UMT

[ESD^{+}17] Chase A. Ehlo, Melody J. Saltzgiver, Thomas E. Dowling, Paul C. Marsh, and Brian R. Kesner. Use of molecular techniques to confirm nonnative fish predation on razorback sucker larvae in Lake Mohave, Arizona and Nevada. Transactions of the American Fisheries Society, 146(2):201–205, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eshenroder:2014:RCC

[Esh14] Randy L. Eshenroder. The role of the Champlain Canal and Erie Canal as putative corridors for colonization of Lake Champlain and Lake Ontario by sea lampreys. Transactions of the American Fisheries Society, 143(3):634–649, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erhardt:2018:JCS

[ETC18] John M. Erhardt, Kenneth F. Tiffan, and William P. Connor. Juvenile Chinook salmon mortality in a Snake River reservoir: Smallmouth bass predation revisited. Transactions of the American Fisheries Society, 147(2):316–328, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:2017:MGR

[Eva17] Thomas M. Evans. Measuring the growth rate in three populations of larval lampreys with mark–recapture techniques. Transactions of the American Fisheries Society, 146(1):147– 159, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eyler:2016:DPI

[EWSR16] Sheila M. Eyler, Stuart A. Welsh, David R. Smith, and Mary M. Rockey. Downstream passage and impact of turbine shutdowns on survival of silver American eels at five hydroelectric dams on the Shenandoah River. Transactions of the American Fisheries Society, 145(5):964–976, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Einfalt:2013:BGH

[EWW13] Lisa M. Einfalt, Douglas B. Wojcieszak, and David H. Wahl. Behavior, growth and habitat selection of hatchery esocids reared with artificial vegetation. Transactions of the American Fisheries Society, 142(2):345–352, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frimpong:2010:CUS

[FA10] Emmanuel A. Frimpong and Paul L. Angermeier. Comparative utility of selected frameworks for regionalizing fish-based bioassessments across the United States. Transactions of the American Fisheries Society, 139(6):1872–1895, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fitzsimons:2013:IDT

[FBB^{+}13] John D. Fitzsimons, Scott B. Brown, Lisa R. Brown, Guy Verreault, Rémi Tardif, Ken G. Drouillard, Scott A. Rush, and Jana R. Lantry. Impacts of diet on thiamine status of Lake Ontario American eels. Transactions of the American Fisheries Society, 142(5):1358–1369, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Falke:2010:SRH

[FBF10] Jeffrey A. Falke, Kevin R. Bestgen, and Kurt D. Fausch. Streamflow reductions and habitat drying affect growth, survival, and recruitment of brassy minnow across a Great Plains riverscape. Transactions of the American Fisheries Society, 139(5):1566–1583, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Friedl:2013:TBM

[FBH^{+}13] Sarah E. Friedl, Jeffrey A. Buckel, Joseph E. Hightower, Fred S. Scharf, and Kenneth H. Pollock. Telemetry-based mortality estimates of juvenile spot in two North Carolina estuarine creeks. Transactions of the American Fisheries Society, 142(2):399–415, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fetzer:2011:OMG

[FBJR11] William W. Fetzer, Thomas E. Brooking, James R. Jackson, and Lars G. Rudstam. Overwinter mortality of gizzard shad: Evaluation of starvation and cold temperature stress. Transactions of the American Fisheries Society, 140(6):1460–1471, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flowers:2011:OAS

[FBP11] H. Jared Flowers, Timothy F. Bonvechio, and Douglas L. Peterson. Observation of Atlantic sturgeon predation by a flathead catfish. Transactions of the American Fisheries Society, 140(2):250–252, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flitcroft:2014:RPA

[FBS^{+}14] R. Flitcroft, K. Burnett, J. Snyder, G. Reeves, and L. Ganio. Riverscape patterns among years of juvenile Coho salmon in midcoastal Oregon: Implications for conservation. Transactions of the American Fisheries Society, 143(1):26–38, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:2019:TFP

[FBWT19] Gregory S. Fraser, Kevin R. Bestgen, Dana L. Winkelman, and Kevin G. Thompson. Temperature — not flow — predicts native fish reproduction with implications for climate change. Transactions of the American Fisheries Society, 148(3):509– 527, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Faulkner:2019:AAF

[FBWZ19] James R. Faulkner, Blane L. Bellerud, Daniel L. Widener, and Richard W. Zabel. Associations among fish length, dam passage history, and survival to adulthood in two at-risk species of Pacific salmon. Transactions of the American Fisheries Society, 148(6):1069–1087, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fang:2017:SRA

[FCG^{+}17] Min Fang, Lu Cai, Yong Gao, Da He, David Johnson, and Yingping Huang. Swimming and recovery abilities of juvenile Chinese sturgeon Acipenser sinensis. Transactions of the American Fisheries Society, 146(6):1186–1192, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fleming:2018:ELC

[FDSB18] B. Paul Fleming, Daniel J. Daugherty, Nathan G. Smith, and Robert K. Betsill. Efficacy of low-cost, side-scan sonar for surveying alligator gars. Transactions of the American Fisheries Society, 147(4):696–703, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farmer:2013:USV

[FDWG13] Troy M. Farmer, Dennis R. DeVries, Russell A. Wright, and Joel E. Gagnon. Using seasonal variation in otolith microchemical composition to indicate largemouth bass and southern flounder residency patterns across an estuarine salinity gradient. Transactions of the American Fisheries Society, 142(5):1415–1429, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraley:2016:SMH

[FFYI16] Kevin M. Fraley, Jeffrey A. Falke, Richard Yanusz, and Sam Ivey. Seasonal movements and habitat use of potamodromous rainbow trout across a complex Alaska riverscape. Transactions of the American Fisheries Society, 145(5):1077–1092, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fritts:2016:SGC

[FGW^{+}16] Mark W. Fritts, Cheryl Grunwald, Isaac Wirgin, Tim L. King, and Douglas L. Peterson. Status and genetic character of Atlantic sturgeon in the Satilla River, Georgia. Transactions of the American Fisheries Society, 145(1):69–82, January 2016.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frimpong:2012:STE

[FH12] E. A. Frimpong and M. L. Henebry. Short-term effects of formalin and ethanol fixation and preservation techniques on weight and size of fish eggs. Transactions of the American Fisheries Society, 141(6):1472–1479, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Furey:2017:BTM

[FH17] N. B. Furey and S. G. Hinch. Bull trout movements match the life history of sockeye salmon: Consumers can exploit seasonally distinct resource pulses. Transactions of the American Fisheries Society, 146(3):450–461, May 2017. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Franklin:2012:ENL

[FHCSN12] Abigail E. Franklin, Alex Haro, Theodore Castro-Santos, and John Noreika. Evaluation of nature-like and technical fishways for the passage of alewives at two coastal streams in New England. Transactions of the American Fisheries Society, 141 (3):624–637, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fang:2012:ICR

[FJJ^{+}12] Xing Fang, Liping Jiang, Peter C. Jacobson, Heinz G. Stefan, Shoeb R. Alam, and Donald L. Pereira. Identifying cisco refuge lakes in Minnesota under future climate scenarios. Transactions of the American Fisheries Society, 141(6):1608–1621, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Favrot:2018:FWM

[FJP18] Scott D. Favrot, Brian C. Jonasson, and James T. Peterson. Fall and winter microhabitat use and suitability for spring Chinook salmon parr in a U.S. Pacific Northwest river. Transactions of the American Fisheries Society, 147(1):151–170, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Favrot:2018:BRE

[FK18] Scott D. Favrot and Thomas J. Kwak. Behavior and reproductive ecology of the sicklefin redhorse: an imperiled Southern Appalachian mountain fish. Transactions of the American Fisheries Society, 147(1):204–222, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischer:2019:GCT

[FKF^{+}19a] Jesse R. Fischer, Thomas J. Kwak, H. Jared Flowers, W. Gregory Cope, Jacob M. Rash, and Douglas A. Besler. Growth, condition, and trophic relations of stocked trout in Southern Appalachian mountain streams. Transactions of the American Fisheries Society, 148(4):771–784, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flowers:2019:BSS

[FKF^{+}19b] H. Jared Flowers, Thomas J. Kwak, Jesse R. Fischer, W. Gregory Cope, Jacob M. Rash, and Douglas A. Besler. Behavior and survival of stocked trout in Southern Appalachian mountain streams. Transactions of the American Fisheries Society, 148(1):3–20, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisk:2015:ERR

[FKH15] J. Michael Fisk II, Thomas J. Kwak, and Ryan J. Heise. Effects of regulated river flows on habitat suitability for the robust redhorse. Transactions of the American Fisheries Society, 144(4):792–806, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fleischer:2016:BRH

[Fle16] Guy W. Fleischer. Book review: Hakes: Biology and Exploitation. Edited by H. Arancibia: Wiley-Blackwell, Chichester, UK. 2015. 376 pages. $209.95 (hardcover), $167.99 (e-book). Transactions of the American Fisheries Society, 145(4):913, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flinders:2017:STC

[FM17] Jon M. Flinders and Daniel D. Magoulick. Spatial and temporal consumption dynamics of trout in catch-and-release areas in Arkansas tailwaters. Transactions of the American Fisheries Society, 146(3):432–449, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fujiwara:2011:ECP

[FMG^{+}11] Masami Fujiwara, Michael S. Mohr, Aaron Greenberg, J. Scott Foott, and Jerri L. Bartholomew. Effects of ceratomyxosis on population dynamics of Klamath fall-run Chinook salmon. Transactions of the American Fisheries Society, 140(5):1380– 1391, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ficke:2011:SJA

[FMJ11] A. D. Ficke, C. A. Myrick, and N. Jud. The swimming and jumping ability of three small Great Plains fishes: Implications for fishway design. Transactions of the American Fisheries Society, 140(6):1521–1531, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frost:2010:ESI

[FMS10] Deborah A. Frost, R. Lynn McComas, and Benjamin P. Sandford. The effects of a surgically implanted microacoustic tag on growth and survival in subyearling fall Chinook salmon. Transactions of the American Fisheries Society, 139(4):1192– 1197, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fox:2019:MMJ

[FP19] Adam G. Fox and Douglas L. Peterson. Movement and out-migration of juvenile Atlantic sturgeon in Georgia, USA. Transactions of the American Fisheries Society, 148(5):952– 962, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischer:2012:FCR

[FPD12] Jason Fischer, Craig Paukert, and Melinda Daniels. Fish community response to habitat alteration: Impacts of sand dredging in the Kansas River. Transactions of the American Fisheries Society, 141(6):1532–1544, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finger:2013:COP

[FPM13] Amanda J. Finger, Steve Parmenter, and Bernie P. May. Conservation of the owens pupfish: Genetic effects of multiple translocations and extirpations. Transactions of the American Fisheries Society, 142(5):1430–1443, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ford:2015:SSE

[FPM15] Michael Ford, Todd N. Pearsons, and Andrew Murdoch. The spawning success of early maturing resident hatchery Chinook salmon in a natural river system. Transactions of the American Fisheries Society, 144(3):539–548, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischer:2018:LSL

[FPR^{+}18] Jason L. Fischer, Jeremy J. Pritt, Edward F. Roseman, Carson G. Prichard, Jaquelyn M. Craig, Gregory W. Kennedy, and Bruce A. Manny. Lake sturgeon, lake whitefish, and walleye egg deposition patterns with response to fish spawning substrate restoration in the St. Clair–Detroit River system. Transactions of the American Fisheries Society, 147(1):79–93, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischer:2010:APL

[FQW^{+}10] Jesse R. Fischer, Michael C. Quist, Skyler L. Wigen, Andrew J. Schaefer, Timothy W. Stewart, and Thomas M. Isenhart. Assemblage and population-level responses of stream fish to riparian buffers at multiple spatial scales. Transactions of the American Fisheries Society, 139(1):185–200, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Feiner:2013:TNI

[FRA13] Zachary S. Feiner, James A. Rice, and D. Derek Aday. Trophic niche of invasive white perch and potential interactions with representative reservoir species. Transactions of the American Fisheries Society, 142(3):628–641, May 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Feiner:2013:TND

[FRBA13] Zachary S. Feiner, James A. Rice, Aaron J. Bunch, and D. Derek Aday. Trophic niche and diet overlap between invasive white perch and resident white bass in a southeastern reservoir. Transactions of the American Fisheries Society, 142 (4):912–919, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fresh:2019:ESE

[Fre19] Kurt L. Fresh. Ecology of salmonids in estuaries around the world: Adaptations, habitats, and conservation. Transactions of the American Fisheries Society, 148(3):683–684, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foley:2018:LPJ

[FRM18] Kevin M. Foley, Amanda Rosenberger, and Franz J. Mueter. Longitudinal patterns of juvenile Coho salmon distribution and densities in headwater streams of the Little Susitna River, Alaska. Transactions of the American Fisheries Society, 147 (2):247–264, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Funes-Rodriguez:2015:IHC

[FRRCGA^{+}15] René Funes-Rodríguez, José A. Ruíz-Chavarría, Rogelio González-Armas, Reginaldo Durazo, and Sergio A. Guzmán del Proó. Influence of hydrographic conditions on the distribution of spiny lobster larvae off the West Coast of Baja California. Transactions of the American Fisheries Society, 144(6):1192–1205, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frenette:2016:NHC

[FS16] Bryan D. Frenette and Richard Snow. Natural habitat conditions in a captive environment lead to spawning of spotted gar. Transactions of the American Fisheries Society, 145(4):835– 838, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forsythe:2013:EAM

[FSC^{+}13] P. S. Forsythe, K. T. Scribner, J. A. Crossman, A. Ragavendran, and E. A. Baker. Experimental assessment of the magnitude and sources of lake sturgeon egg mortality. Transactions of the American Fisheries Society, 142(4):1005–1011, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flannery:2017:OCC

[FSFS17] Joel Flannery, Andrew Stubblefield, Rocco Fiori, and Conor Shea. Observations of channel change from constructed wood jams on a forested gravel-bed stream. Transactions of the American Fisheries Society, 146(1):181–193, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Firman:2011:LMA

[FSJ^{+}11] Julie C. Firman, E. Ashley Steel, David W. Jensen, Kelly M. Burnett, Kelly Christiansen, Blake E. Feist, David P. Larsen, and Kara Anlauf. Landscape models of adult Coho salmon density examined at four spatial extents. Transactions of the American Fisheries Society, 140(2):440–455, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flannery:2013:MAP

[FSN^{+}13] Blair G. Flannery, Robert E. Spangler, Brenda L. Norcross, Cara J. Lewis, and John K. Wenburg. Microsatellite analysis of population structure in Alaska eulachon with application to mixed-stock analysis. Transactions of the American Fisheries Society, 142(4):1036–1048, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Feyrer:2017:PNA

[FSP^{+}17] Frederick Feyrer, Steven B. Slater, Donald E. Portz, Darren Odom, Tara Morgan-King, and Larry R. Brown. Pelagic nekton abundance and distribution in the northern sacramento– san joaquin delta, California. Transactions of the American Fisheries Society, 146(1):128–135, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fukushima:2011:RST

[FSRK11] Michio Fukushima, Hiroto Shimazaki, Peter S. Rand, and Masahide Kaeriyama. Reconstructing Sakhalin taimen Parahucho perryi historical distribution and identifying causes for local extinctions. Transactions of the American Fisheries Society, 140(1):1–13, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fenske:2010:DPA

[FSW10] Kari H. Fenske, David H. Secor, and Michael J. Wilberg. Demographics and parasitism of American eels in the Chesapeake Bay, USA. Transactions of the American Fisheries Society, 139(6):1699–1710, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farley:2018:PDI

[FVK^{+}18] Nicole J. Farley, Adrian A. Vasquez, Richard Kik IV, Solomon R. David, Arjun S. Katailiha, Xavier N. Walker, and Jeffrey L. Ram. Primer designs for identification and Environmental DNA (eDNA) detection of gars. Transactions of the American Fisheries Society, 147(4):687–695, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fetherman:2015:BTR

[FWB^{+}15] Eric R. Fetherman, Dana L. Winkelman, Larissa L. Bailey, George J. Schisler, and K. Davies. Brown trout removal effects on short-term survival and movement of Myxobolus cerebralis-resistant rainbow trout. Transactions of the American Fisheries Society, 144(3):610–626, May 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:2017:TUI

[FWBT17] Gregory S. Fraser, Dana L. Winkelman, Kevin R. Bestgen, and Kevin G. Thompson. Tributary use by imperiled flannelmouth and bluehead suckers in the Upper Colorado River Basin. Transactions of the American Fisheries Society, 146(5): 858–870, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farmer:2010:MCT

[FWD10] Troy M. Farmer, Russell A. Wright, and Dennis R. DeVries. Mercury concentration in two estuarine fish populations across a seasonal salinity gradient. Transactions of the American Fisheries Society, 139(6):1896–1912, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flaherty-Walia:2015:IAC

[FWSW^{+}15] Kerry E. Flaherty-Walia, Theodore S. Switzer, Brent L. Winner, Amanda J. Tyler-Jedlund, and Sean F. Keenan. Improved ability to characterize recruitment of gray snapper in three Florida estuaries along the Gulf of Mexico through targeted sampling of polyhaline seagrass beds. Transactions of the American Fisheries Society, 144(5):911–926, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisher:2014:EOD

[FWT^{+}14] J. P. Fisher, L. A. Weitkamp, D. J. Teel, S. A. Hinton, J. A. Orsi, E. V. Farley, Jr., J. F. T. Morris, M. E. Thiess, R. M. Sweeting, and M. Trudel. Early ocean dispersal patterns of Columbia River Chinook and Coho salmon. Transactions of the American Fisheries Society, 143(2):252–272, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fernandes:2010:SDM

[FZZ^{+}10] Stephen J. Fernandes, Gayle Barbin Zydlewski, Joseph D. Zydlewski, Gail S. Wippelhauser, and Michael T. Kinnison. Seasonal distribution and movements of shortnose sturgeon and Atlantic sturgeon in the Penobscot River Estuary, Maine. Transactions of the American Fisheries Society, 139(5):1436– 1449, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gwinn:2010:EPL

[GA10] Daniel C. Gwinn and Micheal S. Allen. Exploring populationlevel effects of fishery closures during spawning: an example using largemouth bass. Transactions of the American Fisheries Society, 139(2):626–634, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gosselin:2017:CMH

[GA17] J. L. Gosselin and J. J. Anderson. Combining migration history, river conditions, and fish condition to examine cross-life-stage effects on marine survival in Chinook salmon. Transactions of the American Fisheries Society, 146(3):408–421, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gurney:2014:OAS

[GBB^{+}14] W. S. C. Gurney, Louise O. Brennan, P. J. Bacon, K. F. Whelan, Martin O’Grady, Eileen Dillane, and P. McGinnity. Objectively assigning species and ages to salmonid length data from dual-frequency identification sonar. Transactions of the American Fisheries Society, 143(3):573–585, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Groce:2012:ESA

[GBF12] Matthew C. Groce, Larissa L. Bailey, and Kurt D. Fausch. Evaluating the success of Arkansas darter translocations in Colorado: an occupancy sampling approach. Transactions of the American Fisheries Society, 141(3):825–840, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

George:2016:ESD

[GBW16] Scott D. George, Barry P. Baldigo, and Scott M. Wells. Effects of seasonal drawdowns on fish assemblages in sections of an impounded river–canal system in upstate New York. Transactions of the American Fisheries Society, 145(6):1348–1357, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grote:2014:MDS

[GBZ14] Ann B. Grote, Michael M. Bailey, and Joseph D. Zydlewski. Movements and demography of spawning American shad in the Penobscot River, Maine, prior to dam removal. Transactions of the American Fisheries Society, 143(2):552–563, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glaspie:2018:EHD

[GCA^{+}18] Cassandra N. Glaspie, Melissa A. Clouse, Aaron T. Adamack, YoonKyung Cha, Stuart A. Ludsin, Doran M. Mason, Michael R. Roman, Craig A. Stow, and Stephen B. Brandt. Effect of hypoxia on diet of Atlantic bumpers in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 147(4):740–748, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gamble:2018:SGS

[GCG^{+}18] Madilyn M. Gamble, Kristin A. Connelly, Jennifer R. Gardner, Joshua W. Chamberlin, Kenneth I. Warheit, and David A. Beauchamp. Size, growth, and size-selective mortality of subyearling Chinook salmon during early marine residence in Puget Sound. Transactions of the American Fisheries Society, 147(2):370–389, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glass:2012:SSD

[GCM12] W. R. Glass, L. D. Corkum, and N. E. Mandrak. Spring and summer distribution and habitat use by adult threatened spotted gar in Rondeau Bay, Ontario, using radiotelemetry. Transactions of the American Fisheries Society, 141(4):1026– 1035, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ganias:2015:RRA

[GDG^{+}15] Konstantinos Ganias, Jeffrey N. Divino, Katie E. Gherard, Justin P. Davis, Foivos Mouchlianitis, and Eric T. Schultz. A reappraisal of reproduction in anadromous alewives: Determinate versus indeterminate fecundity, batch size, and batch number. Transactions of the American Fisheries Society, 144 (6):1143–1158, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Geist:2010:SGJ

[GDM^{+}10] David R. Geist, Zhiqun Deng, Robert P. Mueller, Steven R. Brink, and James A. Chandler. Survival and growth of juvenile Snake River fall Chinook salmon exposed to constant and fluctuating temperatures. Transactions of the American Fisheries Society, 139(1):92–107, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Geist:2011:EFT

[GDM^{+}11] David R. Geist, Zhiqun Deng, Robert P. Mueller, Valerie Cullinan, Steven Brink, and James A. Chandler. The effect of fluctuating temperatures and ration levels on the growth of juvenile Snake River fall Chinook salmon. Transactions of the American Fisheries Society, 140(1):190–200, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gao:2010:ICV

[GDPW10] Yongwen Gao, David L. Dettman, Kevin R. Piner, and Farron R. Wallace. Isotopic correlation (δ ^{18}O versus δ ^{13}C) of otoliths in identification of groundfish stocks. Transactions of the American Fisheries Society, 139(2):491–501, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gilbert:2016:LAF

[GDST16] Matthew J. H. Gilbert, Caitlyn R. Donadt, Heidi K. Swanson, and Keith B. Tierney. Low annual fidelity and early upstream migration of anadromous Arctic char in a variable environment. Transactions of the American Fisheries Society, 145 (5):931–942, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glover:2010:SPT

[GDWD10] David C. Glover, Dennis R. DeVries, Russell A. Wright, and Donald A. Davis. Sample preparation techniques for determination of fish energy density via bomb calorimetry: an evaluation using largemouth bass. Transactions of the American Fisheries Society, 139(3):671–675, May 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

George:2017:CST

[GGC17] Amy E. George, Tatiana Garcia, and Duane C. Chapman. Comparison of size, terminal fall velocity, and density of bighead carp, silver carp, and grass carp eggs for use in drift modeling. Transactions of the American Fisheries Society, 146(5): 834–843, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garza:2014:PSS

[GGHS^{+}14] John Carlos Garza, Elizabeth A. Gilbert-Horvath, Brian C. Spence, Thomas H. Williams, Heidi Fish, Stephen A. Gough, Joseph H. Anderson, David Hamm, and Eric C. Anderson. Population structure of steelhead in coastal California. Transactions of the American Fisheries Society, 143(1):134–152, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giresi:2015:IDM

[GGP^{+}15] Melissa M. Giresi, R. Dean Grubbs, David S. Portnoy, William B. Driggers III, Lisa Jones, and John R. Gold. Identification and distribution of morphologically conserved smoothhound sharks in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 144(6):1301–1310, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gahagan:2010:EEF

[GGS10] Benjamin I. Gahagan, Katie E. Gherard, and Eric T. Schultz. Environmental and endogenous factors influencing emigration in juvenile anadromous alewives. Transactions of the American Fisheries Society, 139(4):1069–1082, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gilbert-Horvath:2016:HPS

[GHPS^{+}16] Elizabeth A. Gilbert-Horvath, Kerrie A. Pipal, Brian C. Spence, Thomas H. Williams, and John Carlos Garza. Hierarchical phylogeographic structure of Coho salmon in California. Transactions of the American Fisheries Society, 145(5):1122– 1138, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gido:2014:REN

[Gid14] Keith Gido. A review of “Ecology of North American Freshwater Fishes.” By Stephen T. Ross. University of California Press, Oakland. 2013. $75 (hardcover). Transactions of the American Fisheries Society, 143(6):1613, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gross:2013:EPP

[GIW^{+}13] Jackson A. Gross, Kathryn M. Irvine, Siri Wilmoth, Tristany L. Wagner, Patrick A. Shields, and Jeffrey R. Fox. The effects of pulse pressure from seismic water gun technology on northern pike. Transactions of the American Fisheries Society, 142(5):1335–1346, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goertler:2018:EEH

[GJC^{+}18] Pascale Goertler, Kristopher Jones, Jeffery Cordell, Brian Schreier, and Ted Sommer. Effects of extreme hydrologic regimes on juvenile Chinook salmon prey resources and diet composition in a large river floodplain. Transactions of the American Fisheries Society, 147(2):287–299, March 2018. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garvin:2012:MIA

[GKBG12] Justin W. Garvin, Ozge Kureksiz, Philip J. Breczinski, and Mona K. Garvin. A method for an image-analysis-based twodimensional computational fluid dynamics simulation of moving fish. Transactions of the American Fisheries Society, 141 (1):185–198, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gorman:2019:VHU

[GKG^{+}19] A. M. Gorman, R. T. Kraus, L. F. G. Gutowsky, C. S. Vandergoot, Y. Zhao, C. T. Knight, M. D. Faust, T. A. Hayden, and C. C. Krueger. Vertical habitat use by adult walleyes conflicts with expectations from fishery-independent surveys. Transactions of the American Fisheries Society, 148(3):592–604, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galicia:2015:MCD

[GLM^{+}15] David Galicia, Pedro M. Leunda, Rafael Miranda, Javier Madoz, and Steve Parmenter. Morphometric contribution to the detection of introgressive hybridization in the endangered owens tui chub in California. Transactions of the American Fisheries Society, 144(2):431–442, March 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Guan:2019:EDE

[GMB^{+}19] Xin Guan, Emy M. Monroe, Katherine D. Bockrath, Erica L. Mize, Christopher B. Rees, Denise L. Lindsay, Kelly L. Baerwaldt, Leo G. Nico, and Richard F. Lance. Environmental DNA (eDNA) assays for invasive populations of black carp in North America. Transactions of the American Fisheries Society, 148(6):1043–1055, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garvin:2011:DSN

[GMP^{+}11] Michael R. Garvin, Robert W. Marcotte, Katie J. Palof, Rachel J. Riley, Lisa M. Kamin, and Anthony J. Gharrett. Diagnostic single-nucleotide polymorphisms identify Pacific Ocean perch and delineate blackspotted and rougheye rockfish. Transactions of the American Fisheries Society, 140(4):984– 988, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gervelis:2013:AGC

[GN13] Brian J. Gervelis and Lisa J. Natanson. Age and growth of the common thresher shark in the Western North Atlantic Ocean. Transactions of the American Fisheries Society, 142(6):1535– 1545, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grantham:2012:RSL

[GNMM12] Theodore E. Grantham, David A. Newburn, Michael A. Mc-Carthy, and Adina M. Merenlender. The role of streamflow and land use in limiting oversummer survival of juvenile steelhead in California streams. Transactions of the American Fisheries Society, 141(3):585–598, May 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gover:2014:HMC

[GNV^{+}14] Theodore R. Gover, Megan K. Nims, Jason J. Van Tassell, Paris D. Collingsworth, John W. Olesik, Stuart A. Ludsin, and Elizabeth A. Marschall. How much cleaning is needed when processing otoliths from fish larvae for microchemical analysis? Transactions of the American Fisheries Society, 143(3):779– 783, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Golet:2018:BEB

[Gol18] Walt Golet. Biology and ecology of bluefin tuna. Transactions of the American Fisheries Society, 147(5):1011–1012, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:2015:UMS

[Goo15] C. Phillip Goodyear. Understanding maximum size in the catch: Atlantic blue marlin as an example. Transactions of the American Fisheries Society, 144(2):274–282, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:2019:MGC

[Goo19] C. Phillip Goodyear. Modeling growth: Consequences from selecting samples by size. Transactions of the American Fisheries Society, 148(3):528–551, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gido:2012:LTD

[GP12] Keith B. Gido and David L. Propst. Long-term dynamics of native and nonnative fishes in the San Juan River, New Mexico and Utah, under a partially managed flow regime. Transactions of the American Fisheries Society, 141(3):645–659, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gaillard:2018:RSG

[GPBA18] Mélanie Gaillard, Scott A. Pavey, Louis Bernatchez, and Céline Audet. River-specific gene expression patterns associated with habitat selection for key hormone-coding genes in glass eel-stage American eels. Transactions of the American Fisheries Society, 147(5):855–868, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

George:2013:FEP

[GRDO13] Ellen M. George, Edward F. Roseman, Bruce M. Davis, and Timothy P. O’Brien. Feeding ecology of pelagic larval burbot in Northern Lake Huron, Michigan. Transactions of the American Fisheries Society, 142(6):1716–1723, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greenwood:2012:AEN

[Gre12] Marin F. D. Greenwood. Assessing effects of the nonindigenous pike killifish on indigenous fishes in Tampa Bay, Florida, using a weighted-evidence approach. Transactions of the American Fisheries Society, 141(1):84–99, January 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gharrett:2012:GAR

[GRS12] Anthony J. Gharrett, Rachel J. Riley, and Paul D. Spencer. Genetic analysis reveals restricted dispersal of northern rockfish along the continental margin of the Bering Sea and Aleutian Islands. Transactions of the American Fisheries Society, 141(2):370–382, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gorman:2013:UDB

[GS13] Owen T. Gorman and Shawn P. Sitar. Ups and downs of burbot and their predator lake trout in Lake Superior, 1953– 2011. Transactions of the American Fisheries Society, 142(6): 1757–1772, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodchild:2016:ETN

[GS16] Shawn C. Goodchild and Craig A. Stockwell. An experimental test of novel ecological communities of imperiled and invasive species. Transactions of the American Fisheries Society, 145 (2):264–268, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:2019:EGS

[GSA^{+}19] Cory J. Graham, Trent M. Sutton, Milo D. Adkison, Megan V. McPhee, and Philip J. Richards. Evaluation of growth, survival, and recruitment of Chinook salmon in Southeast Alaska Rivers. Transactions of the American Fisheries Society, 148 (2):243–259, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gregalis:2012:EPV

[GSD^{+}12] Kevan C. Gregalis, Lela S. Schlenker, J. Marcus Drymon, John F. Mareska, and Sean P. Powers. Evaluating the performance of vertical longlines to survey reef fish populations in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 141(6):1453–1464, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goetz:2017:RLT

[GSJS17] Frederick Goetz, Shawn Sitar, Andy Jasonowicz, and Michael Seider. Reproduction of lake trout morphotypes at Isle Royale in Northern Lake Superior. Transactions of the American Fisheries Society, 146(2):268–282, March 2017. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goetz:2011:RBS

[GSR^{+}11] Frederick Goetz, Shawn Sitar, Daniel Rosauer, Penny Swanson, Charles R. Bronte, Jon Dickey, and Crystal Simchick. The reproductive biology of Siscowet and lean lake trout in Southern Lake Superior. Transactions of the American Fisheries Society, 140(6):1472–1491, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grzybowski:2010:GVW

[GSVS^{+}10] Michael Grzybowski, Osvaldo J. Sepulveda-Villet, Carol A. Stepien, Daniel Rosauer, Frederick Binkowski, Rebecca Klaper, Brian Shepherd, and Frederick Goetz. Genetic variation of 17 wild yellow perch populations from the Midwest and East Coast analyzed via microsatellites. Transactions of the American Fisheries Society, 139(1):270–287, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garner:2012:UBI

[GvdHBC12] S. B. Garner, J. L. van der Ham, K. M. Boswell, and J. H. Cowan, Jr. Use of bioelectric impedance analysis to assess total-body condition and predict energy density in juvenile Atlantic croaker. Transactions of the American Fisheries Society, 141(4):1117–1125, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gaulke:2015:BPR

[GWB^{+}15] Greg L. Gaulke, John R. Wolfe, Douglas L. Bradley, Penelope E. Moskus, David H. Wahl, and Cory D. Suski. Behavioral and physiological responses of largemouth bass to rain-induced reductions in dissolved oxygen in an urban system. Transactions of the American Fisheries Society, 144(5): 927–941, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grasty:2019:TPR

[GWG^{+}19] Sarah E. Grasty, Carrie C. Wall, John Willis Gray, Jennifer Brizzolara, and Steven Murawski. Temporal persistence of red grouper holes and analysis of associated fish assemblages from towed camera data in the steamboat lumps marine protected area. Transactions of the American Fisheries Society, 148(3):652–660, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gaffney:2018:IME

[GWOD18] L. P. Gaffney, B. M. Watson, J. D. Oakes, and R. H. Devlin. Internal morphological effects of growth hormone transgenesis in Coho salmon. Transactions of the American Fisheries Society, 147(6):1199–1206, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gorsky:2012:CSH

[GZB12] Dimitry Gorsky, Joseph Zydlewski, and David Basley. Characterizing seasonal habitat use and diel vertical activity of lake whitefish in Clear Lake, Maine, as determined with acoustic telemetry. Transactions of the American Fisheries Society, 141(3):761–771, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hangsleben:2013:EEC

[HAG13] Matt A. Hangsleben, Micheal S. Allen, and Daniel C. Gwinn. Evaluation of electrofishing catch per unit effort for indexing fish abundance in Florida lakes. Transactions of the American Fisheries Society, 142(1):247–256, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2011:ODC

[Han11] John Mark Hanson. Ontogenetic diet comparison of Atlantic cod and white hake occurring at historically low population abundances. Transactions of the American Fisheries Society, 140(4):1070–1077, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2018:FIB

[Han18] John Mark Hanson. Feeding interactions between fishes in a coastal ecosystem in the Southern Gulf of St. Lawrence, Atlantic Canada. Transactions of the American Fisheries Society, 147(1):61–78, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:2017:BBB

[Har17] Kyle J. Hartman. Bioenergetics of brown bullhead in a changing climate. Transactions of the American Fisheries Society, 146(4):634–644, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henson:2018:TTS

[HAR18] Mary Nell Henson, D. Derek Aday, and James A. Rice. Thermal tolerance and survival of Nile tilapia and blue tilapia under rapid and natural temperature declination rates. Transactions of the American Fisheries Society, 147(2):278–286, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henson:2018:AIT

[HARL18] Mary Nell Henson, D. Derek Aday, James A. Rice, and Craig A. Layman. Assessing the influence of tilapia on sport fish species in North Carolina reservoirs. Transactions of the American Fisheries Society, 147(2):350–362, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hernandez:2019:QTR

[HAV^{+}19] Jesús M. Rivera Hernández, Noemi Peña Alvarado, Karlen Correa Vélez, Richard Nemeth, Richard Appeldoorn, and Virginia Shervette. Queen triggerfish reproductive biology in U.S. Caribbean waters. Transactions of the American Fisheries Society, 148(1):134–147, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2017:EEM

[HB17] Kyle C. Hanson and James M. Barron. Evaluation of the effects of marking Pacific lamprey ammocoetes with visual implant elastomer, coded wire tags, and passive integrated transponders. Transactions of the American Fisheries Society, 146(4):626–633, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heredia:2018:TEL

[HB18] Nicholas A. Heredia and Phaedra Budy. Trophic ecology of Lahontan cutthroat trout: Historical predator–prey interaction supports native apex predator in a unique desert lake.

Transactions of the American Fisheries Society, 147(5):842– 854, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holbrook:2015:FSP

[HBA^{+}15] Christopher M. Holbrook, Roger Bergstedt, Noah S. Adams, Tyson W. Hatton, and Robert L. McLaughlin. Fine-scale pathways used by adult sea lampreys during riverine spawning migrations. Transactions of the American Fisheries Society, 144(3):549–562, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:2013:ECP

[HBB13] Adam G. Hansen, David A. Beauchamp, and Casey M. Baldwin. Environmental constraints on piscivory: Insights from linking ultrasonic telemetry to a visual foraging model for cutthroat trout. Transactions of the American Fisheries Society, 142(1):300–316, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haskell:2017:TIC

[HBB17] Craig A. Haskell, David A. Beauchamp, and Stephen M. Bollens. Trophic interactions and consumption rates of subyearling Chinook salmon and nonnative juvenile American shad in Columbia River reservoirs. Transactions of the American Fisheries Society, 146(2):291–298, March 2017. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:2015:DHP

[HBC15] Karina C. Hall, Matt K. Broadhurst, and Brian R. Cullis. Deep hooking and postrelease mortality of two Australian native freshwater fishes angled from rivers by using natural baits. Transactions of the American Fisheries Society, 144(4):860– 871, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hovel:2015:DBM

[HBHS15] Rachel A. Hovel, David A. Beauchamp, Adam G. Hansen, and Mark H. Sorel. Development of a bioenergetics model for the threespine stickleback. Transactions of the American Fisheries Society, 144(6):1311–1321, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hessenauer:2012:GPR

[HBS12] Jan-Michael Hessenauer, Mary Tate Bremigan, and Kim T. Scribner. Genetic pedigree reconstruction facilitates lakewide estimates of age-0 largemouth bass dispersal. Transactions of the American Fisheries Society, 141(6):1672–1681, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:2013:VPD

[HBS13] Adam G. Hansen, David A. Beauchamp, and Erik R. Schoen. Visual prey detection responses of piscivorous trout and salmon: Effects of light, turbidity, and prey size. Transactions of the American Fisheries Society, 142(3):854–867, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heggenes:2011:EGD

[HBTT11] Jan Heggenes, Mark Beere, Patrick Tamkee, and Eric B. Taylor. Estimation of genetic diversity within and among populations of Oncorhynchus mykiss in a coastal river experiencing spatially variable hatchery augmentation. Transactions of the American Fisheries Society, 140(1):123–135, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holtrop:2010:ESE

[HCD10] Ann M. Holtrop, Yong Cao, and Chad R. Dolan. Estimating sampling effort required for characterizing species richness and site-to-site similarity in fish assemblage surveys of wadeable Illinois streams. Transactions of the American Fisheries Society, 139(5):1421–1435, September 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hertz:2014:RCE

[HCD^{+}14] Eric Hertz, Danielle C. Claar, Morgan M. Davies, Therese C. Frauendorf, Joel C. B. White, Julia A. Fisher, Abra F. Martin, and Brian Starzomski. A review of “Community Ecology” By Gary G. Mittelbach. Sinauer, Sunderland, Massachusetts. 2012. 400 pages. $64.56 (paper). Transactions of the American Fisheries Society, 143(3):830–831, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hudy:2010:DWS

[HCNL10] Mark Hudy, Jason A. Coombs, Keith H. Nislow, and Benjamin H. Letcher. Dispersal and within-stream spatial population structure of brook trout revealed by pedigree reconstruction analysis. Transactions of the American Fisheries Society, 139(5):1276–1287, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2014:ARC

[HCR14] John Mark Hanson, Michel Comeau, and Amélie Rondeau. Atlantic rock crab, unlike American lobster, is important to ecosystem functioning in Northumberland Strait. Transactions of the American Fisheries Society, 143(5):1266–1279, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howell:2016:LHD

[HCS^{+}16] Philip J. Howell, Michael E. Colvin, Paul M. Sankovich, David V. Buchanan, and Alan R. Hemmingsen. Life histories, demography, and distribution of a fluvial bull trout population. Transactions of the American Fisheries Society, 145(1): 173–194, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hernandez:2014:QAS

[HCW14] Kala Hernandez, Timothy Copeland, and Kristin Wright. Quantitative assessment of scale resorption in migrating and spawning steelhead of the Snake River Basin. Transactions of the American Fisheries Society, 143(6):1562–1568, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogg:2013:ASL

[HCZ13] Robert Hogg, Stephen M. Coghlan, Jr., and Joseph Zydlewski. Anadromous sea lampreys recolonize a Maine coastal river tributary after dam removal. Transactions of the American Fisheries Society, 142(5):1381–1394, September 2013. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogg:2015:FCR

[HCZG15] Robert S. Hogg, Stephen M. Coghlan, Jr., Joseph Zydlewski, and Cory Gardner. Fish community response to a smallstream dam removal in a Maine coastal river tributary. Transactions of the American Fisheries Society, 144(3):467–479, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holsman:2016:NIT

[HD16] Kirstin Holsman and Eric Danner. Numerical integration of temperature-dependent functions in bioenergetics models to avoid overestimation of fish growth. Transactions of the American Fisheries Society, 145(2):334–347, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haponski:2014:GHW

[HDBS14] Amanda E. Haponski, Hillary Dean, Bevin E. Blake, and Carol A. Stepien. Genetic history of walleyes spawning in Lake Erie’s Cattaraugus Creek: a comparison of pre- and poststocking. Transactions of the American Fisheries Society, 143(5): 1295–1307, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huenemann:2012:ITF

[HDF12] Thad W. Huenemann, Eric D. Dibble, and Jonathan P. Fleming. Influence of turbidity on the foraging of largemouth bass. Transactions of the American Fisheries Society, 141(1):107– 111, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Honsey:2016:ALM

[HDH16] Andrew E. Honsey, Steven B. Donabauer, and Tomas O. Höök. An analysis of lake morphometric and land-use characteristics that promote persistence of cisco in Indiana. Transactions of the American Fisheries Society, 145(2):363–373, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:2017:EEO

[HDQ^{+}17] Michelle A. Havey, Andrew H. Dittman, Thomas P. Quinn, Sean C. Lema, and Darran May. Experimental evidence for olfactory imprinting by sockeye salmon at embryonic and smolt stages. Transactions of the American Fisheries Society, 146 (1):74–83, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hostetter:2015:QAP

[HEC^{+}15] Nathan J. Hostetter, Allen F. Evans, Bradley M. Cramer, Ken Collis, Donald E. Lyons, and Daniel D. Roby. Quantifying avian predation on fish populations: Integrating predatorspecific deposition probabilities in tag recovery studies. Transactions of the American Fisheries Society, 144(2):410–422, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hostetter:2011:REF

[HER^{+}11] Nathan J. Hostetter, A. F. Evans, D. D. Roby, K. Collis, M. Hawbecker, B. P. Sandford, D. E. Thompson, and F. J. Loge. Relationship of external fish condition to pathogen prevalence and out-migration survival in juvenile steelhead. Transactions of the American Fisheries Society, 140(5):1158– 1171, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hostetter:2012:SJS

[HERC12] Nathan J. Hostetter, Allen F. Evans, Daniel D. Roby, and Ken Collis. Susceptibility of juvenile steelhead to avian predation: the influence of individual fish characteristics and river conditions. Transactions of the American Fisheries Society, 141(6): 1586–1599, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hitt:2012:DRI

[HEW12] Nathaniel P. Hitt, Sheila Eyler, and John E. B. Wofford. Dam removal increases American eel abundance in distant headwater streams. Transactions of the American Fisheries Society, 141(5):1171–1179, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:2016:RSS

[HGB^{+}16] Adam G. Hansen, Jennifer R. Gardner, David A. Beauchamp, Rebecca Paradis, and Thomas P. Quinn. Recovery of sockeye salmon in the Elwha River, Washington, after dam removal: Dependence of smolt production on the resumption of anadromy by landlocked Kokanee. Transactions of the American Fisheries Society, 145(6):1303–1317, November 2016. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hance:2016:BSV

[HGBE16] Dalton J. Hance, Lisa M. Ganio, Kelly M. Burnett, and Joseph L. Ebersole. Basin-scale variation in the spatial pattern of fall movement of juvenile Coho salmon in the West Fork Smith River, Oregon. Transactions of the American Fisheries Society, 145(5):1018–1034, September 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hintz:2016:SHU

[HGG^{+}16] William D. Hintz, David C. Glover, James E. Garvey, K. Jack Killgore, David P. Herzog, Timothy W. Spier, Robert E. Colombo, and Robert A. Hrabik. Status and habitat use of Scaphirhynchus sturgeons in an important fluvial corridor: Implications for river habitat enhancement. Transactions of the American Fisheries Society, 145(2):386–399, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huynh:2017:ETM

[HGP^{+}17] Quang C. Huynh, Todd Gedamke, Clay E. Porch, John M. Hoenig, John F. Walter, Meaghan Bryan, and Jon Brodziak. Estimating total mortality rates from mean lengths and catch rates in nonequilibrium situations. Transactions of the American Fisheries Society, 146(4):803–815, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:2016:CWU

[HGS^{+}16] John W. Hayes, Eric Goodwin, Karen A. Shearer, Joe Hay, and Lon Kelly. Can weighted useable area predict flow requirements of drift-feeding salmonids? Comparison with a net rate of energy intake model incorporating drift-flow processes. Transactions of the American Fisheries Society, 145(3):589– 609, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hintz:2017:SRL

[HGSG17] William D. Hintz, David C. Glover, Brittany C. Szynkowski, and James E. Garvey. Spatiotemporal reproduction and larval habitat associations of nonnative silver carp and bighead carp. Transactions of the American Fisheries Society, 146(3):422– 431, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hafs:2011:IET

[HH11] Andrew W. Hafs and Kyle J. Hartman. Influence of electrode type and location upon bioelectrical impedance analysis measurements of brook trout. Transactions of the American Fisheries Society, 140(5):1290–1297, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinrichsen:2013:RIT

[HHES13] Richard A. Hinrichsen, Daniel J. Hasselman, Curtis C. Ebbesmeyer, and Barbara A. Shields. The role of impoundments, temperature, and discharge on colonization of the Columbia River Basin, USA, by nonindigenous American shad. Transactions of the American Fisheries Society, 142 (4):887–900, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haus:2017:DSB

[HHJH17] William O. Haus, Kyle J. Hartman, John M. Jacobs, and Reginal M. Harrell. Development of striped bass relative condition models with bioelectrical impedance analysis and associated temperature corrections. Transactions of the American Fisheries Society, 146(5):917–926, September 2017. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harnish:2012:EMP

[HJM^{+}12] Ryan A. Harnish, Gary E. Johnson, Geoffrey A. McMichael, Michael S. Hughes, and Blaine D. Ebberts. Effect of migration pathway on travel time and survival of acoustic-tagged juvenile salmonids in the Columbia River estuary. Transactions of the American Fisheries Society, 141(2):507–519, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2016:EEE

[HJS^{+}16] Julianne E. Harris, Jeffrey C. Jolley, Gregory S. Silver, Henry Yuen, and Timothy A. Whitesel. An experimental evaluation of electrofishing catchability and catch depletion abundance estimates of larval lampreys in a wadeable stream: Use of a hierarchical approach. Transactions of the American Fisheries Society, 145(5):1006–1017, September 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hillyard:2012:TRC

[HK12] Ryan W. Hillyard and Ernest R. Keeley. Temperature-related changes in habitat quality and use by Bonneville cutthroat trout in regulated and unregulated river segments. Transactions of the American Fisheries Society, 141(6):1649–1663, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herrala:2014:HUS

[HKKS14] Jason R. Herrala, Patrick T. Kroboth, Nathan M. Kuntz, and Harold L. Schramm, Jr. Habitat use and selection by adult pallid sturgeon in the Lower Mississippi River. Transactions of the American Fisheries Society, 143(1):153–163, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horwitz:2018:AGR

[HKO^{+}18] Richard J. Horwitz, David H. Keller, Paul F. Overbeck, Steven P. Platania, Robert K. Dudley, and Evan W. Carson. Age and growth of the Rio Grande silvery minnow, an endangered, short-lived cyprinid of the North American Southwest. Transactions of the American Fisheries Society, 147(2):265– 277, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hill:2011:FTZ

[HKP11] Jeffrey E. Hill, Anne R. Kapuscinski, and Tyler Pavlowich. Fluorescent transgenic zebra danio more vulnerable to predators than wild-type fish. Transactions of the American Fisheries Society, 140(4):1001–1005, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanks:2018:CSP

[HKR18] R. Daniel Hanks, Yoichiro Kanno, and Jacob M. Rash. Can single-pass electrofishing replace three-pass depletion for population trend detection? Transactions of the American Fisheries Society, 147(4):729–739, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hager:2014:EAS

[HKW^{+}14] Christian Hager, Jason Kahn, Carter Watterson, Jay Russo, and Kyle Hartman. Evidence of Atlantic sturgeon spawning in the York River system. Transactions of the American Fisheries Society, 143(5):1217–1219, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holbrook:2011:SMA

[HKZ11] Christopher M. Holbrook, Michael T. Kinnison, and Joseph Zydlewski. Survival of migrating Atlantic salmon smolts through the Penobscot River, Maine: a prerestoration assessment. Transactions of the American Fisheries Society, 140(5): 1255–1268, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henningsen:2010:OCB

[HL10] Alan D. Henningsen and Robert T. Leaf. Observations on the captive biology of the southern stingray. Transactions of the American Fisheries Society, 139(3):783–791, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harstad:2014:VMR

[HLB14] Deborah L. Harstad, Donald A. Larsen, and Brian R. Beckman. Variation in minijack rate among hatchery populations of Columbia River Basin Chinook salmon. Transactions of the American Fisheries Society, 143(3):768–778, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hain:2016:CVS

[HLB^{+}16] Ernie F. Hain, Bradley A. Lamphere, Michael J. Blum, Peter B. McIntyre, Stacy A. C. Nelson, and James F. Gilliam. Comparison of visual survey and mark–recapture population estimates of a benthic fish in Hawaii. Transactions of the American Fisheries Society, 145(4):878–887, July 2016. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hill:2014:ART

[HLH14] Jeffrey E. Hill, Larry L. Lawson, Jr., and Scott Hardin. Assessment of the risks of transgenic fluorescent ornamental fishes to the United States using the Fish Invasiveness Screening Kit (FISK). Transactions of the American Fisheries Society, 143 (3):817–829, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horton:2011:MCR

[HLK11] Gregg E. Horton, Benjamin H. Letcher, and William L. Kendall. A multistate capture–recapture modeling strategy to separate true survival from permanent emigration for a passive integrated transponder tagged population of stream fish. Transactions of the American Fisheries Society, 140(2):320– 333, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2012:MVM

[HLR^{+}12] L. N. Harris, T. N. Loewen, J. D. Reist, N. M. Halden, J. A. Babaluk, and R. F. Tallman. Migratory variation in Mackenzie River system broad whitefish: Insights from otolith strontium distributions. Transactions of the American Fisheries Society, 141(6):1574–1585, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:2017:SHH

[HM17] Michael S. Hughes and Andrew R. Murdoch. Spawning habitat of hatchery spring Chinook salmon and possible mechanisms contributing to lower reproductive success. Transactions of the American Fisheries Society, 146(5):1016–1027, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herbst:2013:LWD

[HML13] Seth J. Herbst, J. Ellen Marsden, and Brian F. Lantry. Lake whitefish diet, condition, and energy density in Lake Champlain and the lower four Great Lakes following dreissenid invasions. Transactions of the American Fisheries Society, 142 (2):388–398, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hyle:2014:DVI

[HMO14] A. Reid Hyle, Richard S. McBride, and John E. Olney. Determinate versus indeterminate fecundity in American shad, an anadromous clupeid. Transactions of the American Fisheries Society, 143(3):618–633, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:2018:ETS

[HMP18] Brock M. Huntsman, Roy W. Martin, and Kirk A. Patten. Effects of temperature and spatial scale on Rio Grande cutthroat trout growth and abundance. Transactions of the American Fisheries Society, 147(3):480–496, May 2018. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horreo:2011:ASR

[HMSG^{+}11] J. L. Horreo, G. Machado-Schiaffino, A. M. Griffiths, D. Bright, J. R. Stevens, and E. Garcia-Vazquez. Atlantic salmon at risk: Apparent rapid declines in effective population size in Southern European populations. Transactions of the American Fisheries Society, 140(3):605–610, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haeseker:2012:AFM

[HMTC12] Steven L. Haeseker, Jerry A. McCann, Jack Tuomikoski, and Brandon Chockley. Assessing freshwater and marine environmental influences on life-stage-specific survival rates of Snake River spring–summer Chinook salmon and steelhead. Transactions of the American Fisheries Society, 141(1):121–138, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hickerson:2019:EMN

[HMW19] Brian T. Hickerson, Bryan M. Maitland, and Annika W. Walters. Effects of multiple nonnative fish on an imperiled cyprinid, hornyhead chub. Transactions of the American Fisheries Society, 148(6):1132–1145, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hess:2011:SNP

[HN11] Jon E. Hess and Shawn R. Narum. Single-Nucleotide Polymorphism (SNP) loci correlated with run timing in adult Chinook salmon from the Columbia River Basin. Transactions of the American Fisheries Society, 140(3):855–864, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:2013:SAS

[HN13] Bret C. Harvey and Rodney J. Nakamoto. Seasonal and among-stream variation in predator encounter rates for fish prey. Transactions of the American Fisheries Society, 142 (3):621–627, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:2012:AGS

[HNK^{+}12] Michael J. Hansen, Nancy A. Nate, Charles C. Krueger, Mara S. Zimmerman, Hanna G. Kruckman, and William W. Taylor. Age, growth, survival, and maturity of lake trout morphotypes in Lake Mistassini, Quebec. Transactions of the American Fisheries Society, 141(6):1492–1503, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2019:UDF

[HNRB19] J. Mason Harris, James A. Nelson, Guillaume Rieucau, and Whitney P. Broussard III. Use of drones in fishery science. Transactions of the American Fisheries Society, 148(4):687– 697, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herrmann:2018:MST

[HNV^{+}18] Scott J. Herrmann, Del Wayne R. Nimmo, Brian D. Vanden Heuvel, James S. Carsella, Christopher M. Kennedy, Kevin B. Rogers, John S. Wood, and Lynn M. Herrmann-Hoesing. Mercury and selenium in twelve cutthroat trout tissues from highelevation Colorado lakes, USA: Intraspecific and interspecific comparisons. Transactions of the American Fisheries Society, 147(3):444–458, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoenig:2017:SNM

[Hoe17] John M. Hoenig. Should natural mortality estimators based on maximum age also consider sample size? Transactions of the American Fisheries Society, 146(1):136–146, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2010:CVN

[HOGO10] K. C. Hanson, K. G. Ostrand, A. L. Gannam, and S. L. Ostrand. Comparison and validation of nonlethal techniques for estimating condition in juvenile salmonids. Transactions of the American Fisheries Society, 139(6):1733–1741, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houston:2010:PPT

[HOWS10] Derek D. Houston, T. Heath Ogden, Michael F. Whiting, and Dennis K. Shiozawa. Polyphyly of the pikeminnows (Teleostei: Cyprinidae) inferred using mitochondrial DNA sequences. Transactions of the American Fisheries Society, 139 (1):303–315, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hardy:2013:SKR

[HP13] Ryan Hardy and Vaughn L. Paragamian. A synthesis of Kootenai River burbot stock history and future management goals. Transactions of the American Fisheries Society, 142(6):1662– 1670, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:2014:TPT

[HP14] Kyle J. Hartman and Michael A. Porto. Thermal performance of three rainbow trout strains at above-optimal temperatures. Transactions of the American Fisheries Society, 143(6):1445– 1454, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hale:2016:AEH

[HPF^{+}16] Edward A. Hale, Ian A. Park, Matthew T. Fisher, Richard A. Wong, Michael J. Stangl, and John H. Clark. Abundance estimate for and habitat use by early juvenile Atlantic sturgeon within the Delaware River estuary. Transactions of the American Fisheries Society, 145(6):1193–1201, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hernandez:2010:DEI

[HPG10] Frank J. Hernandez, Jr., Sean P. Powers, and William M. Graham. Detailed examination of ichthyoplankton seasonality from a high-resolution time series in the Northern Gulf of Mexico during 2004–2006. Transactions of the American Fisheries Society, 139(5):1511–1525, September 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Humston:2010:DBT

[HPHB10] Robert Humston, Brant M. Priest, William C. Hamilton, and Paul E. Bugas, Jr. Dispersal between tributary and main-stem rivers by juvenile smallmouth bass evaluated using otolith microchemistry. Transactions of the American Fisheries Society, 139(1):171–184, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howell:2016:HUD

[HPSA16] Penelope T. Howell, Jose J. Pereira, Eric T. Schultz, and Peter J. Auster. Habitat use in a depleted population of winter flounder: Insights into impediments to population recovery. Transactions of the American Fisheries Society, 145(6):1208– 1222, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:2015:FEU

[HR15] Eriek S. Hansen and Frank J. Rahel. Fish energy use among fluctuating and constant thermal regimes simulating winter conditions in rivers. Transactions of the American Fisheries Society, 144(5):990–997, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:2016:LHD

[HRB^{+}16] Jason Hall, Philip Roni, Todd Bennett, John McMillan, Karrie Hanson, Raymond Moses, Mike McHenry, George Pess, and William Ehinger. Life history diversity of steelhead in two coastal Washington watersheds. Transactions of the American Fisheries Society, 145(5):990–1005, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hyun:2012:ATL

[HRG12] Saang-Yoon Hyun, Joel H. Reynolds, and Peter F. Galbreath. Accounting for tag loss and its uncertainty in a mark– recapture study with a mixture of single and double tags. Transactions of the American Fisheries Society, 141(1):11–25, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herbst:2016:WFE

[HRHS16] Seth J. Herbst, Brian M. Roth, Daniel B. Hayes, and Jason D. Stockwell. Walleye foraging ecology in an interconnected chain of lakes influenced by nonnative species. Transactions of the American Fisheries Society, 145(2):319–333, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:2011:DSA

[HRR^{+}11] Michael C. Hayes, Reginald R. Reisenbichler, Stephen P. Rubin, Lisa A. Wetzel, and Anne R. Marshall. Differential survival among sSOD-1* genotypes in Chinook salmon. Transactions of the American Fisheries Society, 140(5):1305–1316, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holecek:2013:CTL

[HS13] Dean E. Holecek and Dennis L. Scarnecchia. Comparison of two life history strategies after impoundment of a historically anadromous stock of Columbia River redband trout. Transactions of the American Fisheries Society, 142(5):1157–1166, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2017:MOW

[HS17] S. Dale Hanson and Craig P. Stafford. Modeling otolith weight using fish age and length: Applications to age determination. Transactions of the American Fisheries Society, 146(4):778– 790, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinrichsen:2016:MLE

[HSA^{+}16] Richard A. Hinrichsen, Craig A. Steele, Michael W. Ackerman, Matthew R. Campbell, Shawn R. Narum, Maureen A. Hess, William P. Young, Barbara A. Shields, and Brian L. Maschhoff. Maximum likelihood estimation of the proportion of hatchery-origin fish on spawning grounds using coded wire tagging and parentage-based tagging. Transactions of the American Fisheries Society, 145(3):671–686, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Homola:2012:GDE

[HSE^{+}12] Jared J. Homola, Kim T. Scribner, Robert F. Elliott, Michael C. Donofrio, Jeannette Kanefsky, Kregg M. Smith, and James N. McNair. Genetically derived estimates of contemporary natural straying rates and historical gene flow among Lake Michigan lake sturgeon populations. Transactions of the American Fisheries Society, 141(5):1374–1388, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinrichsen:2012:PAE

[HSF12] Richard A. Hinrichsen, Rishi Sharma, and Timothy R. Fisher. Precision and accuracy of estimators of the proportion of hatchery-origin spawners. Transactions of the American Fisheries Society, 141(2):437–454, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henery:2010:GMA

[HSG10] Rene E. Henery, Ted R. Sommer, and Charles R. Goldman. Growth and methylmercury accumulation in juvenile Chinook salmon in the Sacramento River and its floodplain, the Yolo Bypass. Transactions of the American Fisheries Society, 139 (2):550–563, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Habicht:2010:SFD

[HSM^{+}10] C. Habicht, L. W. Seeb, K. W. Myers, E. V. Farley, and J. E. Seeb. Summer–fall distribution of stocks of immature sockeye salmon in the Bering Sea as revealed by single-nucleotide polymorphisms. Transactions of the American Fisheries Society, 139(4):1171–1191, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holecek:2012:SIA

[HSM12] Dean E. Holecek, Dennis L. Scarnecchia, and Shannon E. Miller. Smoltification in an impounded, adfluvial redband trout population upstream from an impassable dam: Does it persist? Transactions of the American Fisheries Society, 141 (1):68–75, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hawkes:2017:AEM

[HSS17] James P. Hawkes, Timothy F. Sheehan, and Daniel S. Stich. Assessment of early migration dynamics of river-specific hatchery Atlantic salmon smolts. Transactions of the American Fisheries Society, 146(6):1279–1290, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall-Scharf:2016:OLT

[HSSS16] Brittany J. Hall-Scharf, Theodore S. Switzer, and Christopher D. Stallings. Ontogenetic and long-term diet shifts of a generalist juvenile predatory fish in an urban estuary undergoing dramatic changes in habitat availability. Transactions of the American Fisheries Society, 145(3):502–520, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hasegawa:2013:MEA

[HT13] Koh Hasegawa and Satoru Takahashi. Microscale environments along the seaward migration route of stocked chum salmon fry. Transactions of the American Fisheries Society, 142(5):1232–1237, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hill:2017:RSF

[HTH^{+}17] Jeffrey E. Hill, Quenton M. Tuckett, Scott Hardin, Larry L. Lawson, Jr., Katelyn M. Lawson, Jared L. Ritch, and Lauren Partridge. Risk screen of freshwater tropical ornamental fishes for the conterminous United States. Transactions of the American Fisheries Society, 146(5):927–938, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haskell:2013:EJA

[HTR13] Craig A. Haskell, Kenneth F. Tiffan, and Dennis W. Rondorf. The effects of juvenile American shad planktivory on zooplankton production in Columbia River food webs. Transactions of the American Fisheries Society, 142(3):606–620, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hooley-Underwood:2018:CGI

[HUMG^{+}18] Zachary Hooley-Underwood, Elizabeth G. Mandeville, Paul Gerrity, Joe Deromedi, Kevin Johnson, and Annika W. Walters. Combining genetic, isotopic, and field data to better describe the influence of dams and diversions on burbot movement in the Wind River Drainage, Wyoming. Transactions of the American Fisheries Society, 147(3):606–620, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hooley-Underwood:2019:ISS

[HUSST19] Zachary E. Hooley-Underwood, Summer B. Stevens, Nicholas R. Salinas, and Kevin G. Thompson. An intermittent stream supports extensive spawning of large-river native fishes. Transactions of the American Fisheries Society, 148(2):426–441, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hessenauer:2016:LNA

[HVD^{+}16] Jan-Michael Hessenauer, Jason Vokoun, Justin Davis, Robert Jacobs, and Eileen O’Donnell. Loss of naivety to angling at different rates in fished and unfished populations of largemouth bass. Transactions of the American Fisheries Society, 145(5): 1068–1076, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2014:ADD

[HW14] John Mark Hanson and Tamara Wilson. Abundance, distribution, and diet of a small-bodied ecotype of windowpane. Transactions of the American Fisheries Society, 143(3):650– 659, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hodge:2014:PFB

[HWD14] Brian W. Hodge, Margaret A. Wilzbach, and Walter G. Duffy. Potential fitness benefits of the half-pounder life history in Klamath River steelhead. Transactions of the American Fisheries Society, 143(4):864–875, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hodge:2016:LHD

[HWD^{+}16] Brian W. Hodge, Margaret A. Wilzbach, Walter G. Duffy, Rebecca M. Quiñones, and James A. Hobbs. Life history diversity in Klamath River steelhead. Transactions of the American Fisheries Society, 145(2):227–238, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heintz:2010:IMD

[HWH10] Ron A. Heintz, Mark S. Wipfli, and John P. Hudson. Identification of marine-derived lipids in juvenile Coho salmon and aquatic insects through fatty acid analysis. Transactions of the American Fisheries Society, 139(3):840–854, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Humphrey:2014:ETA

[HWHF14] Jennifer Humphrey, Michael J. Wilberg, Edward D. Houde, and Mary C. Fabrizio. Effects of temperature on age-0 Atlantic menhaden growth in Chesapeake Bay. Transactions of the American Fisheries Society, 143(5):1255–1265, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houde:2015:ECF

[HWN15] Aimee Lee S. Houde, Chris C. Wilson, and Bryan D. Neff. Effects of competition with four nonnative salmonid species on Atlantic salmon from three populations. Transactions of the American Fisheries Society, 144(5):1081–1090, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ivasauskas:2014:DMI

[IB14] Tomas J. Ivasauskas and Phillip W. Bettoli. Development of a multimetric index for fish assemblages in a cold tailwater in Tennessee. Transactions of the American Fisheries Society, 143(2):495–507, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ivasauskas:2012:ESM

[IBH12] Tomas J. Ivasauskas, Phillip W. Bettoli, and Thomas Holt. Effects of suture material and ultrasonic transmitter size on survival, growth, wound healing, and tag expulsion in rainbow trout. Transactions of the American Fisheries Society, 141(1): 100–106, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ivan:2011:LTI

[IHTF11] Lori N. Ivan, Tomas O. Höök, Michael V. Thomas, and David G. Fielder. Long-term and interannual dynamics of walleye and yellow perch in Saginaw Bay, Lake Huron. Transactions of the American Fisheries Society, 140(4):1078–1092, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isaak:2018:GWS

[ILH^{+}18] Daniel J. Isaak, Charles H. Luce, Dona L. Horan, Gwynne L. Chandler, Sherry P. Wollrab, and David E. Nagel. Global warming of salmon and trout rivers in the Northwestern U.S.: Road to ruin or path through purgatory? Transactions of the American Fisheries Society, 147(3):566–587, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Irvine:2014:IDO

[IMO^{+}14] J. R. Irvine, C. J. G. Michielsens, M. O’Brien, B. A. White, and M. Folkes. Increasing dominance of odd-year returning pink salmon. Transactions of the American Fisheries Society, 143(4):939–956, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ivan:2011:IAF

[IRJ11] Lori N. Ivan, Edward S. Rutherford, and Thomas H. Johengen. Impacts of adfluvial fish on the ecology of two Great Lakes tributaries. Transactions of the American Fisheries Society, 140(6):1670–1682, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Irwin:2013:EST

[IWB^{+}13] Brian J. Irwin, Tyler Wagner, James R. Bence, Megan V. Kepler, Weihai Liu, and Daniel B. Hayes. Estimating spatial and temporal components of variation for fisheries count data using negative binomial mixed models. Transactions of the American Fisheries Society, 142(1):171–183, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2013:POP

[JAA^{+}13] Lyndal Johnson, Bernadita Anulacion, Mary Arkoosh, O. Paul Olson, Catherine Sloan, Sean Y. Sol, Julann Spromberg, David J. Teel, Gladys Yanagida, and Gina Ylitalo. Persistent organic pollutants in juvenile Chinook salmon in the Columbia River Basin: Implications for stock recovery. Transactions of the American Fisheries Society, 142(1):21–40, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jackson:2018:FFS

[Jac18] James R. Jackson. Foundations of fisheries science. Transactions of the American Fisheries Society, 147(2):415–416, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:2014:EIW

[JADJ^{+}14] Kim K. Jones, Kara Anlauf-Dunn, Paul S. Jacobsen, Matt Strickland, Lora Tennant, and Sharon E. Tippery. Effectiveness of instream wood treatments to restore stream complexity and winter rearing habitat for juvenile Coho salmon. Transactions of the American Fisheries Society, 143(2):334– 345, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Janak:2012:ENB

[JBC^{+}12] Jill M. Janak, Richard S. Brown, Alison H. Colotelo, Brett D. Pflugrath, John R. Stephenson, Z. Daniel Deng, Thomas J. Carlson, and Adam G. Seaburg. The effects of neutrally buoyant, externally attached transmitters on swimming performance and predator avoidance of juvenile Chinook salmon. Transactions of the American Fisheries Society, 141(5):1424– 1432, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Juncos:2013:MPC

[JBV13] Romina Juncos, David A. Beauchamp, and Pablo H. Vigliano. Modeling prey consumption by native and nonnative piscivorous fishes: Implications for competition and impacts on shared prey in an ultraoligotrophic lake in Patagonia. Transactions of the American Fisheries Society, 142(1):268–281, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:2016:IDG

[JC16] Daniel A. James and Steven R. Chipps. Influence of Didymosphenia geminata blooms on prey composition and associated diet and growth of brown trout. Transactions of the American Fisheries Society, 145(1):195–205, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2012:MRT

[JCK^{+}12] Eric L. Johnson, Christopher C. Caudill, Matthew L. Keefer, Tami S. Clabough, Christopher A. Peery, Michael A. Jepson, and Mary L. Moser. Movement of radio-tagged adult Pacific lampreys during a large-scale fishway velocity experiment. Transactions of the American Fisheries Society, 141 (3):571–579, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jay:2014:EEN

[JCS14] Kathleen Jay, James A. Crossman, and Kim T. Scribner. Estimates of effective number of breeding adults and reproductive success for white sturgeon. Transactions of the American Fisheries Society, 143(5):1204–1216, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2015:EAA

[JDS15] Matthew W. Johnson, Sandra L. Diamond, and Gregory W. Stunz. External attachment of acoustic tags to deepwater reef fishes: an alternate approach when internal implantation affects experimental design. Transactions of the American Fisheries Society, 144(4):851–859, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jordan:2019:MIO

[JHK^{+}19] George R. Jordan, Edward J. Heist, Bernard R. Kuhajda, Gregory R. Moyer, Paul Hartfield, and Matthew S. Piteo. Morphological identification overestimates the number of pallid sturgeon in the Lower Mississippi River due to extensive introgressive hybridization. Transactions of the American Fisheries Society, 148(5):1004–1023, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2012:DLH

[JJC12] June Johnson, Tyler Johnson, and Timothy Copeland. Defining life histories of precocious male parr, minijack, and jack Chinook salmon using scale patterns. Transactions of the American Fisheries Society, 141(6):1545–1556, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2017:WIE

[JJNB17] Clark F. Johnson, Brett M. Johnson, Travis E. Neebling, and Jason C. Burckhardt. Walleye introduction eliminates predation refuge for adfluvial cutthroat trout and rainbow trout. Transactions of the American Fisheries Society, 146(2):252– 267, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jirka:2017:DNW

[JK17] Kurt J. Jirka and Clifford E. Kraft. Diet niche width and individual specialization of brook trout in Adirondack lakes. Transactions of the American Fisheries Society, 146(4):716– 731, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnston:2010:RNN

[JKP10] T. A. Johnston, M. Keir, and M. Power. Response of native and naturalized fish to salmonid cage culture farms in Northern Lake Huron, Canada. Transactions of the American Fisheries Society, 139(3):660–670, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jepsen:2017:PFA

[JLA17] Niels Jepsen, Martin H. Larsen, and Kim Aarestrup. Performance of fast absorbable sutures and histo-glue for closing incisions in brown trout. Transactions of the American Fisheries Society, 146(6):1233–1237, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jager:2010:EMV

[JLV^{+}10] Henriette I. Jager, Ken B. Lepla, Webb Van Winkle, Brad W. James, and Steven O. McAdam. The elusive minimum viable population size for white sturgeon. Transactions of the American Fisheries Society, 139(5):1551–1565, September 2010.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:2014:JRR

[JM14] Michelle K. Jones and Tim Mulligan. Juvenile rockfish recruitment in Trinidad Bay, California. Transactions of the American Fisheries Society, 143(2):543–551, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jacobs:2013:CSF

[JMB^{+}13] Gregory R. Jacobs, Charles P. Madenjian, David B. Bunnell, David M. Warner, and Randall M. Claramunt. Chinook salmon foraging patterns in a changing Lake Michigan. Transactions of the American Fisheries Society, 142(2):362– 372, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Juncos:2011:RRT

[JMM^{+}11] Romina Juncos, Daniela Milano, Patricio J. Macchi, Marcelo F. Alonso, and Pablo H. Vigliano. Response of rainbow trout to different food web structures in Northern Patagonia: Implications for growth, bioenergetics, and invasiveness. Transactions of the American Fisheries Society, 140(2):415–428, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jager:2013:PMA

[JPFB13] Henriette I. Jager, Douglas L. Peterson, Daniel Farrae, and Mark S. Bevelhimer. A population model to assess influences on the viability of the shortnose sturgeon population in the Ogeechee River, Georgia. Transactions of the American Fisheries Society, 142(3):731–746, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jenkins:2018:RPE

[JPG^{+}18] Laura E. Jenkins, Andrew L. Pierce, Neil Graham, Ryan Branstetter, Douglas R. Hatch, and James J. Nagler. Reproductive performance and energy balance in consecutive and skip repeat spawning female steelhead reconditioned in captivity. Transactions of the American Fisheries Society, 147 (5):959–971, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2017:EDD

[JPH17] Brett M. Johnson, William M. Pate, and Adam G. Hansen. Energy density and dry matter content in fish: New observations and an evaluation of some empirical models. Transactions of the American Fisheries Society, 146(6):1262–1278, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2010:AGM

[JPHK10] Matthew W. Johnson, Sean P. Powers, Crystal L. Hightower, and Matthew Kenworthy. Age, growth, mortality, and diet composition of vermilion snapper from the North-Central Gulf of Mexico. Transactions of the American Fisheries Society, 139(4):1136–1149, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2013:CBC

[JPLS13] Donald R. Johnson, Harriet M. Perry, and Joanne Lyczkowski-Shultz. Connections between Campeche Bank and red snapper populations in the Gulf of Mexico via modeled larval transport. Transactions of the American Fisheries Society, 142 (1):50–58, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James-Pirri:2010:MQS

[JPRS10] Mary-Jane James-Pirri, Charles T. Roman, and Jeffrey L. Swanson. A method to quantitatively sample nekton in salt-Marsh ditches and small tidal creeks. Transactions of the American Fisheries Society, 139(2):413–419, March 2010. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2012:UMJ

[JPSM12] Gary E. Johnson, Walter H. Pearson, Susan L. Southard, and Robert P. Mueller. Upstream movement of juvenile Coho salmon in relation to environmental conditions in a culvert test bed. Transactions of the American Fisheries Society, 141 (6):1520–1531, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Janetski:2013:RDA

[JRBC13] David J. Janetski, Carl R. Ruetz III, Yakuta Bhagat, and David F. Clapp. Recruitment dynamics of age-0 yellow perch in a drowned river mouth lake: Assessing synchrony with nearshore Lake Michigan. Transactions of the American Fisheries Society, 142(2):505–514, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2012:PEP

[JRP12] Christopher L. Johnson, Philip Roni, and George R. Pess. Parental effect as a primary factor limiting egg-to-fry survival of spring Chinook salmon in the Upper Yakima River Basin. Transactions of the American Fisheries Society, 141(5):1295– 1309, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jolley:2012:ODL

[JSW12] Jeffrey C. Jolley, Gregory S. Silver, and Timothy A. Whitesel. Occupancy and detection of larval Pacific lampreys and Lampetra spp. in a large river: the Lower Willamette River. Transactions of the American Fisheries Society, 141(2):305– 312, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:2015:CET

[JSW^{+}15] Marissa H. Jones, James E. Seeb, Kenneth I. Warheit, Todd R. Seamons, Thomas P. Quinn, and Lisa W. Seeb. Consequences of emergence timing for the growth and relative survival of steelhead fry from naturally spawning wild and hatchery parents. Transactions of the American Fisheries Society, 144(5): 977–989, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jude:2013:BEL

[JWHJ13] David J. Jude, Yu Wang, Stephen R. Hensler, and John Janssen. Burbot early life history strategies in the Great Lakes. Transactions of the American Fisheries Society, 142 (6):1733–1745, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:2016:SLI

[JZBP16] Alexander J. Jensen, Gayle B. Zydlewski, Sarah Barker, and Michael Pietrak. Sea lice infestations of a wild fish assemblage in the Northwest Atlantic Ocean. Transactions of the American Fisheries Society, 145(1):7–16, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnston:2019:RRR

[JZS^{+}19] Catherine Johnston, Gayle Barbin Zydlewski, Sean Smith, Joseph Zydlewski, and Michael T. Kinnison. River reach restored by dam removal offers suitable spawning habitat for endangered shortnose sturgeon. Transactions of the American Fisheries Society, 148(1):163–175, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kerns:2016:CMW

[KAH16] Janice A. Kerns, Micheal S. Allen, and Joseph E. Hightower. Components of mortality within a black bass high-release recreational fishery. Transactions of the American Fisheries Society, 145(3):578–588, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keefer:2018:CSC

[KBC^{+}18] Matthew L. Keefer, Timothy J. Blubaugh, Tami S. Clabough, Michael A. Jepson, George P. Naughton, and Christopher C. Caudill. Coho salmon colonization of Oregon’s Upper Willamette River Basin. Transactions of the American Fisheries Society, 147(6):1153–1166, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kennedy:2018:OPD

[KBG^{+}18] P. J. Kennedy, T. J. Bartley, D. M. Gillis, K. S. McCann, and M. D. Rennie. Offshore prey densities facilitate similar life history and behavioral patterns in two distinct aquatic apex predators, northern pike and lake trout. Transactions of the American Fisheries Society, 147(5):972–995, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koch:2012:HSM

[KBO^{+}12] Brian Koch, Ronald C. Brooks, Amanda Oliver, David Herzog, James E. Garvey, Robert Hrabik, Robert Colombo, Quinton Phelps, and Timothy Spier. Habitat selection and movement of naturally occurring pallid sturgeon in the Mississippi River. Transactions of the American Fisheries Society, 141(1):112– 120, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kirk:2015:CAP

[KCJ^{+}15] Mark A. Kirk, Christopher C. Caudill, Eric L. Johnson, Matthew L. Keefer, and Tami S. Clabough. Characterization of adult Pacific lamprey swimming behavior in relation to environmental conditions within large-dam fishways. Transactions of the American Fisheries Society, 144(5):998–1012, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keefer:2014:FBR

[KCM14] Matthew L. Keefer, Christopher C. Caudill, and Mary L. Moser. Fishway bottleneck relief models: a case study using radio-tagged Pacific lampreys. Transactions of the American Fisheries Society, 143(4):1049–1060, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kozfkay:2011:IHH

[KCMS11] Christine C. Kozfkay, Matthew R. Campbell, Kevin A. Meyer, and Daniel J. Schill. Influences of habitat and hybridization on the genetic structure of redband trout in the Upper Snake River Basin, Idaho. Transactions of the American Fisheries Society, 140(2):282–295, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katzenmeyer:2019:FGC

[KCS^{+}19] Eric D. Katzenmeyer, Michael E. Colvin, Timothy W. Stewart, Clay L. Pierce, and Scott E. Grummer. Fish growth changes over time in a Midwestern U.S. lake. Transactions of the American Fisheries Society, 148(2):493–506, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kiraly:2014:CTS

[KCZH14] Ian A. Kiraly, Stephen M. Coghlan, Jr., Joseph Zydlewski, and Daniel Hayes. Comparison of two sampling designs for fish assemblage assessment in a large river. Transactions of the American Fisheries Society, 143(2):508–518, March 2014.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:2018:RBP

[KDS18] Sarah M. King, Solomon R. David, and Jeffrey A. Stein. Relative bias and precision of age estimates among calcified structures of spotted gar, shortnose gar, and longnose gar. Transactions of the American Fisheries Society, 147(4):626–638, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaemingk:2012:FAS

[KGC^{+}12] Mark A. Kaemingk, Tracy L. Galarowicz, John A. Clevenger, David F. Clapp, and Herbert L. Lenon. Fish assemblage shifts and population dynamics of smallmouth bass in the Beaver Archipelago, Northern Lake Michigan: a comparison between historical and recent time periods amidst ecosystem changes. Transactions of the American Fisheries Society, 141(2):550– 559, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaemingk:2014:THD

[KGW14] Mark A. Kaemingk, Brian D. S. Graeb, and David W. Willis. Temperature, hatch date, and prey availability influence age-0 yellow perch growth and survival. Transactions of the American Fisheries Society, 143(4):845–855, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kammerer:2013:EST

[KH13] Brittany D. Kammerer and Scott A. Heppell. The effects of semichronic thermal stress on physiological indicators in steelhead. Transactions of the American Fisheries Society, 142(5): 1299–1307, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kneebone:2014:MPS

[KHD^{+}14] Jeff Kneebone, William S. Hoffman, Micah J. Dean, Dewayne A. Fox, and Michael P. Armstrong. Movement patterns and stock composition of adult striped bass tagged in Massachusetts coastal waters. Transactions of the American Fisheries Society, 143(5):1115–1129, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klinard:2018:ESI

[KHF^{+}18] Natalie V. Klinard, Edmund A. Halfyard, Aaron T. Fisk, Thomas J. Stewart, and Timothy B. Johnson. Effects of surgically implanted acoustic tags on body condition, growth, and survival in a small, laterally compressed forage fish. Transactions of the American Fisheries Society, 147(4):749–757, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kinziger:2013:CPS

[KHHG13] Andrew P. Kinziger, Michael Hellmair, David G. Hankin, and John Carlos Garza. Contemporary population structure in Klamath River Basin Chinook salmon revealed by analysis of microsatellite genetic data. Transactions of the American Fisheries Society, 142(5):1347–1357, September 2013. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kazyak:2015:SSM

[KHK^{+}15] David C. Kazyak, Robert H. Hilderbrand, Stephen R. Keller, Mark C. Colaw, Amanda E. Holloway, Raymond P. Morgan II, and Tim L. King. Spatial structure of morphological and neutral genetic variation in brook trout. Transactions of the American Fisheries Society, 144(3):480–490, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kahn:2014:ASA

[KHW^{+}14] J. E. Kahn, C. Hager, J. C. Watterson, J. Russo, K. Moore, and K. Hartman. Atlantic sturgeon annual spawning run estimate in the Pamunkey River, Virginia. Transactions of the American Fisheries Society, 143(6):1508–1514, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:2016:LAS

[KK16] Neala W. Kendall and Kathryn E. Kostow. Length- and ageselective fishing on Chinook salmon of the Columbia River Basin, 2002–2012. Transactions of the American Fisheries Society, 145(5):915–930, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koel:2010:APV

[KKB^{+}10] Todd M. Koel, Billie L. Kerans, Scott C. Barras, Katie C. Hanson, and John S. Wood. Avian piscivores as vectors for Myxobolus cerebralis in the greater Yellowstone ecosystem. Transactions of the American Fisheries Society, 139(4):976– 988, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Korman:2011:EFF

[KKM11] Josh Korman, Matthew Kaplinski, and Theodore S. Melis. Effects of fluctuating flows and a controlled flood on incubation success and early survival rates and growth of age-0 rainbow trout in a large regulated river. Transactions of the American Fisheries Society, 140(2):487–505, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kanno:2015:EFA

[KLR^{+}15] Yoichiro Kanno, Benjamin H. Letcher, Ana L. Rosner, Kyle P. O’Neil, and Keith H. Nislow. Environmental factors affecting brook trout occurrence in headwater stream segments. Transactions of the American Fisheries Society, 144(2):373– 382, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kellogg:2017:ITP

[KLS17] Karen A. Kellogg and Peter Leipzig-Scott. The influence of turbidity on prey consumption in the tessellated darter. Transactions of the American Fisheries Society, 146(3):508– 511, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keenan:2010:ILM

[KM10] Lawrence W. Keenan and Steven J. Miller. An inexpensive, low-maintenance, precision amplifier, dissolved oxygen sensor, and miniature data logger system designed for long-term deployment in riverine systems. Transactions of the American Fisheries Society, 139(2):593–597, March 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kallis:2014:HBS

[KM14] Jahn L. Kallis and Elizabeth A. Marschall. How body size and food availability influence first-winter growth and survival of a stocked piscivore. Transactions of the American Fisheries Society, 143(6):1434–1444, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaller:2013:BFE

[KMKS13] Michael D. Kaller, Catherine E. Murphy, William E. Kelso, and Mark R. Stead. Basins for fish and ecoregions for macroinvertebrates: Different spatial scales are needed to assess Louisiana wadeable streams. Transactions of the American Fisheries Society, 142(3):767–782, May 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Killgore:2011:FER

[KMM^{+}11] K. Jack Killgore, L. E. Miranda, Catherine E. Murphy, Douglas M. Wolff, Jan Jeffrey Hoover, Thomas M. Keevin, Steven T. Maynord, and Mark A. Cornish. Fish entrainment rates through towboat propellers in the Upper Mississippi and Illinois Rivers. Transactions of the American Fisheries Society, 140(3):570–581, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kennedy:2018:SRS

[KMS^{+}18] Patrick A. Kennedy, Kevin A. Meyer, Daniel J. Schill, Matthew R. Campbell, and Ninh V. Vu. Survival and reproductive success of hatchery YY male brook trout stocked in Idaho streams. Transactions of the American Fisheries Society, 147(3):419–430, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kuehne:2016:EDO

[KO16] Lauren M. Kuehne and Julian D. Olden. Environmental drivers of occupancy and detection of Olympic mudminnow. Transactions of the American Fisheries Society, 145(1):17–26, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kirsch:2014:MSA

[KP14] Joseph E. Kirsch and James T. Peterson. A multi-scaled approach to evaluating the fish assemblage structure within Southern Appalachian streams. Transactions of the American Fisheries Society, 143(5):1358–1371, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kozfkay:2019:PVS

[KPS^{+}19] Christine C. Kozfkay, Mike Peterson, Benjamin P. Sandford, Eric L. Johnson, and Paul Kline. The productivity and viability of Snake River sockeye salmon hatchery adults released into Redfish Lake, Idaho. Transactions of the American Fisheries Society, 148(2):308–323, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kohler:2012:NES

[KPZ^{+}12] Andre E. Kohler, Todd N. Pearsons, Joseph S. Zendt, Matthew G. Mesa, Christopher L. Johnson, and Patrick J. Connolly. Nutrient enrichment with salmon carcass analogs in the Columbia River Basin, USA: a stream food web analysis. Transactions of the American Fisheries Society, 141(3):802– 824, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:2011:LAT

[KQ11] Neala W. Kendall and Thomas P. Quinn. Length and age trends of Chinook salmon in the Nushagak River, Alaska, related to commercial and recreational fishery selection and exploitation. Transactions of the American Fisheries Society, 140(3):611–622, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kimmerer:2018:IBM

[KR18] Wim J. Kimmerer and Kenneth A. Rose. Individual-based modeling of delta smelt population dynamics in the Upper San Francisco Estuary III. Effects of entrainment mortality and changes in prey. Transactions of the American Fisheries Society, 147(1):223–243, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:2011:IPM

[KRM11] Damon M. Krueger, Edward S. Rutherford, and Doran M. Mason. Influence of predation mortality on survival of Chinook salmon parr in a Lake Michigan tributary. Transactions of the American Fisheries Society, 140(1):147–163, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:2013:MIP

[KRM13] Damon M. Krueger, Edward S. Rutherford, and Doran M. Mason. Modeling the influence of parr predation by walleyes and brown trout on the long-term population dynamics of Chinook salmon in Lake Michigan: a stage matrix approach. Transactions of the American Fisheries Society, 142(4):1101–1113, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kirk:2018:ETO

[KRRW18] Mark A. Kirk, Anna N. Rosswog, Kirsten N. Ressel, and Scott A. Wissinger. Evaluating the trade-offs between invasion and isolation for native brook trout and nonnative brown trout in Pennsylvania streams. Transactions of the American Fisheries Society, 147(5):806–817, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kocovsky:2011:IDM

[KS11] Patrick M. Kocovsky and Martin A. Stapanian. Influence of dreissenid mussels on catchability of benthic fishes in bottom trawls. Transactions of the American Fisheries Society, 140 (6):1565–1573, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klobucar:2016:LLC

[KSB16] Stephen L. Klobucar, W. Carl Saunders, and Phaedra Budy. A Lota lota consumption: Trophic dynamics of nonnative burbot in a valuable sport fishery. Transactions of the American Fisheries Society, 145(6):1386–1398, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kapuscinski:2013:GPS

[KSF13] Kevin L. Kapuscinski, Brian L. Sloss, and John M. Farrell. Genetic population structure of muskellunge in the Great Lakes. Transactions of the American Fisheries Society, 142(4):1075– 1089, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelly-Stormer:2017:GTR

[KSSK^{+}17] Amanda Kelly-Stormer, Virginia Shervette, Kevin Kolmos, David Wyanski, Tracey Smart, Chris McDonough, and Marcel J. M. Reichert. Gray triggerfish reproductive biology, age, and growth off the Atlantic Coast of the Southeastern USA. Transactions of the American Fisheries Society, 146(3):523– 538, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keefer:2012:URD

[KST^{+}12] Matthew L. Keefer, Robert J. Stansell, Sean C. Tackley, William T. Nagy, Karrie M. Gibbons, Christopher A. Peery, and Christopher C. Caudill. Use of radiotelemetry and direct observations to evaluate sea lion predation on adult Pacific salmonids at Bonneville Dam. Transactions of the American Fisheries Society, 141(5):1236–1251, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kellock:2014:SIL

[KTE^{+}14] Kristen A. Kellock, Brittany E. Trushel, Patrick C. Ely, Cecil A. Jennings, and Robert B. Bringolf. Survey of intersex largemouth bass from impoundments in Georgia USA. Transactions of the American Fisheries Society, 143(3):565–572, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koetsier:2012:DTW

[KU12] Peter Koetsier and Alexander N. Urquhart. Desiccation tolerance in a wild population of the invasive oriental weatherfish Misgurnus anguillicaudatus in Idaho, USA. Transactions of the American Fisheries Society, 141(2):365–369, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kock:2019:AOS

[KVA^{+}19] Tobias J. Kock, Nicholas E. Verretto, Nicklaus K. Ackerman, Russell W. Perry, John W. Beeman, Michael C. Garello, and Scott D. Fielding. Assessment of operational and structural factors influencing performance of fish collectors in forebays of high-head dams. Transactions of the American Fisheries Society, 148(2):464–479, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kappenman:2010:DUT

[KWCI10] Kevin M. Kappenman, Molly A. H. Webb, Elijah S. Cureton, and Jason Ilgen. Determination of upper temperature tolerance in June sucker larvae: Is the transition to Utah Lake temperatures a recruitment bottleneck? Transactions of the American Fisheries Society, 139(5):1386–1398, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kepler:2014:CBM

[KWS14] Megan V. Kepler, Tyler Wagner, and John A. Sweka. Comparative bioenergetics modeling of two lake trout morphotypes. Transactions of the American Fisheries Society, 143(6):1592– 1604, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kowalski:2014:FAC

[KWW14] Kurt P. Kowalski, Michael J. Wiley, and Douglas A. Wilcox. Fish assemblages, connectivity, and habitat rehabilitation in a diked Great Lakes coastal wetland complex. Transactions of the American Fisheries Society, 143(5):1130–1142, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Korman:2017:TRT

[KYK17] Josh Korman, Michael D. Yard, and Theodore A. Kennedy. Trends in rainbow trout recruitment, abundance, survival, and growth during a boom-and-bust cycle in a tailwater fishery. Transactions of the American Fisheries Society, 146(5):1043– 1057, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Landress:2016:FAA

[Lan16] Chad M. Landress. Fish assemblage associations with floodplain connectivity following restoration to benefit an endangered catostomid. Transactions of the American Fisheries Society, 145(1):83–93, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lincoln:2016:SMM

[LAR16] Kelsey J. Lincoln, D. Derek Aday, and James A. Rice. Seasonal mortality and movement patterns of white bass in a Southeastern U.S. reservoir. Transactions of the American Fisheries Society, 145(5):1035–1046, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lowery:2015:TOF

[LB15] Erin D. Lowery and David A. Beauchamp. Trophic ontogeny of fluvial bull trout and seasonal predation on Pacific salmon in a riverine food web. Transactions of the American Fisheries Society, 144(4):724–741, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Letvin:2017:ECC

[LBBW17] Alexander P. Letvin, Michael L. Brown, Katie N. Bertrand, and Michael J. Weber. Effects of common carp on trophic dynamics of sport fishes in shallow South Dakota water bodies. Transactions of the American Fisheries Society, 146(2):331– 340, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larsen:2010:ECB

[LBC10] Donald A. Larsen, Brian R. Beckman, and Kathleen A. Cooper. Examining the conflict between smolting and precocious male maturation in spring (stream-type) Chinook salmon. Transactions of the American Fisheries Society, 139 (2):564–578, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liss:2016:MTR

[LBD^{+}16] Stephanie A. Liss, Richard S. Brown, Katherine A. Deters, Ricardo W. Walker, Z. Daniel Deng, M. Brad Eppard, Richard L. Townsend, and Adam G. Seaburg. Mortality, transmitter retention, growth, and wound healing in juvenile salmon injected with micro acoustic transmitters. Transactions of the American Fisheries Society, 145(5):1047–1058, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lutz-Carrillo:2018:HGS

[LCHB^{+}18] Dijar J. Lutz-Carrillo, Martin Husemann, Preston T. Bean, Jánaye C. Williamson, Marcos J. De Jesús, and Jesse W. Ray. Hybridization and genetic structure in phenotypic spotted bass in Texas. Transactions of the American Fisheries Society, 147(5):891–905, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liu:2015:PSL

[LCO^{+}15] Chunlong Liu, Yifeng Chen, Julian D. Olden, Dekui He, Xiaoyun Sui, and Chengzhi Ding. Phenotypic shifts in life history traits influence invasion success of goldfish in the Yarlung Tsangpo River, Tibet. Transactions of the American Fisheries Society, 144(3):602–609, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lucey:2013:RSC

[LDA^{+}13] Sean M. Lucey, William J. Duffy, Erin Adams, Heather Marshall, Owen Nichols, and Katherine Thompson. A review of “Scientific Communication for Natural Resource Professionals” Edited by C. A. Jennings, T. E. Lauer, and B. Vondracek. American Fisheries Society, Bethesda, Maryland. 2012. 180 pages. $35.00 ($25 for American Fisheries Society members) (paper). Transactions of the American Fisheries Society, 142 (4):1154–1155, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larochelle:2015:VEC

[LDLH15] Mélissa Larochelle, Pierre Dumont, Claude Lavoie, and Daniel Hatin. Varying effects of common reed invasion on early life history of northern pike. Transactions of the American Fisheries Society, 144(1):196–210, January 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lieschke:2019:EEE

[LDP19] Jason A. Lieschke, Jan C. Dean, and Andrew Pickworth. Extending the effectiveness of electrofishing to estuarine habitats: Laboratory and field assessments. Transactions of the American Fisheries Society, 148(3):584–591, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lindley:2011:ETG

[LEM^{+}11] Steven T. Lindley, Daniel L. Erickson, Mary L. Moser, Greg Williams, Olaf P. Langness, Barry W. McCovey, Jr., Michael Belchik, Dave Vogel, William Pinnix, John T. Kelly, Joseph C. Heublein, and A. Peter Klimley. Electronic tagging of green sturgeon reveals population structure and movement among estuaries. Transactions of the American Fisheries Society, 140 (1):108–122, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Levings:2018:ABO

[Lev18] Colin D. Levings. All the boats on the ocean: How government subsidies led to global overfishing. Transactions of the American Fisheries Society, 147(2):414, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lefebvre:2015:RCL

[LF15] Lyndsey S. Lefebvre and John C. Field. Reproductive complexity in a long-lived deepwater fish, the blackgill rockfish. Transactions of the American Fisheries Society, 144(2):383– 399, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lew:2015:UNG

[LFB^{+}15] Ryan M. Lew, Amanda J. Finger, Melinda R. Baerwald, Alisha Goodbla, Bernie May, and Mariah H. Meek. Using nextgeneration sequencing to assist a conservation hatchery: a single-nucleotide polymorphism panel for the genetic management of endangered delta smelt. Transactions of the American Fisheries Society, 144(4):767–779, July 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lehman:2017:RBC

[LHHL17] Brendan Lehman, David D. Huff, Sean A. Hayes, and Steven T. Lindley. Relationships between Chinook salmon swimming performance and water quality in the San Joaquin River, California. Transactions of the American Fisheries Society, 146(2):349–358, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larsen:2013:ELH

[LHS^{+}13] Donald A. Larsen, Deborah L. Harstad, Charles R. Strom, Mark V. Johnston, Curtis M. Knudsen, David E. Fast, Todd N. Pearsons, and Brian R. Beckman. Early life history variation in hatchery- and natural-origin spring Chinook salmon in the Yakima River, Washington. Transactions of the American Fisheries Society, 142(2):540–555, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lister:2014:NPS

[Lis14] D. Brent Lister. Natural productivity in steelhead populations of natural and hatchery origin: Assessing hatchery spawner influence. Transactions of the American Fisheries Society, 143 (1):1–16, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lister:2015:NPS

[Lis15] D. Brent Lister. Natural productivity in steelhead populations of natural and hatchery origin — assessing hatchery spawner influence: Response to comment. Transactions of the American Fisheries Society, 144(6):1298–1300, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [CCH15].

Legler:2010:WTP

[LJHL10] Nicholas D. Legler, Timothy B. Johnson, Daniel D. Heath, and Stuart A. Ludsin. Water temperature and prey size effects on the rate of digestion of larval and early juvenile fish. Transactions of the American Fisheries Society, 139(3):868–875, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lerner:2013:SVD

[LKP^{+}13] Justin D. Lerner, David W. Kerstetter, Eric D. Prince, Liana Talaue-McManus, Eric S. Orbesen, Arthur Mariano, Derke Snodgrass, and Gary L. Thomas. Swordfish vertical distribution and habitat use in relation to diel and lunar cycles in the Western North Atlantic. Transactions of the American Fisheries Society, 142(1):95–104, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lacroix:2012:DUA

[LKS^{+}12] Gilles L. Lacroix, Derek Knox, Timothy F. Sheehan, Mark D. Renkawitz, and Meredith L. Bartron. Distribution of U.S. Atlantic salmon postsmolts in the Gulf of Maine. Transactions of the American Fisheries Society, 141(4):934–942, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Love:2019:SBC

[LLWW19] Seth A. Love, Nathan J. Lederman, Tristan Widloe, and Gregory W. Whitledge. Sources of bighead carp and silver carp found in Chicago urban fishing program ponds. Transactions of the American Fisheries Society, 148(2):417–425, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larsen:2017:UMP

[LMD^{+}17] Donald A. Larsen, Mollie A. Middleton, Jon T. Dickey, Ryan S. Gerstenberger, Chris V. Brun, and Penny Swanson. Use of morphological and physiological indices to characterize life history diversity in juvenile hatchery winter-run steelhead. Transactions of the American Fisheries Society, 146(4):663– 679, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ladago:2016:EFL

[LME16] Bret J. Ladago, J. Ellen Marsden, and Allison N. Evans. Early feeding by lake trout fry. Transactions of the American Fisheries Society, 145(1):1–6, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Landsman:2018:NVS

[LMPvdH18] Sean J. Landsman, Nic McLellan, Jonathan Platts, and Michael R. van den Heuvel. Nonsalmonid versus salmonid passage at nature-like and pool-and-weir fishways in Atlantic Canada, with special attention to rainbow smelt. Transactions of the American Fisheries Society, 147(1):94–110, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Love:2018:ENS

[LN18] Joseph W. Love and Joshua J. Newhard. Expansion of northern snakehead in the Chesapeake Bay watershed. Transactions of the American Fisheries Society, 147(2):342–349, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Love:2011:HSI

[Lov11] Joseph W. Love. Habitat suitability index for largemouth bass in tidal rivers of the Chesapeake Bay watershed. Transactions of the American Fisheries Society, 140(4):1049–1059, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liermann:2017:RRC

[LPM^{+}17] Martin Liermann, George Pess, Mike McHenry, John McMillan, Mel Elofson, Todd Bennett, and Raymond Moses. Relocation and recolonization of Coho salmon in two tributaries to the Elwha River: Implications for management and monitoring. Transactions of the American Fisheries Society, 146(5): 955–966, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laske:2012:DIT

[LRH12] Sarah M. Laske, Frank J. Rahel, and Wayne A. Hubert. Differential interactions of two introduced piscivorous salmonids with a native cyprinid in lentic systems: Implications for conservation of roundtail chub. Transactions of the American Fisheries Society, 141(2):495–506, March 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leduc:2010:APR

[LRM^{+}10] Antoine O. H. C. Leduc, Ellie Roh, Camille J. Macnaughton, Frederic Benz, Jordan Rosenfeld, and Grant E. Brown. Ambient pH and the response to chemical alarm cues in juvenile Atlantic salmon: Mechanisms of reduced behavioral responses. Transactions of the American Fisheries Society, 139(1):117– 128, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lyons:2015:TRP

[LRR^{+}15] John Lyons, Andrew L. Rypel, Paul W. Rasmussen, Thomas E. Burzynski, Bradley T. Eggold, Jared T. Myers, Tammie J. Paoli, and Peter B. McIntyre. Trends in the reproductive phenology of two Great Lakes fishes. Transactions of the American Fisheries Society, 144(6):1263–1274, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lokteff:2013:DBD

[LRW13] Ryan L. Lokteff, Brett B. Roper, and Joseph M. Wheaton. Do beaver dams impede the movement of trout? Transactions of the American Fisheries Society, 142(4):1114–1125, July 2013.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Long:2010:VOI

[LS10a] James M. Long and David R. Stewart. Verification of otolith identity used by fisheries scientists for aging channel catfish. Transactions of the American Fisheries Society, 139(6):1775– 1779, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lunn:2010:EBR

[LS10b] Brianne D. Lunn and Geoffrey B. Steinhart. Effect of brood reduction on nest abandonment of smallmouth bass. Transactions of the American Fisheries Society, 139(2):586–592, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Long:2016:ODO

[LS16] James M. Long and Richard A. Snow. Ontogenetic development of otoliths in alligator gar. Transactions of the American Fisheries Society, 145(3):537–544, May 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Long:2018:PDO

[LS18] James M. Long and Richard A. Snow. Posthatch development of otoliths and daily ring genesis in age-0 spotted gars. Transactions of the American Fisheries Society, 147(6):1146–1152, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ley:2014:UMA

[LSD^{+}14] Guillermo Ley, Melody J. Saltzgiver, Thomas E. Dowling, Abraham P. Karam, Brian R. Kesner, and Paul C. Marsh. Use of a molecular assay to detect predation on an endangered fish species. Transactions of the American Fisheries Society, 143 (1):49–54, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liao:2013:QEA

[LSJN13] Hongsheng Liao, Alexei F. Sharov, Cynthia M. Jones, and Gary A. Nelson. Quantifying the effects of aging bias in Atlantic striped bass stock assessment. Transactions of the American Fisheries Society, 142(1):193–207, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leclerc:2011:DFS

[LSPB11] Véronique Leclerc, Pascal Sirois, Dolors Planas, and Pierre Bérubé. Diet and feeding success of fast-growing yellow perch larvae and juveniles in perturbed boreal lakes. Transactions of the American Fisheries Society, 140(5):1193–1205, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loke-Smith:2010:UME

[LSSYL10] Kerri A. Loke-Smith, Michael A. Sundberg, Kelly A. Young, and Christopher G. Lowe. Use of morphology and endocrinology to predict sex in California sheephead: Evidence of altered timing of sex change at Santa Catalina Island, California. Transactions of the American Fisheries Society, 139(6): 1742–1750, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loomis:2013:ESA

[LSV^{+}13] John H. Loomis, Harold L. Schramm, Jr., Bruce Vondracek, Patrick D. Gerard, and Christopher J. Chizinski. Effects of simulated angler capture and live-release tournaments on walleye survival. Transactions of the American Fisheries Society, 142(3):868–875, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lennox:2015:DCR

[LUC^{+}15] Robert J. Lennox, Ingebrigt Uglem, Steven J. Cooke, Tor F. Næsje, Frederick G. Whoriskey, Torgeir B. Havn, Eva M. Ulvan, Øyvind Solem, and Eva B. Thorstad. Does catch-and-release angling alter the behavior and fate of adult Atlantic salmon during upriver migration? Transactions of the American Fisheries Society, 144(2):400–409, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loppnow:2014:SSS

[LV14] Grace L. Loppnow and Paul A. Venturelli. Stage-structured simulations suggest that removing young of the year is an effective method for controlling invasive smallmouth bass. Transactions of the American Fisheries Society, 143(5):1341–1347, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Li:2010:GDV

[LYX^{+}10] Si-Fa Li, Qin-Ling Yang, Jia-Wei Xu, Cheng-Hui Wang, Duane C. Chapman, and Guoqing Lu. Genetic diversity and variation of mitochondrial DNA in native and introduced bighead carp. Transactions of the American Fisheries Society, 139(4):937–946, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

May:2012:MHS

[MAH^{+}12] Cassandra J. May, D. Derek Aday, R. Scott Hale, Jonathan C. S. Denlinger, and Elizabeth A. Marschall. Modeling habitat selection of a top predator: Considering growth and physical environments in a spatial context. Transactions of the American Fisheries Society, 141(1):215–223, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2012:DGM

[MAK12] L. E. Miranda, J. N. Aycock, and K. J. Killgore. A directgradient multivariate index of biotic condition. Transactions of the American Fisheries Society, 141(6):1637–1648, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mueller:2012:LTM

[MAN^{+}12] Robert Mueller, Evan Arntzen, Marc Nabelek, Ben Miller, Katherine Klett, and Ryan Harnish. Laboratory testing of a modified electroshocking system designed for deepwater juvenile lamprey sampling. Transactions of the American Fisheries Society, 141(3):841–845, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacDonald:2010:AES

[MAS10] Alyssa M. MacDonald, Charles F. Adams, and Kevin D. E. Stokesbury. Abundance estimates of skates (Rajidae) on the Continental Shelf of the Northeastern United States using a video survey. Transactions of the American Fisheries Society, 139(5):1415–1420, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matta:2018:AGF

[Mat18] Mary Elizabeth Matta. Age and growth of fishes: Principles and techniques. Transactions of the American Fisheries Society, 147(6):1215, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mueller:2010:TBP

[MBBM10] Anna-Maria Mueller, Deborah L. Burwen, Kevin M. Boswell, and Tim Mulligan. Tail-beat patterns in dual-frequency identification sonar echograms and their potential use for species identification and bioenergetics studies. Transactions of the American Fisheries Society, 139(3):900–910, May 2010. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McClure-Baker:2010:GSH

[MBEvdB^{+}10] Sherri A. McClure-Baker, Anthony A. Echelle, Ronald A. van den Bussche, Alice F. Echelle, Dean A. Hendrickson, and Gary P. Garrett. Genetic status of headwater catfish in Texas and New Mexico: a perspective from mtDNA and morphology. Transactions of the American Fisheries Society, 139(6): 1780–1791, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:2010:NSA

[MBG10] Charles P. Madenjian, David B. Bunnell, and Owen T. Gorman. Ninespine stickleback abundance in Lake Michigan increases after invasion of dreissenid mussels. Transactions of the American Fisheries Society, 139(1):11–20, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mullen:2011:DFA

[MBG^{+}11] Jason A. Mullen, Robert G. Bramblett, Christopher S. Guy, Alexander V. Zale, and David W. Roberts. Determinants of fish assemblage structure in Northwestern Great Plains streams. Transactions of the American Fisheries Society, 140 (2):271–281, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:2012:SPW

[MBH12] Paul H. Michaletz, Przemyslaw G. Bajer, and Robert S. Hayward. Seasonal patterns in white crappies’ consumption and growth: Influences of varying water temperatures and prey availability. Transactions of the American Fisheries Society, 141(3):681–696, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNatt:2016:RMJ

[MBH16] Regan A. McNatt, Daniel L. Bottom, and Susan A. Hinton. Residency and movement of juvenile Chinook salmon at multiple spatial scales in a tidal marsh of the Columbia River Estuary. Transactions of the American Fisheries Society, 145 (4):774–785, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morley:2012:CMP

[MBL12] James W. Morley, Jeffrey A. Buckel, and Thomas E. Lankford, Jr. Comparing multiple predictors of energy content in juvenile bluefish. Transactions of the American Fisheries Society, 141 (4):1109–1116, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mora:2018:EAS

[MBL^{+}18] Ethan A. Mora, Ryan D. Battleson, Steven T. Lindley, Michael J. Thomas, Russ Bellmer, Liam J. Zarri, and A. Peter Klimley. Estimating the annual spawning run size and population size of the southern distinct population segment of green sturgeon. Transactions of the American Fisheries Society, 147(1):195–203, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKenna:2010:SSW

[MBM10] James E. McKenna, Jr., Ryan S. Butryn, and Richard P. Mc-Donald. Summer stream water temperature models for Great Lakes streams: New York. Transactions of the American Fisheries Society, 139(5):1399–1414, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mattei:2011:EST

[MBP^{+}11] J. H. Mattei, M. A. Beekey, H. R. Potter, C. S. Bond, A. R. Woronik, J. A. Roberts, and K. A. Smith. Estimation of short-term tag-induced mortality in horseshoe crabs Limulus polyphemus. Transactions of the American Fisheries Society, 140(4):954–958, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDougall:2013:MPS

[MBPA13] C. A. McDougall, P. J. Blanchfield, S. J. Peake, and W. G. Anderson. Movement patterns and size-class influence entrainment susceptibility of lake sturgeon in a small hydroelectric reservoir. Transactions of the American Fisheries Society, 142 (6):1508–1521, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moncrief:2018:AGR

[MBPP18] Trevor Moncrief, Nancy J. Brown-Peterson, and Mark S. Peterson. Age, growth, and reproduction of vermilion snapper in the North-Central Gulf of Mexico. Transactions of the American Fisheries Society, 147(5):996–1010, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:2010:EMS

[MBT10] Megan E. Moore, Barry A. Berejikian, and Eugene P. Tezak. Early marine survival and behavior of steelhead smolts through Hood Canal and the Strait of Juan de Fuca. Transactions of the American Fisheries Society, 139(1):49–61, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muir:2014:EDL

[MBZ^{+}14] A. M. Muir, C. R. Bronte, M. S. Zimmerman, H. R. Quinlan, J. D. Glase, and C. C. Krueger. Ecomorphological diversity of lake trout at Isle Royale, Lake Superior. Transactions of the American Fisheries Society, 143(4):972–987, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martens:2014:JAS

[MC14] Kyle D. Martens and Patrick J. Connolly. Juvenile anadromous salmonid production in Upper Columbia River side channels with different levels of hydrological connection. Transactions of the American Fisheries Society, 143(3):757–767, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:2012:SGJ

[MCC^{+}12] Matthew G. Mesa, Elizabeth S. Copeland, Helena E. Christiansen, Jacob L. Gregg, Sean R. Roon, and Paul K. Hershberger. Survival and growth of juvenile Pacific lampreys tagged with Passive Integrated Transponders (PIT) in freshwater and seawater. Transactions of the American Fisheries Society, 141(5):1260–1268, September 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:2017:MGI

[MCD17] Paul C. Marsh, Robert W. Clarkson, and Thomas E. Dowling. Molecular genetics informs spatial segregation of two desert stream Gila species. Transactions of the American Fisheries Society, 146(1):47–59, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mussen:2013:AJC

[MCH^{+}13] Timothy D. Mussen, Dennis Cocherell, Zachary Hockett, Ali Ercan, Hossein Bandeh, M. Levent Kavvas, Joseph J. Cech, Jr., and Nann A. Fangue. Assessing juvenile Chinook salmon behavior and entrainment risk near unscreened water diversions: Large flume simulations. Transactions of the American Fisheries Society, 142(1):130–142, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murry:2010:LEA

[MCO^{+}10] Brent A. Murry, Michael J. Connerton, Robert O’Gorman, Donald J. Stewart, and Neil H. Ringler. Lakewide estimates of alewife biomass and Chinook salmon abundance and consumption in Lake Ontario, 1989–2005: Implications for prey fish sustainability. Transactions of the American Fisheries Society, 139(1):223–240, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mussen:2015:MWD

[MCP^{+}15] Timothy D. Mussen, Dennis E. Cocherell, Oliver Patton, Denisse Jauregui, Ali Ercan, Hossein Bandeh, Dan Meier, Steve Thomas, M. Levent Kavvas, Joseph J. Cech, Jr., and Nann A. Fangue. Modified water diversion structures can behaviorally deter juvenile Chinook salmon from entrainment.

Transactions of the American Fisheries Society, 144(5):1070– 1080, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCauley:2014:DBW

[MCSF14] M. M. McCauley, R. M. Cerrato, M. Sclafani, and M. G. Frisk. Diel behavior in white perch revealed using acoustic telemetry. Transactions of the American Fisheries Society, 143(5):1330– 1340, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Munsch:2017:QCD

[MCT17] Stuart H. Munsch, Jeffery R. Cordell, and Jason D. Toft. A quantitative chronology of diurnal feeding in juvenile Pacific salmon. Transactions of the American Fisheries Society, 146 (2):222–229, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mahlum:2014:EBA

[MCW^{+}14] Shad Mahlum, David Cote, Yolanda F. Wiersma, Dan Kehler, and Keith D. Clarke. Evaluating the barrier assessment technique derived from FishXing software and the upstream movement of brook trout through road culverts. Transactions of the American Fisheries Society, 143(1):39–48, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miehls:2011:FIH

[MD11] Scott M. Miehls and John M. Dettmers. Factors influencing habitat shifts of age-0 yellow perch in Southwestern Lake Michigan. Transactions of the American Fisheries Society, 140(5):1317–1329, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muhlfeld:2016:GSC

[MDD^{+}16] Clint C. Muhlfeld, Vincent S. D’Angelo, Christopher Downs, John Powell, Stephen Amish, Gordon Luikart, Ryan Kovach, Matthew Boyer, and Steven Kalinowski. Genetic status and conservation of westslope cutthroat trout in glacier national park. Transactions of the American Fisheries Society, 145(5): 1093–1109, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:2012:EAE

[MDP12] Charles P. Madenjian, Solomon R. David, and Steven A. Pothoven. Effects of activity and energy budget balancing algorithm on laboratory performance of a fish bioenergetics model. Transactions of the American Fisheries Society, 141 (5):1328–1337, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyers:2010:CCI

[MDT10] Erica M. Meyers, Bridget Dobrowski, and Christina L. Tague. Climate change impacts on flood frequency, intensity, and timing may affect trout species in Sagehen Creek, California. Transactions of the American Fisheries Society, 139(6):1657– 1664, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDermid:2012:VAT

[MFAS^{+}12] Jenni L. McDermid, Friedrich A. Fischer, Mohammed Al-Shamlih, William N. Sloan, Nicholas E. Jones, and Chris C. Wilson. Variation in acute thermal tolerance within and among hatchery strains of brook trout. Transactions of the American Fisheries Society, 141(5):1230–1235, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Middaugh:2013:LLS

[MFH13] Christopher R. Middaugh, Carolyn J. Foley, and Tomas O. Höök. Local and lake-scale habitat effects on abundance, lengths, and diets of age-0 largemouth bass and bluegill in Indiana temperate lakes. Transactions of the American Fisheries Society, 142(6):1576–1589, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Monzyk:2015:IJS

[MFR15] Fred R. Monzyk, Thomas A. Friesen, and Jeremy D. Romer. Infection of juvenile salmonids by Salmincola californiensis (Copepoda: Lernaeopodidae) in reservoirs and streams of the Willamette River Basin, Oregon. Transactions of the American Fisheries Society, 144(5):891–902, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Medeiros:2018:PKI

[MGK^{+}18] Lea R. Medeiros, Peter F. Galbreath, Curtis M. Knudsen, Chad A. Stockton, Ilana J. Koch, William J. Bosch, Shawn R. Narum, James J. Nagler, and Andrew L. Pierce. Plasma 11ketotestosterone in individual age-1 spring Chinook salmon males accurately predicts age-2 maturation status. Transactions of the American Fisheries Society, 147(6):1042–1051, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martinez:2011:MGS

[MGP11] Andrés Martínez, John Carlos Garza, and Devon E. Pearse. A microsatellite genome screen identifies chromosomal regions under differential selection in steelhead and rainbow trout. Transactions of the American Fisheries Society, 140(3):829– 842, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morrison:2013:ESD

[MGW13] Phillip R. Morrison, Erick P. Groot, and David W. Welch. The effect of short-duration seawater exposure and acoustic tag implantation on the swimming performance and physiology of presmolt juvenile Coho salmon. Transactions of the American Fisheries Society, 142(3):783–792, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maranto:2010:IPR

[MGWP10] Christina J. Maranto, Thomas P. Good, Francis K. Wiese, and Julia K. Parrish. Impact of the potholes reservoir Caspian tern breeding colony on out-migrating juvenile salmonids in the Mid-Columbia River. Transactions of the American Fisheries Society, 139(2):362–381, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Melnychuk:2011:MED

[MH11] Michael C. Melnychuk and Stephen J. Hausch. Method for estimating detection probabilities of nonmigrant tagged fish: Applications for quantifying residualization rates. Transactions of the American Fisheries Society, 140(6):1613–1628, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:2012:ESC

[MHE12] Kevin A. Meyer, Brett High, and F. Steven Elle. Effects of stocking catchable-sized hatchery rainbow trout on wild rainbow trout abundance, survival, growth, and recruitment. Transactions of the American Fisheries Society, 141(1):224– 237, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:2012:TOV

[MHF12] Jennifer M. Marsh, Nicola Hillgruber, and Robert J. Foy. Temporal and ontogenetic variability in trophic role of four groundfish species — walleye pollock, Pacific cod, arrowtooth flounder, and Pacific halibut — around Kodiak Island in the Gulf of Alaska. Transactions of the American Fisheries Society, 141 (2):468–486, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McBride:2010:SRB

[MHHH10] Richard S. McBride, Julianne E. Harris, A. Reid Hyle, and Jay C. Holder. The spawning run of blueback herring in the St. Johns River, Florida. Transactions of the American Fisheries Society, 139(2):598–609, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matala:2011:RAD

[MHN11] Andrew P. Matala, Jon E. Hess, and Shawn R. Narum. Resolving adaptive and demographic divergence among Chinook salmon populations in the Columbia River Basin. Transactions of the American Fisheries Society, 140(3):783–807, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murawski:2014:PES

[MHPB14] Steven A. Murawski, William T. Hogarth, Ernst B. Peebles, and Luiz Barbeiri. Prevalence of external skin lesions and polycyclic aromatic hydrocarbon concentrations in Gulf of Mexico fishes, post-Deepwater Horizon. Transactions of the American Fisheries Society, 143(4):1084–1097, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:2010:OSA

[Mic10] Paul H. Michaletz. Overwinter survival of age-0 gizzard shad in Missouri reservoirs spanning a productivity gradient: Roles of body size and winter severity. Transactions of the American Fisheries Society, 139(1):241–256, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Milner:2019:BTB

[Mil19] Nigel Milner. Brown trout: Biology, ecology, and management. Transactions of the American Fisheries Society, 148(3):682, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mitro:2016:BTB

[Mit16] Matthew G. Mitro. Brook trout, brown trout, and ectoparasitic copepods Salmincola edwardsii: Species interactions as a proximate cause of brook trout loss under changing environmental conditions. Transactions of the American Fisheries Society, 145(6):1223–1233, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miralles:2014:ISB

[MJGV14] L. Miralles, F. Juanes, and E. Garcia-Vazquez. Interoceanic sex-biased migration in bluefish. Transactions of the American Fisheries Society, 143(5):1308–1315, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miehls:2017:EGE

[MJH17] Scott M. Miehls, Nicholas S. Johnson, and Alex Haro. Electrical guidance efficiency of downstream-migrating juvenile sea lampreys decreases with increasing water velocity. Transactions of the American Fisheries Society, 146(2):299–307, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCann:2018:CSL

[MJHP18] Erin L. McCann, Nicholas S. Johnson, Peter J. Hrodey, and Kevin L. Pangle. Characterization of sea lamprey stream entry using dual-frequency identification sonar. Transactions of the American Fisheries Society, 147(3):514–524, May 2018.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCartin:2019:NPA

[MJS^{+}19] Kellie McCartin, Adrian Jordaan, Matthew Sclafani, Robert Cerrato, and Michael G. Frisk. A new paradigm in alewife migration: Oscillations between spawning grounds and estuarine habitats. Transactions of the American Fisheries Society, 148 (3):605–619, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2010:UEE

[MK10] L. E. Miranda and R. H. Kidwell. Unintended effects of electrofishing on nongame fishes. Transactions of the American Fisheries Society, 139(5):1315–1321, September 2010. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Myrvold:2015:VJS

[MK15] Knut Marius Myrvold and Brian P. Kennedy. Variation in juvenile steelhead density in relation to instream habitat and watershed characteristics. Transactions of the American Fisheries Society, 144(3):577–590, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miyakoshi:2016:OES

[MK16] Yasuyuki Miyakoshi and Shuichi Kitada. Overdispersion in the estimation of salmon escapements by mark–recapture surveys and its related factors. Transactions of the American Fisheries Society, 145(2):269–276, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Midway:2010:HSC

[MKA10] Stephen R. Midway, Thomas J. Kwak, and D. Derek Aday. Habitat suitability of the Carolina madtom, an imperiled, endemic stream fish. Transactions of the American Fisheries Society, 139(2):325–338, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macdonald:2010:SSR

[MKH10] J. Stevenson Macdonald, Cher A. King, and Herb Herunter. Sediment and salmon: The role of spawning sockeye salmon in annual bed load transport characteristics in small, interior streams of British Columbia. Transactions of the American Fisheries Society, 139(3):758–767, May 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2013:FAB

[MKH13] L. E. Miranda, K. J. Killgore, and J. J. Hoover. Fish assemblages in borrow-pit lakes of the Lower Mississippi River. Transactions of the American Fisheries Society, 142(3):596– 605, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:2017:PYC

[MKHC17] Kevin A. Meyer, Patrick Kennedy, Brett High, and Matthew R. Campbell. Purifying a Yellowstone cutthroat trout stream by removing rainbow trout and hybrids via electrofishing. Transactions of the American Fisheries Society, 146(6):1193–1203, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:2015:DPD

[MKK15] Brett A. Miller, William E. Kelso, and Michael D. Kaller. Diet partitioning in a diverse centrarchid assemblage in the Atchafalaya River Basin, Louisiana. Transactions of the American Fisheries Society, 144(4):780–791, July 2015. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mollenhauer:2018:QSD

[MLB18] Robert Mollenhauer, Daniel Logue, and Shannon K. Brewer. Quantifying seining detection probability for fishes of Great Plains sand-bed rivers. Transactions of the American Fisheries Society, 147(2):329–341, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mychek-Londer:2013:UDR

[MLBS^{+}13] Justin G. Mychek-Londer, David B. Bunnell, Wendylee Stott, James S. Diana, John R. P. French III, and Margret A. Chriscinske. Using diets to reveal overlap and egg predation among benthivorous fishes in Lake Michigan. Transactions of the American Fisheries Society, 142(2):492–504, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morin:2013:BRV

[MLC13] Roderick Morin, Sophie G. LeBlanc, and Steven E. Campana. Bomb radiocarbon validates age and long-term growth declines in American plaice in the Southern Gulf of St. Lawrence. Transactions of the American Fisheries Society, 142(2):458– 470, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McMillan:2013:USN

[MLS^{+}13] John R. McMillan, Martin C. Liermann, James Starr, George R. Pess, and Xan Augerot. Using a stream network census of fish and habitat to assess models of juvenile salmonid distribution. Transactions of the American Fisheries Society, 142(4):942–956, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNicol:2010:AUS

[MM10] Richard E. McNicol and Shayne E. MacLellan. Accuracy of using scales to age mixed-stock Chinook salmon of hatchery origin. Transactions of the American Fisheries Society, 139 (3):727–734, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marklevitz:2017:HUA

[MM17] Stephen A. C. Marklevitz and Yolanda E. Morbey. Habitat use and arrival timing of hatchery and naturalized Chinook salmon spawning in a Great Lakes tributary. Transactions of the American Fisheries Society, 146(4):567–583, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:2012:ISC

[MMY12] Loren M. Miller, Steven W. Mero, and Jerry A. Younk. The impact of stocking on the current ancestry in twenty native and introduced muskellunge populations in Minnesota. Transactions of the American Fisheries Society, 141(5):1411–1423, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDougall:2018:EFI

[MNB18] C. A. McDougall, P. A. Nelson, and C. C. Barth. Extrinsic factors influencing somatic growth of lake sturgeon. Transactions of the American Fisheries Society, 147(3):459–479, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDougall:2017:HQR

[MNM^{+}17] C. A. McDougall, P. A. Nelson, D. Macdonald, D. Kroeker, K. Kansas, C. C. Barth, and D. S. MacDonell. Habitat quantity required to support self-sustaining lake sturgeon populations: an alternative hypothesis. Transactions of the American Fisheries Society, 146(6):1137–1155, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matala:2019:EOM

[MNS^{+}19] Andrew P. Matala, Shawn R. Narum, Brian P. Saluskin, Mark V. Johnston, Jeremiah E. Newell, David E. Fast, and Peter F. Galbreath. Early observations from monitoring a reintroduction program: Return of sockeye salmon to a nursery lake of historical importance. Transactions of the American Fisheries Society, 148(2):271–288, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murauskas:2013:RBA

[MOS13] Joshua G. Murauskas, Alexei M. Orlov, and Kevin A. Siwicke. Relationships between the abundance of Pacific lamprey in the Columbia River and their common hosts in the marine environment. Transactions of the American Fisheries Society, 142(1):143–155, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCargo:2010:EIS

[MP10] Jeremy W. McCargo and James T. Peterson. An evaluation of the influence of seasonal base flow and geomorphic stream characteristics on coastal plain stream fish assemblages. Transactions of the American Fisheries Society, 139(1):29–48, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macdonald:2010:MIE

[MPHG10] J. Stevenson Macdonald, David A. Patterson, Merran J. Hague, and Ian C. Guthrie. Modeling the influence of environmental factors on spawning migration mortality for sockeye salmon fisheries management in the Fraser River, British Columbia. Transactions of the American Fisheries Society, 139(3):768–782, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McMillan:2014:DUF

[MPM^{+}14] John R. McMillan, George R. Pess, Michael L. McHenry, Raymond Moses, and Thomas P. Quinn. Documentation of unusual, fall spawning by coastal cutthroat trout in the Elwha River System, Washington. Transactions of the American Fisheries Society, 143(6):1605–1611, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murauskas:2011:RDR

[MR11] J. G. Murauskas and R. A. Rulifson. Reproductive development and related observations during the spawning migration of hickory shad. Transactions of the American Fisheries Society, 140(4):1035–1048, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Midwood:2018:IIT

[MRB^{+}18] Jonathan D. Midwood, David T. Reddick, Jill L. Brooks, Christine Boston, Susan E. Doka, and Steven J. Cooke. Intracoelomic implantation of transmitters in longnose gar. Transactions of the American Fisheries Society, 147(4):704–710, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mandeville:2019:IFA

[MRPW19] Caitlin P. Mandeville, Frank J. Rahel, Lindsay S. Patterson, and Annika W. Walters. Integrating fish assemblage data, modeled stream temperatures, and thermal tolerance metrics to develop thermal guilds for water temperature regulation: Wyoming case study. Transactions of the American Fisheries Society, 148(4):739–754, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mudrak:2012:PEL

[MS12] Peter A. Mudrak and Stephen T. Szedlmayer. Proximity effects of larger resident fishes on recruitment of age-0 red snapper in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 141(2):487–494, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:2013:TSB

[MS13] Paul H. Michaletz and Michael J. Siepker. Trends and synchrony in black bass and crappie recruitment in Missouri reservoirs. Transactions of the American Fisheries Society, 142(1): 105–118, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Massie:2018:SVM

[MSB^{+}18] Danielle L. Massie, Geoffrey D. Smith, Timothy F. Bonvechio, Aaron J. Bunch, David O. Lucchesi, and Tyler Wagner. Spatial variability and macro-scale drivers of growth for native and introduced flathead catfish populations. Transactions of the American Fisheries Society, 147(3):554–565, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:2010:MVB

[MSBT10] Kyle E. Martin, Craig A. Steele, Joseph P. Brunelli, and Gary H. Thorgaard. Mitochondrial variation and biogeographic history of Chinook salmon. Transactions of the American Fisheries Society, 139(3):792–802, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDonald:2018:CTB

[MSD18] Dusty L. McDonald, J. Warren Schlechte, and Daniel J. Daugherty. Comparison of two biometric methods for nonlethal sex determination of alligator gars. Transactions of the American Fisheries Society, 147(4):711–715, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:2019:CSG

[MSD^{+}19] Douglas J. Martin, Alice A. Shelly, Robert J. Danehy, Emily D. Lang, and Jonathon Hvozda. Coho salmon growth in relation to natural turbidity regimes in a coastal stream of Northern California. Transactions of the American Fisheries Society, 148(4):817–831, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Myers:2014:SSS

[MSH^{+}14] Ransom A. Myers, Matthew W. Smith, John M. Hoenig, Neil Kmiecik, Mark A. Luehring, Melissa T. Drake, Patrick J.

Schmalz, and Greg G. Sass. Size- and sex-specific capture and harvest selectivity of walleyes from tagging studies. Transactions of the American Fisheries Society, 143(2):438–450, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matsubu:2017:JSL

[MSH17] William Matsubu, Charles A. Simenstad, and Gregg E. Horton. Juvenile steelhead locate coldwater refugia in an intermittently closed estuary. Transactions of the American Fisheries Society, 146(4):680–695, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marion:2015:MEI

[MSK15] Cathy A. Marion, Mark C. Scott, and Kevin M. Kubach. Multiscale environmental influences on fish assemblage structure of South Atlantic coastal plain streams. Transactions of the American Fisheries Society, 144(5):1040–1057, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyer:2012:PPP

[MSP12] G. R. Moyer, J. A. Sweka, and D. L. Peterson. Past and present processes influencing genetic diversity and effective population size in a natural population of Atlantic sturgeon. Transactions of the American Fisheries Society, 141(1):56–67, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGlauflin:2011:SHG

[MSS^{+}11] Molly T. McGlauflin, Daniel E. Schindler, Lisa W. Seeb, Christian T. Smith, Christopher Habicht, and James E. Seeb. Spawning habitat and geography influence population structure and juvenile migration timing of sockeye salmon in the Wood River Lakes, Alaska. Transactions of the American Fisheries Society, 140(3):763–782, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muttray:2017:DCN

[MSS^{+}17] A. F. Muttray, D. Sakhrani, J. L. Smith, I. Nakayama, W. S. Davidson, L. Park, and R. H. Devlin. Deletion and copy number variation of Y-chromosomal regions in Coho salmon, chum salmon, and pink salmon populations. Transactions of the American Fisheries Society, 146(2):240–251, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGlauflin:2010:HRM

[MSW^{+}10] Molly T. McGlauflin, Matt J. Smith, Jonathan T. Wang, Sewall F. Young, Neng Chen, Yong C. Lee, Carita Pascal, Lisa W. Seeb, Junko Stevens, and James E. Seeb. Highresolution melting analysis for the discovery of novel singlenucleotide polymorphisms in rainbow and cutthroat trout for species identification. Transactions of the American Fisheries Society, 139(3):676–684, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mayfield:2014:SPP

[MSW^{+}14] M. P. Mayfield, L. D. Schultz, L. A. Wyss, B. J. Clemens, and C. B. Schreck. Spawning patterns of Pacific lamprey in tributaries to the Willamette River, Oregon. Transactions of the American Fisheries Society, 143(6):1544–1554, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muttray:2015:LVC

[MSWD15] Annette F. Muttray, Dionne Sakhrani, Ruth E. Withler, and Robert H. Devlin. Low variation in a Y-chromosomal growth hormone pseudogene relative to its functional autosomal progenitor gene in Chinook salmon. Transactions of the American Fisheries Society, 144(5):1029–1039, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDermid:2010:ELH

[MSWS10] Jenni L. McDermid, William N. Sloan, Chris C. Wilson, and Brian J. Shuter. Early life history variation among hatcheryand wild-origin lake trout reared in a hatchery environment. Transactions of the American Fisheries Society, 139(1):21–28, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:2019:LGV

[MT19] Daniel L. Martin and Timothy E. Targett. Latitudinal growth variation in young tautogs: a field, laboratory, and growing degree-day assessment of potential intrinsic differences in growth responses to temperature. Transactions of the American Fisheries Society, 148(4):698–709, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muir:2012:PPM

[MVK12] A. M. Muir, P. Vecsei, and C. C. Krueger. A perspective on perspectives: Methods to reduce variation in shape analysis of digital images. Transactions of the American Fisheries Society, 141(4):1161–1170, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Montie:2015:SSP

[MVP15] Eric W. Montie, Steven Vega, and Michael Powell. Seasonal and spatial patterns of fish sound production in the May River, South Carolina. Transactions of the American Fisheries Society, 144(4):705–716, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:2013:DEB

[MWC^{+}13] Matthew G. Mesa, Lisa K. Weiland, Helena E. Christiansen, Sally T. Sauter, and David A. Beauchamp. Development and evaluation of a bioenergetics model for bull trout. Transactions of the American Fisheries Society, 142(1):41–49, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moran:2018:KMF

[MWG18] Clinton J. Moran, David L. Ward, and Alice C. Gibb. Key morphological features favor the success of nonnative fish species under reduced turbidity conditions in the Lower Colorado River Basin. Transactions of the American Fisheries Society, 147(5):948–958, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mollenhauer:2013:FEW

[MWKS13] Robert Mollenhauer, Tyler Wagner, Megan V. Kepler, and John A. Sweka. Fall and early winter movement and habitat use of wild brook trout. Transactions of the American Fisheries Society, 142(5):1167–1178, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:2010:ACR

[MWS10] Aaron E. Martin, Mark S. Wipfli, and Robert E. Spangler. Aquatic community responses to salmon carcass analog and wood bundle additions in restored floodplain habitats in an Alaskan stream. Transactions of the American Fisheries Society, 139(6):1828–1845, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDermid:2013:IDT

[MWSS13] Jenni L. McDermid, Chris C. Wilson, William N. Sloan, and Brian J. Shuter. Intraspecific differences in thermal biology among inland lake trout populations. Transactions of the American Fisheries Society, 142(3):756–766, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Midway:2014:HCO

[MWT14] Stephen R. Midway, Tyler Wagner, and Bryn H. Tracy. A hierarchical community occurrence model for North Carolina stream fish. Transactions of the American Fisheries Society, 143(5):1348–1357, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Malcolm:2011:IVG

[MYS^{+}11] I. A. Malcolm, A. F. Youngson, C. Soulsby, C. Imholt, and R. J. Fryer. Is interstitial velocity a good predictor of salmonid embryo survival? Transactions of the American Fisheries Society, 140(4):898–904, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nutile:2013:EEE

[NAG13] Sam Nutile, Jon J. Amberg, and Reuben R. Goforth. Evaluating the effects of electricity on fish embryos as a potential strategy for controlling invasive cyprinids. Transactions of the American Fisheries Society, 142(1):1–9, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neeson:2012:TPD

[NAW12] Thomas M. Neeson, Sara A. Adlerstein, and Michael J. Wiley. Towards a process domain-sensitive substrate habitat model for sea lampreys in Michigan rivers. Transactions of the American Fisheries Society, 141(2):313–326, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Natanson:2014:AGB

[NAWM14] Lisa J. Natanson, Douglas H. Adams, Megan V. Winton, and Jasmine R. Maurer. Age and growth of the bull shark in the Western North Atlantic Ocean. Transactions of the American Fisheries Society, 143(3):732–743, May 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Negus:2012:RIT

[NB12] Mary T. Negus and Roger A. Bergstedt. Rates of intraperitoneal temperature change in lake trout implanted with archival tags. Transactions of the American Fisheries Society, 141(2):383–391, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neville:2013:CGS

[NB13] H. M. Neville and L. Bernatchez. Coding gene single nucleotide polymorphism population genetics of nonnative brook trout: The ghost of introductions past. Transactions of the American Fisheries Society, 142(5):1215–1231, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neville:2015:PSA

[NBC15] Chrys M. Neville, Richard J. Beamish, and Cedar M. Chittenden. Poor survival of acoustically-tagged juvenile Chinook salmon in the Strait of Georgia, British Columbia, Canada. Transactions of the American Fisheries Society, 144(1):25–33, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Novak:2017:CFH

[NCW^{+}17] Ashleigh J. Novak, Amy E. Carlson, Carolyn R. Wheeler, Gail S. Wippelhauser, and James A. Sulikowski. Critical foraging habitat of Atlantic sturgeon based on feeding habits, prey distribution, and movement patterns in the Saco River Estuary, Maine. Transactions of the American Fisheries Society, 146(2):308–317, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nannini:2014:ALT

[NDC^{+}14] Michael A. Nannini, Matthew J. Diana, Julie E. Claussen, David P. Philipp, and David H. Wahl. Assessing the long-term reproductive contribution of stocked fish to largemouth bass populations. Transactions of the American Fisheries Society, 143(3):606–612, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neville:2016:MDG

[NDP16] Helen Neville, Dan Dauwalter, and Mary Peacock. Monitoring demographic and genetic responses of a threatened inland trout to habitat reconnection. Transactions of the American Fisheries Society, 145(3):610–626, May 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norris:2010:CEH

[NDW10] Alicia J. Norris, Dennis R. DeVries, and Russell A. Wright. Coastal estuaries as habitat for a freshwater fish species: Exploring population-level effects of salinity on largemouth bass. Transactions of the American Fisheries Society, 139(2):610– 625, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:2014:BRB

[Nel14a] Gary A. Nelson. Book review: Biology and Management of Inland Striped Bass and Hybrid Striped Bass: Edited by James S. Bulak, Charles C. Coutant, and James A. Rice. American Fisheries Society, Symposium 80, Bethesda, Maryland. 2013. 588 pages. $79.00 ($55.30 for American Fisheries Society members) (hardcover). Transactions of the American Fisheries Society, 143(4):1112–1113, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:2014:CSP

[Nel14b] Gary A. Nelson. Cluster sampling: a pervasive, yet little recognized survey design in fisheries research. Transactions of the American Fisheries Society, 143(4):926–938, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Natanson:2013:RBC

[NG13] Lisa J. Natanson and Brian J. Gervelis. The reproductive biology of the common thresher shark in the Western North Atlantic Ocean. Transactions of the American Fisheries Society, 142(6):1546–1562, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Narum:2010:EGL

[NHM10] Shawn R. Narum, Jon E. Hess, and Andrew P. Matala. Examining genetic lineages of Chinook salmon in the Columbia River Basin. Transactions of the American Fisheries Society, 139(5):1465–1477, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neff:2012:GSM

[NJ12] Margaret R. Neff and Donald A. Jackson. Geology as a structuring mechanism of stream fish communities. Transactions of the American Fisheries Society, 141(4):962–974, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nobriga:2013:CPR

[NLF13] Matthew L. Nobriga, Erik Loboschefsky, and Frederick Feyrer. Common predator, rare prey: Exploring juvenile striped bass predation on delta smelt in California’s San Francisco Estuary. Transactions of the American Fisheries Society, 142(6):1563– 1575, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neely:2012:CED

[NMH12] Kathleen G. Neely, James M. Myers, and Jeffrey J. Hard. A comparison of early development between a domesticated stock of Coho salmon and its parental stock. Transactions of the American Fisheries Society, 141(6):1504–1509, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nannini:2012:DBS

[NPW12] Michael A. Nannini, Joseph Parkos III, and David H. Wahl. Do behavioral syndromes affect foraging strategy and risk-taking in a juvenile fish predator? Transactions of the American Fisheries Society, 141(1):26–33, January 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nobriga:2016:PDE

[NR16] Matthew L. Nobriga and Jonathan A. Rosenfield. Population dynamics of an estuarine forage fish: Disaggregating forces driving long-term decline of longfin smelt in California’s San Francisco Estuary. Transactions of the American Fisheries Society, 145(1):44–58, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neher:2013:EER

[NRZ^{+}13] Tammy D. Hoem Neher, Amanda E. Rosenberger, Christian E. Zimmerman, Coowe M. Walker, and Steven J. Baird. Estuarine environments as rearing habitats for juvenile Coho salmon in contrasting south-central Alaska watersheds. Transactions of the American Fisheries Society, 142(6):1481–1494, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nohner:2018:IAM

[NTHR18] Joel K. Nohner, William W. Taylor, Daniel B. Hayes, and Brian M. Roth. Influence of aquatic macrophytes on age-0 largemouth bass growth and diets. Transactions of the American Fisheries Society, 147(4):758–769, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nims:2014:CSF

[NW14] Megan K. Nims and Benjamin D. Walther. Contingents of southern flounder from subtropical estuaries revealed by otolith chemistry. Transactions of the American Fisheries Society, 143(3):721–731, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nguyen:2019:EAM

[NYCC19] Phong L. Nguyen, Benjamin A. Yefimov, Allison Coffin, and Cynthia Cooper. Electroimmobilization as an alternative method to chemical immobilization of larval fish. Transactions of the American Fisheries Society, 148(4):725–738, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Narum:2011:CGM

[NZF^{+}11] Shawn R. Narum, Joseph S. Zendt, Chris Frederiksen, Nate Campbell, Andrew Matala, and William R. Sharp. Candidate genetic markers associated with anadromy in Oncorhynchus mykiss of the Klickitat River. Transactions of the American Fisheries Society, 140(3):843–854, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ostberg:2011:GMD

[ODG^{+}11] C. O. Ostberg, J. J. Duda, J. H. Graham, S. Zhang, K. P. Haywood III, B. Miller, and T. L. Lerud. Growth, morphology, and developmental instability of rainbow trout, Yellowstone cutthroat trout, and four hybrid generations. Transactions of the American Fisheries Society, 140(2):334–344, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Overton:2015:CLT

[OGM^{+}15] Anthony S. Overton, Jennifer C. Griffin, F. Joseph Margraf, Eric B. May, and Kyle J. Hartman. Chronicling long-term predator responses to a shifting forage base in Chesapeake Bay: an energetics approach. Transactions of the American Fisheries Society, 144(5):956–966, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Odenkirk:2016:TAN

[OI16] John S. Odenkirk and Mike W. Isel. Trends in abundance of northern snakeheads in Virginia tributaries of the Potomac River. Transactions of the American Fisheries Society, 145 (4):687–692, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OMalley:2013:TPF

[OMA13] Kathleen G. O’Malley, Douglas F. Markle, and William R. Ardren. Timing of population fragmentation in a vulnerable minnow, the umpqua chub, and the role of nonnative predators. Transactions of the American Fisheries Society, 142(2): 447–457, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Obedzinski:2018:EFR

[OPHD18] Mariska Obedzinski, Sarah Nossaman Pierce, Gregg E. Horton, and Mathew J. Deitch. Effects of flow-related variables on oversummer survival of juvenile Coho salmon in intermittent streams. Transactions of the American Fisheries Society, 147 (3):588–605, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olin:2014:PES

[OPS^{+}14] Jill A. Olin, Gregg R. Poulakis, Philip W. Stevens, Jacquelyn A. DeAngelo, and Aaron T. Fisk. Preservation effects on stable isotope values of archived elasmobranch fin tissue: Comparisons between frozen and ethanol-stored samples. Transactions of the American Fisheries Society, 143(6):1569–1576, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

ONeal:2011:PMR

[OS11] Sarah L. O’Neal and Jack A. Stanford. Partial migration in a robust brown trout population of a Patagonian river. Transactions of the American Fisheries Society, 140(3):623–635, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ogburn:2017:ARH

[OSA^{+}17] Matthew B. Ogburn, Jason Spires, Robert Aguilar, Michael R. Goodison, Keira Heggie, Eva Kinnebrew, William McBurney, Kimberly D. Richie, Paige M. Roberts, and Anson H. Hines. Assessment of river herring spawning runs in a Chesapeake Bay coastal plain stream using imaging sonar. Transactions of the American Fisheries Society, 146(1):22–35, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osborne:2018:VRS

[OSDT18] Megan J. Osborne, Alyssa V. Sanchez, Thomas E. Dowling, and Thomas F. Turner. Variance in reproductive success is driven by environmental factors, not mating system, in bonytails. Transactions of the American Fisheries Society, 147(6): 1100–1114, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OMalley:2017:UHD

[OSS^{+}17] Michael B. O’Malley, Rory Saunders, Justin R. Stevens, J. Michael Jech, and Timothy F. Sheehan. Using hydroacoustics to describe pelagic fish distribution in the Penobscot Estuary, Maine. Transactions of the American Fisheries Society, 146(5):817–833, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OBrien:2014:EFA

[OTR^{+}14] Timothy P. O’Brien, William W. Taylor, Edward F. Roseman, Charles P. Madenjian, and Stephen C. Riley. Ecological factors affecting rainbow smelt recruitment in the main basin of Lake Huron, 1976–2010. Transactions of the American Fisheries Society, 143(3):784–795, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oplinger:2011:EPG

[OWN11] Randall W. Oplinger, David H. Wahl, and Michael A. Nannini. Effect of predators on growth and reproductive investment in bluegills: Comparison of populations with differing size and age of maturation. Transactions of the American Fisheries Society, 140(6):1532–1539, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oplinger:2013:PIS

[OWP13] Randall W. Oplinger, David H. Wahl, and David P. Philipp. Parental influence on the size and age at maturity of bluegills. Transactions of the American Fisheries Society, 142(4):1067– 1074, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:2017:TOG

[OWP^{+}17] Jeffrey B. Olsen, John K. Wenburg, Scott A. Pavey, Joe L. Miller, and Troy R. Hamon. The time of origin and genetic diversity of three isolated Kokanee populations in Southwest Alaska. Transactions of the American Fisheries Society, 146 (6):1212–1222, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:2016:RTV

[OWS16] Zachary Olsen, Tom Wagner, and Glen Sutton. Regional and temporal variation in size at maturity of female blue crabs in Texas coastal waters. Transactions of the American Fisheries Society, 145(5):1059–1067, September 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Purtlebaugh:2010:RAG

[PA10] Caleb H. Purtlebaugh and Micheal S. Allen. Relative abundance, growth, and mortality of five age-0 estuarine fishes in relation to discharge of the Suwannee River, Florida. Transactions of the American Fisheries Society, 139(4):1233–1246, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pope:2014:FR

[PAA14] Kevin L. Pope, Craig R. Allen, and David G. Angeler. Fishing for resilience. Transactions of the American Fisheries Society, 143(2):467–478, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearson:2011:MFI

[PAB^{+}11] Mark S. Pearson, Ted R. Angradi, David W. Bolgrien, Terri M. Jicha, Debra L. Taylor, Mary F. Moffett, and Brian H. Hill. Multimetric fish indices for midcontinent (USA) great rivers. Transactions of the American Fisheries Society, 140(6):1547– 1564, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:2011:UME

[PAG11] Catherine T. Phillips, Mara L. Alexander, and Alyson M. Gonzales. Use of macrophytes for egg deposition by the endangered fountain darter. Transactions of the American Fisheries Society, 140(5):1392–1397, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:2019:CTE

[Pag19] Lawrence M. Page. Cutthroat trout: Evolutionary biology and taxonomy. Transactions of the American Fisheries Society, 148(4):869–870, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2013:ASS

[PB13] Douglas L. Peterson and Michael S. Bednarski. Abundance and size structure of shortnose sturgeon in the Altamaha River, Georgia. Transactions of the American Fisheries Society, 142(5):1444–1452, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:2015:EDW

[PB15] Timothy M. Parker and Michael E. Barnes. Effects of different water velocities on the hatchery rearing performance and recovery from transportation of rainbow trout fed two different rations. Transactions of the American Fisheries Society, 144(5):882–890, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pothoven:2016:SBC

[PB16] Steven A. Pothoven and David B. Bunnell. A shift in bloater consumption in Lake Michigan between 1993 and 2011 and its effects on Diporeia and Mysis prey. Transactions of the American Fisheries Society, 145(1):59–68, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petre:2017:DHR

[PB17] Sally J. Petre and Scott A. Bonar. Determination of habitat requirements for Apache trout. Transactions of the American Fisheries Society, 146(1):1–15, January 2017. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perry:2015:ETR

[PBB^{+}15] Russell W. Perry, Patricia L. Brandes, Jon R. Burau, Philip T. Sandstrom, and John R. Skalski. Effect of tides, river flow, and gate operations on entrainment of juvenile salmon into the Interior Sacramento–San Joaquin River Delta. Transactions of the American Fisheries Society, 144(3):445–455, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pflugrath:2012:MNB

[PBC12] Brett D. Pflugrath, Richard S. Brown, and Thomas J. Carlson. Maximum neutral buoyancy depth of juvenile Chinook salmon: Implications for survival during hydroturbine passage. Transactions of the American Fisheries Society, 141(2): 520–525, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Panther:2011:IIL

[PBG^{+}11] Jennifer L. Panther, Richard S. Brown, Greg L. Gaulke, Katherine A. Deters, Christa M. Woodley, and M. Brad Eppard. Influence of incision location on transmitter loss, healing, survival, growth, and suture retention of juvenile Chinook salmon. Transactions of the American Fisheries Society, 140 (6):1492–1503, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pothoven:2012:EDB

[PBM^{+}12] Steven A. Pothoven, David B. Bunnell, Charles P. Madenjian, Owen T. Gorman, and Edward F. Roseman. Energy density of bloaters in the Upper Great Lakes. Transactions of the American Fisheries Society, 141(3):772–780, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Padilla:2016:DMD

[PBW16] Andrew J. Padilla, Randy J. Brown, and Matthew J. Wooller. Determining the movements and distribution of anadromous Bering ciscoes by use of otolith strontium isotopes. Transactions of the American Fisheries Society, 145(6):1374–1385, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2013:DTH

[PCC^{+}13] Rodney B. Pierce, Andrew J. Carlson, Bruce M. Carlson, Dallas Hudson, and David F. Staples. Depths and thermal habitat used by large versus small northern pike in three Minnesota lakes. Transactions of the American Fisheries Society, 142(6): 1629–1639, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peoples:2017:DBE

[PCFH17] Brandon K. Peoples, Pearce Cooper, Emmanuel A. Frimpong, and Eric M. Hallerman. DNA barcoding elucidates cyprinid reproductive interactions in a Southwest Virginia stream. Transactions of the American Fisheries Society, 146(1):84–91, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Plumb:2012:EPC

[PCT^{+}12] John M. Plumb, William P. Connor, Kenneth F. Tiffan, Christine M. Moffitt, Russell W. Perry, and Noah S. Adams. Estimating and predicting collection probability of fish at dams using multistate modeling. Transactions of the American Fisheries Society, 141(5):1364–1373, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powers:2018:DAC

[PDH^{+}18] Sean P. Powers, J. Marcus Drymon, Crystal L. Hightower, Trey Spearman, George S. Bosarge, and Amanda Jefferson. Distribution and age composition of red snapper across the Inner Continental Shelf of the North-Central Gulf of Mexico. Transactions of the American Fisheries Society, 147(5): 791–805, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pritt:2014:SLF

[PDM^{+}14] Jeremy J. Pritt, Mark R. DuFour, Christine M. Mayer, Edward F. Roseman, and Robin L. DeBruyne. Sampling little fish in big rivers: Larval fish detection probabilities in two Lake Erie tributaries and implications for sampling effort and abundance indices. Transactions of the American Fisheries Society, 143(4):1011–1027, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2012:SBM

[PDP12] Ron Pierce, Mike Davidson, and Craig Podner. Spawning behavior of mountain whitefish and co-occurrence of Myxobolus cerebralis in the Blackfoot River Basin, Montana. Transactions of the American Fisheries Society, 141(3):720–730, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pampoulie:2011:NNG

[PDSP^{+}11] Christophe Pampoulie, Anna Kristín Daníelsdóttir, Marie Storr-Paulsen, Holger Hovgård, Einar Hjörleifsson, and Björn Ævarr Steinarsson. Neutral and nonneutral genetic markers revealed the presence of inshore and offshore stock components of Atlantic cod in Greenland waters. Transactions of the American Fisheries Society, 140(2):307–319, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierson:2014:REF

[PE14] Katherine J. Pierson and David B. Eggleston. Response of estuarine fish to large-scale oyster reef restoration. Transactions of the American Fisheries Society, 143(2):273–288, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2015:RRF

[Pet15] William Peterson. A review of “Regional Fisheries Oceanography of the California Current System: The CalCOFI Program”, by Sam McClatchie. Springer, New York. 2014. 235 pages. $189 (hardcover), $149 (e-book). Transactions of the American Fisheries Society, 144(2):443–444, March 2015. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prystay:2017:CBC

[PET^{+}17] Tanya S. Prystay, Chris K. Elvidge, William M. Twardek, James M. Logan, Connor H. Reid, Shannon H. Clarke, Jordann G. Foster, Emma L. L. Cooke, and Steven J. Cooke. Comparison of the behavioral consequences and recovery patterns of largemouth bass exposed to MS-222 or electrosedation. Transactions of the American Fisheries Society, 146 (3):556–566, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peoples:2011:API

[PF11a] Brandon K. Peoples and Emmanuel A. Frimpong. Amongpass, interregional, and single- versus multiple-season comparisons of detection probabilities of stream fishes. Transactions of the American Fisheries Society, 140(1):67–83, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2011:EMD

[PF11b] Douglas L. Peterson and Daniel J. Farrae. Evidence of metapopulation dynamics in shortnose sturgeon in the southern part of their range. Transactions of the American Fisheries Society, 140(6):1540–1546, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pennock:2019:FEE

[PFG19] Casey A. Pennock, Michael A. Farrington, and Keith B. Gido. Feeding ecology of early life stage razorback sucker relative to other sucker species in the San Juan River, Utah. Transactions of the American Fisheries Society, 148(5):938–951, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearse:2011:PSG

[PGJ11] Devon E. Pearse, Stephanie L. Gunckel, and Steven E. Jacobs. Population structure and genetic divergence of coastal rainbow and redband trout in the Upper Klamath Basin. Transactions of the American Fisheries Society, 140(3):587–597, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Price:2013:PSF

[PGR13] Michael H. H. Price, Barry W. Glickman, and John D. Reynolds. Prey selectivity of Fraser River sockeye salmon during early marine migration in British Columbia. Transactions of the American Fisheries Society, 142(4):1126–1133, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Price:2013:ASR

[PGS13] Michael H. H. Price, Nick Gayeski, and Jack A. Stanford. Abundance of Skeena River chum salmon during the early rise of commercial fishing. Transactions of the American Fisheries Society, 142(4):989–1004, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Poytress:2015:STD

[PGVG15] William R. Poytress, Joshua J. Gruber, Joel P. Van Eenennaam, and Mark Gard. Spatial and temporal distribution of spawning events and habitat characteristics of Sacramento River green sturgeon. Transactions of the American Fisheries Society, 144(6):1129–1142, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2011:SSP

[PGW^{+}11] Landon L. Pierce, Brian D. S. Graeb, David W. Willis, Jason S. Sorensen, and Mark A. Pegg. Stocking success of paddlefish in Lake Francis case, South Dakota: Population characteristics and sport fishery potential. Transactions of the American Fisheries Society, 140(5):1359–1369, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2015:EEE

[PGWS15] Landon L. Pierce, Brian D. S. Graeb, David W. Willis, and Jason S. Sorensen. Evaluating effects of exploitation on annual apparent mortality rates of paddlefish using mark–recapture data. Transactions of the American Fisheries Society, 144(2): 337–344, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Preston:2017:USM

[PHFN17] Daniel L. Preston, Jeremy S. Henderson, Landon P. Falke, and Mark Novak. Using survival models to estimate invertebrate prey identification times in a generalist stream fish. Transactions of the American Fisheries Society, 146(6):1303–1314, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2016:CSE

[PHG^{+}16] Mark S. Peterson, Jeanne-Marie Havrylkoff, Paul O. Grammer, Paul F. Mickle, and William T. Slack. Consistent SpatioTemporal estuarine habitat use during emigration or immigration of a western population of Gulf sturgeon. Transactions of the American Fisheries Society, 145(1):27–43, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petty:2012:BTM

[PHHM12] J. Todd Petty, Jeff L. Hansbarger, Brock M. Huntsman, and Patricia M. Mazik. Brook trout movement in response to temperature, flow, and thermal refugia within a complex Appalachian Riverscape. Transactions of the American Fisheries Society, 141(4):1060–1073, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2017:DHP

[PHM17] Matthew L. Peterson, David G. Hankin, and Kaitlyn Manishin. Decline in the half-pounder life history among Trinity River, California, steelhead. Transactions of the American Fisheries Society, 146(6):1245–1261, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perkins:2011:FBD

[PJ11] T. Alex Perkins and Henrïette I. Jager. Falling behind: Delayed growth explains life-history variation in Snake River fall Chinook salmon. Transactions of the American Fisheries Society, 140(4):959–972, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pennington:2011:PFL

[PK11] Kelly M. Pennington and Anne R. Kapuscinski. Predation and food limitation influence fitness traits of growth-enhanced transgenic and wild-type fish. Transactions of the American Fisheries Society, 140(2):221–234, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pess:2011:IBS

[PKL^{+}11] G. R. Pess, P. M. Kiffney, M. C. Liermann, T. R. Bennett, J. H. Anderson, and T. P. Quinn. The influences of body size, habitat quality, and competition on the movement and survival of juvenile Coho salmon during the early stages of stream recolonization. Transactions of the American Fisheries Society, 140(4):883–897, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paller:2017:DIB

[PKP^{+}17] Michael H. Paller, Ely Kosnicki, Blair A. Prusha, Dean E. Fletcher, Stephen A. Sefick, and Jack W. Feminella. Development of an index of biotic integrity for the sand hills ecoregion of the Southeastern United States. Transactions of the American Fisheries Society, 146(1):112–127, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pribyl:2011:RFD

[PKPS11] Alena L. Pribyl, Michael L. Kent, Steven J. Parker, and Carl B. Schreck. The response to forced decompression in six species of Pacific rockfish. Transactions of the American Fisheries Society, 140(2):374–383, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pratt:2013:HUS

[PL13] Anne E. Pratt and Thomas E. Lauer. Habitat use and separation among congeneric darter species. Transactions of the American Fisheries Society, 142(2):568–577, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pracheil:2010:PDL

[PM10] Brenda M. Pracheil and Patrick M. Muzzall. Population dynamics of larval trematodes in juvenile bluegills from Three Lakes II, Michigan, and the potential for overwinter parasiteinduced host mortality. Transactions of the American Fisheries Society, 139(3):652–659, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perez:2013:VBC

[PM13] K. O. Perez and S. B. Munch. Validating back-calculation models using population data. Transactions of the American Fisheries Society, 142(1):82–94, January 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Penney:2014:PCE

[PM14] Zachary L. Penney and Christine M. Moffitt. Proximate composition and energy density of stream-maturing adult steelhead during upstream migration, sexual maturity, and kelt emigration. Transactions of the American Fisheries Society, 143(2):399–413, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Plumb:2015:RET

[PM15] John M. Plumb and Christine M. Moffitt. Re-estimating temperature-dependent consumption parameters in bioenergetics models for juvenile Chinook salmon. Transactions of the American Fisheries Society, 144(2):323–330, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peng:2019:PPW

[PM19] Dao-Min Peng and Yong-Tong Mu. Predicting the production of the World’s cephalopod fisheries by means of differences in level of development and production trends. Transactions of the American Fisheries Society, 148(2):260–270, March 2019.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pautzke:2010:SUN

[PMF^{+}10] Sarah M. Pautzke, Martha E. Mather, John T. Finn, Linda A. Deegan, and Robert M. Muth. Seasonal use of a New England estuary by foraging contingents of migratory striped bass. Transactions of the American Fisheries Society, 139(1):257– 269, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Polansky:2019:UMD

[PMN19] Leo Polansky, Lara Mitchell, and Ken B. Newman. Using multistage design-based methods to construct abundance indices and uncertainty measures for delta smelt. Transactions of the American Fisheries Society, 148(4):710–724, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pavey:2010:CES

[PNM^{+}10] Scott A. Pavey, Jennifer L. Nielsen, Renae H. MacKas, Troy R. Hamon, and Felix Breden. Contrasting ecology shapes juvenile lake-type and riverine sockeye salmon. Transactions of the American Fisheries Society, 139(5):1584–1594, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prchalova:2012:CGN

[PNMH^{+}12] Marie Prchalová, J. Wesley Neal, Milton Muñoz-Hincapié, Tomáš Jůza, Milan Říha, Jiří Peterka, and Jan Kubečka. Comparison of gill nets and fixed-frame trawls for sampling threadfin shad in tropical reservoirs. Transactions of the American Fisheries Society, 141(4):1151–1160, July 2012. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Poesch:2014:DSM

[Poe14] Mark S. Poesch. Developing standardized methods for sampling freshwater fishes with multiple gears: Effects of sampling order versus sampling method. Transactions of the American Fisheries Society, 143(2):353–362, March 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2013:RWT

[PPC13] Ron Pierce, Craig Podner, and Kellie Carim. Response of wild trout to stream restoration over two decades in the Blackfoot River Basin, Montana. Transactions of the American Fisheries Society, 142(1):68–81, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Puchala:2016:PSE

[PPD16] Elizabeth A. Puchala, Donna L. Parrish, and Therese M. Donovan. Predicting the stability of endangered stonecats in the LaPlatte River, Vermont. Transactions of the American Fisheries Society, 145(4):903–912, July 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perry:2013:CET

[PPF^{+}13] Russell W. Perry, John M. Plumb, Scott D. Fielding, Noah S. Adams, and Dennis W. Rondorf. Comparing effects of transmitters within and among populations: Application to swimming performance of juvenile Chinook salmon. Transactions of the American Fisheries Society, 142(4):901–911, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paller:2016:FIS

[PPF^{+}16] Michael H. Paller, Blair A. Prusha, Dean E. Fletcher, Ely Kosnicki, Stephen A. Sefick, Miller S. Jarrell, Sean C. Sterrett, Andrew M. Grosse, Tracey D. Tuberville, and Jack W. Feminella. Factors influencing stream fish species composition and functional properties at multiple spatial scales in the sand hills of the Southeastern United States. Transactions of the American Fisheries Society, 145(3):545–562, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perry:2015:ULB

[PPH15] Russell W. Perry, John M. Plumb, and Charles W. Huntington. Using a laboratory-based growth model to estimate massand temperature-dependent growth parameters across populations of juvenile Chinook salmon. Transactions of the American Fisheries Society, 144(2):331–336, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2015:LTI

[PPJ15] Ron Pierce, Craig Podner, and Leslie Jones. Long-term increases in trout abundance following channel reconstruction, instream wood placement, and livestock removal from a spring creek in the Blackfoot Basin, Montana. Transactions of the American Fisheries Society, 144(1):184–195, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2014:IHR

[PPMJ14] Ron Pierce, Craig Podner, Laurie Marczak, and Leslie Jones. Instream habitat restoration and stream temperature reduction in a whirling disease-positive spring creek in the Blackfoot River Basin, Montana. Transactions of the American Fisheries Society, 143(5):1188–1198, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pinto:2013:LLD

[PPP^{+}13] Madalena C. Pinto, John R. Post, Andrew J. Paul, Fiona D. Johnston, Craig J. Mushens, and Jim D. Stelfox. Lateral and longitudinal displacement of stream-rearing juvenile bull trout in response to upstream migration of spawning adults. Transactions of the American Fisheries Society, 142(6):1590–1601, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2014:WCT

[PPW^{+}14] Ron Pierce, Craig Podner, Tracy Wendt, Ron Shields, and Kellie Carim. Westslope cutthroat trout movements through restored habitat and Coanda Diversions in the Nevada Spring Creek Complex, Blackfoot Basin, Montana. Transactions of the American Fisheries Society, 143(1):230–239, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pratt:2014:RBS

[Pra14] Harold “Wes” Pratt. A review of “The Biology of Sharks and Rays”. By A. Peter Klimley. University of Chicago Press, Chicago, Illinois. 2013. 512 pages. $40.00 (hardcover). Transactions of the American Fisheries Society, 143(6):1612, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prichard:2016:LSC

[PRK^{+}16] Carson G. Prichard, Edward F. Roseman, Kevin M. Keeler, Timothy P. O’Brien, and Stephen C. Riley. Large-scale changes in bloater growth and condition in Lake Huron. Transactions of the American Fisheries Society, 145(6):1241–1251, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peacock:2010:ESU

[PRW^{+}10] Mary M. Peacock, Morgan L. Robinson, Timothy Walters, Heather A. Mathewson, and Ray Perkins. The evolutionarily significant unit concept and the role of translocated populations in preserving the genetic legacy of Lahontan cutthroat trout. Transactions of the American Fisheries Society, 139 (2):382–395, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Piraino:2014:FSM

[PS14] Maria N. Piraino and Stephen T. Szedlmayer. Fine-scale movements and home ranges of red snapper around artificial reefs in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 143(4):988–998, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Potoka:2016:MOM

[PSB16] Kathryn M. Potoka, Colin P. Shea, and Phillip W. Bettoli. Multispecies occupancy modeling as a tool for evaluating the status and distribution of darters in the Elk River, Tennessee. Transactions of the American Fisheries Society, 145(5):1110– 1121, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perkin:2013:FDL

[PSGB13] Joshuah S. Perkin, Zachary R. Shattuck, Joseph E. Gerken, and Timothy H. Bonner. Fragmentation and drought legacy correlate with distribution of burrhead chub in subtropical streams of North America. Transactions of the American Fisheries Society, 142(5):1287–1298, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pendleton:2019:ATB

[PSH^{+}19] Richard M. Pendleton, Christopher R. Standley, Amanda L. Higgs, Gregg H. Kenney, Patrick J. Sullivan, Suresh A. Sethi, and Bradley P. Harris. Acoustic telemetry and benthic habitat mapping inform the spatial ecology of shortnose sturgeon in the Hudson River, New York, USA. Transactions of the American Fisheries Society, 148(1):35–47, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peer:2012:SMA

[PSM12] A. C. Peer, G. M. Selckmann, and T. J. Miller. A standardized method and analytical approach for predicting female reproductive stage in teleosts by using ovary color and female characteristics. Transactions of the American Fisheries Society, 141(4):1036–1044, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearsons:2010:CRT

[PT10] Todd N. Pearsons and Gabriel M. Temple. Changes to rainbow trout abundance and salmonid biomass in a Washington watershed as related to hatchery salmon supplementation. Transactions of the American Fisheries Society, 139(2):502– 520, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phelps:2010:HUD

[PTG^{+}10] Quinton E. Phelps, Sara J. Tripp, James E. Garvey, David P. Herzog, David E. Ostendorf, Joseph W. Ridings, Jason W. Crites, and Robert A. Hrabik. Habitat use during early life history infers recovery needs for shovelnose sturgeon and pallid sturgeon in the Middle Mississippi River. Transactions of the American Fisheries Society, 139(4):1060–1068, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2010:DCA

[PTPS10] Rodney B. Pierce, Cynthia M. Tomcko, Donald L. Pereira, and David F. Staples. Differing catchability among lakes: Influences of lake basin morphology and other factors on gill-net catchability of northern pike. Transactions of the American Fisheries Society, 139(4):1109–1120, July 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pease:2011:MEI

[PTWK11] Allison A. Pease, Jason M. Taylor, Kirk O. Winemiller, and Ryan S. King. Multiscale environmental influences on fish assemblage structure in central Texas streams. Transactions of the American Fisheries Society, 140(5):1409–1427, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pregler:2015:UMO

[PVJH15] Kasey C. Pregler, Jason C. Vokoun, Timothy Jensen, and Neal Hagstrom. Using multimethod occupancy estimation models to quantify gear differences in detection probabilities: Is backpack electrofishing missing occurrences for a species of concern? Transactions of the American Fisheries Society, 144 (1):89–95, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parkos:2010:IIV

[PW10a] Joseph J. Parkos III and D. H. Wahl. Intra- and intersystem variation in largemouth bass recruitment: Reproduction, prey availability, and the timing of year-class establishment. Transactions of the American Fisheries Society, 139(5):1372–1385, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parkos:2010:IBS

[PW10b] Joseph J. Parkos III and David H. Wahl. Influence of body size and prey type on the willingness of age-0 fish to forage under predation risk. Transactions of the American Fisheries Society, 139(4):969–975, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parkos:2013:CAP

[PWP13] Joseph J. Parkos III, D. H. Wahl, and D. P. Philipp. Comment: Assessing the potential for the angling of nesting males to influence largemouth bass recruitment should consider the quality of nesting individuals. Transactions of the American Fisheries Society, 142(2):416–418, March 2013. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [ARC^{+}13].

Page:2017:STG

[PZS17] Kevin S. Page, Richard D. Zweifel, and Wendylee Stott. Spatial and temporal genetic analysis of walleyes in the Ohio River. Transactions of the American Fisheries Society, 146 (6):1168–1185, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2013:JCS

[QHS^{+}13] Thomas P. Quinn, Nicole Harris, J. Anne Shaffer, Chris Byrnes, and Patrick Crain. Juvenile Coho salmon in the Elwha River Estuary prior to dam removal: Seasonal occupancy, size distribution, and comparison to nearby Salt Creek. Transactions of the American Fisheries Society, 142(4):1058–1066, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2012:SIC

[QSJ12] Thomas P. Quinn, Todd R. Seamons, and Susan P. Johnson. Stable isotopes of carbon and nitrogen indicate differences in marine ecology between wild and hatchery-produced steelhead. Transactions of the American Fisheries Society, 141 (2):526–532, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2011:EGR

[QSVD11] Thomas P. Quinn, Todd R. Seamons, Leif Asbjørn Vøllestad, and Elisabeth Duffy. Effects of growth and reproductive history on the egg size–fecundity trade-off in steelhead. Transactions of the American Fisheries Society, 140(1):45–51, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2016:BRM

[Qui16] Thomas P. Quinn. Book review: Migration Ecology of Marine Fishes. By David Hallock Secor: Johns Hopkins University Press, Baltimore, Maryland. 2015. 292 pages. $99.95. Transactions of the American Fisheries Society, 145(5):1152, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2019:OEP

[Qui19] Thomas P. Quinn. The ocean ecology of Pacific salmon and trout. Transactions of the American Fisheries Society, 148(1): 232, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riley:2010:LTT

[RA10] Stephen C. Riley and Jean V. Adams. Long-term trends in habitat use of offshore demersal fishes in Western Lake Huron suggest large-scale ecosystem change. Transactions of the American Fisheries Society, 139(5):1322–1334, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:2018:DFS

[RAB^{+}18] Clinton R. Robertson, Karim Aziz, David L. Buckmeier, Nathan G. Smith, and Nolan Raphelt. Development of a flowspecific floodplain inundation model to assess alligator gar recruitment success. Transactions of the American Fisheries Society, 147(4):674–686, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rutherford:2016:DSW

[RAR^{+}16] E. S. Rutherford, J. Allison, C. R. Ruetz III, J. R. Elliott, J. K. Nohner, M. R. DuFour, R. P. O’Neal, D. J. Jude, and S. R. Hensler. Density and survival of walleye eggs and larvae in a Great Lakes tributary. Transactions of the American Fisheries Society, 145(3):563–577, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raabe:2012:QSH

[RB12] Joshua K. Raabe and Michael A. Bozek. Quantity, structure, and habitat selection of natural spawning reefs by walleyes in a North temperate lake: a multiscale analysis. Transactions of the American Fisheries Society, 141(4):1097–1108, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Recsetar:2013:SAT

[RB13] Matthew S. Recsetar and Scott A. Bonar. Survival of Apache trout eggs and alevins under static and fluctuating temperature regimes. Transactions of the American Fisheries Society, 142(2):373–379, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rohan:2018:TNS

[RB18] Sean K. Rohan and Troy W. Buckley. Trophic niche separation between sympatric sibling flatfishes in relation to gill raker morphology. Transactions of the American Fisheries Society, 147(3):431–443, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rudershausen:2010:FEB

[RBE^{+}10] Paul J. Rudershausen, Jeffrey A. Buckel, Jason Edwards, Damon P. Gannon, Christopher M. Butler, and Tyler W. Averett. Feeding ecology of blue marlins, dolphinfish, yellowfin tuna, and wahoos from the North Atlantic Ocean and comparisons with other oceans. Transactions of the American Fisheries Society, 139(5):1335–1359, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberts:2011:EHC

[RBF^{+}11] James J. Roberts, Stephen B. Brandt, David Fanslow, Stuart A. Ludsin, Steven A. Pothoven, Donald Scavia, and Tomas O. Höök. Effects of hypoxia on consumption, growth, and RNA:DNA ratios of young yellow perch. Transactions of the American Fisheries Society, 140(6):1574–1586, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Recsetar:2014:GSA

[RBF14] Matthew S. Recsetar, Scott A. Bonar, and Olin G. Feuerbacher. Growth and survival of Apache trout under static and fluctuating temperature regimes. Transactions of the American Fisheries Society, 143(5):1247–1254, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roni:2012:FAM

[RBH^{+}12] Philip Roni, Todd Bennett, Ranae Holland, George Pess, Karrie Hanson, Raymond Moses, Mike McHenry, William Ehinger, and Jason Walter. Factors affecting migration timing, growth, and survival of juvenile Coho salmon in two coastal Washington watersheds. Transactions of the American Fisheries Society, 141(4):890–906, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rudstam:2011:ACR

[RBK^{+}11] L. G. Rudstam, T. E. Brooking, S. D. Krueger, J. R. Jackson, and L. Wetherbee. Analysis of compensatory responses in land-locked alewives to walleye predation: a tale of two lakes. Transactions of the American Fisheries Society, 140(6):1587– 1603, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reinhold:2017:SDR

[RBZ^{+}17] Ann Marie Reinhold, Robert G. Bramblett, Alexander V. Zale, Geoffrey C. Poole, and David W. Roberts. Spatially dependent responses of a large-river fish assemblage to bank stabilization and side channels. Transactions of the American Fisheries Society, 146(5):967–982, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raposa:2016:CBL

[RC16] Kenneth B. Raposa and Marnita Chintala. Comparison of bottomless lift nets and breder traps for sampling salt-Marsh nekton. Transactions of the American Fisheries Society, 145 (1):163–172, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raby:2013:RPS

[RCC^{+}13] Graham D. Raby, Steven J. Cooke, Katrina V. Cook, Sarah H. McConnachie, Michael R. Donaldson, Scott G. Hinch, Charlotte K. Whitney, S. Matthew Drenner, David A. Patterson, Timothy D. Clark, and Anthony P. Farrell. Resilience of pink salmon and chum salmon to simulated fisheries capture stress incurred upon arrival at spawning grounds. Transactions of the American Fisheries Society, 142(2):524–539, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Redman:2011:OTR

[RCD^{+}11] Rebecca A. Redman, Sergiusz J. Czesny, John M. Dettmers, Michael J. Weber, and Daniel Makauskas. Old tales in recent context: Current perspective on yellow perch recruitment in Lake Michigan. Transactions of the American Fisheries Society, 140(5):1277–1289, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reeves:2019:TES

[RCM^{+}19] David B. Reeves, Edward J. Chesney, Ryan T. Munnelly, Donald M. Baltz, and Kanchan Maiti. Trophic ecology of sheepshead and stone crabs at oil and gas platforms in the Northern Gulf of Mexico’s hypoxic zone. Transactions of the American Fisheries Society, 148(2):324–338, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ridenour:2011:HAS

[RDH11] Clayton J. Ridenour, Wyatt J. Doyle, and Tracy D. Hill. Habitats of age-0 sturgeon in the Lower Missouri River. Transactions of the American Fisheries Society, 140(5):1351–1358, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roegner:2010:JSU

[RDR^{+}10] G. Curtis Roegner, Earl W. Dawley, Micah Russell, Allan Whiting, and David J. Teel. Juvenile salmonid use of reconnected tidal freshwater wetlands in Grays River, Lower Columbia River Basin. Transactions of the American Fisheries Society, 139(4):1211–1232, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robichaud:2017:AMA

[REN17] David Robichaud, Karl K. English, and Troy C. Nelson. Annual movements of acoustic-tagged white sturgeon in the lower reaches of the Fraser River and its tributaries. Transactions of the American Fisheries Society, 146(4):611–625, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reynolds:2016:SRR

[Rey16] James B. Reynolds. Spheres, rings, and rods as electrodes in electrofishing: Their effects on system resistance and electrical fields. Transactions of the American Fisheries Society, 145 (2):239–248, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rous:2015:OPF

[RFM^{+}15] Andrew M. Rous, Alexandra Forrest, Elisabeth Hart McKittrick, George Letterio, Jordan Roszell, Taylor Wright, and Steven J. Cooke. Orientation and position of fish affects recovery time from electrosedation. Transactions of the American Fisheries Society, 144(4):820–828, July 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:2015:DPL

[RG15] Stewart B. Reid and Damon H. Goodman. Detectability of Pacific lamprey occupancy in western drainages: Implications for distribution surveys. Transactions of the American Fisheries Society, 144(2):315–322, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:2016:PLC

[RG16] Stewart B. Reid and Damon H. Goodman. Pacific lamprey in coastal drainages of California: Occupancy patterns and contraction of the southern range. Transactions of the American Fisheries Society, 145(4):703–711, July 2016. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rice:2011:ASO

[RGM^{+}11] Casimir A. Rice, Correigh M. Greene, Paul Moran, David J. Teel, David R. Kuligowski, Reginald R. Reisenbichler, Eric M. Beamer, James R. Karr, and Kurt L. Fresh. Abundance, stock origin, and length of marked and unmarked juvenile Chinook salmon in the surface waters of Greater Puget Sound. Transactions of the American Fisheries Society, 140(1):170– 189, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rountos:2017:ODS

[RGP17] Konstantine J. Rountos, Christopher J. Gobler, and Ellen K. Pikitch. Ontogenetic differences in swimming behavior of fish exposed to the harmful dinoflagellate Cochlodinium polykrikoides. Transactions of the American Fisheries Society, 146(5):1081–1091, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raabe:2014:ASM

[RH14] Joshua K. Raabe and Joseph E. Hightower. American shad migratory behavior, weight loss, survival, and abundance in a North Carolina River following dam removals. Transactions of the American Fisheries Society, 143(3):673–688, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raabe:2019:EFP

[RHEF19] Joshua K. Raabe, Joseph E. Hightower, Timothy A. Ellis, and Joseph J. Facendola. Evaluation of fish passage at a naturelike rock ramp fishway on a large coastal river. Transactions of the American Fisheries Society, 148(4):798–816, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ray:2012:GAR

[RHKD12] Jesse W. Ray, Martin Husemann, Ryan S. King, and Patrick D. Danley. Genetic analysis reveals dispersal of Florida bass haplotypes from reservoirs to rivers in central Texas. Transactions of the American Fisheries Society, 141(5):1269– 1273, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robinson:2015:IPV

[RHR^{+}15] Kendra A. Robinson, Scott G. Hinch, Graham D. Raby, Michael R. Donaldson, Dave Robichaud, David A. Patterson, and Steven J. Cooke. Influence of postcapture ventilation assistance on migration success of adult sockeye salmon following capture and release. Transactions of the American Fisheries Society, 144(4):693–704, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Railsback:2015:ESE

[RHW15] Steven F. Railsback, Bret C. Harvey, and Jason L. White. Effects of spatial extent on modeled relations between habitat and anadromous salmonid spawning success. Transactions of the American Fisheries Society, 144(6):1220–1236, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rothschild:2011:CUF

[RJ11] Brian J. Rothschild and Yue Jiao. Characterizing uncertainty in fish stock assessments: the case of the Southern New England–Mid-Atlantic winter flounder. Transactions of the American Fisheries Society, 140(3):557–569, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rothschild:2012:SSB

[RJH12] Brian J. Rothschild, Yue Jiao, and Saang-Yoon Hyun. Simulation study of biological reference points for summer flounder. Transactions of the American Fisheries Society, 141(2):426– 436, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robinson:2010:IVI

[RJWK10] Jason M. Robinson, Daniel C. Josephson, Brian C. Weidel, and Clifford E. Kraft. Influence of variable interannual summer water temperatures on brook trout growth, consumption, reproduction, and mortality in an unstratified Adirondack lake. Transactions of the American Fisheries Society, 139(3):685–699, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:2013:IBMa

[RKEB13a] Kenneth A. Rose, Wim J. Kimmerer, Karen P. Edwards, and William A. Bennett. Individual-based modeling of delta smelt population dynamics in the Upper San Francisco Estuary: I. Model description and baseline results. Transactions of the American Fisheries Society, 142(5):1238–1259, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:2013:IBMb

[RKEB13b] Kenneth A. Rose, Wim J. Kimmerer, Karen P. Edwards, and William A. Bennett. Individual-based modeling of delta smelt population dynamics in the Upper San Francisco Estuary: II. Alternative baselines and good versus bad years. Transactions of the American Fisheries Society, 142(5):1260–1272, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rohde:2013:PMD

[RKF^{+}13] Jessica Rohde, Anna N. Kagley, Kurt L. Fresh, Fred A. Goetz, and Thomas P. Quinn. Partial migration and diel movement patterns in Puget Sound Coho salmon. Transactions of the American Fisheries Society, 142(6):1615–1628, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ralph:2014:CHS

[RL14] Gina M. Ralph and Romuald N. Lipcius. Critical habitats and stock assessment: Age-specific bias in the Chesapeake Bay blue crab population survey. Transactions of the American Fisheries Society, 143(4):889–898, July 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rundio:2019:DVS

[RL19] David E. Rundio and Steven T. Lindley. Diet variability of steelhead/rainbow trout in a coastal basin in Central California: Relative importance of seasonal, spatial, and ontogenetic variation. Transactions of the American Fisheries Society, 148 (1):88–105, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rudershausen:2019:SHY

[RLL^{+}19] Paul J. Rudershausen, Laura M. Lee, Steven M. Lombardo, Jeffery H. Merrell, and Jeffrey A. Buckel. Survival and habitat of yellow-phase American eels in North Carolina tidal creeks. Transactions of the American Fisheries Society, 148(5):978– 990, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reeves:2017:SCM

[RMC^{+}17] David B. Reeves, Ryan T. Munnelly, Edward J. Chesney, Donald M. Baltz, and Brian D. Marx. Stone crab Menippe spp. populations on Louisiana’s nearshore oil and gas platforms: Higher density and size at maturity on a sand shoal. Transactions of the American Fisheries Society, 146(3):371–383, May 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:2011:ATM

[RMDA11] R. Anne Richards, Joshua Moser, Bridget Dunnigan, and Larry A. Alade. Archival tagging methods for monkfish. Transactions of the American Fisheries Society, 140(3):582– 586, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robillard:2011:DHU

[RMM11] Melissa M. Robillard, Robert L. McLaughlin, and Robert W. Mackereth. Diversity in habitat use and trophic ecology of brook trout in Lake Superior and tributary streams revealed through stable isotopes. Transactions of the American Fisheries Society, 140(4):943–953, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Renan:2016:ASA

[RMMGPB16] Ximena Renán, Jorge Montero-Muñoz, Joaquin R. Garza-Pérez, and Thierry Brulé. Age and stock analysis using otolith shape in gags from the Southern Gulf of Mexico. Transactions of the American Fisheries Society, 145(6):1252–1265, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodger:2016:PFR

[RMW16] Anthony W. Rodger, Kevin B. Mayes, and Kirk O. Winemiller. Preliminary findings for a relationship between instream flow and shoal chub recruitment in the Lower Brazos River, Texas. Transactions of the American Fisheries Society, 145(5):943–950, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rhody:2013:ARC

[RNG^{+}13] N. R. Rhody, C. L. Neidig, H. J. Grier, K. L. Main, and H. Migaud. Assessing reproductive condition in captive and wild common snook stocks: a comparison between the wet mount technique and histological preparations. Transactions of the American Fisheries Society, 142(4):979–988, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rubenson:2016:SSP

[RO16] Erika S. Rubenson and Julian D. Olden. Spatiotemporal spawning patterns of smallmouth bass at its upstream invasion edge. Transactions of the American Fisheries Society, 145 (4):693–702, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reilly:2014:PGA

[RPC14] Jessica R. Reilly, Cynthia A. Paszkowski, and David W. Coltman. Population genetics of Arctic grayling distributed across large, unobstructed river systems. Transactions of the American Fisheries Society, 143(3):802–816, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ray:2014:UCM

[RPSB14] R. Adam Ray, Russell W. Perry, Nicholas A. Som, and Jerri L. Bartholomew. Using cure models for analyzing the influence of pathogens on salmon survival. Transactions of the American Fisheries Society, 143(2):387–398, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robillard:2015:BPS

[RPVS15] Megan M. Reese Robillard, Laura M. Payne, Robert R. Vega, and Gregory W. Stunz. Best practices for surgically implanting acoustic transmitters in spotted seatrout. Transactions of the American Fisheries Society, 144(1):81–88, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richard:2013:WBT

[RR13] Jordan C. Richard and Andrew L. Rypel. Water body type influences climate-growth relationships of freshwater drum. Transactions of the American Fisheries Society, 142(5):1308– 1320, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rebenack:2015:EEJ

[RRA^{+}15] Jennifer J. Rebenack, Seth Ricker, Colin Anderson, Michael Wallace, and Darren M. Ward. Early emigration of juvenile Coho salmon: Implications for population monitoring. Transactions of the American Fisheries Society, 144(1):163–172, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosset:2017:TPM

[RRG^{+}17] Julianne Rosset, Allison H. Roy, Benjamin I. Gahagan, Andrew R. Whiteley, Michael P. Armstrong, John J. Sheppard, and Adrian Jordaan. Temporal patterns of migration and spawning of river herring in coastal Massachusetts. Transactions of the American Fisheries Society, 146(6):1101–1114, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawding:2014:GBE

[RSB14] Daniel J. Rawding, Cameron S. Sharpe, and Scott M. Blankenship. Genetic-based estimates of adult Chinook salmon spawner abundance from carcass surveys and juvenile outmigrant traps. Transactions of the American Fisheries Society, 143(1):55–67, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roseman:2014:ACP

[RSBF14] Edward F. Roseman, Jeffrey S. Schaeffer, Ethan Bright, and David G. Fielder. Angler-caught piscivore diets reflect fish community changes in Lake Huron. Transactions of the American Fisheries Society, 143(6):1419–1433, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Renkawitz:2012:SDB

[RSG12] Mark D. Renkawitz, Timothy F. Sheehan, and Graham S. Goulette. Swimming depth, behavior, and survival of Atlantic salmon postsmolts in Penobscot Bay, Maine. Transactions of the American Fisheries Society, 141(5):1219–1229, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reeves:2011:SCH

[RSL11] Gordon H. Reeves, Jack D. Sleeper, and Dirk W. Lang. Seasonal changes in habitat availability and the distribution and abundance of salmonids along a stream gradient from headwaters to mouth in coastal Oregon. Transactions of the American Fisheries Society, 140(3):537–548, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rincon-Sandoval:2019:FAC

[RSVAB19] Luis A. Rincón-Sandoval, Iván Velázquez-Abunader, and Thierry Brulé. Factors affecting catchability in longline fishing of red grouper in the Southeastern Gulf of Mexico. Transactions of the American Fisheries Society, 148(4):857–868, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roegner:2014:DCS

[RT14] G. Curtis Roegner and David J. Teel. Density and condition of subyearling Chinook salmon in the Lower Columbia River and estuary in relation to water temperature and genetic stock of origin. Transactions of the American Fisheries Society, 143(5): 1161–1176, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roach:2011:DTA

[RW11] Katherine A. Roach and Kirk O. Winemiller. Diel turnover of assemblages of fish and shrimp on sandbanks in a temperate floodplain river. Transactions of the American Fisheries Society, 140(1):84–90, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rypel:2013:RFT

[Ryp13] Andrew L. Rypel. A review of “Fisheries Techniques, third edition.” Edited by A. V. Zale, D. L. Parrish, and T. M. Sutton. American Fisheries Society, Bethesda, Maryland. 2012. 1,009 pages. $104.00 ($72.80 for American Fisheries Society members). Transactions of the American Fisheries Society, 142(4):1153, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Recsetar:2012:RBF

[RZW^{+}12] Matthew S. Recsetar, Matthew P. Zeigler, David L. Ward, Scott A. Bonar, and Colleen A. Caldwell. Relationship between fish size and upper thermal tolerance. Transactions of the American Fisheries Society, 141(6):1433–1438, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaw:2014:LST

[SA14] S. L. Shaw and M. S. Allen. Localized spatial and temporal variation in reproductive effort of Florida bass. Transactions of the American Fisheries Society, 143(1):85–96, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaw:2016:QRC

[SA16] Stephanie L. Shaw and Micheal S. Allen. Quantifying recruitment compensation in Florida largemouth bass, with implications for fisheries. Transactions of the American Fisheries Society, 145(3):462–475, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sbragaglia:2019:ESS

[SAF^{+}19] Valerio Sbragaglia, Josep Alós, Kim Fromm, Christopher T. Monk, Carlos Díaz-Gil, Silva Uusi-Heikkilä, Andrew E. Honsey, Alexander D. M. Wilson, and Robert Arlinghaus. Experimental size-selective harvesting affects behavioral types of a social fish. Transactions of the American Fisheries Society, 148(3):552–568, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranko:2014:FBM

[SAH^{+}14] Scott A. Stranko, Matthew J. Ashton, Robert H. Hilderbrand, Sara L. Weglein, David C. Kazyak, and Jay V. Kilian. Fish and benthic macroinvertebrate densities in small streams with and without American eels. Transactions of the American Fisheries Society, 143(3):700–708, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sackett:2013:VPM

[SARC13] Dana K. Sackett, D. Derek Aday, James A. Rice, and W. Gregory Cope. Validation of a predictive model for fish tissue mercury concentrations. Transactions of the American Fisheries Society, 142(2):380–387, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Slagle:2017:FIN

[SAS17] Z. Slagle, M. S. Allen, and S. L. Shaw. Factors influencing nest survival in Florida bass. Transactions of the American Fisheries Society, 146(4):696–702, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Streich:2017:CSS

[SAW^{+}17] Matthew K. Streich, Matthew J. Ajemian, Jennifer J. Wetz, Jason A. Williams, J. Brooke Shipley, and Gregory W. Stunz. A comparison of size structure, age, and growth of red snapper from artificial and natural habitats in the Western Gulf of Mexico. Transactions of the American Fisheries Society, 146 (4):762–777, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoenebeck:2012:DAA

[SB12a] Casey W. Schoenebeck and Michael L. Brown. Does anaerobic activity differ seasonally or between sexes in yellow perch populations? Transactions of the American Fisheries Society, 141(1):199–203, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simmons:2012:ADG

[SB12b] Bryan R. Simmons and Daniel W. Beckman. Age determination, growth, and population structure of the striped shiner and duskystripe shiner. Transactions of the American Fisheries Society, 141(3):846–854, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schreier:2016:EPE

[SBC^{+}16] Brian M. Schreier, Melinda R. Baerwald, J. Louise Conrad, Gregg Schumer, and Bernie May. Examination of predation on early life stage delta smelt in the San Francisco Estuary using DNA diet analysis. Transactions of the American Fisheries Society, 145(4):723–733, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strohm:2017:MWO

[SBC17] Deanna D. Strohm, Phaedra Budy, and Todd A. Crowl. Matching watershed and otolith chemistry to establish natal origin of an endangered desert lake sucker. Transactions of the American Fisheries Society, 146(4):732–743, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Small:2011:DLC

[SBDW11] Maureen P. Small, Dave Burgess, Cheryl Dean, and Kenneth I. Warheit. Does Lower Crab Creek in the Eastern Washington desert have a native population of Chinook salmon? Transactions of the American Fisheries Society, 140(3):808–821, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stepien:2012:TSG

[SBMH12] Carol A. Stepien, Jo Ann Banda, Douglas M. Murphy, and Amanda E. Haponski. Temporal and spatial genetic consistency of walleye spawning groups. Transactions of the American Fisheries Society, 141(3):660–672, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoen:2012:QLI

[SBO12] Erik R. Schoen, David A. Beauchamp, and Nathanael C. Overman. Quantifying latent impacts of an introduced piscivore: Pulsed predatory inertia of lake trout and decline of Kokanee. Transactions of the American Fisheries Society, 141(5):1191– 1206, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:2014:DLP

[SBW14] Daniel E. Shoup, Kevin M. Boswell, and David H. Wahl. Diel littoral–pelagic movements by juvenile bluegills in a small lake. Transactions of the American Fisheries Society, 143(3):796– 801, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Su:2013:ESD

[SC13] Zhenming Su and David Clapp. Evaluation of sample design and estimation methods for Great Lakes angler surveys. Transactions of the American Fisheries Society, 142(1):234– 246, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Switzer:2015:HUJ

[SCB15] Theodore S. Switzer, Edward J. Chesney, and Donald M. Baltz. Habitat use by juvenile red snapper in the Northern Gulf of Mexico: Ontogeny, seasonality, and the effects of hypoxia. Transactions of the American Fisheries Society, 144 (2):300–314, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Satterthwaite:2017:OSC

[SCC17] William H. Satterthwaite, Stephanie M. Carlson, and Anne Criss. Ocean size and corresponding life history diversity among the four run timings of California Central Valley Chinook salmon. Transactions of the American Fisheries Society, 146(4):594–610, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sloychuk:2016:JLS

[SCF16] Janelle R. Sloychuk, Douglas P. Chivers, and Maud C. O. Ferrari. Juvenile lake sturgeon go to school: Life-skills training for hatchery fish. Transactions of the American Fisheries Society, 145(2):287–294, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sebring:2013:RVP

[SCL^{+}13] Scott H. Sebring, Melissa C. Carper, Richard D. Ledgerwood, Benjamin P. Sandford, Gene M. Matthews, and Allen F. Evans. Relative vulnerability of PIT-Tagged subyearling fall Chinook salmon to predation by Caspian terns and doublecrested cormorants in the Columbia River estuary. Transactions of the American Fisheries Society, 142(5):1321–1334, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaw:2013:ISR

[SCW^{+}13] Stephanie L. Shaw, Steven R. Chipps, Steve K. Windels, Molly A. H. Webb, and Darryl T. McLeod. Influence of sex and reproductive status on seasonal movement of lake sturgeon in Namakan Reservoir, Minnesota–Ontario. Transactions of the American Fisheries Society, 142(1):10–20, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saul:2012:IBM

[SDBB12] Steven Saul, David Die, Elizabeth N. Brooks, and Karen Burns. An individual-based model of ontogenetic migration in reef fish using a biased random walk. Transactions of the American Fisheries Society, 141(6):1439–1452, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sato:2010:HIL

[SDK^{+}10] Takuya Sato, Teruko Demise, Hitoshi Kubota, Makoto Nagoshi, and Katsutoshi Watanabe. Hybridization, isolation, and low genetic diversity of kirikuchi char, the southernmost populations of the genus Salvelinus. Transactions of the American Fisheries Society, 139(6):1758–1774, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smiley:2011:AEG

[SDOK11] Jeffrey E. Smiley, Mark A. Drawbridge, Mark S. Okihiro, and Ronald S. Kaufmann. Acute effects of gas supersaturation on juvenile cultured white seabass. Transactions of the American Fisheries Society, 140(5):1269–1276, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shephard:2019:SCI

[SDOK19] Samuel Shephard, Karen Delanty, Martin O’Grady, and Fiona Kelly. Salmonid conservation in an invaded lake: Changing outcomes of predator removal with introduction of nonnative prey. Transactions of the American Fisheries Society, 148(1): 219–231, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2018:MRA

[SDSB18] Nathan G. Smith, Daniel J. Daugherty, J. Warren Schlechte, and David L. Buckmeier. Modeling the responses of alligator gar populations to harvest under various length-based regulations: Implications for conservation and management. Transactions of the American Fisheries Society, 147(4):665– 673, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simcox:2015:MCP

[SDW15] Brandon L. Simcox, Dennis R. DeVries, and Russell A. Wright. Migratory characteristics and passage of paddlefishat two Southeastern U.S. lock-and-dam systems. Transactions of the American Fisheries Society, 144(3):456–466, May 2015.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schumann:2017:EPI

[SDW^{+}17a] David A. Schumann, David Deslauriers, Matthew D. Wagner, Katie N. Bertrand, and Brian D. S. Graeb. Effects of passive integrated transponder tags on survival, growth, and swimming performance of age-0 shovelnose sturgeon. Transactions of the American Fisheries Society, 146(2):230–239, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sweka:2017:FWS

[SDW17b] John A. Sweka, Lori A. Davis, and Tyler Wagner. Fall and winter survival of brook trout and brown trout in a North-Central Pennsylvania watershed. Transactions of the American Fisheries Society, 146(4):744–752, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2011:PRI

[SE11] Christian T. Smith and Rod Engle. Persistent reproductive isolation between sympatric lineages of fall Chinook salmon in White Salmon River, Washington. Transactions of the American Fisheries Society, 140(3):699–715, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Secor:2019:SE

[Sec19] David H. Secor. Seascape ecology. Transactions of the American Fisheries Society, 148(4):871–872, July 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schloesser:2015:RAP

[SF15] Ryan W. Schloesser and Mary C. Fabrizio. Relationships among proximate components and energy density of juvenile Atlantic estuarine fishes. Transactions of the American Fisheries Society, 144(5):942–955, September 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schloesser:2017:CIS

[SF17] Ryan W. Schloesser and Mary C. Fabrizio. Condition indices as surrogates of energy density and lipid content in juveniles of three fish species. Transactions of the American Fisheries Society, 146(5):1058–1069, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaefer:2016:DNC

[SFB16] Jacob Schaefer, Nkrumah Frazier, and Jonathan Barr. Dynamics of near-coastal fish assemblages following the Deepwater Horizon oil spill in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 145(1):108–119, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stiansen:2010:HVO

[SFF^{+}10] Stian Stiansen, Anders Fernö, Dag Furevik, Terje Jørgensen, and Svein Løkkeborg. Horizontal and vertical odor plume trapping of red king crabs explains the different efficiency of topand side-entrance pot designs. Transactions of the American Fisheries Society, 139(2):483–490, March 2010. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Straight:2014:PAM

[SFF14] Carrie A. Straight, Byron J. Freeman, and Mary C. Freeman. Passive acoustic monitoring to detect spawning in large-bodied catostomids. Transactions of the American Fisheries Society, 143(3):595–605, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2015:FSA

[SFH15] Joseph A. Smith, H. Jared Flowers, and Joseph E. Hightower. Fall spawning of Atlantic sturgeon in the Roanoke River, North Carolina. Transactions of the American Fisheries Society, 144(1):48–54, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2014:GCW

[SFLH14] Christian T. Smith, Rod French, Jens Lovtang, and David Hand. Genetic composition of the Warm Springs River Chinook salmon population maintained following eight generations of hatchery production. Transactions of the American Fisheries Society, 143(5):1280–1294, September 2014. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevens:2013:FAO

[SGB13] Philip W. Stevens, Marin F. D. Greenwood, and David A. Blewett. Fish assemblages in the oligohaline stretch of a Southwest Florida river during periods of extreme freshwater inflow variation. Transactions of the American Fisheries Society, 142 (6):1644–1658, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Syslo:2016:FEN

[SGK16] John M. Syslo, Christopher S. Guy, and Todd M. Koel. Feeding ecology of native and nonnative salmonids during the expansion of a nonnative apex predator in Yellowstone Lake, Yellowstone National Park. Transactions of the American Fisheries Society, 145(3):476–492, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schrandt:2016:DPA

[SGP16] Meagan N. Schrandt, Kevan C. Gregalis, and Sean P. Powers. Differential predictors of adult versus juvenile distributions of two coastal pelagic, surf zone fishes along the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 145(6):1358–1373, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schueller:2010:SLS

[SH10] Amy M. Schueller and Daniel B. Hayes. Sensitivity of lake sturgeon population dynamics and genetics to demographic parameters. Transactions of the American Fisheries Society, 139(2):521–534, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2012:ELH

[SH12] Joseph A. Smith and Joseph E. Hightower. Effect of lowhead lock-and-dam structures on migration and spawning of American shad and striped bass in the Cape Fear River, North Carolina. Transactions of the American Fisheries Society, 141 (2):402–413, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sorel:2016:PNP

[SHC^{+}16] Mark H. Sorel, Adam G. Hansen, Kristin A. Connelly, Andrew C. Wilson, Erin D. Lowery, and David A. Beauchamp.

Predation by northern pikeminnow and tiger muskellunge on juvenile salmonids in a high-head reservoir: Implications for anadromous fish reintroductions. Transactions of the American Fisheries Society, 145(3):521–536, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sorel:2016:TFR

[SHCB16] Mark H. Sorel, Adam G. Hansen, Kristin A. Connelly, and David A. Beauchamp. Trophic feasibility of reintroducing anadromous salmonids in three reservoirs on the North Fork Lewis River, Washington: Prey supply and consumption demand of resident fishes. Transactions of the American Fisheries Society, 145(6):1331–1347, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Singer:2019:BRJ

[SHH^{+}19] Gabriel P. Singer, Matthew J. Hansen, Kristina V. Ho, Katie W. Lee, Dennis E. Cocherell, A. Peter Klimley, Andrew L. Rypel, and Nann A. Fangue. Behavioral response of juvenile Chinook salmon to surgical implantation of microacoustic transmitters. Transactions of the American Fisheries Society, 148(2):480–492, March 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shiffman:2013:RBS

[Shi13] David Shiffman. A review of “Biology of Sharks and Their Relatives” second edition. Edited by Jeffrey C. Carrier, John A. Musick, and Michael R. Heithaus. CRC Press, Boca Raton, Florida. 2012. 666 pages. $99.95. Transactions of the American Fisheries Society, 142(6):1773, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Satterthwaite:2012:SDM

[SHM^{+}12] William H. Satterthwaite, Sean A. Hayes, Joseph E. Merz, Susan M. Sogard, Danielle M. Frechette, and Marc Mangel. State-dependent migration timing and use of multiple habitat types in anadromous salmonids. Transactions of the American Fisheries Society, 141(3):781–794, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevenson:2019:ISA

[SHP^{+}19] Christine F. Stevenson, Scott G. Hinch, Aswea D. Porter, Erin L. Rechisky, David W. Welch, Stephen J. Healy, Andrew G. Lotto, and Nathan B. Furey. The influence of smolt age on freshwater and early marine behavior and survival of migrating juvenile sockeye salmon. Transactions of the American Fisheries Society, 148(3):636–651, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Slavik:2012:BTS

[SHR^{+}12] Ondřej Slavík, Pavel Horký, Tomáš Randák, Pavel Balvín, and Michal Bílý. Brown trout spawning migration in fragmented central European headwaters: Effect of isolation by artificial obstacles and the Moon phase. Transactions of the American Fisheries Society, 141(3):673–680, May 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snobl:2017:RSE

[SIKR17] Zachary R. Snobl, Daniel A. Isermann, Ryan P. Koenigs, and Joshua K. Raabe. Relative sampling efficiency and movements of subadult lake sturgeon in the Lower Wolf River, Wisconsin. Transactions of the American Fisheries Society, 146(5):1070– 1080, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schweizer:2011:MRV

[SJ11] Peter E. Schweizer and Henriette I. Jager. Modeling regional variation in riverine fish biodiversity in the Arkansas–White– Red River Basin. Transactions of the American Fisheries Society, 140(5):1227–1239, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swink:2014:GSS

[SJ14] William D. Swink and Nicholas S. Johnson. Growth and survival of sea lampreys from metamorphosis to spawning in Lake Huron. Transactions of the American Fisheries Society, 143 (2):380–386, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Straight:2015:DPT

[SJFF15] Carrie A. Straight, C. Rhett Jackson, Byron J. Freeman, and Mary C. Freeman. Diel patterns and temporal trends in spawning activities of robust redhorse and river redhorse in Georgia, assessed using passive acoustic monitoring. Transactions of the American Fisheries Society, 144(3):563–576, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schmucker:2016:GES

[SJGL16] Andrew K. Schmucker, Nicholas S. Johnson, Heather S. Galbraith, and Weiming Li. Glass-eel-stage American eels respond to conspecific odor as a function of concentration. Transactions of the American Fisheries Society, 145(4):712–722, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sitar:2014:ESS

[SJMG14] Shawn P. Sitar, Andrew J. Jasonowicz, Cheryl A. Murphy, and Frederick W. Goetz. Estimates of skipped spawning in lean and Siscowet lake trout in Southern Lake Superior: Implications for stock assessment. Transactions of the American Fisheries Society, 143(3):660–672, May 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sather:2016:STF

[SJT^{+}16] Nichole K. Sather, Gary E. Johnson, David J. Teel, Adam J. Storch, John R. Skalski, and Valerie I. Cullinan. Shallow tidal freshwater habitats of the Columbia River: Spatial and temporal variability of fish communities and density, size, and genetic stock composition of juvenile Chinook salmon. Transactions of the American Fisheries Society, 145(4):734–753, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoby:2011:HRS

[SK11] Gregory P. Schoby and Ernest R. Keeley. Home range size and foraging ecology of bull trout and westslope cutthroat trout in the Upper Salmon River Basin, Idaho. Transactions of the American Fisheries Society, 140(3):636–645, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stein:2018:CCA

[SKBS18] Jeffrey A. Stein, Sarah M. King, David L. Buckmeier, and Nathan G. Smith. Comment: The Challenge of Age Estimation in Gars Lepisosteus spp. Transactions of the American Fisheries Society, 147(4):649–652, July 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [BSSP18].

Schroder:2010:BBS

[SKP^{+}10] Steven L. Schroder, Curtis M. Knudsen, Todd N. Pearsons, Todd W. Kassler, Sewall F. Young, Edward P. Beall, and David E. Fast. Behavior and breeding success of wild and first-generation hatchery male spring Chinook salmon spawning in an artificial stream. Transactions of the American Fisheries Society, 139(4):989–1003, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stewart:2016:UHB

[SL16] David R. Stewart and James M. Long. Using hierarchical Bayesian multispecies mixture models to estimate tandem hoop-net-based habitat associations and detection probabilities of fishes in reservoirs. Transactions of the American Fisheries Society, 145(3):450–461, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spangenberg:2014:EVR

[SLG^{+}14] Dina Spangenberg, Donald A. Larsen, Ryan Gerstenberger, Chris Brun, and Brian R. Beckman. The effects of variation in rearing conditions on growth, smolt development, and minijack rate in yearling Chinook salmon: a hatchery scale experiment. Transactions of the American Fisheries Society, 143(5): 1220–1230, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simmons:2010:IHR

[SLM10] Rachel E. Simmons, Phil Lavretsky, and Bernie May. Introgressive hybridization of redband trout in the Upper McCloud River watershed. Transactions of the American Fisheries Society, 139(1):201–213, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siepker:2013:EIS

[SM13a] Michael J. Siepker and Paul H. Michaletz. Exploring the influence of stock-recruitment relationships and environmental variables on black bass and crappie recruitment dynamics in Missouri reservoirs. Transactions of the American Fisheries Society, 142(1):119–129, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stapanian:2013:ISS

[SM13b] Martin A. Stapanian and Charles P. Madenjian. Introduction to a special section: Ecology, culture, and management of burbot. Transactions of the American Fisheries Society, 142 (6):1659–1661, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sommer:2011:LTS

[SMH^{+}11] Ted Sommer, Francine Mejia, Kathryn Hieb, Randall Baxter, Erik Loboschefsky, and Frank Loge. Long-term shifts in the lateral distribution of age-0 striped bass in the San Francisco Estuary. Transactions of the American Fisheries Society, 140 (6):1451–1459, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2013:RAF

[Smi13] Tim D. Smith. A review of “All the Fish in the Sea: Maximum Sustainable Yield and the Failure of Fisheries Management”, by Carmel Finley. University of Chicago Press, Chicago. 2011. 224 pages. $35.00. Transactions of the American Fisheries Society, 142(3):885–886, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaffler:2014:STV

[SMJ14] Jason J. Schaffler, Thomas J. Miller, and Cynthia M. Jones. Spatial and temporal variation in otolith chemistry of juvenile Atlantic menhaden in the Chesapeake Bay. Transactions of the American Fisheries Society, 143(4):1061–1071, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schreier:2013:PPS

[SMM13] A. Drauch Schreier, B. Mahardja, and B. May. Patterns of population structure vary across the range of the white sturgeon. Transactions of the American Fisheries Society, 142(5): 1273–1286, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Switzer:2012:RJG

[SMMK12] Theodore S. Switzer, Timothy C. MacDonald, Robert H. McMichael, Jr., and Sean F. Keenan. Recruitment of juvenile gags in the Eastern Gulf of Mexico and factors contributing to observed spatial and temporal patterns of estuarine occupancy. Transactions of the American Fisheries Society, 141 (3):707–719, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steinhorst:2010:USB

[SMN^{+}10] Kirk Steinhorst, Deborah Milks, George P. Naughton, Mark Schuck, and Bill Arnsberg. Use of statistical bootstrapping to calculate confidence intervals for the fall Chinook salmon run reconstruction to Lower Granite Dam. Transactions of the American Fisheries Society, 139(6):1792–1801, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Satterthwaite:2014:UGS

[SMO^{+}14] William H. Satterthwaite, Michael S. Mohr, Michael R. O’Farrell, Eric C. Anderson, Michael A. Banks, Sarah J. Bates, M. Renee Bellinger, Lisa A. Borgerson, Eric D. Crandall, John Carlos Garza, Brett J. Kormos, Peter W. Lawson, and Melodie L. Palmer-Zwahlen. Use of genetic stock identification data for comparison of the ocean spatial distribution, size at age, and fishery exposure of an untagged stock and its indicator: California Coastal versus Klamath River Chinook salmon. Transactions of the American Fisheries Society, 143 (1):117–133, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sogard:2012:CHC

[SMS^{+}12] Susan M. Sogard, Joseph E. Merz, William H. Satterthwaite, Michael P. Beakes, David R. Swank, Erin M. Collins, Robert G. Titus, and Marc Mangel. Contrasts in habitat characteristics and life history patterns of Oncorhynchus mykiss in California’s Central Coast and Central Valley. Transactions of the American Fisheries Society, 141(3):747–760, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sestrich:2011:IFN

[SMY11] Clint M. Sestrich, Thomas E. McMahon, and Michael K. Young. Influence of fire on native and nonnative salmonid populations and habitat in a Western Montana Basin. Transactions of the American Fisheries Society, 140(1):136–146, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schneider:2010:TWS

[SNC^{+}10] Kristal N. Schneider, Raymond M. Newman, Virginia Card, Sanford Weisberg, and Donald L. Pereira. Timing of walleye spawning as an indicator of climate change. Transactions of the American Fisheries Society, 139(4):1198–1210, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stephenson:2013:SDS

[SNI^{+}13] Sarah M. Stephenson, Matthew D. Neufeld, Susan C. Ireland, Shawn Young, Ryan S. Hardy, and Pete Rust. Survival and dispersal of sonic-tagged, hatchery-reared burbot released into the Kootenay River. Transactions of the American Fisheries Society, 142(6):1671–1679, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seitz:2010:FSI

[SNP^{+}10] Andrew C. Seitz, Brenda L. Norcross, John C. Payne, Anna N. Kagley, Buck Meloy, Jacob L. Gregg, and Paul K. Hershberger. Feasibility of surgically implanting acoustic tags into Pacific herring. Transactions of the American Fisheries Society, 139(5):1288–1291, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simpson:2012:EEB

[SO12] William G. Simpson and Kenneth G. Ostrand. Effects of entrainment and bypass at screened irrigation canals on juvenile steelhead. Transactions of the American Fisheries Society, 141 (3):599–609, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schmitt:2015:FRP

[SO15] Joseph D. Schmitt and Donald J. Orth. First record of pughead deformity in blue catfish. Transactions of the American Fisheries Society, 144(6):1111–1116, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schueller:2010:ARJ

[SP10] Paul Schueller and Douglas L. Peterson. Abundance and recruitment of juvenile Atlantic sturgeon in the Altamaha River, Georgia. Transactions of the American Fisheries Society, 139 (5):1526–1535, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sanft:2018:VJB

[SPC^{+}18] Eric Sanft, Joseph J. Parkos III, Scott F. Collins, Anthony P. Porreca, and David H. Wahl. Vulnerability of juvenile bighead and silver carps to predation by largemouth bass. Transactions of the American Fisheries Society, 147(6):1207–1214, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stockwell:2011:ESP

[SPCJ11] Craig A. Stockwell, Kevin M. Purcell, Michael L. Collyer, and John Janovy. Effects of salinity on Physa acuta, the intermediate host for the parasite Posthodiplostomum minimum: Implications for the translocation of the protected white sands pupfish. Transactions of the American Fisheries Society, 140(5): 1370–1374, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spurgeon:2015:THC

[SPH^{+}15] Jonathan J. Spurgeon, Craig P. Paukert, Brian D. Healy, Melissa Trammell, Dave Speas, and Emily Omana-Smith. Translocation of humpback chub into tributary streams of the Colorado River: Implications for conservation of large-river fishes. Transactions of the American Fisheries Society, 144 (3):502–514, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Suski:2019:INS

[SPH19] Cory D. Suski, Madison A. Philipp, and Caleb T. Hasler. Influence of nutritional status on carbon dioxide tolerance and avoidance behavior in a freshwater teleost. Transactions of the American Fisheries Society, 148(5):914–925, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spies:2012:LGR

[Spi12] Ingrid Spies. Landscape genetics reveals population subdivision in Bering Sea and Aleutian Islands Pacific cod. Transactions of the American Fisheries Society, 141(6):1557–1573, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saglam:2017:GAR

[SPM^{+}17] Ismail K. Saglam, Daniel J. Prince, Mariah Meek, Omar A. Ali, Michael R. Miller, Mary Peacock, Helen Neville, Alisha Goodbla, Chad Mellison, William Somer, Bernie May, and Amanda J. Finger. Genomic analysis reveals genetic distinctiveness of the Paiute cutthroat trout Oncorhynchus clarkii seleniris. Transactions of the American Fisheries Society, 146 (6):1291–1302, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simonin:2016:IBH

[SPR^{+}16] Paul W. Simonin, Donna L. Parrish, Lars G. Rudstam, Bernard Pientka, and Patrick J. Sullivan. Interactions between hatch dates, growth rates, and mortality of age-0 native rainbow smelt and nonnative alewife in Lake Champlain. Transactions of the American Fisheries Society, 145(3):649– 656, May 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sable:2010:PYP

[SR10] Shaye E. Sable and Kenneth A. Rose. Predicting yellow perch population responses using a density-dependent agestructured matrix projection model: How many annual data points are needed? Transactions of the American Fisheries Society, 139(6):1857–1871, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2015:OHU

[SR15] M. Chad Smith and Roger A. Rulifson. Overlapping habitat use of multiple anadromous fish species in a restricted coastal watershed. Transactions of the American Fisheries Society, 144(6):1173–1183, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seegert:2014:HDO

[SRMB^{+}14] Sarah E. Zahn Seegert, Emma J. Rosi-Marshall, Colden V. Baxter, Theodore A. Kennedy, Robert O. Hall, Jr., and Wyatt F. Cross. High diet overlap between native small-bodied fishes and nonnative fathead minnow in the Colorado River, Grand Canyon, Arizona. Transactions of the American Fisheries Society, 143(4):1072–1083, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seegert:2015:TAF

[SRMB^{+}15] Sarah E. Zahn Seegert, Emma J. Rosi-Marshall, Colden V. Baxter, Theodore A. Kennedy, Robert O. Hall, and Wyatt F. Cross. Transactions of the American Fisheries Society: Guide for authors. Transactions of the American Fisheries Society, 144(1):228–233, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simonin:2018:PDS

[SRP^{+}18] Paul W. Simonin, Lars G. Rudstam, Donna L. Parrish, Bernard Pientka, and Patrick J. Sullivan. Piscivore diet shifts and trophic level change after alewife establishment in Lake Champlain. Transactions of the American Fisheries Society, 147(5):939–947, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simmons:2011:RAG

[SS11] Carrie M. Simmons and Stephen T. Szedlmayer. Recruitment of age-0 gray triggerfish to benthic structured habitat in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 140(1):14–20, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Silbernagel:2013:DFL

[SS13] Justin J. Silbernagel and Peter W. Sorensen. Direct field and laboratory evidence that a combination of egg and larval predation controls recruitment of invasive common carp in many lakes of the Upper Mississippi River Basin. Transactions of the American Fisheries Society, 142(4):1134–1140, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siwicke:2015:ILA

[SS15a] Kevin A. Siwicke and Andrew C. Seitz. Interpreting lamprey attacks on Pacific cod in the Eastern Bering Sea. Transactions of the American Fisheries Society, 144(6):1249–1262, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stahr:2015:AWW

[SS15b] Kristopher J. Stahr and Daniel E. Shoup. American water willow mediates survival and antipredator behavior of juvenile largemouth bass. Transactions of the American Fisheries Society, 144(5):903–910, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sullivan:2015:TPG

[SS15c] Timothy J. Sullivan and Carol A. Stepien. Temporal population genetic structure of yellow perch spawning groups in the Lower Great Lakes. Transactions of the American Fisheries Society, 144(1):211–226, January 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sass:2018:WPR

[SS18a] Greg G. Sass and Stephanie L. Shaw. Walleye population responses to experimental exploitation in a Northern Wisconsin lake. Transactions of the American Fisheries Society, 147(5): 869–878, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siwicke:2018:SDD

[SS18b] Kevin A. Siwicke and Andrew C. Seitz. Spatial differences in the distributions of Arctic and Pacific lampreys in the Eastern Bering Sea. Transactions of the American Fisheries Society, 147(1):111–127, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seeb:2011:SNP

[SSH^{+}11a] Lisa W. Seeb, James E. Seeb, Christopher Habicht, Edward V. Farley, Jr., and Fred M. Utter. Single-nucleotide polymorphic genotypes reveal patterns of early juvenile migration of sockeye salmon in the Eastern Bering Sea. Transactions of the American Fisheries Society, 140(3):734–748, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shippentower:2011:WYM

[SSH11b] Gene E. Shippentower, Carl B. Schreck, and Scott A. Heppell. Who’s your momma? Recognizing maternal origin of juvenile steelhead using injections of strontium chloride to create transgenerational marks. Transactions of the American Fisheries Society, 140(5):1330–1339, September 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smylie:2016:AGR

[SSM16] Meredith Smylie, Virginia Shervette, and Chris McDonough. Age, growth, and reproduction in two coastal populations of longnose gars. Transactions of the American Fisheries Society, 145(1):120–135, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sotola:2017:GEI

[SSR^{+}17] V. Alex Sotola, Aaron W. Schrey, Alexandria K. Ragsdale, Gregory W. Whitledge, Les Frankland, Eric K. Bollinger, and Robert E. Colombo. Genetic evidence of isolation by distance and impact of impoundments on genetic diversity of riverine channel catfish. Transactions of the American Fisheries Society, 146(6):1204–1211, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaeffer:2012:EFF

[SSS^{+}12] Travis W. Schaeffer, Daniel E. Spengler, Casey W. Schoenebeck, Michael L. Brown, and Steven R. Chipps. Effect of feeding– fasting cycles on oxygen consumption and bioenergetics of female yellow perch. Transactions of the American Fisheries Society, 141(6):1480–1491, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:2012:CFC

[STK12] Charles E. Stanley, Jason M. Taylor, and Ryan S. King. Coupling fish community structure with instream flow and habitat connectivity between two hydrologically extreme years. Transactions of the American Fisheries Society, 141(4):1000–1015, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:2010:OES

[Sto10] Dennis M. Stone. Overriding effects of species-specific turbidity thresholds on hoop-net catch rates of native fishes in the Little Colorado River, Arizona. Transactions of the American Fisheries Society, 139(4):1150–1170, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strange:2010:UTL

[Str10] Joshua S. Strange. Upper thermal limits to migration in adult Chinook salmon: Evidence from the Klamath River Basin. Transactions of the American Fisheries Society, 139(4):1091– 1108, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strange:2012:MSA

[Str12] Joshua S. Strange. Migration strategies of adult Chinook salmon runs in response to diverse environmental conditions in the Klamath River Basin. Transactions of the American Fisheries Society, 141(6):1622–1636, November 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sato:2014:GSD

[STS^{+}14] Shunpei Sato, William D. Templin, Lisa W. Seeb, James E. Seeb, and Shigehiko Urawa. Genetic structure and diversity of Japanese chum salmon populations inferred from singlenucleotide polymorphism markers. Transactions of the American Fisheries Society, 143(5):1231–1246, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stuby:2018:CLH

[Stu18] Lisa Stuby. Contributions to the life history of Kuskokwim River inconnu. Transactions of the American Fisheries Society, 147(5):879–890, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shepard:2013:IVE

[STZ13] Bradley B. Shepard, Mark L. Taper, and Alexander V. Zale. Improved variance estimates of biomass for stream-dwelling fish calculated using removal estimators. Transactions of the American Fisheries Society, 142(3):841–853, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stark:2018:EVE

[SVKK18] Eric J. Stark, Dmitri T. Vidergar, Christine C. Kozfkay, and Paul A. Kline. Egg viability and egg-to-fry survival of captivereared Chinook salmon released to spawn naturally. Transactions of the American Fisheries Society, 147(1):128–138, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:2011:BSF

[SW11] Daniel E. Shoup and David H. Wahl. Body size, food, and temperature affect overwinter survival of age-0 bluegills. Transactions of the American Fisheries Society, 140(5):1298–1304, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simpson:2016:TRA

[SWB^{+}16] Cara A. Simpson, Michael J. Wilberg, Hongsheng Bi, Amy M. Schueller, Geneviève M. Nesslage, and Harvey J. Walsh. Trends in relative abundance and early life survival of Atlantic menhaden during 1977–2013 from long-term ichthyoplankton programs. Transactions of the American Fisheries Society, 145(5):1139–1151, September 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stapanian:2013:TCS

[SWC13] Martin A. Stapanian, Larry D. Witzel, and Andy Cook. Temporal changes and sexual differences in spatial distribution of burbot in Lake Erie. Transactions of the American Fisheries Society, 142(6):1724–1732, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sellheim:2017:VFR

[SWH^{+}17] K. Sellheim, M. Willmes, J. A. Hobbs, J. J. G. Glessner, Z. J. Jackson, and J. E. Merz. Validating fin ray microchemistry as a tool to reconstruct the migratory history of white sturgeon. Transactions of the American Fisheries Society, 146(5): 844–857, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sullivan:2018:FRY

[SWP^{+}18] Christopher J. Sullivan, Michael J. Weber, Clay L. Pierce, David H. Wahl, Quinton E. Phelps, Carlos A. Camacho, and Robert E. Colombo. Factors regulating year-class strength of silver carp throughout the Mississippi River Basin. Transactions of the American Fisheries Society, 147(3):541–553, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steen:2010:PFC

[SWS10] Paul J. Steen, Michael J. Wiley, and Jeffrey S. Schaeffer. Predicting future changes in Muskegon River watershed game fish distributions under future land cover alteration and climate change scenarios. Transactions of the American Fisheries Society, 139(2):396–412, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaffler:2015:OCD

[SYH^{+}15] Jason J. Schaffler, Shawn P. Young, Steve Herrington, Travis Ingram, and Josh Tannehill. Otolith chemistry to determine within-river origins of Alabama shad in the Apalachicola– Chattahoochee–Flint River Basin. Transactions of the American Fisheries Society, 144(1):1–10, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sandford:2012:ELM

[SZGS12] Benjamin P. Sandford, Richard W. Zabel, Lyle G. Gilbreath, and Steven G. Smith. Exploring latent mortality of juvenile salmonids related to migration through the Columbia River hydropower system. Transactions of the American Fisheries Society, 141(2):343–352, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spencer:2010:MUG

[SZZ10] Randall C. Spencer, Joseph Zydlewski, and Gayle Zydlewski. Migratory urge and gill Na^{+}, K^{+}-ATPase activity of hatcheryreared Atlantic salmon smolts from the Dennys and Penobscot River stocks, Maine. Transactions of the American Fisheries Society, 139(4):947–956, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tremain:2012:MGS

[TA12] Derek M. Tremain and Douglas H. Adams. Mercury in groupers and sea basses from the Gulf of Mexico: Relationships with size, age, and feeding ecology. Transactions of the American Fisheries Society, 141(5):1274–1286, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:2013:VBC

[TA13] Josh L. Taylor and Steven K. A. Arndt. Variability in burbot cohort abundance at juvenile and adult stages in Columbia Lake, British Columbia. Transactions of the American Fisheries Society, 142(6):1705–1715, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trippel:2015:IRS

[TAM15] Nicholas A. Trippel, Micheal S. Allen, and Richard S. McBride. Importance of resident and seasonally transient prey to largemouth bass in the St. Johns River, Florida. Transactions of the American Fisheries Society, 144(1):140–149, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Terrazas:2017:EDT

[TASC17] Marc M. Terrazas, Jennifer R. Adams, Ponnerassery S. Sudheesh, and Kenneth D. Cain. Effects of diel temperature fluctuation on growth, stress response, and immune function of burbot. Transactions of the American Fisheries Society, 146 (5):996–1007, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:2014:SSM

[TB14] Jamie N. Thompson and David A. Beauchamp. Size-selective mortality of steelhead during freshwater and marine life stages related to freshwater growth in the Skagit River, Washington. Transactions of the American Fisheries Society, 143(4):910– 925, July 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:2016:FDF

[TB16] Neil F. Thompson and Michael S. Blouin. Family dominance at the fry stage weakly influences mean family body length at smolting in hatchery-reared steelhead. Transactions of the American Fisheries Society, 145(6):1282–1289, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trushenski:2013:IRU

[TBC^{+}13] J. T. Trushenski, J. D. Bowker, S. J. Cooke, D. Erdahl, T. Bell, J. R. MacMillan, R. P. Yanong, J. E. Hill, M. C. Fabrizio, J. E. Garvey, and S. Sharon. Issues regarding the use of sedatives in fisheries and the need for immediate-release options. Transactions of the American Fisheries Society, 142(1):156– 170, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trushenski:2012:CEA

[TBGM12] Jesse T. Trushenski, James D. Bowker, Brian R. Gause, and Bonnie L. Mulligan. Chemical and electrical approaches to sedation of hybrid striped bass: Induction, recovery, and physiological responses to sedation. Transactions of the American Fisheries Society, 141(2):455–467, March 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tabor:2017:PBS

[TBL^{+}17] Roger A. Tabor, Alex T. C. Bell, Daniel W. Lantz, Chris N. Gregersen, Hans B. Berge, and Denise K. Hawkins. Phototaxic behavior of subyearling salmonids in the nearshore area of two urban lakes in Western Washington State. Transactions of the American Fisheries Society, 146(4):753–761, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trotter:2012:MCS

[TBTS12] Alexis A. Trotter, David A. Blewett, Ronald G. Taylor, and Philip W. Stevens. Migrations of common snook from a tidal river with implications for skipped spawning. Transactions of the American Fisheries Society, 141(4):1016–1025, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffan:2011:DBN

[TC11] Kenneth F. Tiffan and William P. Connor. Distinguishing between natural and hatchery Snake River fall Chinook salmon subyearlings in the field using body morphology. Transactions of the American Fisheries Society, 140(1):21–30, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trested:2011:SMM

[TCBI11] Drew G. Trested, Matthew D. Chan, William C. Bridges, and J. Jeffery Isely. Seasonal movement and mesohabitat usage of adult and juvenile lake sturgeon in the Grasse River, New York. Transactions of the American Fisheries Society, 140 (4):1006–1014, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thibault:2013:IER

[TD13] Isabel Thibault and Julian Dodson. Impacts of exotic rainbow trout on habitat use by native juvenile salmonid species at an early invasive stage. Transactions of the American Fisheries Society, 142(4):1141–1150, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffan:2014:PAC

[TES14] Kenneth F. Tiffan, John M. Erhardt, and Scott J. St. John. Prey availability, consumption, and quality contribute to variation in growth of subyearling Chinook salmon rearing in riverine and reservoir habitats. Transactions of the American Fisheries Society, 143(1):219–229, January 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tuckey:2013:ISD

[TF13] Troy D. Tuckey and Mary C. Fabrizio. Influence of survey design on fish assemblages: Implications from a study in Chesapeake Bay tributaries. Transactions of the American Fisheries Society, 142(4):957–973, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tibbles:2018:ISA

[TFM^{+}18] Marguerite Tibbles, Jeffrey A. Falke, Andrew R. Mahoney, Martin D. Robards, and Andrew C. Seitz. An Interferometric Synthetic Aperture Radar (InSAR) habitat suitability model to identify overwinter conditions for coregonine whitefishes in Arctic lagoons. Transactions of the American Fisheries Society, 147(6):1167–1178, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trebitz:2015:CWS

[TH15] Anett S. Trebitz and Joel C. Hoffman. Coastal wetland support of Great Lakes fisheries: Progress from concept to quantification. Transactions of the American Fisheries Society, 144 (2):352–372, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Then:2015:CTL

[THGA15] Amy Y. Then, John M. Hoenig, Todd Gedamke, and Jerald S. Ault. Comparison of two length-based estimators of total mortality: a simulation approach. Transactions of the American Fisheries Society, 144(6):1206–1219, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tronstad:2010:ILT

[THKG10] Lusha M. Tronstad, Robert O. Hall, Jr., Todd M. Koel, and Ken G. Gerow. Introduced lake trout produced a fourlevel trophic cascade in Yellowstone Lake. Transactions of the American Fisheries Society, 139(5):1536–1550, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trippel:2017:AII

[THL^{+}17] Nicholas A. Trippel, John S. Hargrove, Erin H. Leone, James D. Austin, and Micheal S. Allen. Angling-induced impacts on recruitment and contributions to reproduction in Florida bass. Transactions of the American Fisheries Society, 146(5):871–887, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thoni:2014:ESD

[THNG14] Ryan Thoni, Jordan Holcomb, Robert Nichols, and Michael M. Gangloff. Effects of small dams on sunfish assemblages in North Carolina Piedmont and coastal plain streams. Transactions of the American Fisheries Society, 143(1):97–103, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tsehaye:2014:CSC

[TJB^{+}14] Iyob Tsehaye, Michael L. Jones, Travis O. Brenden, James R. Bence, and Randall M. Claramunt. Changes in the salmonine community of Lake Michigan and their implications for predator–prey balance. Transactions of the American Fisheries Society, 143(2):420–437, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffan:2012:DMF

[TKC^{+}12] Kenneth F. Tiffan, Tobias J. Kock, William P. Connor, Frank Mullins, and R. Kirk Steinhorst. Downstream movement of fall Chinook salmon juveniles in the Lower Snake River reservoirs during winter and early spring. Transactions of the American Fisheries Society, 141(2):285–293, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffan:2018:MBP

[TKC^{+}18] Kenneth F. Tiffan, Tobias J. Kock, William P. Connor, Marshall C. Richmond, and William A. Perkins. Migratory behavior and physiological development as potential determinants of life history diversity in fall Chinook salmon in the Clearwater River. Transactions of the American Fisheries Society, 147 (2):400–413, March 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:2014:DRH

[TL14] Sara M. Turner and Karin E. Limburg. Determination of river herring natal origin using otolith chemical markers: Accuracy as a function of spatial scale and choice of markers. Transactions of the American Fisheries Society, 143(6):1530–1543, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:2015:DDG

[TL15] S. M. Turner and K. E. Limburg. Does daily growth affect the rate of manganese uptake in juvenile river herring otoliths? Transactions of the American Fisheries Society, 144(5):873– 881, September 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:2017:TLS

[TL17] Andrew D. Taylor and Matthew K. Litvak. Timing and location of spawning based on larval capture and ultrasonic telemetry of Atlantic sturgeon in the Saint John River, New Brunswick. Transactions of the American Fisheries Society, 146(2):283–290, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Torre:2017:NES

[TLKT17] Michael P. Torre, Danielle M. Lifavi, William P. Kaestner, and Timothy E. Targett. No evidence of site fidelity in Atlantic croakers or spot in the shore zone of Lower Delaware Bay. Transactions of the American Fisheries Society, 146(6):1093– 1100, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trushenski:2019:SSG

[TLM^{+}19] Jesse T. Trushenski, Donald A. Larsen, Mollie A. Middleton, Michelle Jakaitis, Eric L. Johnson, Christine C. Kozfkay, and Paul A. Kline. Search for the smoking gun: Identifying and addressing the causes of postrelease morbidity and mortality of hatchery-reared Snake River sockeye salmon smolts. Transactions of the American Fisheries Society, 148(5):875–895, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thorn:2019:SUG

[TM19] Michael W. Thorn and Yolanda E. Morbey. Survival under the gravel: the effects of egg size and parental identity on the in situ hatching success of Chinook salmon. Transactions of the American Fisheries Society, 148(5):991–1003, September 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teel:2011:ISS

[TNOU11] David J. Teel, Shawn R. Narum, Jeffrey B. Olsen, and Fred M. Utter. Introduction to a special section: Genetic adaptation of natural salmonid populations. Transactions of the American Fisheries Society, 140(3):659–664, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trumbo:2014:RSV

[TNS^{+}14] Bradly A. Trumbo, Keith H. Nislow, Jonathan Stallings, Mark Hudy, Eric P. Smith, Dong-Yun Kim, Bruce Wiggins, and Charles A. Dolloff. Ranking site vulnerability to increasing temperatures in Southern Appalachian brook trout streams in Virginia: an exposure-sensitivity approach. Transactions of the American Fisheries Society, 143(1):173–187, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thorson:2011:BCC

[TP11] James T. Thorson and Michael H. Prager. Better catch curves: Incorporating age-specific natural mortality and logistic selectivity. Transactions of the American Fisheries Society, 140 (2):356–366, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:2014:ESI

[TPA14] Brandon C. Thompson, Wesley Porak, and Micheal S. Allen. Effects of surgically implanting radio transmitters in juvenile largemouth bass. Transactions of the American Fisheries Society, 143(2):346–352, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Torterotot:2014:IFR

[TPBB14] Jean-Baptiste Torterotot, Charles Perrier, Normand E. Bergeron, and Louis Bernatchez. Influence of forest road culverts and waterfalls on the fine-scale distribution of brook trout genetic diversity in a boreal watershed. Transactions of the American Fisheries Society, 143(6):1577–1591, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:2016:URC

[TPLA16] Brandon C. Thompson, Wesley F. Porak, Erin H. Leone, and Micheal S. Allen. Using radiotelemetry to compare the initial behavior and mortality of hatchery-reared and wild juvenile Florida bass. Transactions of the American Fisheries Society, 145(2):374–385, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Torres-Rojas:2014:FHV

[TRHHOGSJ14] Y. E. Torres-Rojas, A. Hernández-Herrera, S. Ortega-García, and M. F. Soto-Jiménez. Feeding habits variability and trophic position of dolphinfish in waters south of the Baja California Peninsula, Mexico. Transactions of the American Fisheries Society, 143(2):528–542, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tattam:2013:BSG

[TRLG13] Ian A. Tattam, James R. Ruzycki, Hiram W. Li, and Guillermo R. Giannico. Body size and growth rate influence emigration timing of Oncorhynchus mykiss. Transactions of the American Fisheries Society, 142(5):1406–1414, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tattam:2015:LCW

[TRMC15] Ian A. Tattam, James R. Ruzycki, Josh L. McCormick, and Richard W. Carmichael. Length and condition of wild Chinook salmon smolts influence age at maturity. Transactions of the American Fisheries Society, 144(6):1237–1248, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:2010:SBH

[TRW10] Jessica S. Thompson, James A. Rice, and D. Scott Waters. Striped bass habitat selection rules in reservoirs without suitable summer habitat offer insight into consequences for growth. Transactions of the American Fisheries Society, 139(5):1450–1464, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Topping:2013:UUT

[TS13] Darin T. Topping and Stephen T. Szedlmayer. Use of ultrasonic telemetry to estimate natural and fishing mortality of red snapper. Transactions of the American Fisheries Society, 142 (4):1090–1100, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Torre:2017:FBW

[TT17] Michael P. Torre and Timothy E. Targett. Feeding by bluefish and weakfish along riprap-hardened shorelines: Comparisons with adjacent sandy beach in Delaware Bay, USA. Transactions of the American Fisheries Society, 146(2):341–348, March 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tucker:2015:CDC

[TTM^{+}15] S. Tucker, M. E. Thiess, J. F. T. Morris, D. Mackas, W. T. Peterson, J. R. Candy, T. D. Beacham, E. M. Iwamoto, D. J. Teel, M. Peterson, and M. Trudel. Coastal distribution and consequent factors influencing production of endangered Snake River sockeye salmon. Transactions of the American Fisheries Society, 144(1):107–123, January 2015. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tucker:2011:LHS

[TTW^{+}11] S. Tucker, M. Trudel, D. W. Welch, J. R. Candy, J. F. T. Morris, M. E. Thiess, C. Wallace, and T. D. Beacham. Life history and seasonal stock-specific ocean migration of juvenile Chinook salmon. Transactions of the American Fisheries Society, 140(4):1101–1119, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:2012:SSH

[TVL^{+}12] Lisa C. Thompson, Jenna L. Voss, Royce E. Larsen, William D. Tietje, Ryan A. Cooper, and Peter B. Moyle. Southern steelhead, hard woody debris, and temperature in a California Central Coast watershed. Transactions of the American Fisheries Society, 141(2):275–284, March 2012. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Timm:2014:INA

[TW14] Raymond K. Timm and Robert C. Wissmar. Influence of natural and anthropogenic disturbances on spawning sockeye salmon distributions in the Cedar River, Washington. Transactions of the American Fisheries Society, 143(3):709–720, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Uthe:2019:ECR

[UACS^{+}19] P. Uthe, R. Al-Chokhachy, B. B. Shepard, A. V. Zale, and J. L. Kershner. Effects of climate-related stream factors on patterns of individual summer growth of cutthroat trout. Transactions of the American Fisheries Society, 148(1):21–34, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Urquhart:2014:LTT

[UK14] Alexander N. Urquhart and Peter Koetsier. Low-temperature tolerance and critical thermal minimum of the invasive oriental weatherfish Misgurnus anguillicaudatus in Idaho, USA. Transactions of the American Fisheries Society, 143(1):68–76, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Urabe:2010:EHQ

[UNTA10] Hirokazu Urabe, Miyuki Nakajima, Mitsuru Torao, and Tomoya Aoyama. Evaluation of habitat quality for stream salmonids based on a bioenergetics model. Transactions of the American Fisheries Society, 139(6):1665–1676, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Usvyatsov:2013:TED

[UPT^{+}13] Sima Usvyatsov, Jeffrey Picka, Andrew Taylor, James Watmough, and Matthew Kenneth Litvak. Timing and extent of drift of shortnose sturgeon larvae in the Saint John River, New Brunswick, Canada. Transactions of the American Fisheries Society, 142(3):717–730, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Usvyatsov:2012:APS

[UWL12a] Sima Usvyatsov, James Watmough, and Matthew K. Litvak. Age and population size estimates of overwintering shortnose sturgeon in the Saint John River, New Brunswick, Canada. Transactions of the American Fisheries Society, 141(4):1126– 1136, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Usvyatsov:2012:MEE

[UWL12b] Sima Usvyatsov, James Watmough, and Matthew K. Litvak. Modeling the effect of environmental parameters on feeding ecology of the shortnose sturgeon in the Saint John River, New Brunswick. Transactions of the American Fisheries Society, 141(1):238–256, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vari:2013:RHB

[Var13] Richard P. Vari. A review of “Historical Biogeography of Neotropical Freshwater Fishes” Edited by J. S. Albert and R. R. Reis. University of California Press, Berkeley. 388 pages. $85.00. Transactions of the American Fisheries Society, 142 (4):1151–1152, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vandergoot:2014:SVP

[VB14] Christopher S. Vandergoot and Travis O. Brenden. Spatially varying population demographics and fishery characteristics of Lake Erie walleyes inferred from a long-term tag recovery study. Transactions of the American Fisheries Society, 143(1): 188–204, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Volk:2010:RJC

[VBJS10] Eric C. Volk, Daniel L. Bottom, Kim K. Jones, and Charles A. Simenstad. Reconstructing juvenile Chinook salmon life history in the Salmon River Estuary, Oregon, using otolith microchemistry and microstructure. Transactions of the American Fisheries Society, 139(2):535–549, March 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Van:2014:LSG

[VCP^{+}14] Wishingrad Van, Meghan K. Carr, Michael S. Pollock, Maud C. O. Ferrari, and Douglas P. Chivers. Lake sturgeon geographic range, distribution, and migration patterns in the Saskatchewan River. Transactions of the American Fisheries Society, 143(6):1555–1561, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDeValk:2016:FYD

[VFB^{+}16] Anthony J. VanDeValk, John L. Forney, Thomas E. Brooking, James R. Jackson, and Lars G. Rudstam. First-year density and growth as they relate to recruitment of white perch to the adult stock in Oneida Lake, New York, 1968–2011. Transactions of the American Fisheries Society, 145(2):416– 426, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanLandeghem:2013:IGA

[VFFP13] Matthew M. VanLandeghem, Mukhtar Farooqi, Bobby Farquhar, and Reynaldo Patiño. Impacts of golden alga Prymnesium parvum on fish populations in reservoirs of the Upper Colorado River and Brazos River Basins, Texas. Transactions of the American Fisheries Society, 142(3):581–595, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDoornik:2015:GPS

[VHJ^{+}15] Donald M. Van Doornik, Maureen A. Hess, Marc A. Johnson, David J. Teel, Thomas A. Friesen, and James M. Myers. Genetic population structure of Willamette River steelhead and the influence of introduced stocks. Transactions of the American Fisheries Society, 144(1):150–162, January 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vine:2019:IEC

[VHPP19] Joshua R. Vine, S. Chad Holbrook, William C. Post, and Brandon K. Peoples. Identifying environmental cues for Atlantic sturgeon and shortnose sturgeon spawning migrations in the Savannah River. Transactions of the American Fisheries Society, 148(3):671–681, May 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vidal:2017:UVS

[VIW^{+}17] Tiffany E. Vidal, Brian J. Irwin, Tyler Wagner, Lars G. Rudstam, James R. Jackson, and James R. Bence. Using variance structure to quantify responses to perturbation in fish catches. Transactions of the American Fisheries Society, 146(4):584– 593, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Valiente:2011:IRT

[VJGV11] A. G. Valiente, F. Juanes, and E. Garcia-Vazquez. Increasing regional temperatures associated with delays in Atlantic salmon sea-run timing at the southern edge of the European distribution. Transactions of the American Fisheries Society, 140(2):367–373, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanPoorten:2012:FBB

[vPWT12] Brett T. van Poorten, Carl J. Walters, and Nathan G. Taylor. A field-based bioenergetics model for estimating time-varying food consumption and growth. Transactions of the American Fisheries Society, 141(4):943–961, July 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanSickle:2010:CMY

[vS10] John van Sickle. Correlated metrics yield multimetric indices with inferior performance. Transactions of the American Fisheries Society, 139(6):1802–1817, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Valenca-Silva:2014:LDA

[VSBB^{+}14] Graziela Valença-Silva, Mariana G. Braz, Rodrigo E. Barreto, Daisy M. F. Salvadori, and Gilson L. Volpato. Low dose of the anesthetic propofol does not induce genotoxic or mutagenic effects in Nile tilapia. Transactions of the American Fisheries Society, 143(2):414–419, March 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vizza:2013:EEP

[VSBC13] Carmella Vizza, Beth L. Sanderson, Douglas G. Burrows, and Holly J. Coe. The effects of ethanol preservation on fish fin stable isotopes: Does variation in C:N ratio and body size matter? Transactions of the American Fisheries Society, 142(5): 1469–1476, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vue:2015:BIA

[VSC15] Sherr Vue, Kurt M. Samways, and Richard A. Cunjak. Bioelectrical impedance analysis to estimate lipid content in Atlantic salmon parr as influenced by temperature, PIT tags, and instrument precision and application in field studies. Transactions of the American Fisheries Society, 144(2):235–245, March 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanVrancken:2010:EHK

[vVO10] Jeffrey van Vrancken and Martin O’Connell. Effects of Hurricane Katrina on freshwater fish assemblages in a small coastal tributary of Lake Pontchartrain, Louisiana. Transactions of the American Fisheries Society, 139(6):1723–1732, November 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanLandeghem:2013:TVP

[VWS13] Matthew M. VanLandeghem, David H. Wahl, and Cory D. Suski. Temporal variation in the physiological responses in largemouth bass following small club angling tournaments. Transactions of the American Fisheries Society, 142(1):257– 267, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanZwol:2012:ENS

[vZNW12] Jessica A. van Zwol, Bryan D. Neff, and Chris C. Wilson. The effect of nonnative salmonids on social dominance and growth of juvenile Atlantic salmon. Transactions of the American Fisheries Society, 141(4):907–918, July 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weinheimer:2017:MCI

[WAD^{+}17] Joshua Weinheimer, Joseph H. Anderson, Mark Downen, Mara Zimmerman, and Thom Johnson. Monitoring climate impacts: Survival and migration timing of summer chum salmon in Salmon Creek, Washington. Transactions of the American Fisheries Society, 146(5):983–995, September 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2010:CRB

[WAFL10] Gregory D. Williams, Kelly S. Andrews, Deborah A. Farrer, and Phillip S. Levin. Catch rates and biological characteristics of bluntnose sixgill sharks in Puget Sound. Transactions of the American Fisheries Society, 139(1):108–116, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wellband:2012:FSP

[WAKH12] Kyle W. Wellband, Dana Y. Atagi, Rachel A. Koehler, and Daniel D. Heath. Fine-scale population genetic structure and dispersal of juvenile steelhead in the Bulkley–Morice River, British Columbia. Transactions of the American Fisheries Society, 141(2):392–401, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilzbach:2012:MRR

[WAR12] Margaret A. Wilzbach, Mark J. Ashenfelter, and Seth J. Ricker. Movement of resident rainbow trout transplanted below a barrier to anadromy. Transactions of the American Fisheries Society, 141(2):294–304, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waples:2011:PGS

[WAW11] Robin S. Waples, Paul B. Aebersold, and Gary A. Winans. Population genetic structure and life history variability in Oncorhynchus nerka from the Snake River Basin. Transactions of the American Fisheries Society, 140(3):716–733, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weber:2013:DDE

[WB13a] Michael J. Weber and Michael L. Brown. Density-dependence and environmental conditions regulate recruitment and firstyear growth of common carp in shallow lakes. Transactions of the American Fisheries Society, 142(2):471–482, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weber:2013:SVJ

[WB13b] Michael J. Weber and Michael L. Brown. Spatiotemporal variation of juvenile common carp foraging patterns as inferred from stable isotope analysis. Transactions of the American Fisheries Society, 142(5):1179–1191, September 2013. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walsworth:2015:INS

[WB15a] Timothy E. Walsworth and Phaedra Budy. Integrating nonnative species in niche models to prioritize native fish restoration activity locations along a desert river corridor. Transactions of the American Fisheries Society, 144(4):667–681, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winters:2015:ECT

[WB15b] Lisa K. Winters and Phaedra Budy. Exploring crowded trophic niche space in a novel reservoir fish assemblage: How many is too many? Transactions of the American Fisheries Society, 144(6):1117–1128, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wipf:2014:LTV

[WBD14] Matthew M. Wipf, Michael E. Barnes, and Dan Durben. Lack of temporal variation in egg size in landlocked fall Chinook salmon from Lake Oahe, South Dakota. Transactions of the American Fisheries Society, 143(2):289–293, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitney:2017:STV

[WBD17] Emily J. Whitney, Anne H. Beaudreau, and Douglas H. Duncan. Spatial and temporal variation in the diets of Pacific staghorn sculpins related to hydrological factors in a glacially influenced estuary. Transactions of the American Fisheries Society, 146(6):1156–1167, November 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitledge:2010:LET

[WBH10] Gregory W. Whitledge, Przemyslaw G. Bajer, and Robert S. Hayward. Laboratory evaluation of two bioenergetics models for brown trout. Transactions of the American Fisheries Society, 139(4):929–936, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winans:2017:GCO

[WBM^{+}17] Gary A. Winans, Jon Baker, Mike McHenry, Larry Ward, and Jim Myers. Genetic characterization of Oncorhynchus mykiss prior to dam removal with implications for recolonization of the Elwha River watershed, Washington. Transactions of the American Fisheries Society, 146(1):160–172, January 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winner:2010:RAD

[WBMG10] Brent L. Winner, David A. Blewett, Robert H. McMichael, Jr., and Cameron B. Guenther. Relative abundance and distribution of common snook along shoreline habitats of Florida estuaries. Transactions of the American Fisheries Society, 139 (1):62–79, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warnock:2011:EPC

[WBR11] Will G. Warnock, Jason K. Blackburn, and Joseph B. Rasmussen. Estimating proportional contributions of migratory bull trout from hierarchical populations to mixed-stock recreational fisheries using genetic and trapping data. Transactions of the American Fisheries Society, 140(2):345–355, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wisniewski:2013:MFR

[WBW^{+}13] Jason M. Wisniewski, Katherine D. Bockrath, John P. Wares, Andrea K. Fritts, and Matthew J. Hill. The mussel–fish relationship: a potential new twist in North America? Transactions of the American Fisheries Society, 142(3):642–648, May 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wehrly:2012:LBC

[WBWSK12] Kevin E. Wehrly, James E. Breck, Lizhu Wang, and Lidia Szabo-Kraft. A landscape-based classification of fish assemblages in sampled and unsampled lakes. Transactions of the American Fisheries Society, 141(2):414–425, March 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wesner:2011:GPT

[WCD^{+}11] Jeff Scott Wesner, James Wes Cornelison, Christopher D. Dankmeyer, Peter F. Galbreath, and Thomas H. Martin. Growth, pH tolerance, survival, and diet of introduced northern-strain and native southern-strain Appalachian brook trout. Transactions of the American Fisheries Society, 140 (1):37–44, March 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weigel:2013:CSN

[WCMP13] Dana E. Weigel, Patrick J. Connolly, Kyle D. Martens, and Madison S. Powell. Colonization of steelhead in a natal stream after barrier removal. Transactions of the American Fisheries Society, 142(4):920–930, July 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weinstein:2019:EEB

[WCN^{+}19] Spencer Y. Weinstein, Jason A. Coombs, Keith H. Nislow, Chris Riley, Allison H. Roy, and Andrew R. Whiteley. Evaluating the effects of barriers on slimy sculpin movement and population connectivity using novel sibship-based and traditional genetic metrics. Transactions of the American Fisheries Society, 148(6):1117–1131, November 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitlock:2018:UGP

[WCQD18] Steven L. Whitlock, Matthew R. Campbell, Michael C. Quist, and Andrew M. Dux. Using genetic and phenotypic comparisons to evaluate apparent segregation among Kokanee spawning groups. Transactions of the American Fisheries Society, 147(1):43–60, January 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:2018:PAC

[WCYH18] Maryann S. Watson, Katrina V. Cook, Nathan Young, and Scott G. Hinch. Perceptions and actions of commercial fishers in response to conservation measures in Canadian Pacific salmon fisheries. Transactions of the American Fisheries Society, 147(5):906–918, September 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:2018:DRF

[WCZ^{+}18] Jonathan M. Watson, Stephen M. Coghlan, Jr., Joseph Zydlewski, Daniel B. Hayes, and Ian A. Kiraly. Dam removal and fish passage improvement influence fish assemblages in the Penobscot River, Maine. Transactions of the American Fisheries Society, 147(3):525–540, May 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wathen:2011:HSO

[WCZT11] Gus Wathen, Stephen M. Coghlan, Jr., Joseph Zydlewski, and Joan G. Trial. Habitat selection and overlap of Atlantic salmon and smallmouth bass juveniles in nursery streams. Transactions of the American Fisheries Society, 140(5):1145–1157, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:2013:LSE

[WDDS13] Tyler Wagner, Jefferson T. Deweber, Jason Detar, and John A. Sweka. Landscape-scale evaluation of asymmetric interactions between brown trout and brook trout using twospecies occupancy models. Transactions of the American Fisheries Society, 142(2):353–361, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warren:2014:GVE

[WDHW14] Dana R. Warren, Jason B. Dunham, and David Hockman-Wert. Geographic variability in elevation and topographic constraints on the distribution of native and nonnative trout in the great basin. Transactions of the American Fisheries Society, 143(1):205–218, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walrath:2016:ISC

[WDR16] John D. Walrath, Daniel C. Dauwalter, and Drew Reinke. Influence of stream condition on habitat diversity and fish assemblages in an impaired Upper Snake River Basin watershed. Transactions of the American Fisheries Society, 145(4):821– 834, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weber:2011:GSA

[WDW11] Michael J. Weber, John M. Dettmers, and David H. Wahl. Growth and survival of age-0 yellow perch across habitats in Southwestern Lake Michigan: Early life history in a large freshwater environment. Transactions of the American Fisheries Society, 140(5):1172–1185, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weber:2010:EPP

[WDWC10] Michael J. Weber, John M. Dettmers, David H. Wahl, and Sergiusz J. Czesny. Effects of predator–prey interactions and benthic habitat complexity on selectivity of a foraging generalist. Transactions of the American Fisheries Society, 139 (4):1004–1013, July 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weitkamp:2010:MDC

[Wei10] Laurie A. Weitkamp. Marine distributions of Chinook salmon from the West Coast of North America determined by coded wire tag recoveries. Transactions of the American Fisheries Society, 139(1):147–170, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wahl:2012:EEP

[WEW12] David H. Wahl, Lisa M. Einfalt, and Douglas B. Wojcieszak. Effect of experience with predators on the behavior and survival of muskellunge and tiger muskellunge. Transactions of the American Fisheries Society, 141(1):139–146, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2011:DEF

[WF11] Branson D. Williams and Mary C. Fabrizio. Detectability of estuarine fishes in a beach seine survey of tidal tributaries of Lower Chesapeake Bay. Transactions of the American Fisheries Society, 140(5):1340–1350, September 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weber:2019:REM

[WF19] Michael J. Weber and Mark Flammang. Relative effects of mortality versus emigration on a reservoir walleye population. Transactions of the American Fisheries Society, 148(1):123– 133, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wipfli:2010:SCI

[WHC^{+}10] Mark S. Wipfli, John P. Hudson, John P. Caouette, Nicole L. Mitchell, Joanna L. Lessard, Ron A. Heintz, and Dominic T. Chaloner. Salmon carcasses increase stream productivity more than inorganic fertilizer pellets: a test on multiple trophic levels in streamside experimental channels. Transactions of the American Fisheries Society, 139(3):824–839, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:2012:QCE

[WHF^{+}12] Annika W. Walters, Damon M. Holzer, James R. Faulkner, Charles D. Warren, Patrick D. Murphy, and Michelle M. Mc-Clure. Quantifying cumulative entrainment effects for Chinook salmon in a heavily irrigated watershed. Transactions of the American Fisheries Society, 141(5):1180–1190, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2012:MDS

[Whi12] Matthew M. White. Mitochondrial DNA sequence variation in saugers. Transactions of the American Fisheries Society, 141 (3):795–801, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitney:2014:POW

[WHP14] C. K. Whitney, S. G. Hinch, and D. A. Patterson. Population origin and water temperature affect development timing in embryonic sockeye salmon. Transactions of the American Fisheries Society, 143(5):1316–1329, September 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitney:2019:LTC

[WHSK19] James E. Whitney, Joshua A. Holloway, Derek T. Scholes, and Alexandra D. King. Long-term change of fish communities in a polluted watershed: Does cleaner water “Act” on fishes? Transactions of the American Fisheries Society, 148(1):191– 206, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weaver:2013:AES

[WK13] Daniel M. Weaver and Thomas J. Kwak. Assessing effects of stocked trout on nongame fish assemblages in Southern Appalachian mountain streams. Transactions of the American Fisheries Society, 142(6):1495–1507, November 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2018:HCG

[WKC18] Matthew M. White, Brooks A. Kohli, and Paul E. Converse. Historical and contemporary gene flow and the genetic structure of muskellunge in the Ohio River drainage. Transactions of the American Fisheries Society, 147(6):1067–1077, November 2018. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weathers:2019:NGP

[WKS^{+}19] T. Casey Weathers, David C. Kazyak, Jay R. Stauffer, Jr., Matt A. Kulp, Stephen E. Moore, Tim L. King, and John E. Carlson. Neutral genetic and phenotypic variation within and among isolated headwater populations of brook trout. Transactions of the American Fisheries Society, 148(1):58–72, January 2019. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Welsh:2013:ECU

[WL13] Stuart A. Welsh and Heather L. Liller. Environmental correlates of upstream migration of yellow-phase American eels in the Potomac River drainage. Transactions of the American Fisheries Society, 142(2):483–491, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weller:2016:SSD

[WLLCF16] J. Daniel Weller, John Paul Leblanc, Arunas Liskauskas, and Patricia Chow-Fraser. Spawning season distribution in subpopulations of muskellunge in Georgian Bay, Lake Huron. Transactions of the American Fisheries Society, 145(4):795– 809, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilson:2016:PGS

[WLW16] Chris C. Wilson, Arunas P. Liskauskas, and Kristyne M. Wozney. Pronounced genetic structure and site fidelity among native muskellunge populations in Lake Huron and Georgian Bay. Transactions of the American Fisheries Society, 145(6): 1290–1302, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warren:2012:QSD

[WM12] Dana R. Warren and Michelle M. McClure. Quantifying salmon-derived nutrient loads from the mortality of hatcheryorigin juvenile Chinook salmon in the Snake River Basin. Transactions of the American Fisheries Society, 141(5):1287– 1294, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wirgin:2014:CSS

[WMG^{+}14] Isaac Wirgin, Lorraine Maceda, Cheryl Grunwald, Nirmal K. Roy, and John R. Waldman. Coastwide stock structure of winter flounder using nuclear DNA analyses. Transactions of the American Fisheries Society, 143(1):240–251, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:2015:EWT

[WMS15] David L. Ward and Rylan Morton-Starner. Effects of water temperature and fish size on predation vulnerability of juvenile humpback chub to rainbow trout and brown trout. Transactions of the American Fisheries Society, 144(6):1184–1191, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:2016:DUT

[WMV^{+}16] Tyler Wagner, Stephen R. Midway, Tiffany Vidal, Brian J. Irwin, and James R. Jackson. Detecting unusual temporal patterns in fisheries time series data. Transactions of the American Fisheries Society, 145(4):786–794, July 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wirgin:2012:SOM

[WMW^{+}12] Isaac Wirgin, Lorraine Maceda, John R. Waldman, Sierra Wehrell, Michael Dadswell, and Tim King. Stock origin of migratory Atlantic sturgeon in Minas Basin, Inner Bay of Fundy, Canada, determined by microsatellite and mitochondrial DNA analyses. Transactions of the American Fisheries Society, 141(5):1389–1398, September 2012. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:2010:REG

[WNCF10] Darren M. Ward, Keith H. Nislow, Celia Y. Chen, and Carol L. Folt. Rapid, efficient growth reduces mercury concentrations in stream-dwelling Atlantic salmon. Transactions of the American Fisheries Society, 139(1):1–10, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winship:2014:FHE

[WOM14] Arliss J. Winship, Michael R. O’Farrell, and Michael S. Mohr. Fishery and hatchery effects on an endangered salmon population with low productivity. Transactions of the American Fisheries Society, 143(4):957–971, July 2014. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitlock:2014:IHC

[WQD14] Steven L. Whitlock, Michael C. Quist, and Andrew M. Dux. Influence of habitat characteristics on shore-spawning Kokanee. Transactions of the American Fisheries Society, 143(6): 1404–1418, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watkins:2014:EST

[WR14] Katherine Shepard Watkins and Kenneth A. Rose. The effects of spatial and temporal resolution in simulating fish movement in individual-based models. Transactions of the American Fisheries Society, 143(5):1143–1160, September 2014.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wirth:2012:RSG

[WRP^{+}12] Lisa Wirth, Amanda Rosenberger, Anupma Prakash, Rudiger Gens, F. Joseph Margraf, and Toshihide Hamazaki. A remote-sensing, GIS-Based approach to identify, characterize, and model spawning habitat for fall-run chum salmon in a sub-Arctic, glacially fed river. Transactions of the American Fisheries Society, 141(5):1349–1363, September 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watkins:2017:RFP

[WRQH17] Carson J. Watkins, Tyler J. Ross, Michael C. Quist, and Ryan S. Hardy. Response of fish population dynamics to mitigation activities in a large regulated river. Transactions of the American Fisheries Society, 146(4):703–715, July 2017. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wippelhauser:2015:SSA

[WS15] Gail S. Wippelhauser and Thomas S. Squiers, Jr. Shortnose sturgeon and Atlantic sturgeon in the Kennebec River System, Maine: a 1977–2001 retrospective of abundance and important habitat. Transactions of the American Fisheries Society, 144 (3):591–601, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waterhouse:2014:RAW

[WSI14] Matthew D. Waterhouse, Brian L. Sloss, and Daniel A. Isermann. Relationships among walleye population characteristics and genetic diversity in Northern Wisconsin lakes. Transactions of the American Fisheries Society, 143(3):744–756, May 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wingate:2011:SPM

[WSK11] Rebecca L. Wingate, David H. Secor, and Richard T. Kraus. Seasonal patterns of movement and residency by striped bass within a subestuary of the Chesapeake Bay. Transactions of the American Fisheries Society, 140(6):1441–1450, November 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watkins:2015:PFA

[WSQ^{+}15] Carson J. Watkins, Bryan S. Stevens, Michael C. Quist, Bradley B. Shepard, and Susan C. Ireland. Patterns of fish assemblage structure and habitat use among main- and sidechannel environments in the Lower Kootenai River, Idaho. Transactions of the American Fisheries Society, 144(6):1340– 1355, November 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wissmar:2010:RDD

[WTB10] Robert C. Wissmar, Raymond K. Timm, and Mason D. Bryant. Radar-derived digital elevation models and fieldsurveyed variables to predict distributions of juvenile Coho salmon and Dolly Varden in remote streams of Alaska. Transactions of the American Fisheries Society, 139(1):288–302, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:2012:BPD

[WvPC12] Carl J. Walters, Brett T. van Poorten, and Lewis G. Coggins. Bioenergetics and population dynamics of flannelmouth sucker and bluehead sucker in grand canyon as evidenced by tag recapture observations. Transactions of the American Fisheries Society, 141(1):158–173, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:2014:BIU

[WZC^{+}14] David M. Walters, Robert E. Zuellig, Harry J. Crockett, James F. Bruce, Paul M. Lukacs, and Ryan M. Fitzpatrick. Barriers impede upstream spawning migration of flathead chub. Transactions of the American Fisheries Society, 143 (1):17–25, January 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wathen:2012:ESB

[WZCT12] Gus Wathen, Joseph Zydlewski, Stephen M. Coghlan, Jr., and Joan G. Trial. Effects of smallmouth bass on Atlantic salmon habitat use and diel movements in an artificial stream. Transactions of the American Fisheries Society, 141(1):174–184, January 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wippelhauser:2015:SSG

[WZK^{+}15] Gail S. Wippelhauser, Gayle B. Zydlewski, Micah Kieffer, James Sulikowski, and Michael T. Kinnison. Shortnose sturgeon in the Gulf of Maine: Use of spawning habitat in the Kennebec System and response to dam removal. Transactions of the American Fisheries Society, 144(4):742–752, July 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yasumiishi:2016:BRH

[Yas16] Ellen Martinson Yasumiishi. Book review: A History of Sockeye Salmon Research, Karluk River System, Alaska, 1880– 2010. By Richard Gard and Richard Lee Bottorff: NOAA Technical Memorandum NMFS-F/SPO-125. 2014. 413 pages. Available for free at the National Marine Fisheries Service, Scientific Publication Office, 7600 Sand Point Way Northeast, Seattle, Washington 98115, USA. Transactions of the American Fisheries Society, 145(6):1399, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yard:2011:TPC

[YCB^{+}11] Michael D. Yard, Lewis G. Coggins, Jr., Colden V. Baxter, Glenn E. Bennett, and Josh Korman. Trout piscivory in the Colorado River, grand canyon: Effects of turbidity, temperature, and fish prey availability. Transactions of the American Fisheries Society, 140(2):471–486, May 2011. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2010:MMA

[YHWH10] Roger G. Young, John W. Hayes, Jeremy Wilkinson, and Joe Hay. Movement and mortality of adult brown trout in the Motupiko River, New Zealand: Effects of water temperature, flow, and flooding. Transactions of the American Fisheries Society, 139(1):137–146, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2012:PSA

[YITI12] Shawn P. Young, Travis R. Ingram, Josh E. Tannehill, and J. Jeffery Isely. Passage of spawning Alabama shad at Jim Woodruff Lock and Dam, Apalachicola River, Florida. Transactions of the American Fisheries Society, 141(4):881–889, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yu:2011:CRS

[YJW11] Hao Yu, Yan Jiao, and Andreas Winter. Catch-rate standardization for yellow perch in Lake Erie: a comparison of the spatial generalized linear model and the generalized additive model. Transactions of the American Fisheries Society, 140 (4):905–918, July 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2010:CEF

[YSC10] Paciencia S. Young, Christina Swanson, and Joseph J. Cech, Jr. Close encounters with a fish screen III: Behavior, performance, physiological stress responses, and recovery of adult delta smelt exposed to two-vector flows near a fish screen. Transactions of the American Fisheries Society, 139(3):713– 726, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yun:2014:SBA

[YSW^{+}14] Sang-Seon Yun, Andras Szeitz, Andrew Wildbill, Michael Siefkes, and David Close. Sulfated bile acids as putative sex pheromone components in Pacific lamprey. Transactions of the American Fisheries Society, 143(6):1455–1459, November 2014. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2016:SSF

[YYA^{+}16] Joy M. Young, Beau G. Yeiser, Erick R. Ault, James A. Whittington, and Jynessa Dutka-Gianelli. Spawning site fidelity, catchment, and dispersal of common snook along the East Coast of Florida. Transactions of the American Fisheries Society, 145(2):400–415, March 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zeigler:2013:UTT

[ZBC^{+}13] Matthew P. Zeigler, Stephen F. Brinkman, Colleen A. Caldwell, Andrew S. Todd, Matthew S. Recsetar, and Scott A. Bonar. Upper thermal tolerances of Rio Grande cutthroat trout under constant and fluctuating temperatures. Transactions of the American Fisheries Society, 142(5):1395–1405, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zelasko:2016:ELT

[ZBHW16] Koreen A. Zelasko, Kevin R. Bestgen, John A. Hawkins, and Gary C. White. Evaluation of a long-term predator removal program: Abundance and population dynamics of invasive northern pike in the Yampa River, Colorado. Transactions of the American Fisheries Society, 145(6):1153–1170, November 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zelasko:2010:SRM

[ZBW10] Koreen A. Zelasko, Kevin R. Bestgen, and Gary C. White. Survival rates and movement of hatchery-reared razorback suckers in the Upper Colorado River Basin, Utah and Colorado. Transactions of the American Fisheries Society, 139(5): 1478–1499, September 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:2013:EES

[ZC13] Yuying Zhang and Steve X. Cadrin. Estimating effective sample size for monitoring length distributions: a comparative study of Georges Bank groundfish. Transactions of the American Fisheries Society, 142(1):59–67, January 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ziskin:2011:ICO

[ZHWR11] Gabriel L. Ziskin, Patrick J. Harris, David M. Wyanski, and Marcel J. M. Reichert. Indications of continued overexploitation of speckled hind along the Atlantic Coast of the Southeastern United States. Transactions of the American Fisheries Society, 140(2):384–398, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:2015:ASL

[ZKB^{+}15] Mara S. Zimmerman, Clayton Kinsel, Eric Beamer, Edward J. Connor, and David E. Pflug. Abundance, survival, and life history strategies of juvenile Chinook salmon in the Skagit River, Washington. Transactions of the American Fisheries Society, 144(3):627–641, May 2015. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zuray:2012:SCI

[ZKH12] Stanley Zuray, Richard Kocan, and Paul Hershberger. Synchronous cycling of ichthyophoniasis with Chinook salmon density revealed during the annual Yukon River spawning migration. Transactions of the American Fisheries Society, 141 (3):615–623, May 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zweifel:2010:DEB

[ZLHS10] Richard D. Zweifel, Andrew M. Gascho Landis, R. Scott Hale, and Roy A. Stein. Development and evaluation of a bioenergetics model for saugeye. Transactions of the American Fisheries Society, 139(3):855–867, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:2010:SBB

[ZM10] Chang Ik Zhang and Bernard A. Megrey. A simple biomassbased length-cohort analysis for estimating biomass and fishing mortality. Transactions of the American Fisheries Society, 139(3):911–924, May 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:2016:FIS

[ZRM^{+}16] Hongyan Zhang, Edward S. Rutherford, Doran M. Mason, Jason T. Breck, Marion E. Wittmann, Roger M. Cooke, David M. Lodge, John D. Rothlisberger, Xinhua Zhu, and Timothy B. Johnson. Forecasting the impacts of silver and bighead carp on the Lake Erie food web. Transactions of the American Fisheries Society, 145(1):136–162, January 2016. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:2013:SLH

[ZSVK13] Christian E. Zimmerman, Heidi K. Swanson, Eric C. Volk, and Adam J. R. Kent. Species and life history affect the utility of otolith chemical composition for determining natal stream of origin for Pacific salmon. Transactions of the American Fisheries Society, 142(5):1370–1380, September 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zeigler:2012:ERC

[ZTC12] Matthew P. Zeigler, Andrew S. Todd, and Colleen A. Caldwell. Evidence of recent climate change within the historic range of Rio Grande cutthroat trout: Implications for management and future persistence. Transactions of the American Fisheries Society, 141(4):1045–1059, July 2012. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zydlewski:2011:REE

[Zyd11] Gayle Barbin Zydlewski. A review of “Eels at the Edge: Science, Status, and Conservation Concerns” Edited by John M. Casselman and David K. Cairns. American Fisheries Society, Bethesda, Maryland. 2009. 449 pages. $69 ($48 for AFS members). Transactions of the American Fisheries Society, 140 (2):526–527, May 2011. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zydlewski:2010:DII

[ZZD10] Joseph Zydlewski, Gayle Zydlewski, and G. Russell Danner. Descaling injury impairs the osmoregulatory ability of Atlantic salmon smolts entering seawater. Transactions of the American Fisheries Society, 139(1):129–136, January 2010. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhu:2013:LFA

[ZZMC13] Xinhua Zhu, Yingming Zhao, Alastair Mathers, and Lynda D. Corkum. Length frequency age estimations of American eel recruiting to the Upper St. Lawrence River and Lake Ontario. Transactions of the American Fisheries Society, 142(2):333– 344, March 2013. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).